A Complete Bibliography of Publications in the Journal of Aquaculture Research & Development

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

10 April 2024 Version 1.13

Title word cross-reference

17α [85, 59]. 17β [62]. 20β [85]. ^{+} [365]. ^{3} [448]. _{2} [81, 130]. _{96h} [156]. a [77]. α [34, 145]. β [148, 153, 145]. ω3 [555]. ω6 [555]. × [347, 323, 119].

-ATPase [365]. -Carvon [127]. -Deoxynivalenol [448]. -Dihydroxyprogesterone [85]. -encapsulated [237]. -Estradiol [62]. -Glucan [148, 153]. -Methyl [59].

/-Carvon [127]. /K [365].

08^{o} [516].

1 [97, 145, 357]. 13-Cis-Retinoic [266]. 16S [177, 49]. 17- [34].

2 [136]. 25.327’E [516].

3 [145, 242]. 31.912’N [516]. 318 [516].

4-Trizole [88].

71 [346]. 78^{o} [516].

Abaya [226]. Abdominal [590, 514]. Aberrations [565]. Ablation [60, 332]. Abnormal [312]. Abnormalities [344]. Absorption [153]. Abundance [461, 504, 319, 293]. Accelerate [227]. Acclimation [386]. Acclimatization [374]. Accumulation [416]. Acephate [95]. Acid [257, 359, 1, 163, 37, 466, 473, 437, 6, 84, 400, 266, 483, 47, 79, 193, 660, 555, 237, 14, 12]. Acid-Induced [266]. Acidolysis [427]. Acidolysis-Oxidized [427]. Acids [385, 437, 12]. Acipenser [81, 85, 62, 76, 108, 210]. acrassipes [429]. Acting [18]. Action [608]. Activate [175]. Activated [368]. Active [418]. Activities [80, 103, 191, 26, 743, 131, 274]. Activity [115, 335, 573, 114, 338, 105, 483, 194, 365]. acts [9]. Acute [375, 708, 170]. Adamawa [442]. Adaptability [25]. Adaptations [609]. Added [54, 176]. Additional [543, 223]. Additions [388]. Additive [148, 708, 97]. Additives [113, 235, 717, 111]. Administration [531, 418, 9]. Adult [555, 133, 288]. Adults [469]. Advance [101]. Advanced [264]. Advancement [117]. Advancements [674, 670]. Advances [135, 652]. Advantages [643, 592]. Aegean [317]. Aerated [130]. Aeromonas [456, 537, 544, 406, 225, 33, 459, 677, 261, 398, 121, 294, 432, 105, 24, 49, 237]. aeruginosa [225, 571]. aestivum [527]. aestuarii [96]. Afar [297]. Affect [600]. Affecting [256, 226]. affects [526]. Aflatoxin [148]. African [370, 140, 503, 459, 397, 456, 137, 350, 436, 434, 395, 35, 717]. After [494, 171, 279]. Against [459, 1, 282, 398, 431, 148, 32, 175, 204, 292, 501, 488, 274, 237]. Age [168]. Ageing [496]. Agent [282, 315]. Agents [725]. AgNPs [521]. Agusan [505]. AI [651]. Ailments [612]. Air [391, 279]. akaara [441]. Al [439]. Al-Faw [439]. Alanine [88]. Albertine [554]. Alfalfa [71]. Algae [392, 191, 275, 322, 99]. Algal [396, 593, 100, 631]. alginolyticus [175]. Alkaline [149]. Allometric [326]. Allowed [362]. Alma [77]. Aloe [333]. Along [563, 232]. Alpha [204, 1]. Alpha-Tocopherol [204]. alpinus [115, 114]. Alterations [383, 346]. Alternative [535, 492, 522, 110]. Alum [370]. alvarezii [715, 454, 431]. AMAs [471]. Amblypharyngodon [156]. Ameiurus [124]. American [573, 274]. americanum [60]. Amhara [174]. Amines [78]. Amino [257, 359, 466, 473, 14, 483, 193, 385]. Ammonia [368, 176, 714]. Ammonia-Oxidizing [714]. amoA [177]. among [509, 80, 236, 125, 356]. Amphiprion [366]. Amplified [236]. AMPs [522]. Amylase [115]. Anabas [346]. Anaemia [11]. Anaerobic [410]. Anal [173]. Analysing [697]. Analysis [80, 120, 239, 5, 616, 265, 634, 385, 446, 327, 138, 180, 627, 264, 305, 380, 321, 443, 281, 177, 645, 650, 377, 203, 144]. Analytical [606]. Analyze [216]. Analyzing [611, 500, 494]. Anatomo [727]. Anatomo-Pathological [727]. andersonii [59]. Andhra [149, 525]. Anemonefish [366]. Anesthesia [127]. Anesthetic [8, 366]. Angel [329]. Angler [305]. Angling [703]. anguillicaudatus [485]. Anhydride [605]. Animal [539, 31, 620, 521, 627, 68, 646, 40]. Animals [682, 603, 482]. Anomalous [154]. Anterior [284]. Anti [367]. Anti-pangas [367]. Antibacterial [282, 274]. Antibiofilm [353]. Antibiotic [692, 689, 501, 24]. Antibiotics [522, 206]. Antigenic [317]. Antimicrobial [17, 33, 191, 522, 121, 136, 131]. Antimicrobials [155]. Antioxidant [335, 4, 437, 338, 194, 485]. Antioxidants [468, 191]. antiquata [344]. Antiviral [725]. Aphanomyces [294]. Apostichopus [327]. Apparent [157, 359, 306]. Applicability [368]. Application [201, 506, 440, 41, 438, 394]. Applications [521, 672, 684, 199, 723, 532]. Approach [283, 606, 625, 648, 380, 440]. Approaches [617, 622, 615, 229]. Aqua [190, 342, 101, 524]. aqua-chemicals [524]. Aquaculture [531, 296, 545, 599, 692, 155, 453, 120, 303, 535, 662, 90, 640, 651, 17, 608, 639, 623, 512, 673, 606, 725, 728, 167, 655, 637, 358, 491, 720, 550, 141, 513, 689, 498, 304, 601, 636, 724, 722, 166, 569, 418, 598, 10, 82, 369, 341, 378, 679, 628, 649, 107, 117, 285, 474, 627, 721, 614, 604, 709, 435, 172, 731, 20, 730, 46, 618, 26, 688, 629, 661, 529, 551, 671, 199, 510, 124, 654, 726, 486, 729, 423, 611, 723, 681, 676, 286, 694, 438, 660, 394, 600, 631, 351, 494, 64, 665, 532, 680, 607]. Aquaculture [701, 658, 633, 635, 508, 229, 630, 471]. Aquacultured [208]. Aquaculturists [159]. Aquafeed [479]. aquamarina [99]. Aquaponic [228]. Aquaponics [554, 549, 407]. Aquariums [599]. Aquatic [352, 58, 596, 608, 620, 621, 580, 581, 568, 582, 521, 682, 683, 627, 740, 595, 202, 88, 615, 696, 607, 482]. Aquatic(R) [292]. Aqueous [32]. Arabian [134]. Arabinoxylan [137]. Arabinoxylan-Oligosaccharides [137]. Arachidonic [163, 400, 12]. arameters [8]. Arapaima [179]. Arctic [115, 114]. arcuata [458]. Area [120, 334, 192, 45, 216, 547]. Areas [471, 446, 475, 477, 525]. areolata [102, 234, 233]. Argo [686]. Argopecten [84]. Argyrosomus [328, 134, 437]. Aristaeomorpha [361]. armatus [337]. aromatica [487]. Arsedi [389]. Arsenic [605, 625, 648]. Artemia [61, 164, 275]. Arthrospira [457]. Artificial [392, 543, 25, 710, 350, 214, 71, 36]. Artificially [562]. Artrazine [288]. Asa [565]. Ascanius [44, 29]. Asian [425, 31, 91, 720, 28, 27, 483, 36, 40]. asiatica [432]. Aspect [160]. Aspects [34, 641, 666]. Assam [236]. Assay [467]. Assemblages [122]. Assessed [236]. Assessing [28, 411]. Assessment [157, 503, 451, 604, 538, 399, 361, 174]. Associated [409, 727]. Astaxanthin [247]. Asterales [506]. Asteroids [103]. Aswan [200]. Atlantic [403, 39, 167, 250, 360]. Atmospheric [374]. ATPase [365]. Attenuation [275]. Attractor [73]. Attributes [208]. Atyidae [430]. Atypical [23]. auranticum [318]. aurata [590, 495]. auratus [37, 8, 138, 161, 246]. Austin [300]. Australia [20]. Australian [19].

Austrocambarus [241]. Automatic [239, 316]. Available [126, 557, 172]. Azolla [452, 71]. azoricus [374].

B1 [148]. Babylon [102, 234, 233]. Babylonia [102, 234, 233]. bachiata [286]. Bacillus [13, 529, 551, 526, 300, 50, 330, 466, 473, 38]. Background [508]. Backwater [348]. Backwaters [118]. Bacteria [410, 624, 137, 336, 247, 84, 294, 714, 177, 274, 41, 269, 660]. Bacterial [367, 150, 411, 422]. Bacteriological [446, 291]. Bagrid [188]. Bagridae [287]. Bagrus [258]. Bai [307]. baker [9]. Balance [141, 437]. Balanced [311]. Balancing [671]. Ballan [44]. Baltic [686]. Baluchestan [401]. Bamboo [176]. Ban [399]. Bangiales [206]. Bangladesh [515, 326, 358, 261, 446, 408, 267, 393, 43, 381, 399, 449, 423, 280, 371, 500, 542, 524]. Banyuwangi [504]. Barb [320, 278]. barila [16]. Barilius [16]. Bark [189]. Barley [558]. Barramundi [371]. Based [313, 157, 404, 6, 707, 719, 306, 380, 346, 553, 486, 168]. Basic [525]. Basil [338]. basilicum [338]. basin [505]. Basra [382]. Basrah [439]. Bass [590, 248, 73, 5, 27, 249, 317]. Batch [137]. Bathymodiolus [374]. Bati [174]. Batie [476]. batrachus [31, 91, 279, 36, 40]. Bay [340, 450, 422]. bayad [258]. be [198]. Bean [313, 263, 271]. Bean-Based [313]. bebe [511]. Bee [430]. Beel [515]. Beetroot [594]. Before [494]. Behavior [443]. Behaviour [330]. Beluga [14]. Benefits [726, 681, 530, 588]. Benthic [363]. bergylta [44, 29]. Besar [334]. Best [512]. Beta [95, 182, 594]. Beta-cypermethrin [95]. between [77, 245, 45, 467]. Bheri [186]. bicolor [255]. BIG [541]. Bigelowiella [732]. Bigeye [488]. Binders [472]. binghamiae [175]. Bio [2, 238, 606]. Bio-indicator [238]. Bio-Inoculation [2]. BioAcid [364]. Bioactive [353, 28, 131]. Bioaugmentation [315]. Biochar [723, 694]. Biochar-Enhanced [694]. Biochemical [468, 396, 225, 700, 74, 94, 338, 71, 208, 70, 111, 364, 330]. Biochemistry [425, 383, 189]. Biocides [292]. Biodiesel [130]. Biodiversity [643, 25, 732, 704, 293, 739]. Bioeconomic [221]. Biofilms [411]. Biofloc [224]. Biogenic [78]. Biogeochemical [169]. Biological [113, 112, 268, 134, 263, 94, 632, 379, 695, 356, 186]. Biologically [165]. Biology [586, 252, 337, 433, 163, 106, 302]. Biomarkers [468]. biomass [631]. Biomedicine [17]. Biometric [79]. Biomolecule [377]. Biosecurity [491]. Biosensor [303]. Biota [595]. Biotechnology [67, 690, 676]. Bira [174]. Bitter [252]. Bivalve [458]. Bivalves [165]. Bizerte [344, 340]. BKD [150]. Black [309, 291, 501, 311, 147, 168, 305]. Black-Lipped [501]. Black-Motley [305]. Bleeker [505]. Blended [31]. Bloch [27, 371]. Blood [323, 79, 299]. Bloom [593]. Blooms [198]. Blue [667, 666, 68]. Bluespotted [219]. board [381]. Bodies [575, 578, 595]. Body [425, 31, 74, 324, 267, 63, 71, 87, 143, 246, 15]. Bone [313]. Bonylip [278]. Boon [706, 718]. Boone [214, 99, 30]. boost [541]. Borne [167]. borneensis [505]. Both [520, 183, 603, 482]. Bottom [103]. Boulanger [572]. Boulenger [700, 306]. Box [242]. Brachyura [259]. Brackish [214, 260, 184, 561]. Brain [507]. Branches [616]. Brand [321]. Brazil [388, 265, 273]. Bread [527]. Bream [590, 495]. Breathing [279]. Breeding [188, 91, 679, 35, 541, 381, 184, 36]. Breeds [612]. Brief [586, 534]. British [7]. Broad [263]. Bronopol [292]. Brooding [184]. Broodstock [217, 257, 256, 55, 163, 227, 289]. Broodstocks [386]. Brook [150]. Brycinus [511]. Buch [154]. Buguma [260]. Bulls [305]. Buoyancy [472]. Burchell [397, 121, 35, 189, 70, 251]. Burkholderia [409]. Businesse [635]. Butuan [505]. By-Product [234, 299]. Bycatch [716]. Byproducts [54].

C [61, 21, 37, 138, 145]. C. [138]. C/EBP [145]. Ca [313]. Ca/P [313]. Cadmium [468, 355]. Cage [389, 528, 512, 711, 523, 722, 267, 488]. Cages [455, 301]. cainii [300]. Cake [243]. calcarifer [425, 27, 483, 371]. calcitrans [269]. Calcium [62, 74]. Caledonia [169]. California [510]. Caligus [256, 223, 144]. Calimere [83]. Calorimetric [536]. Cambarellus [463]. Camellia [425]. Cameroon [476, 442, 727]. Camphor [366]. Campus [496]. can [198]. Canada [209]. Canals [372]. Candidate [285, 64]. Canola [19, 27]. Cans [69]. cantonensis [430]. Capacity [283, 4, 485]. capillaceaas [222]. Captive [188, 184]. Captivity [92, 572]. Carassius [8, 37, 138, 161]. Carbohydrate [211, 324, 143, 181]. Carbohydrate/Lipid [324]. Carbon [368, 473, 391, 667]. Carbonate [281, 307]. Carboxymethyl [153]. Carcass [700]. Carcinus [96]. Cardinal [184]. Caridina [430]. Caring [482]. carota [594]. Carotenoids [347, 323]. Carp [148, 526, 140, 211, 28, 659, 284, 290, 71, 322, 127, 496]. carpio [148, 526, 140, 413, 322, 127, 237]. Carrageenan [715]. Carrassius [246]. Carrot [594]. Carrying [283]. Cartilage [160]. Carvon [127]. Case [509, 417, 564]. Cases [515]. Caspase [105]. Caspase-3 [105]. Caspian [53, 139]. caspica [414, 139]. Castelnau [59]. Cat [236, 294]. Catarina [388]. Catch [103, 703, 272]. Catching [482]. Categorization [524]. Catfish [367, 80, 370, 140, 257, 287, 188, 197, 347, 31, 503, 91, 459, 397, 456, 253, 137, 224, 196, 19, 121, 350, 436, 434, 395, 35, 136, 572, 124, 36, 40, 229]. Catla [460, 151, 460, 154]. Caudal [173]. Caught [73, 482, 489]. Causative [282]. Caused [88]. Causes [7, 663]. Caviar [108]. cDNAs [203]. Ceara [265]. Cell [5]. Cellana [163]. Cells [357]. Cellulose [153]. Centella [432]. Center [631]. Central [388, 489, 122, 363, 464]. cepacia [409]. cephalus [383]. cereus [466, 473]. cerevisiae [9]. Cestoda [134]. Chabahar [401]. Chaetoceros [269]. Chain [354, 254]. Challenge [398]. Challenged [294]. Challenges [491, 720, 721, 685, 699]. Change [736, 358, 667, 272]. Changes [562, 225, 62, 457, 279, 221, 143, 615, 139]. Chankanay [111]. Channa [219]. Channel [258, 224, 229]. Channidae [219]. Chanos [445, 556]. Characterisation [317]. Characteristic [270]. Characteristics [563, 462, 634, 700, 497, 632, 618, 26, 147, 542]. Characterization [158, 409, 65, 373, 430, 571, 408, 136, 203, 144]. Charcoal [499, 176]. Charr [114, 115]. Charybdis [194, 193]. Chemical [462, 700, 192, 420, 111, 661, 174]. Chemicals [429, 524]. Chenab [312].

Cherax [300, 126, 65]. Chi [321]. Chile [628, 649]. Chilling [391, 208]. China [455, 444, 440, 426, 635, 356]. Chinook [710]. Chitin [203]. Chloramine [292]. Chloramine-T [292]. Chlorella [4, 447]. Chloride [468, 298, 355]. Chlorophyll [77]. Chlorophyta [497]. Choice [342]. Chromolaena [282, 506]. Chronic [708, 94, 170, 355]. Chum [351]. Chytridiomyces [23]. Cichlasoma [323]. Cichlidae [442]. Cichlids [478]. Cinnamomum [487]. Cirrhinus [460, 645, 650]. Cis [266]. Citalopram [75]. Citharinus [308, 309]. citharus [308, 309]. citrinellum [323]. Citrus [318]. City [439, 382, 445, 505, 509, 321]. Clam [363]. Clarias [271, 370, 313, 140, 217, 257, 288, 218, 197, 347, 31, 91, 459, 397, 231, 700, 279, 121, 350, 436, 434, 35, 189, 717, 87, 70, 572, 251, 36, 40]. Clariids [565]. Clarius [395]. Class [555]. Classification [619]. Clea [92]. Cleaning [450]. Climate [736, 358, 272, 500]. Climatic [667]. Clinical [150]. Clinostomum [226]. Closed [726]. Closed-Loop [726]. Closer [245]. Clove [8, 366]. Clown [366]. Clupisoma [380]. Co [212, 417, 237, 130]. Co-Infection [212]. Co-management [417]. Coagulation [661]. Coast [103, 240, 86, 165, 89, 106, 149, 83, 104, 516, 64, 101, 123, 293, 168, 403, 207, 272]. Coastal [120, 7, 334, 283, 326, 738, 513, 523, 446, 475, 477, 735, 10, 622, 272, 618, 510, 702, 371, 678, 524, 542]. Coasts [317]. Coated [69]. Cocos [288]. Cocultured [454]. Cod [39]. Cohabitation [133]. coioides [171]. Cold [189]. coli [377]. Collagen [462]. Collapse [245]. Collected [340, 516]. Color [327]. Coloration [247]. Colouration [161]. Columbia [7]. Column [173]. columnare [354]. Combination [569]. Combined [330, 87]. Commercial [80, 235, 499, 570, 555]. Common [148, 526, 140, 224, 659, 322, 124, 127, 663]. Communities [513, 363, 202, 702, 530, 422]. Community [334, 314, 499, 373, 245, 18, 118, 508]. Comparative [308, 81, 39, 232, 446, 372, 348, 255, 436, 305, 478, 177, 210, 645, 650]. Comparison [318, 124, 504, 467, 133]. Compartmental [739]. Competency [353]. Complex [344, 173, 483]. Component [12]. Components [527, 269]. Composition [476, 425, 244, 4, 192, 437, 102, 324, 267, 350, 63, 71, 412, 87, 118, 123, 555, 74]. Compositions [416]. Compost [100]. Compounds [353, 224, 131]. Computer [443]. Concentration [77, 112, 268, 473, 165, 78]. Concentrations [605, 340, 189, 365]. Conceptual [623, 307]. Concomitant [256]. Concreate [493]. Concrete [219, 631]. Condition [511, 188, 186, 357, 464, 504, 438]. Conditions [590, 183, 102, 338, 346, 600, 181]. Conferred [459]. Confinement [366]. Congener [73]. Connection [702]. Consequences [742]. Conservation [220, 433, 732, 240, 53, 685, 669]. Consideration [508]. Constant [434]. Constituents [361]. Construction [90, 658]. Consumption [509, 42, 28]. Containing [225, 167, 19, 97, 479]. contamination [291]. Content [63, 447, 412, 715, 143, 286, 485]. Contents [211, 123, 299]. Context [449]. Contracaecum [226]. Contribute [145]. Contributing [731].

Contribution [679]. Control [535]. Controlled [91, 490]. Controlling [429]. Controverted [550]. Conventional [211, 318, 560]. Cooking [233]. Coordination [245]. Copepoda [223]. Copepods [502]. Copper [375, 39, 279, 215]. Copra [93]. Copra-Derived [93]. Coral [643, 12, 669, 739]. Core [281, 356]. Coreau [265]. Corn [157, 102, 6, 143]. Correlation [422]. Cost [52, 311]. costata [414]. Costs [216]. Cottonseed [536, 55]. Count [241]. County [528]. Coupled [207]. Course [608, 365]. COVID19 [494]. COX [400]. CpG [237]. CpG-oligonucleotide [237]. Crab [162, 332, 227, 289, 540, 400, 266, 193, 259, 123]. Crablet [162]. Crabs [194]. crassipes [428]. Crayfish [277, 241, 65, 541]. Creeks [523]. crenulata [32]. Crescent [86]. crispus [333]. Critically [380, 320]. Croaker [385, 455]. crocea [385, 455]. Cross [291, 168]. Cross-contamination [291]. Crucial [209]. Crude [447, 472]. Crustacea [96, 516]. Crustacean [64]. Crustaceans [480]. Cryobanking [608]. Cryobiology [733, 617, 652, 679]. Cryogenic [391]. Cryopreservation [617, 652]. Cryoprotectants [652]. Ctenopharyngodon [284, 290, 71]. cuchia [451]. Cucumber [25, 327]. Cucumbers [412]. Cuddalore [706, 718, 89, 106]. Cultivated [672]. Cultivation [164, 506, 435, 490]. Cultural [508]. Culture [389, 130, 66, 283, 642, 492, 404, 711, 137, 497, 192, 401, 275, 180, 735, 94, 240, 267, 452, 214, 379, 465, 48, 43, 504, 496, 525, 269, 438, 712, 633, 670]. Cultured [113, 39, 390, 146, 56, 455, 446, 214, 727, 23, 426, 537]. Cultures [100]. Cum [699]. cumini [559]. curcas [98]. Current [528, 331, 141, 336]. Curries [69]. Customer [482]. Cuvier [520, 571, 366]. cuvieri [138]. Cyanobacterial [377]. Cycle [332, 259]. Cycling [714]. cypermethrin [95]. Cyprinidae [278]. Cyprinus [148, 526, 140, 413, 322, 127, 237]. Cytochrome [138]. Cytotoxicity [191].

D [237]. Dam [442, 322, 174]. Dam/Duhok [322]. damselae [256]. Danio [145]. Daniorerio [396]. Daphnia [262]. Darbhanga [496]. Data [239, 494]. Date [230]. Daucus [594]. Days [399]. Debate [550]. Decapod [516]. Decapoda [96, 516, 430, 259]. Deccan [47]. Decreases [95]. Decreasing [241]. Deep [455, 516]. Deep-Water [455]. Deepwater [705]. Deformity [590]. Degradation [475, 477]. Delegate [607]. Delineation [395]. Delta [260, 372]. Densities [404, 92, 94, 490]. Density [512, 56, 267, 128]. Deoxynivalenol [448]. Dependent [205]. Deployment [190]. Deposition [427, 149]. Depuration [458, 3]. Derived [93, 131]. Design [563, 190, 435]. Detailed [589, 588, 593]. Detection [367, 207, 195, 354, 467, 658, 229]. Determinants [20, 564]. Determination [519, 211, 168]. Determining [108]. Detoxification [55]. Developed [590]. Developing [209]. Development [367, 471, 662, 201, 492, 404, 91, 623, 620, 253, 42, 498, 659, 119, 173, 435, 132, 316, 29, 311, 685, 251, 187, 206, 668, 484, 508, 369, 378]. Developments [545, 67, 474]. Deviations [525]. DHA [42]. Diabetic [539]. diabolicus [374]. Diagnositic [11]. dibrugarensis [188]. Dicentrarchus [368, 590, 73, 5, 256, 249, 317]. Dicer [696]. Diet [313, 300, 511, 157, 396, 66, 92, 225, 620, 231, 163, 724, 6, 560, 214, 71, 70, 27, 147, 47, 485]. Dietary [115, 343, 313, 526, 425, 287, 662, 277, 427, 347, 31, 335, 700, 499, 19, 37, 205, 74, 222, 457, 102, 233, 324, 333, 604, 298, 717, 48, 143, 572, 483, 210, 98, 40, 181]. Diets [520, 329, 708, 387, 235, 518, 244, 553, 19, 97, 102, 234, 233, 68, 63, 172, 306, 556, 271, 30, 555, 142]. Differences [455]. Different [126, 295, 559, 347, 404, 383, 162, 250, 499, 37, 74, 323, 324, 463, 72, 68, 215, 434, 63, 71, 447, 35, 87, 364, 386, 210, 22, 15, 104, 555, 142, 91]. Differentiating [360]. Differentiation [138]. Digestibility [115, 157, 359, 243, 430, 306, 181]. Digestive [430, 114, 151, 483, 210]. Digeston [97]. Digeston-1 [97]. Dihydroxyprogesterone [85]. Diketopiperazines [38, 84]. Diluted [35]. Dimension [371]. Dimensions [622, 653]. Dioxide [391]. Direction [528]. Discharge [41, 99, 269]. Discriminant [380]. Discriminate [356]. Disease [535, 150, 282, 506, 333, 96, 161, 529, 551, 431]. Diseased [406, 121]. Diseases [612, 486, 663]. Disruptor [32]. Dissemination [513]. Dissolved [639]. Distillation [318]. Distribution [476, 122, 448, 185, 83, 104, 319, 422, 524]. Distributional [103]. District [120, 334, 552, 706, 718, 558, 525, 154, 358, 504]. Diversification [668]. Diversity [638, 552, 697, 451, 740, 741, 496, 645, 650, 104, 13]. DLEC [5]. Dloop [395]. DNA [395]. docmac [258]. Documentation [516]. Does [384]. Domestic [14]. Domesticated [73]. Domestication [219]. Done [198]. Donor [659]. Dorsal [173, 63]. Dosages [434]. Dose [227, 205]. Doses [383, 289, 35]. Drain [200]. Draining [312]. Dried [309]. Drinking [462]. Drinks [462]. Dropsy [514]. Drug [531]. Drugs [418]. Dry [359]. due [612]. Duhok [322]. Dunaliella [130]. Duration [459]. During [374, 384, 289, 133, 179, 710, 42, 487]. Dust [384]. Dynamic [606].

E. [119]. EA [219]. Early [91, 42, 324, 484]. Earth [626]. Earthen [520, 56]. East [471, 334, 220, 86, 412, 165, 89, 106, 525, 83, 118, 64, 101, 123, 293, 177, 216]. eastern [524, 207, 486]. EBP [145]. Eco [429, 675, 435]. Eco-Friendly [429, 435]. Ecological [134, 681, 607]. Ecology [641, 348]. Economic [109, 512, 725, 523, 742, 267, 444, 671, 449, 737, 695, 216, 508, 541]. Economics [515, 358, 43]. Economy [597]. Ecosystem [638, 621]. Ecosystems [643, 675, 735, 13, 510, 104, 680, 669, 608]. Ectoine [262]. Ectoparasite [414]. Ecuador [57]. Ecuadorian [498]. Eddy [634]. Edible [165, 266, 259]. Editor [369, 378]. Editorial [579, 583, 567, 584, 534]. Education [631]. Edwardsiella [432]. Eel [337, 573, 451]. Eels [319]. Effect [392, 115, 425, 389, 287, 85, 248, 396, 171, 390, 347, 241, 92, 397, 458, 677, 244, 442, 224, 163, 421, 74, 75, 437, 228, 415, 479, 72, 294, 363, 350, 709, 298, 71, 208, 357, 487, 48, 105, 149, 495, 200, 22, 47, 286, 187, 431, 485, 555, 98, 182, 365, 181, 546]. Effective [178, 52, 693, 311, 602, 362, 678]. Effectiveness [514, 417]. Effects [8, 343, 262, 61, 476, 375, 468, 368, 217, 257, 60, 238, 374, 657, 66, 225, 204, 557, 56, 231, 328, 480, 19, 37, 205, 457, 94, 102, 323, 324, 591, 683, 333, 644, 267, 215, 612, 434, 63, 447, 26, 717, 472, 366, 465, 488, 176, 529, 551, 128, 572, 127, 364, 483, 170, 325, 611, 663, 79, 88, 469, 299]. Efficacy [148, 32, 601]. Efficiency [543, 368, 557, 328, 52, 636, 159, 440, 98, 40, 635]. Efficient [206]. Effluent [21, 233, 379, 286]. Effluents [169]. Egg [241, 685]. Eggs [292, 434, 83]. Egypt [252, 258, 397, 372, 125, 200]. Egyptian [458, 223]. Eichhorni [429]. Eichhornia [428]. Ekiti [159]. El-Sail [200]. Elasmobranch [122]. Electrolyzed [199]. Electron [284, 290]. Elemental [356]. Eleutheronema [507, 517]. Eleyele [420]. Eliminate [3]. Elimination [502]. elongatus [256]. Embryogenesis [42]. Embryonic [253, 187]. Embryos [76]. Emerging [304]. Emphasis [433, 320]. EM(R) [368]. Emulating [66]. Emulsion [127]. encapsulated [237]. Encoding [203]. Endangered [380, 320]. Endemic [164]. Endocrine [32, 517]. Endogenous [727]. Endoparasites [445]. Energy [359, 700, 63, 15, 30]. Enhanced [694]. Enhancement [335, 728, 466, 473, 202]. Enhancing [691, 703, 644, 688, 680]. Enriched [61, 275, 145, 47]. Enrichment [184]. Ensures [512]. Enteromorpha [392]. Environment [596, 25, 336, 240, 495, 202, 546]. Environmental [283, 265, 397, 401, 475, 477, 245, 713, 628, 649, 125, 671, 726, 615, 600, 422, 696, 10]. Environmentally [141, 605]. Environments [348, 740]. Enzymatic [210, 460]. Enzyme [543, 430, 114, 151, 483]. Enzymes [70]. Epalzeorhynchos [255]. Epinephelus [171, 324, 119, 441, 270]. Episode [96]. Epizootiological [248]. Equipped [303]. erectus [365]. Escape [249]. Escherichia [377]. esocinus [312]. Essential [359, 330, 466, 473, 318, 245, 366, 487]. Establish [177]. Establishment [249]. Estimate [283]. Estimates [639]. Estradiol [62]. Estuary [265, 706, 718, 165, 393, 657]. Ethiopia [489, 389, 140, 297, 218, 226, 527, 174, 564, 428]. Eugenia [487]. Eukaryotic [373]. Eurasian [115, 114, 15]. European [368, 590, 248, 5, 256, 249]. Eutrophicated [422]. Evaluation [309, 31, 183, 710, 93, 620, 112, 606, 594, 345, 292, 330, 230, 263, 222, 3, 628, 649, 68, 360, 132, 316, 18, 254, 631]. Evidence [209]. Evolution [730]. Ex [250]. Ex-vivo [250]. Examining [695]. Excel [346]. Exchange [56, 167, 94, 45, 315]. Excretion [448]. Existence [577, 488]. Exogenous [217]. Expenditures [15]. Experimental [66, 706, 718]. Exploring [738, 690, 705, 672, 682, 741, 674, 663]. Export [291]. Exposed [288, 383, 189, 346]. Exposure [383, 250, 605, 279, 170, 355]. Expressed [136]. Expression [343, 5, 179, 427, 479, 507, 105, 136, 203, 133]. Expressions [424]. Extension [391]. Extensive [642, 504]. External [360]. Extract [32, 288, 175, 434, 556, 100]. Extracted [191]. Extracts [352, 308, 462, 91, 189, 394]. Extruded [536, 63]. Eyed [292]. Eyestalk [60, 332]. Eyestalks [266].

f [559]. Fab. [120]. Fabricius [295, 282, 506, 52, 291, 194, 259, 101]. Facilitates [212]. Facilities [633]. Factor [511, 248, 464]. Factors [708, 640, 179, 145, 731, 727, 517, 695, 422]. Facultative [573]. Fall [710].

Family [390]. Farm [651, 169, 379, 128, 423, 260]. Farmed [2, 382, 108, 45, 360, 544, 317]. farmers [541, 301]. Farming [331, 265, 642, 691, 706, 718, 52, 523, 583, 342, 713, 629, 481, 699, 315, 613, 371, 694, 668, 216, 101, 482, 303, 524]. Farming/Catching [482]. Farms [7, 503, 486, 502, 482]. fasciata [392]. Fast [391]. Fat [63, 447]. Fate [448]. Fatty [37, 437, 6, 14, 47, 12, 555]. Fauna [575]. Faw [439]. Feasibility [523, 180, 76, 69]. Features [134]. Fecundity [259, 505]. Fed [520, 313, 300, 126, 186, 235, 553, 700, 250, 19, 59, 556, 87, 271, 70, 147, 210, 555, 142, 427]. Feed [392, 543, 148, 708, 58, 235, 536, 557, 328, 506, 230, 421, 211, 97, 457, 324, 610, 298, 472, 161, 111, 488, 176, 362, 676, 47, 36, 485, 98, 142, 40]. Feeder [316]. Feeding [61, 389, 258, 337, 433, 52, 330, 28, 263, 151, 322, 89, 246, 47, 460, 278, 64, 101, 181]. Feeds [665]. Feedstuffs [359]. Female [257, 287, 60, 241, 223, 251, 339]. female-sign [339]. Females [244]. Fennel [330]. Fenneropenaeus [51]. Fermentation [137, 466, 473]. Fermented [27]. Fertilization [14]. fertilized [631]. Fertilizer [476, 558]. Fertilizers [350]. Fiber [230, 485]. Field [183, 332, 307, 346, 325]. Fields [618, 541]. Fiji [450, 481]. filiculoides [71]. Fillet [146, 63]. Fillets [254]. Filter [28, 72]. Filter-Feeding [28]. Fin [173, 83]. Fin-fish [83]. Final [163]. Finfish [452, 221]. Fingerling [79]. Fingerlings [425, 375, 34, 31, 384, 383, 231, 700, 230, 74, 97, 263, 222, 68, 87, 271, 572, 364]. Fingerprint [356]. Fingerprinting [49]. First [217, 257, 276, 209, 96, 151, 47, 154]. fischeri [207]. Fish [410, 489, 352, 21, 308, 309, 539, 382, 599, 329, 462, 528, 238, 239, 519, 218, 640, 220, 396, 17, 277, 186, 492, 33, 156, 518, 152, 310, 480, 250, 236, 141, 523, 636, 625, 648, 415, 11, 583, 6, 318, 534, 223, 45, 107, 291, 610, 272, 264, 612, 727, 443, 132, 416, 316, 357, 379, 322, 129, 161, 311, 354, 592, 464, 273, 432, 69, 551, 16, 213, 411, 147, 128, 685, 495, 699, 202, 274, 613, 663, 371, 88, 18, 615, 588, 694, 301, 564, 216, 154, 607, 133, 482, 585, 299, 548, 294, 83, 12]. Fish-farmers [301]. Fisheries [623, 693, 734, 721, 272, 614, 417, 402, 716, 729, 202, 702, 737, 530, 619, 547]. Fishermen [515]. Fishers [449, 702, 500, 542]. Fishery [509, 209]. Fishes [439, 297, 109, 738, 409, 195, 707, 719, 280, 79, 630]. Fishing [584, 399, 530]. Fishmeal [557, 480, 19, 234, 233, 27]. Fitness [478]. Five [158]. flavescens [209]. Flavobacterium [292, 354]. Flavor [224, 229]. flaxusa [392]. Flexibility [693]. Floating [512, 488, 654, 301]. Floating-net-cages [301]. Floats [686]. Flocculation [661]. Flora [575]. Flounder [479, 465]. Flow [57, 228, 3, 465, 658]. Flow-Through [3]. fluviatilis [115, 114, 15]. foliacea [361]. Folic [79]. Following [355]. fontinalis [150]. Food [258, 337, 722, 558, 581, 574, 672, 407, 350, 311, 89, 64, 101, 665]. Foods [322]. Foot [344]. Formation [413]. Forms [347]. Formulated [520, 329, 102, 234, 233, 362, 47, 142]. Formulating [68]. Formulation [724, 560]. Formulations [620]. Forsskal [445, 86]. Fortification [437]. fossilis [294]. Four [250]. Fourier [108]. Fraction [367]. Fracture [281, 307]. Framing [587]. fransiscana [61]. Free [69]. Freezing [652].

Frequencies [181]. Frequency [389, 328, 246]. Fresh [392, 439, 195, 391, 189, 161, 266, 154]. Freshwater [331, 638, 201, 738, 146, 458, 196, 332, 601, 446, 6, 451, 13, 311, 400, 486, 259, 319, 123, 635]. Friendly [429, 435]. Fries [553]. Frontier [440]. Frozen [2, 321]. Fructooligosaccharide [335]. Fry [352, 368, 171, 330, 698, 350, 470, 128, 495, 47, 392]. FT [108]. FT-IR [108]. Fuel [672]. fulvidraco [136]. Function [60, 696]. Functional [634, 153, 419, 361, 658]. Functions [682]. Fungal [23]. Fungi [294]. Fungicide [88]. Fusarium [559]. fuscoguttatus [119]. Future [601, 82, 360, 46].

Gadus [39]. galilaeus [355]. Galilee [355]. Gariepinus [271, 370, 313, 140, 288, 218, 197, 347, 459, 397, 231, 121, 350, 436, 434, 395, 35, 189, 717, 87, 70, 251]. Garlic [343, 459, 330]. garua [380]. Gastropoda [403, 362]. Gastropods [232]. Gelatinized [157, 6, 143]. Gender [80]. Gene [239, 5, 427, 345, 138, 479, 145, 136, 177, 144, 133]. Gene2Path [239]. General [614, 647]. Generating [244]. Genes [343, 239, 135]. Genetic [201, 158, 673, 236, 304, 724, 138, 209, 451, 645, 650]. Genitalia [403]. Genomics [545]. Genotoxicity [510]. Genotype [527]. Genotypic [409, 13]. Genus [276, 24]. Geoffery [308, 309]. Geology [684]. Geotextile [152]. Ghana [471]. Ghanas [453]. Giant [331, 201, 446, 324, 119, 270, 461, 484]. gibelio [37]. GIFT [94]. gigas [179]. Gill [563, 375, 579, 393, 170, 365]. Gills [268, 279]. Gilthead [708, 590, 487, 495]. Ginger [677]. giuris [302]. given [30]. Gland [91, 434]. Glandless [536]. Global [245, 665]. Globulin [367]. Globulin-HRPO [367]. Glossogobius [302]. Glucan [148, 153, 182]. Glucomannan [427]. Glucosamine [31]. Glutaraldehyde [292]. glycolic [237]. Glycoprotein [144]. Glyphosate [468, 346]. Glyphosate-Based [346]. Goals [601]. Godavari [525]. Gol [77]. Gold [161]. Golden [391]. Goldfish [37, 659, 246]. Golestan [139]. Gonad [119, 251, 362]. Gonadal [12]. Gonado [288]. Gonado-hepatosomatic [288]. Gonadosomatic [505]. Gondar [489]. Gorgan [509]. Gorge [511]. Gorripudi [525]. Gossypol [55]. Gourami [68, 484]. gouramy [484]. Government [216]. Grass [284, 290, 71]. Grasshopper [231]. Greece [248]. Green [425, 287, 184]. Greenhouse [426]. Grey [383]. Grounds [399]. Grouper [240, 324, 119, 441, 270]. Grow [41, 40]. Grow-Out [41, 40]. Growing [234, 233, 572]. Growout [438]. Growth [392, 343, 262, 520, 543, 61, 370, 313, 425, 389, 329, 396, 66, 597, 179, 126, 326, 39, 295, 387, 427, 390, 347, 31, 235, 146, 258, 335, 557, 56, 345, 553, 328, 700, 330, 19, 74, 493, 558, 59, 437, 457, 415, 568, 102, 318, 324, 145, 534, 628, 649, 333, 339, 538, 267, 215, 214, 298, 349, 71, 556, 447, 715, 717, 87, 271, 70, 161, 465, 488, 143, 176, 470, 529, 551, 454, 671, 246, 128, 572, 124, 364, 729, 483, 210, 170, 504, 22, 47, 160, 15, 431, 600, 469, 555, 98, 142]. Growth [40, 299, 181, 9]. Growth-Related [179]. GSI [505]. GTPase [419]. Gulf [516, 510]. Gully [228]. Gunther [16, 493]. Guppy [255, 215]. Gurdaspur [154]. Gut [540]. Gwatr [401]. Gymnarchus [520]. Gymnocypris [356].

H [448]. Habit [337]. Habitat [644, 139]. Habits [258, 89, 64, 101]. haemastoma [403]. Haemato [526]. Haemato-immunological [526]. Haematological [225, 294, 189, 70]. Haemocyte [182]. Haemolymph [353]. Halomonas [99]. Ham [381, 9, 154]. Hamilton [310, 151, 380, 302, 399, 320]. Handling [213]. Hapalindole [377]. Hapalindole-T [377]. Hard [194]. Harvest [192, 213]. Harvesting [363, 213]. harveyi [282, 431]. Hassan [307]. Hatchability [55]. Hatcheries [492, 408]. Hatchery [178, 390, 150, 152, 570, 102, 234, 233, 72, 184, 411, 99]. Hatchery-Raised [150]. Hatchery-Reared [234, 233]. Hatching [29]. Hatchlings [434]. Hawaiian [163]. Hawassa [631, 564, 519]. Hayke [140]. HCN [571]. Healing [398]. Health [155, 675, 491, 521, 534, 688, 646, 588, 548]. Heat [413]. Heavy [238, 519, 268, 416, 165]. Heckel [312]. Height [528]. heinii [134]. helena [92]. Heller [116]. Helminthes [140]. Hemato [330]. Hemato-biochemical [330]. Hematological [8, 562, 347, 677, 421, 94, 170]. Hematologichal [364]. Hematopoietic [21]. Hemibagrus [287]. Hemorrhagic [21]. Hepatic [427, 204]. hepatosomatic [288]. Herbal [333]. Herbal-mixed [333]. Herbicide [346]. Herbst [193, 123]. Herklots [461]. Herpes [467]. Herpes-Virus [467]. Herring [304]. Heteropnuestes [294]. High [673, 149, 490]. High-Yield [673]. Highlights [623]. Highly [37, 573]. Hilaire [308, 309]. Hilgendorf [110, 187]. Hilsa [592, 393, 381, 449]. Himalayan [312]. Hindgut [137]. Hippocampus [365]. Histochemical [1]. Histological [565, 457, 111, 106]. Histology [375, 189, 170]. Histopathological [562, 256, 1, 279, 346]. Histopathology [398, 121]. History [92]. HNF [145]. HNF-1 [145]. HNF-3 [145]. hoferi [195]. Holothuria [25]. Holothuroidea [412]. Holstein [305]. Home [240]. Hora [389]. Hora-Arsedi [389]. Hordeum [558]. Hormone [81, 34, 345, 227, 289, 145, 266]. Hormone-Induced [145]. Hormones [167, 436]. Hormozgan [232]. Horse [576]. Host [336, 133]. HPLC [394]. HRPO [367]. hsp70 [424]. HUFA [61]. Hulls [263]. Human [733, 605, 3, 743]. Humane [213]. Humans [603]. Huso [14]. Hyacinth [429, 428]. Hybrid [90, 347, 119, 333]. Hybrids [339]. Hydro [186]. hydrodroma [123]. Hydrodynamics [344, 633]. Hydrogen [292]. hydrophila [406, 225, 459, 677, 261, 398, 121, 294, 432, 105, 49, 237]. Hydroponic [228]. Hydroponics [569, 338]. Hyperglycemia [266]. Hyperglycemic [266]. Hyperopisus [511]. Hyperpycnites [705]. Hyperspectral [654]. hypophthalmus [176].

IAR [190]. IAR-Kick [190]. Ibadan [420, 80]. Ibeshe [109]. Ice [78]. Ichthyophonus [195]. Ichthyophthirius [205]. idae [116]. idella [187, 187, 284, 290, 71]. idellaidella [110]. Identification [353, 640, 345, 540, 49, 320, 206, 631]. IHC [467]. IHNV [21]. II [290]. Ijede [416]. Ikere [511]. ilisha [393, 381, 399]. Images [654]. Imidacloprid [709]. Immediate [363]. Immune [343, 60, 135, 175, 424, 205, 479, 529, 551, 483, 51, 182, 237]. Immune-related [343]. Immunity [300, 175, 335]. Immunoconjugate [367]. Immunological [396, 676, 526]. Immunomodulation [50]. immunomodulator [9]. Immunomodulatory [374]. Immunostimulants [205]. Impact [329, 155, 662, 733, 652, 55, 167, 358, 513, 401, 686, 475, 477, 735, 707, 719, 664, 14, 399, 743, 510, 665, 428]. Impacts [153, 275, 125, 139]. Implications [602, 320]. Importance [599, 17, 616, 623, 598, 667, 647, 646, 674, 613, 533, 494]. Important [592, 165, 273]. Imported [439]. Improvement [235, 673]. Improving [596, 498, 723, 548]. in-vitro [432]. In-Vivo [88]. Inactivate [21]. Incidence [2, 224]. Incidences [486]. Incidental [103]. Incimaxx [292]. Inclusion [115, 98]. Incorporated [396, 421]. Increase [175]. Index [505]. Indexes [288]. India [154, 563, 220, 103, 552, 312, 236, 601, 86, 23, 165, 89, 106, 486, 149, 496, 525, 83, 104, 516, 64, 101, 123, 293]. Indian [2, 276, 156, 196, 310, 507, 571, 517, 202, 51]. Indicative [42]. Indicator [268, 238]. Indices [31, 330, 74, 94, 338, 70, 246, 51]. indicus [51]. Indigenous [156, 69, 99]. Indonesia [120, 185, 180, 435, 715, 464, 301]. Indoor [438]. Induced [32, 188, 204, 1, 332, 145, 35, 266]. Induction [130, 112, 441, 325]. Inductions [344]. Industry [453, 498, 474, 199, 494, 712, 532, 482]. Infected [140, 562, 225, 409, 195, 133]. Infection [367, 7, 677, 212, 134, 419, 354, 23, 105, 537, 544]. Infections [256, 456, 442]. Infectious [21, 212, 612, 486, 11]. Infestation [125]. Influence [157, 689, 6, 151, 366, 111, 470, 213, 246]. Influences [482]. Influencing [517]. Information [513, 223]. Infrared [108]. Ingredient [230]. Ingredients [557, 418, 142]. Inhabits [252]. Inhibiting [258]. Inhibition [432]. Inhibitor [419, 88]. Inhibitors [400]. Inhibits [38]. iniae [273]. Initial [387]. Initially [133]. Initiate [523]. Initiative [213]. Inkofele [558]. Inland [578]. Inlet [7]. Innovation [668]. Innovations [720, 712]. Innovative [734, 716, 676]. Inoculation [2]. Inorganic [350]. Insight [500]. Insights [341, 619]. Integrated [598, 465, 699]. Intelligence [494]. Intense [661]. Intensities [447]. Intensive [66, 52, 689, 315, 504, 269, 438]. Interact [655]. Interactions [37, 628, 649]. Intercrystalline [281]. Internal [489, 140, 218]. International [77, 177]. Internet [651]. Intersection [690]. Interstitial [160]. Intertidal [232, 363]. Intestine [250, 284, 290]. Intramuscular [398]. Introduction [177]. invadans [294]. Invasive [160]. Inventory [388]. Investigating [675]. Investigation [4, 380]. Investigations [125, 355]. Invitation [550]. Invitro [4]. Involved [135]. Ion [365]. IR [108]. Iran [509, 414, 77, 232, 501, 139]. Iranian [168]. Iraq [439, 382, 322, 307]. Iron [55]. Irradiation [411]. ISAV [11, 11]. Island [249]. Islands [363, 351]. Isolated [406, 382, 33, 409, 501, 177, 13]. Isolates [559, 121, 377, 317]. Isolation [65, 318, 571]. Issue [550, 304, 369, 378].

Issues [141]. Italy [668]. Items [211].

J&K [312]. jaensis [700, 572]. Jammu [312]. Japan [12]. japonica [192, 203]. japonicus [25, 328, 327]. jarbua [86]. Jatropha [98]. Java [464]. Jewel [184]. Jiangsu [426]. Job [26]. Johnius [505]. Journal [369, 378]. Jumbo [101]. Juvenile [313, 277, 126, 39, 359, 243, 102, 234, 233, 707, 719, 184, 98, 365, 299]. Juveniles [520, 370, 295, 448, 94, 324, 717, 70, 469].

Kappaphycus [715, 454, 431]. Kaptai [267]. kauderni [184]. Kayamkulam [348]. Kei [334]. Kenya [528, 513, 395]. Kenyan [523]. Kerala [563, 348, 405]. Kernel [98]. khudree [47]. KHV [467]. Kibuyuni [528]. Kick [190]. Kidney [150]. Kiln [132]. Kilograms [146]. Klawing [464]. Kneiss [363]. Koi [467]. Kolkata [177, 186]. Kongodu [525]. Konjac [427]. Kopasanda [282]. Kottakudi [118]. Koumanns [184]. kubo [276]. Kunju [276]. Kupang [715]. Kurdistan [322]. Kuril [351]. kuwaitensis [223]. Kwale [528].

L [385, 237]. L-amino [385]. L-lactide-co-glycolic [237]. L. [148, 389, 5, 383, 677, 243, 282, 506, 250, 330, 398, 558, 318, 338, 267, 127, 495, 527, 525, 142, 537, 317]. Labakkang [120]. Labeo [32, 157, 225, 384, 514, 6, 143, 105, 460, 645, 650, 9]. Laboratory [183, 346, 278, 181]. labrax [368, 590, 73, 5, 256, 249, 317]. Labrus [44, 29]. Lactic [660]. lactide [237]. Lagoon [344, 109]. Lagoons [10]. Lagos [109, 416]. Lake [185, 348, 267, 302, 226, 489, 389, 140, 519, 218, 268, 209, 495, 301, 564, 428, 356]. Lakes [252]. Lakshmipur [449]. Lanceolatus [119, 324, 270]. Land [120]. Landlocked [710]. Lanka [417, 530]. Large [385, 224, 455, 549, 658]. Large-Scale [658]. Larimichthys [385]. Larva [278]. Larvae [61, 44, 282, 437, 151, 173, 717, 29, 83]. Larval [253, 42, 275, 436, 381, 386, 36]. Latency [91, 434]. Lates [425, 27, 483, 371]. latipinna [311]. Lead [383, 204]. Lead-Induced [204]. Leadwort [189]. Leaf [384, 549, 70]. LEAP [136]. LEAP-2 [136]. Learned [150]. Leaves [32, 282, 506]. Leg [66, 93, 214, 95]. Lemna [71]. Length [511, 109, 693, 228, 86, 461, 464, 16, 505]. Lepeophtheirus [133]. Lesions [727]. Lessons [150]. Lethal [156, 383]. Lettuce [228]. Leucocytes [383]. leucospilota [25]. Level [287, 314, 163, 572]. Levels [343, 462, 81, 297, 130, 241, 499, 62, 37, 74, 59, 215, 63, 71, 472, 70, 143, 364, 210, 181]. LHRH [81]. LHRH-A [81]. Lice [7]. licheniformis [330]. Life [92, 384, 391, 579, 574, 576, 88]. Light [284, 290, 447]. Lima [271]. Limnology [656]. Limpet [163]. Line [5]. Lined [365]. Link [245]. Linkage [26]. Linn [413]. Linnaeus [326, 235, 518, 442, 470, 325, 36, 15, 40]. Linne [22]. Linneaus [450]. Linseed [243]. Lionfish [89, 106]. Lipid [130, 427, 244, 211, 233, 479, 324, 87, 361, 555]. Lipids [191, 40]. Lipoic [1].

Lipped [501]. Liquid [233]. Lithophaga [340]. Litopenaeus [66, 175, 424, 706, 718, 212, 57, 275, 214, 510, 149, 504, 99, 269, 438, 431, 30, 490, 64, 95, 182, 181]. Live [520, 578, 184, 47, 36]. Livelihood [449, 500, 542]. Livelihoods [515, 475, 477]. Liver [375, 6, 479, 145, 136]. Liver-Enriched [145]. Liver-Expressed [136]. llamasi [277, 241]. LNMU [496]. Loach [485]. Load [171, 84]. Loading [185, 48]. Lobster [435]. Local [513, 216]. Locally [439, 557, 172]. Loci [65]. Logs [281]. Long [374, 492, 497, 37]. Long-Term [37, 374, 492, 497]. longirostris [361]. Loop [726]. Loss [612]. Louse [144]. Low [424, 212, 438, 95, 365]. Lower [563, 505]. Loyalty [321]. lucifera [194]. lucius [219]. Lugo [140]. lunatus [271]. LWR [505].

M [516]. Macro [392]. Macroalgal [198]. Macrobenthos [104]. Macrobrachium [60, 331, 201, 404, 446, 461, 23, 110, 187, 41, 116]. macrocephalus [217, 257, 347]. macrolepidotus [511]. Macronutrient [210]. Magdalena [657]. magna [262]. Mahseer [47]. Maintaining [595]. Maitum [319]. Major [471, 90, 624, 621, 591, 627, 618, 399, 646, 613, 12]. Malawi [655, 172, 444, 440]. Malaysia [331, 197]. Male [217, 110, 339]. male-sign [339]. Males [81]. Malformations [403]. Maluku [334]. Man [41]. Manage [666]. Management [471, 220, 651, 491, 498, 693, 521, 10, 209, 734, 610, 402, 320, 695, 712, 678, 417]. Managing [729]. Manchar [302]. Mangrove [162, 348, 169, 104, 293]. Mangroves [667]. Mania [412]. Manifested [605]. Maninjau [301, 185]. Mannar [516]. Manner [141]. Mannooligosaccharides [93]. manure [631]. Manures [22]. Many [623]. Manzala [268]. MAPE [442]. Mapping [684]. margaritifera [390, 501, 450, 454, 481]. Mariculture [513, 644, 221]. marigala [645, 650]. Marine [388, 547, 439, 624, 67, 283, 738, 191, 641, 697, 711, 732, 704, 690, 735, 574, 522, 247, 84, 591, 742, 72, 223, 684, 714, 743, 315, 741, 280, 695, 712, 658, 633, 670, 546]. Mariotteya [125]. maritimum [317]. Markers [201, 158, 236, 451, 645, 650]. Market [564]. Markets [197]. Marking [183, 360]. Marphysa [465, 469]. Marron [300]. Mart [429]. Martius [428]. Masculinization [415]. Mass [125, 116]. Massive [198, 249]. Mastacembelus [337]. Material [304, 54]. Materials [185]. Mating [110]. Matter [657, 359]. Maturation [257, 85, 163, 57, 227, 289, 119]. Maturing [167]. Maturity [146, 108, 362]. Mauremys [414, 139]. Mazandaran [139]. Maze [73]. MBG [182]. Meager [134]. Meagre [437]. Meal [536, 55, 231, 328, 19, 234, 479, 698, 452, 70, 27, 495, 98, 299, 518]. Measures [491]. Meat [637]. Media [406]. Median [156]. mediated [266]. Medicinal [394]. Medicine [617, 647, 603, 664, 602, 646]. Mediterranean [122, 340, 363, 403, 283, 45]. Meghna [393]. Meiliang [422]. melas [124]. Melatonin [217]. Mentha [338]. Mera [346]. Mesocosms [66]. Metabolic [126, 568]. Metabolism [427, 479]. Metabolites [274, 88]. Metal [340, 416, 165]. Metals [297, 238, 519, 268]. Metamorphosis [29]. metcalfei [165]. Method [642, 10, 407, 502]. Methods [535, 609, 183, 418, 318, 72, 360]. Methyl [34, 59, 127]. Methyltestosterone [698, 325]. Metropolis [80]. Mexico [510]. MIB [229]. Micro [28, 275, 707, 719]. Micro-Particles [28]. Microalgae [637, 672, 131, 269]. Microalgal [492]. Microbes [658]. Microbial [309, 439, 224, 373, 740, 670]. Microbiological [197, 294, 208, 254]. Microbiology [376, 624, 533]. Microbiomes [655]. Micrococcus [540]. Microcrustacean [164]. Microcystis [225]. Microemulsified [347]. Microsatellite [158, 65]. Microsatellites [201]. Microscopic [284, 290]. Microscopy [207]. Microsystems [656]. Microwave [318]. Migratory [573]. Milkfish [556]. Mindanao [445]. Mineral [271, 149, 123]. Mineralization [313, 173]. Minerals [74]. Minh [321]. Mint [366]. Misgurnus [485]. Mitchill [150]. Mitigating [736]. Mitigation [716]. Mitochondrial [395]. Mixed [493, 333]. Miyadiella [276]. MLS [156]. mmunerelated [5]. Mode [625, 648]. Model [357, 653]. Modelling [283, 307, 633]. Models [651, 730]. Modern [623, 664]. Modiolus [165]. Modulation [5]. Mogalipalem [525]. mola [156]. Molecular [606, 270, 136, 645, 650, 203, 144]. Molecule [419]. Mollusc [388]. Mollusca [362]. Mollusks [668]. Molly [311]. Molt [332]. Monitoring [190, 499, 249, 615]. Mono [408]. Mono-sex [408]. monodon [295, 359, 282, 506, 52, 275, 291, 48, 101]. Monomeric [137]. Monopterus [451]. Monosex [34, 493, 267, 325]. Montagu [469]. montezumae [463]. Monthly [186]. mookalee [571]. Moon [308, 309]. Moon-Fish [308]. morhua [39]. Moringa [370, 70]. Morphological [609, 559, 223]. Morphology [232]. Morphometric [326, 264]. Morphometrics [187]. Morphotypes [110]. Mortalities [125]. mossambica [197]. mossambicus [58]. Most [501]. Mosul [322]. Mother [454]. Motley [305]. Motorized [132]. Mouth [184]. Mozambique [58]. Mrigala [460, 460]. Mt [698]. Mud [227, 289, 540, 451, 35]. Mudflat [363]. muelarianus [288]. Muess [258]. Mugil [383]. Mullet [383]. Mulloway [328]. Multi [598, 129]. Multi-Use [129]. multifiliis [205]. Multiple [426]. Munidopsidae [516]. Munidopsis [516]. Muricidae [403]. Muscle [179, 247, 555]. Muscles [268]. Mushroom [182]. Muthupettai [83, 293]. mycoides [300]. mykiss [146, 448, 345, 74, 75, 538, 63, 208, 111, 78, 364, 210, 79, 555]. Myonecrosis [212]. Mystus [188, 236, 405, 464]. Myxosporean [442].

naccarii [210]. Nadu [104, 706, 718, 118]. Nanoparticles [514, 521]. Nari [118]. Naringi [32]. natans [732]. natator [193]. National [564]. Natural [539, 368, 511, 66, 332, 522, 350, 436, 322, 259, 645, 650, 502, 356]. Naturally [195]. Navigating [721]. Near [706, 718]. Necrosis [21]. Need [42, 242]. Negative [81, 275]. Nematode [489, 226]. nemurus [287]. Net [563, 39, 488, 393, 301]. Networks [190]. Neurobiology [609]. Neuropeptides [507]. Neurosteroids [507]. Neutrality [666]. Nicotiana [384]. Nigella [50]. Niger [260, 243, 168]. Nigeria [565, 80, 511, 109, 475, 477, 416, 478, 26, 420, 159, 260, 461, 699, 216, 508]. Nigerian [251]. Nile [489, 343, 389, 375, 468, 140, 519, 562, 387, 518, 335, 50, 677, 557, 55, 594, 243, 553, 230, 97, 222, 457, 415, 408, 349, 379, 470, 128, 325, 200, 98, 142, 537, 544, 299, 254]. Nilem [462]. Nilotic [372]. niloticus [113, 489, 406, 343, 520, 389, 375, 140, 34, 519, 218, 387, 158, 235, 518, 335, 268, 204, 50, 677, 557, 594, 442, 243, 553, 330, 336, 230, 421, 398, 97, 263, 457, 94, 342, 698, 125, 339, 408, 267, 54, 349, 379, 470, 226, 128, 325, 200, 18, 467, 142, 537, 544, 299, 254]. Nitrifying [177, 41, 269]. Nitrite [170]. Nitrogen [112, 57, 558, 666, 48, 714]. NkhataBay [172]. NNV [171]. Noakhali [358]. Non [406, 157, 347, 335, 211, 6, 560, 143]. Non-Conventional [211, 560]. Non-Gelatinized [6, 143, 157]. Non-microemulsified [347]. Non-Selectivity [406]. Non-specific [335]. Normal [35]. North [334, 515, 414, 220]. North-East [334]. North-Western [515]. Northeast [140]. Northern [344, 340, 307, 405]. Note [589, 583, 567, 369, 378, 588, 593, 586]. Notes [405]. Novobiocin [501]. NTT [715]. nucifera [288]. Nucleotides [357]. Nuggets [153]. Nursery [566, 18, 431]. Nusa [412]. Nutrient [157, 636, 568, 666, 306, 447, 717, 48, 286, 100]. Nutrients [308, 77, 299]. Nutrition [620, 673, 141, 636, 581, 610, 548]. Nutritional [126, 359, 458, 604, 172, 412, 416, 478, 147]. Nutritionally [311]. Nutritious [141]. Nutritive [97]. nxrA [177]. Nylon [39].

O. [594]. OBA [478]. Objectives [619]. Obtained [197, 565]. occidentalis [511, 274]. Ocean [657, 567, 622, 687]. Oceanic [705, 249]. Oceanographic [634, 704, 622, 653]. Oceanography [616, 626, 632, 674]. Oceanology [684]. ocellaris [366]. Ocimum [338]. ODN [237]. odorata [282, 506]. Off [224, 229, 516]. Off-Flavor [224, 229]. officinale [677]. ohita [6]. Oil [8, 396, 277, 153, 330, 102, 129, 127, 307]. Oils [318, 366, 487]. Old [328]. Oleic [84]. oleifera [370, 70]. Olfactory [385]. Oligofructose [137]. oligonucleotide [237]. Oligosaccharides [113, 137]. olivacea [162]. olivaceus [465]. Olive [479, 465, 320]. Oman [168]. Omega [242]. Omega-3 [242]. Ompok [433]. On-board [381]. Oncorhynchus [146, 448, 345, 74, 75, 538, 63, 208, 111, 78, 364, 210, 79, 555]. Oocyte [85]. Operated [563]. Operation [642, 57]. Operations [639]. Opportunity [668]. Optimization [52]. Optimize [229]. Optimizing [724]. Optimum [313, 227]. Oral [9]. Orange [318]. Oreochromic [18]. Oreochromis [113, 489, 406, 343, 389, 375, 140, 58, 34, 519, 218, 387, 158, 235, 518, 335, 204, 50, 677, 557, 442, 243, 553, 330, 336, 230, 421, 398, 493, 59, 97, 263, 457, 94, 342, 698, 125, 333, 339, 408, 267, 54, 298, 349, 379, 470, 226, 128, 325, 200, 467, 142, 537, 544, 299, 254]. Oreochromisniloticus [543]. Organic [706, 718, 185, 166, 350, 46, 22]. Organism [504]. Organisms [352, 697, 580, 567, 578, 582]. Organization [255]. Organochlorine [351]. Organogenesis [44]. Organs [565]. Origanum [487]. Origin [40]. Ornamental [329, 220, 396, 33, 195, 161, 311, 630]. Oromia [558, 527]. Orthologous [239]. Osmotic [365]. Osphronemidae [484]. Osphronemus [484]. Osteochilus [278]. Ostreopsis [207]. Osun [461]. Otolith [168, 356]. Outbreaks [248, 150]. outcomes [548]. Ovaprim [91, 35]. Ovarian [652, 227, 289, 108]. ovata [207]. Oven [309]. Oven-Dried [309]. Overview [545, 626, 614, 647, 632, 449, 587]. Ovicell [244]. Oxidase [138]. Oxidative [325]. Oxidized [427]. Oxidizing [714]. Oxygen [639]. oxysporum [559]. Oxytetracycline [398]. Oyo [80, 216]. Oyster [90, 501, 629, 454, 668]. Oysters [3]. Oziothelphusa [332, 400, 266, 259].

P [144, 313]. P-Glycoprotein [144]. P. [215]. pabda [433]. Pacific [687, 373, 734, 431, 64, 182]. Packing [171]. Padang [484]. paddy [541]. Pagrus [223]. PAHs [510]. Pakistan [302]. Palaemon [662]. Palm [277, 717]. Pandalopsis [203]. Panels [90]. Pangas [367]. Pangasius [367, 253, 310, 176]. Pangi [319]. Papain [543]. Paracentrotuslividus [183]. Paradise [326]. paradiseus [326]. parahaemolyticus [38, 84]. Paralichythys [465]. Parameter [180]. Parameters [262, 526, 326, 424, 171, 347, 397, 677, 345, 37, 205, 71, 260, 78, 364, 170, 525, 79]. Parangipettai [86, 101, 123]. Parapenaeus [361]. Parasite [205, 502]. Parasites [489, 218, 226]. Parasitic [140, 125]. Parastromateus [168]. Park [283]. Parrot [323]. Part [489, 524]. Partial [277, 31, 518, 557, 102, 495]. Participation [296]. Particle [657]. Particles [28]. Partitioning [30]. Parts [623, 63]. Passive [486]. pastoris [728]. Patenting [304]. Pathogen [250, 3, 84, 273]. Pathogenic [559, 432, 274]. Pathogenicity [336]. Pathogens [410, 426]. Pathological [727, 355]. Pathology [640, 582, 627, 612, 607]. Pathway [135]. Pathways [239]. Pattern [169, 173, 136]. Patterns [33]. PCR [11, 49, 467]. Pearl [390, 735, 501, 454]. Pecaron [240]. Peel [318]. pelagicus [353]. Pelteobagrus [136]. Penaeus [295, 359, 56, 282, 506, 52, 275, 291, 48, 426, 101]. Penaeusmonodon [120]. Pennar [149]. Pens [39]. People [509]. Pepper [309]. Peppermint [338]. Peptide [136]. Peptides [17, 522]. Perca [115, 114, 209, 15]. Perceptions [482]. Perch [114, 209, 86, 184, 15, 115]. Perciformes [484]. Perfect [512]. Performance [392, 343, 520, 543, 61, 370, 389, 368, 287, 554, 427, 390, 347, 235, 146, 335, 557, 56, 328, 480, 19, 493, 437, 549, 102, 333, 339, 538, 350, 436, 71, 556, 132, 316, 717, 70, 488, 470, 529, 551, 572, 124, 364, 495, 483, 504, 527, 269, 98, 142, 299, 246]. Performances [94, 215]. Period [192, 473, 434]. Periods [91, 383]. Permeability [281]. Perna [165]. Peroxide [292]. Persian [81, 85, 62, 76]. persicus [81, 85, 62, 76]. Personality [321]. Perspective [632]. Perspectives [67, 498, 129]. Pesticides [351, 95, 1]. Petalonia [175]. Pethia [156]. Pets [482]. pH [430]. Pharmaceutic [418]. Pharmaceutical [609]. Pharmaceuticals [672]. Phase [387, 570, 41, 431]. Phaseolus [271]. Phenotypic [409, 13]. Philippi [92]. Philippines [445, 12, 505]. phoenicis [717]. Phosphorus [19, 48]. Photobacteriosis [248, 45]. Photobacterium [256]. Photoperiods [295]. Photosynthetic [567]. Phronima [164]. Phyllanthus [288]. Phylogenetic [327]. Physical [153]. Physico [462, 420, 174]. Physico-Chemical [462, 420, 174]. Physicochemical [260, 78]. Physiological [262, 126, 682]. Physiology [615]. Phytase [313, 271]. Phytase-Supplemented [271]. Phyto [717]. Phyto-Additives [717]. Phytochemical [4]. Phytogenic [708]. Phytoplankton [736, 322, 118, 293, 631]. Pichia [728]. Pigmentation [347, 323]. Pilot [356]. Pinctada [390, 501, 450, 454, 481]. Pineapple [556]. Pink [361, 351]. piperita [338]. Pirarucu [179]. Pisces [328]. piscicida [256]. Pisciculture [597]. Pituitary [91, 434]. Placed [73]. Placobdella [414]. Plan [209]. Plankton [687, 589, 504, 496, 18]. Plans [498]. Plant [352, 21, 31, 518, 457, 291, 452, 306, 495, 30, 394, 40]. Plants [228, 549, 172, 71, 595]. Plasma [365]. Plasmids [376, 533]. Plastic [546]. Plastics [707, 719]. Platanus [274]. Plateau [356]. platensis [343, 112, 457]. Plerocercoid [134]. PLGA [237]. Plumbago [189]. PmRab7 [419]. Poecilancistrum [134]. Poecilia [255, 311]. Point [83]. Politics [245]. Pollution [238, 268, 591, 742, 125]. Poly [437, 237]. Poly-Unsaturated [437]. Polychaete [465, 469]. Polyculture [211, 488, 460]. Polymer [69]. Polymerase [354]. Polymorphic [236, 65]. Polymorphisms [345]. Polynemus [326]. Polypeptide [367]. Polyploidy [413]. Polysaccharides [387, 112]. Pomfret [168]. Pompano [566, 571]. Pond [503, 655, 506, 211, 172, 43, 124, 139, 631]. Ponds [520, 476, 599, 197, 56, 373, 493, 10, 496, 426, 460, 525, 490, 154]. Pontederiaceae [428]. Popular [738]. Population [34, 236, 628, 649, 615, 259, 645, 650, 139]. Populations [183, 209, 395, 482]. Porosity [281]. Porphyra [206]. Portunidae [96]. Portunus [353]. Positive [81]. Possible [7, 10]. Post [700, 282, 213]. Post-Harvest [213]. Posterior [290]. Potassium [74]. Potential [410, 58, 282, 722, 521, 407, 172, 161, 249, 351, 64]. Potentials [416]. Poultry [234, 299, 631]. ppa [424]. Practical [485]. Practices [491, 417, 213]. Pradesh [149, 525]. Prawn [331, 295, 164, 446, 461, 110]. Prawns [201]. Pre [308, 309, 572]. Pre-growing [572]. Pre-Treated [308, 309]. Prebiotic [148, 93, 153]. Precise [494]. Predatory [352]. Predictive [730]. Preference [462]. Preferences [509, 460]. Preliminary [219, 278]. prenanti [427]. Preparation [54]. Preparations [418]. Presence [499, 507]. Present [272, 280]. Preservation [585]. Preserved [78]. Preserving [704]. Pressure [374, 365]. Prevalence [489, 140, 218]. Prevalent [426]. Prevention [692, 506]. Prey [184]. Primary [516]. Principal [248]. Probiotic [526, 250, 294, 529, 551, 99]. Probiotics [368, 404, 688, 470, 611]. Procambarus [277, 241]. Process [11]. Processed [717]. Processes [167, 705, 622]. Processing [21, 80, 291, 54]. Processors [80]. Produced [439, 34, 390, 38, 84, 291, 350, 436]. Producing [571]. Product [234, 213, 299]. Production [130, 186, 620, 512, 164, 224, 153, 637, 52, 19, 108, 627, 715, 444, 440, 184, 386, 423, 116, 701, 254]. Productive [283, 495]. Productivity [503, 594, 338, 99]. Products [2, 509, 480]. Profile [421, 6, 151, 538, 111, 210, 47, 394]. Profiles [14, 193]. Profiling [367, 460]. Profitability [594, 444]. Progenies [55]. Progeny [659]. Programs [679].

Progranulin [145]. Prokaryotic [373]. Prolactin [158]. Prominence [608]. Promising [522, 285]. Promoter [318, 9]. Promoters [708, 487]. Propagation [350, 18]. Properties [536, 153, 169]. Prophenoloxidase [51]. propo [424]. Proposed [491]. Prospective [331]. Prospects [699]. Protease [540]. protected [547]. Protecting [739]. Protection [459]. Protective [32, 204, 1]. Protein [367, 287, 359, 31, 536, 518, 250, 211, 457, 233, 560, 538, 68, 452, 71, 472, 143, 572, 495, 30]. Proteomic [377]. Proteomics [605, 625, 648, 627]. Provide [530]. Providing [141]. Province [232, 258, 715, 426, 401, 185, 319]. Province-Indonesia [185]. Provinces [139]. Proximate [102, 123]. Proximity [697]. przewalskii [356]. Pseudomonas [409, 571, 20]. Pseudosciaena [455]. psychrophilum [292]. Pterapogon [184]. Pterocladia [222]. Pterois [89, 106]. Pterophyllum [329]. Puberty [217, 441]. Public [155]. Pufas [242]. Puffer [280]. pugilis [362]. Pulp [466, 473]. Punjab [154]. Punten [413]. Puntius [320]. purpuratus [84].

Qaroun [495]. Qinghai [356]. Qinghai-Tibet [356]. quadricarinatus [126, 65]. Qualitative [496]. Qualities [308]. Quality [439, 638, 197, 171, 146, 152, 656, 455, 499, 180, 372, 713, 72, 291, 366, 743, 213, 78, 723, 525, 174]. Quantity [434]. quelen [729].

R [406]. R-S [406]. Rabbitfish [392, 488, 170]. Race [413]. Radiation [21, 90]. Radular [232]. Rafinesque [124]. Raft [654]. Raijua [412]. Rainbow [146, 448, 345, 292, 205, 74, 75, 391, 247, 538, 63, 208, 111, 78, 364, 79]. Raised [150, 493, 229]. Ramgati [449]. Ramsar [177]. Randomly [236]. RAPD [236, 451]. Rapid [367, 354]. Rate [115, 543, 241, 167, 228, 436, 349, 71, 447, 484, 98]. Rates [56, 558, 556]. Ratio [313, 14, 461]. Ratios [59, 324, 555]. rDNA [49]. rDNA-PCR [49]. Re [242]. Re-Think [242]. Reaction [354]. Real [11]. Real-time [11]. Reared [370, 424, 92, 328, 234, 233, 124, 278]. Rearing [590, 162, 566, 381, 36]. received [175]. Receiving [169]. Receptor [345]. Recirculating [178, 146, 57, 124]. Recirculation [167, 438]. Record [276, 312, 154]. Records [103]. Recreational [584]. Red [499, 304, 430, 333, 255, 441, 89, 106, 386, 361, 431]. Red-Tail [255]. Redclaw [65]. Reduce [84, 411]. Reducing [75]. Reduction [153, 176]. Reef [12]. Reefs [643, 739]. Reference [256, 272]. Refugia [18]. Regency [180, 412, 715, 334]. Regime [450]. Regimes [162, 151]. Region [471, 297, 554, 312, 727, 322, 174, 351, 293]. Regional [564, 635]. Regions [455]. regius [437]. regulated [424]. Regulation [135, 419, 400]. Related [179, 427, 343]. Relation [186, 390, 143]. Relations [463]. Relationship [511, 336, 86, 461, 464, 16, 78, 505]. Relationships [77, 109]. Release [703, 266]. Relevant [605]. Remarkable [738]. Remediation [152]. Remote [639, 654]. Removal [368]. Renal [204]. rendalli [172, 306].

Replace [27]. Replacement [31, 518, 557, 19, 102, 234, 495]. Replacing [480, 233]. Replication [212]. Report [582]. Reproduction [733, 59, 463, 400, 729]. Reproductive [287, 60, 608, 146, 617, 337, 433, 163, 215, 517, 106, 302, 259]. Requirements [604]. rerio [396, 145]. Research [609, 276, 675, 686, 703, 369, 341, 378, 631]. Researcher [26]. Researches [135]. Reservoir [297, 552, 478, 420, 307]. Reservoirs [471, 402, 281, 202]. Resilience [725]. Resistance [61, 692, 175, 33, 689, 333, 731, 20, 161, 529, 551, 24, 431]. Resistant [677, 377]. Resource [129]. Resources [220, 695]. Respiration [579]. Responding [81]. Response [396, 126, 175, 112, 553, 700, 558, 529, 551, 483, 696, 182, 237]. Responses [256, 189, 40]. Responsibility [713]. Responsiveness [294]. Restoration [683]. Restrictions [693]. Result [462]. Resuspension [10]. reticulata [255, 161]. reticulate [215]. Retinoic [266]. Return [216]. Returns [512, 267]. Reuse [379]. Reveals [385]. Reversal [698]. Reversed [470]. Review [453, 535, 331, 519, 42, 196, 337, 433, 474, 407, 436, 452, 709, 202, 428]. Reviewing [498]. Revolution [711]. RFLP [451]. Rhamdia [729]. Rheological [536]. Rhodophyta [206]. Rhynchophorus [717]. Rice [332, 699, 18]. Rice-Cum-Fish [699]. Rice/Fish [18]. Ridged [193]. rigida [222]. Ringtest [11]. Rising [728]. Risk [248, 612, 678]. River [565, 657, 265, 326, 312, 416, 322, 461, 393, 464, 149, 505, 319, 96, 461, 200]. Riverine [16]. rivoliana [173]. rivulatus [392, 252, 170]. Rockfish [147]. Rockworm [465, 469]. Roe [14]. rogercresseyi [144]. rohita [32, 157, 225, 384, 514, 143, 105, 460, 645, 650, 9]. Rohu [384, 105, 460]. Role [113, 410, 376, 624, 597, 736, 711, 1, 728, 621, 732, 704, 419, 339, 627, 603, 618, 400, 714, 595, 184, 646, 723, 613, 701, 548]. root [189]. rosenbergii [331, 201, 404, 446, 23, 41]. Rotenone [352]. Rotenone-Yielding [352]. Round [467]. Roxb [286]. rRNA [177]. RT [11]. RT-PCR [11]. rubens [458]. Rural [550, 460]. Russian [244]. ruthenus [85].

s [377, 406]. Sabu [412]. Saccharina [192]. Saccharomyces [9]. Safe [141]. Sail [200]. Saint [209, 308, 309]. Saint-Hilaire [308, 309]. Saint-Pierre [209]. Saki [216]. Saki-East [216]. salar [167, 250, 133]. Salicornia [286]. Salicylate [127]. salina [130, 164, 275]. Saline [543, 35]. Salinities [424, 386]. Salinity [156, 212, 162, 339, 438, 469, 95, 365]. Salmo [167, 250, 133]. Salmon [7, 710, 167, 250, 11, 507, 360, 517, 588, 502, 351, 144, 680]. Salmonids [178]. salmonis [133]. Salt [298]. Salvelinus [115, 150, 114]. Sampled [478]. sandwicensis [163]. sanguinea [465, 469]. Sanitization [292]. Santa [388]. sarana [320]. Sarangani [319]. Sarotherodon [355]. Saskatchewan [314]. sativa [50]. Saussure [463]. Save [384]. Savusavu [450]. scalare [329]. Scale [314, 550, 722, 240, 716, 423, 737, 530, 658, 417, 542]. Scallop [72]. Scanning [284, 290]. Schedule [278]. Schizothoraichthys [312].

Schizothorax [427]. schlegeli [147]. Sciaenidae [328]. Science [617, 690, 626, 474, 619, 101]. Scientific [245]. Scorpaena [285]. Screw [22]. scrofa [285]. Scylla [162, 227, 289, 540]. Sea [590, 7, 248, 73, 122, 5, 25, 183, 455, 327, 340, 579, 576, 577, 363, 412, 27, 249, 495, 516, 317, 686, 134, 53, 168]. Seabass [425, 368, 256, 483, 681]. Seabream [61, 708, 487]. Seafloor [684]. Seafood [141, 321, 199, 532, 701]. Seagrass [586, 334]. Seahorse [365]. Search [239]. Season [442, 693, 214, 260]. Seasonal [704, 62, 340, 169]. Seasonality [248]. Seawater [72]. Seaweed [392, 161, 431]. Seaweeds [596, 222, 454]. Sebastes [147]. Sections [168]. Security [722, 407, 665]. Sediment [297, 344, 10, 422]. Sediments [169]. Seed [370, 243, 566, 116, 715]. seenghala [196]. Selected [352, 429, 565, 5, 359, 466, 473, 395, 715, 416, 478]. Selection [528, 724, 133]. Selectivity [406, 393]. Selenium [204, 421]. SEM [207]. Semen [305]. Semi [52, 493]. Semi-Intensive [52]. semisulcatus [56]. senex [332, 400, 266, 259, 332, 266, 400, 259]. Senggaringan [464]. Sensing [639, 654]. Sensitive [501, 377]. Sensor [190]. Sensors [606]. Sensory [308, 153, 208]. Sentinels [303]. Separately [330]. Septicemia [21, 537]. Sequences [177]. Seriola [173]. Serological [261]. serrata [227, 289, 540]. Serum [367, 425, 383, 62, 94, 189]. Sewage [186]. Sex [81, 493, 59, 463, 698, 461, 470, 408]. Sexual [257, 119]. Sexually [167]. Shampoo [32]. Shanigaram [552]. Shark [255]. Sharkia [258]. Sharptooth [456]. Shed [621]. Shelf [391]. Shelf-Life [391]. Shell [390, 96, 194]. Shellfish [69, 741]. Shewanella [99]. shiranus [298]. SHK [357]. SHK-1 [357]. Shocking [413]. Short [279]. Showing [131]. Shrimp [2, 692, 120, 535, 66, 265, 642, 135, 175, 424, 691, 536, 56, 725, 212, 153, 282, 506, 498, 373, 401, 180, 430, 713, 291, 169, 214, 731, 48, 671, 587, 510, 149, 315, 504, 426, 51, 525, 99, 269, 438, 431, 490, 203, 64, 101, 95, 182, 181]. Shrimps [93, 361]. Siberian [137]. sic [38]. Sicilian [285]. Sicily [207]. Sidama [564]. Siddipet [552]. Side [325]. Siderophore [571]. Siganus [392, 252, 170]. sign [339]. Significance [596, 641, 291, 630, 452, 629, 741, 737]. Silage [553]. Silage-based [553]. Silite [226]. Silver [514, 566, 521]. Simmental [305]. Simple [178]. Simplification [609]. Simulated [300, 171]. Simulations [704]. Sindh [302]. sinensis [425]. singaringan [464]. Single [467]. Single-Round [467]. Siphonostomatoida [223]. Sistan [401]. Sistan-Baluchestan [401]. Sitakunda [272]. Site [528, 401, 177]. Situbondo [240]. Sizes [15]. Skeletal [590, 179, 173]. Skin [655, 255, 147]. Skins [462]. Slaughter [487]. Small [156, 550, 722, 549, 419, 417, 716, 423, 737, 530, 542, 226]. Small-Scale [550, 722, 716, 737, 530, 417, 542]. Smart [303]. Smart-farming [303]. Smoking [132]. Snakehead [219]. Social [523, 644, 301]. Socio [515, 358, 449]. Socio-Economic [449]. Socio-Economics [358]. sod [424]. Sodium [298]. Soft [340, 194]. Solid [466, 473]. Solms [429]. Solute [633]. Solutions [716]. Some [8, 429, 297, 109, 232, 409, 191, 268, 74, 70, 519, 478, 260]. Son [128]. Source [536, 233, 452, 447, 100]. Sources [250, 457, 418, 247, 560, 68, 87, 48, 495]. South [80, 120, 511, 334, 446, 86, 478, 165, 89, 106, 83, 118, 101, 123, 293, 524, 475, 477, 501, 64]. South-East [334]. south-eastern [524]. South-West [511]. South-Western [80, 446]. Southeast [149, 103, 720]. Southeastern [558]. Southern [388, 122]. Southwestern [489, 461]. Sovereignty [407]. Soy [466, 473]. Soya [313]. Soybean [243, 480, 479, 698]. Sp [134, 71, 409, 164, 333]. sp. [559]. Sparus [590, 495]. Spathopsis [458]. Spatial [314, 687, 422]. Spawn [659]. Spawning [241, 710, 399]. Spearmint [127]. Special [256, 272]. Species [388, 489, 406, 565, 382, 90, 519, 218, 25, 158, 156, 621, 581, 568, 38, 707, 719, 20, 727, 416, 405, 165, 249, 645, 650, 516, 600, 696, 562]. Speciesspecific [354]. Specific [620, 645, 650, 335]. Specifications [563]. Spectroscopy [108]. Speed [528]. Sperata [196]. Sperm [37]. Spices [308]. spiked [11]. Spiny [337]. spiralis [22]. Spiralothelphusa [123]. Spirulina [457, 343, 112]. Spoilage [585]. Sport [584]. Spot [2, 135]. Spotted [102, 234, 233, 441]. Spp [24, 544, 33, 497, 13]. Sri [417, 530]. Stability [527]. Stable [206]. Stage [217]. Stages [577, 517]. Starch [115, 157, 6]. Starving [357]. State [388, 80, 265, 312, 336, 466, 473, 159, 564, 216]. State/SC [388]. Statistical [229]. Status [471, 528, 331, 334, 491, 272, 420, 699, 280, 301]. Steam [318]. Steel [69]. Steelhead [555]. Steindachner [134]. Step [209]. Sterile [497]. Sterlet [85]. Steroid [81, 88]. Steroid-Inhibitor [88]. Stichopus [25]. Stimulant [385]. Stirring [710]. Stochastic [440]. Stock [303, 492, 659, 380, 395, 320]. Stocking [404, 92, 512, 56, 224, 566, 94, 267, 128, 490]. Stocks [687, 53, 356]. Storage [308, 309, 492]. Stored [69]. Story [423]. Strain [662, 484]. Strains [158, 345]. Stramonita [403]. Strategies [498, 402, 110, 669, 678]. Stream [125, 653]. Streptococcus [273]. Stress [61, 708, 75, 94, 325, 696]. Stretches [563]. Strobilanthes [333]. Strombus [362]. Structural [536, 255]. Structure [380, 395, 18, 118]. Studies [2, 308, 1, 261, 76, 284, 290, 436, 478]. Study [509, 81, 219, 239, 186, 295, 656, 4, 282, 523, 372, 279, 463, 647, 349, 517, 106, 226, 486, 210, 325, 88, 278, 101]. Sturgeon [81, 85, 90, 244, 137, 62, 76, 108, 53]. Sub [288, 383]. Sub-adult [288]. Subchronic [262]. Subcontinent [310, 196]. Sublethal [189]. subsp [256]. Substitution [277, 299]. Substrate [41]. Substrates [126]. Subsurface [658]. subtilis [526, 50]. Subtitute [164]. Suburban [186]. Success [399, 423]. Successful [188]. Sugars [137]. Suitability [120]. Sulawesi [120]. Sulfate [215, 483]. Sulfated [387, 112]. Sulphate [279]. Sumatra [185]. Sun [4]. Sundarban [156]. Super [391, 315, 269, 438]. Super-Intensive [438]. Supermarkets [321]. Suplementation [289]. suppa [164]. Supplement [161]. Supplementation [343, 526, 425, 427, 225, 227, 457, 568, 333, 298]. Supplemented [300, 708, 55, 556, 271]. Supplements [19, 222]. Supplies [411]. Supply [42, 564]. Surface [656]. Surrogate [651, 659]. Surveillance [486, 607]. Survey [523, 443]. Survivability [520]. Survival [262, 543, 61, 389, 329, 300, 39, 171, 700, 437, 436, 298, 71, 556, 454, 128, 364, 47, 187, 431, 484, 469, 98, 40]. Survive [580]. Susceptibility [205, 121]. Sustainability [691, 720, 688, 671, 702, 694, 500]. Sustainable [52, 141, 129, 729, 723, 676, 530, 660, 701, 508]. Swelling [590]. Swimming [193]. sycamore [274]. Sykes [196, 47]. Sympatric [209]. Syndrome [2, 135, 96]. synspilum [323]. Synthase [203]. Synthesizing [540]. Synthetic [247]. Synthetics [436]. System [389, 66, 190, 404, 512, 224, 52, 570, 622, 3, 407, 267, 214, 465, 488, 124, 315, 99, 269, 438, 206, 186]. Systems [146, 152, 673, 167, 228, 726, 504].

T [73, 292, 377]. T-Maze [73]. tabacum [384]. Tackle [304, 245]. Tag [356]. Tagging [183]. Tail [253, 255]. Takalar [180]. Tamil [706, 718, 118, 104]. Tana [489, 428]. Tandanus [19]. tanganicae [493]. Tank [219]. Tanks [328]. TaqMan(R) [11]. tarda [432]. Target [377]. Taxonomic [476, 276, 405]. TCBS [406]. Tea [425]. Technical [265, 159, 440]. Technique [654]. Techniques [606, 264, 585, 670]. Technologies [554, 734]. Technology [492, 224, 605, 474, 664, 41]. Telangana [552]. Teleost [279]. Temperature [248, 415, 580, 538, 413, 441, 555]. Temperatures [91, 187, 15]. Temporal [314, 687]. Ten [359]. Tenacibaculum [317]. Tendaho [297]. Tenggara [412]. Tenualosa [393, 381, 399]. Terapon [86]. Term [37, 279, 374, 492, 497]. Terrestrial [596]. Tesabela [715]. Test [207, 356]. Testicular [1]. Testing [2, 381]. Testosterone [34, 59]. testudineus [346]. tetradactylum [507, 517]. Textile [41]. Texture [344]. their [155, 91, 641, 55, 137, 336, 686, 457, 568, 707, 719, 607, 739]. Therapy [602]. there [578]. these [482]. Things [651]. Think [242]. Thoothukudi [516]. Threadfin [326]. Threatened [310, 669]. Three [235, 327, 360, 653, 439]. Thriving [712]. Thyroxine [227, 289]. Tibet [356]. ticto [156]. Tidal [528, 653]. Tiger [120, 295, 164, 56, 282, 506, 119, 291, 48, 101]. Tigris [322]. Tilapia [489, 543, 389, 375, 468, 140, 519, 197, 158, 235, 518, 50, 557, 55, 594, 243, 655, 553, 499, 336, 230, 421, 97, 263, 222, 457, 415, 94, 342, 698, 333, 267, 54, 349, 379, 444, 440, 43, 470, 128, 386, 325, 200, 355, 142, 537, 544, 299, 254, 343, 58, 562, 387, 677, 408, 98, 197, 268, 172, 306, 335]. Time [365, 11]. Tin [69]. Tire [337]. Tire-track [337]. Tissue [652, 448, 340]. Tissues [297, 11]. Toasted [271]. Tobacco [384]. Tocopherol [204]. Tolerance [95]. Tono [471]. Tool [239, 508]. Tools [634]. Tor [47]. Total [277, 495]. Toxic [612]. Toxicity [352, 148, 375, 32, 204, 1, 625, 648]. Toxicological [675]. Trace [297, 340]. Trachinotus [571]. track [337]. Tract [151]. Trade [291, 482]. Trading [482]. Traits [92]. Transcription [145, 144]. Transcriptome [385]. Transcripts [5]. Transform [108]. Transformation [206]. Transformative [683]. Transforming [711]. transmontanus [108]. Transport [300, 171, 384, 633]. Transportation [75]. Transpose [607]. Trawling [103]. Treated [308, 309, 370, 398, 698]. Treatment [167, 263, 514, 79, 694]. Treatments [379]. Trematode [226]. Trends [341]. Trevally [488]. Trewavas [298]. Trial [381]. Trials [18]. Trichogaster [68]. trichopterus [68]. trifolia [333]. Triggering [266]. Triticum [527]. Trizole [88]. Trophic [471]. Tropic [598]. Tropical [6, 12]. Trout [146, 150, 448, 345, 292, 205, 74, 75, 391, 247, 63, 208, 111, 78, 364, 79, 555, 538]. Trypanorhyncha [134]. Tuna [102, 233]. Tuna-Cooking [233]. Tunisia [344, 340]. Turbines [653]. Turkey [166, 317]. Turtle [139]. Twenty [399]. Twenty-Two [399]. Two [308, 90, 158, 156, 455, 328, 372, 318, 71, 399, 496, 95, 471, 565, 510]. Type [476, 488, 160]. Types [624, 610, 71, 142]. Typing [49]. Tyrrhenian [122].

UAV [711]. Uganda [554, 128, 18]. Ultra [391, 364]. Ultra-Fast [391]. Ultrastructural [609, 346]. Ultrastructure [151]. Ultraviolet [21, 411]. Ulva [392, 497, 222, 161]. Understanding [625, 648, 542, 377]. Unfamiliar [73]. Unfavorable [424]. Uninfected [133]. Units [18]. University [631]. Unlocking [722]. Unsaturated [37, 437]. Unusual [273]. Up-regulated [424]. Upon [707, 719]. Upper [471]. Uptake [286]. Urchin [183, 577]. Urgent [242]. urolepis [339, 339]. Use [392, 21, 155, 238, 314, 233, 360, 715, 366, 129, 27, 41, 99, 561]. used [73, 211, 524]. Uses [637]. Using [429, 34, 651, 190, 404, 207, 152, 153, 560, 108, 451, 395, 184, 611, 645, 650, 536, 91, 11, 659, 698, 35, 143, 494]. Utility [411]. Utilization [392, 543, 557, 55, 457, 324, 298, 717, 271, 660, 485, 206, 98, 299]. Utilizing [651]. UVC [21].

Valenciennes [173]. Validation [11, 419]. Vallisneria [22, 22]. Value [359, 458, 97, 54, 172]. Vaname [504, 490]. vannamei [66, 175, 424, 706, 718, 212, 57, 275, 214, 510, 149, 426, 525, 99, 269, 438, 431, 30, 490, 64, 95, 182, 181]. vannameiei [504]. Variability [687, 559, 395]. Varians [662]. Variation [186, 236, 192]. Variations [327, 340, 260, 528]. Varieties [558, 142]. Various [130, 624, 616, 409, 641, 345, 700, 622, 618]. Varying [59, 70]. Vea [471]. Vellar [718, 706, 165]. Vembanad [563]. Venericardia [344]. Ventral [63]. Venus [344]. vera [333]. verrucosa [344]. Versatile [672]. Vertebral [173]. Vertical [41]. Veterinary [647, 603, 664, 602, 646]. VHSV [21]. via [391]. Viability [2]. Vibrio [406, 382, 374, 175, 207, 282, 3, 38, 84, 431, 562]. Vibrionaceae [501]. Vibriosis [248]. Vietnam [221]. Villalobos [241]. Viral [21, 212]. viridis [165]. Virus [2, 21, 135, 212, 11, 467]. Virus-spiked [11]. Vision [711, 443]. Visual [73]. Vital [702, 701]. Vitamin [61, 37]. Vitellogenin [270]. Vitex [333]. Vitrification [76]. vitro [137, 85, 430, 432]. vittatus [236, 278]. Vivo [88, 250]. Vol [369]. volitans [89, 106]. vollenhovenii [461]. Volturno [96]. Volume [378]. Vugy [281]. vulgare [558, 487]. vulgaris [594, 447]. Vulnerabilities [542]. Vulnerable [500]. vulnificus [3].

Walbaum [538, 208]. Walking [257]. Waste [19, 742, 610, 54]. Wastewater [370, 661, 200]. Water [429, 297, 519, 288, 314, 276, 171, 241, 268, 152, 656, 56, 167, 621, 655, 455, 499, 195, 180, 372, 415, 575, 578, 94, 713, 223, 538, 214, 441, 189, 366, 260, 161, 595, 743, 184, 199, 266, 411, 149, 495, 315, 386, 504, 723, 525, 694, 41, 99, 269, 174, 516, 712, 154, 428, 658, 254, 561]. Water-Borne [167]. Waters [412, 633]. Watershed [43]. Waterside [109]. Waves [528]. Web [574]. Weed [58]. Weevil [717]. Weight [511, 109, 390, 86, 461, 464, 16, 505]. Welfare [708, 107, 487, 602, 213]. Well [411, 281, 245]. Wereda [174]. West [511, 475, 477, 185, 727]. Western [476, 80, 554, 445, 446, 478, 515]. Wet [197, 260]. Wetland [77, 658, 177]. Wetlands [563, 445]. Wheat [527]. White [2, 66, 135, 175, 424, 93, 373, 214, 51, 99, 269, 438, 431, 64, 95, 182, 181]. White-Leg [66, 93, 95]. Whole [74, 208]. Wild [73, 45, 360, 14, 482]. Will [530]. Winter [7, 214]. Wireless [190]. within [80, 146]. Without [392, 386]. Wood [499]. World [738, 623, 550]. Wound [398]. Wrasse [44]. WSSV [2, 419].

year [328]. year-Old [328]. Yeast [27, 9]. Yeast-Fermented [27]. Yellow [253, 385, 455, 209, 136]. Yellowfin [61]. yezoensis [206]. Yield [673, 558, 267, 527, 286]. Yielding [352]. Young [509, 223]. Youth [296]. yunnanensis [540].

Zambia [491, 474]. Zamboanga [445]. Zebrafish [609, 145]. Zeolites [368, 111]. Zero [94, 315, 41, 99, 269]. zeylanica [189]. zeylanicum [487]. Zinc [257, 483]. Zingiber [677]. Ziway [218]. Zone [226]. Zones [363]. Zooplankton [476, 552, 397, 499, 322]. ZWD [269].

References

Gawish:2010:PRA

[1] Azza M. Gawish. The protective role of alpha lipoic acid against pesticides induced testicular toxicity — histopathological and histochemical studies. Journal of Aquaculture Research & Development, 1(1): ??, ???? 2010. CODEN JARDCR. ISSN 2155-9546. URL https:

     //www.walshmedicalmedia.com/open-access/the-protective-role-
     of-alpha-lipoic-acid-against-pesticides-induced-testicular-toxicity-
     histopathological-and-histochemi-26026.html.

A:2010:IWS

[2] Devivaraprasad Reddy A, G. Jeyasekaran, and Jeya Shakila R. Incidence of white spot syndrome virus (WSSV) in Indian farmed frozen shrimp products and testing for viability through bio-inoculation studies. Journal of Aquaculture Research & Development, 1(1):1–5, ???? 2010. CODEN

     JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
     com/open-access/incidence-of-white-spot-syndrome-virus-wssv-

in-indian-farmed-frozen-shrimp-products-and-testing-for-viability-
through-bi-26027.html.

Lewis:2010:EFT

[3] Matthew Lewis, Scott Rikard, and Covadonga R. Arias. Evaluation of a flow-through depuration system to eliminate the human pathogen Vibrio vulnificus from oysters. Journal of Aquaculture Research & Development, 1(2):??, ???? 2010. CODEN JARDCR. ISSN 2155-9546. URL

https://www.walshmedicalmedia.com/open-access/evaluation-of-
a-flowthrough-depuration-system-to-eliminate-the-human-pathogen-
vibrio-vulnificus-from-oysters-26084.html.

Geetha:2010:IAC

[4] Bose Vijaya Geetha, Rajendran Navasakthi, and Ekambaram Padmini. Investigation of antioxidant capacity and phytochemical composition of Sun Chlorella — an invitro study. Journal of Aquaculture Research & Development, 1(2):1–7, ???? 2010. CODEN JARDCR. ISSN 2155-9546. URL

https://www.walshmedicalmedia.com/open-access/investigation-
of-antioxidant-capacity-and-phytochemical-composition-of-sun-
chlorella-an-invitro-study-26085.html.

Buonocore:2011:AEM

[5] Francesco Buonocore, Elisa Randelli, Niels Lorenzen, Katja Einer-Jensen, and Giuseppe Scapigliati. Analysis of the expression and modulation of selected mmunerelated gene transcripts in the DLEC cell line from European sea bass (Dicentrarchus labrax L.). Journal of Aquaculture Research & Development, 2(1):??, ???? 2011. CODEN JARDCR. ISSN 2155-9546.

URL https://www.walshmedicalmedia.com/open-access/analysis-
of-the-expression-and-modulation-of-selected-mmunerelated-gene-
transcripts-in-the-dlec-cell-line-from-european--26021.html.

Kumar:2011:GNG

[6] Vikas Kumar, N. P. Sahu, A. K. Pal, K. K. Jain, Shivendra Kumar, Vidya Sagar, Amit K. Sinha, and Jayant Ranjan. Gelatinized and nongelatinized corn starch based diet influence the fatty acid profile in the liver of tropical freshwater fish, Labeo ohita. Journal of Aquaculture Research & Development, 2(1):??, ???? 2011. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
gelatinized-and-nongelatinized-corn-starch-based-diet-influence-
the-fatty-acid-profile-in-the-liver-of-tropical-freshwat-26022.
html.

Beamish:2011:WIS

[7] R. Beamish, E. Gordon, J. Wade, B. Pennell, C. Neville, K. Lange, R. Sweeting, and S. Jones. The winter infection of sea lice on salmon in farms in a coastal inlet in British Columbia and possible causes. Journal of Aquaculture Research & Development, 2(1):??, ???? 2011. CODEN JARDCR. ISSN 2155-9546. URL https:/

/www.walshmedicalmedia.com/open-access/the-winter-infection-
of-sea-lice-on-salmon-in-farms-in-a-coastal-inlet-in-british-
columbia-and-possible-causes-26023.html.

Abdolazizi:2011:ECO

[8] Shahedah Abdolazizi, Edris Ghaderi, Noshin Naghdi, and Barzan Bahrami Kamangar. Effects of clove oil as an anesthetic on some hematological arameters of Carassius auratus. Journal of Aquaculture Research & Development, 2(1):??, ???? 2011. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
effects-of-clove-oil-as-an-anesthetic-on-some-hematological-
arameters-of-emcarassius-auratusem-26024.html.

Tewary:2011:OAB

[9] Arup Tewary and Bidhan C. Patra. Oral administration of baker’s yeast (Saccharomyces cerevisiae) acts as a growth promoter and immunomodulator in Labeo rohita (Ham.). Journal of Aquaculture Research & Development, 2(1):??, ???? 2011. CODEN JARDCR. ISSN 2155-

9546. URL https://www.walshmedicalmedia.com/open-access/oral-
administration-of-bakers-yeast-saccharomyces-cerevisiae-acts-
as-a-growth-promoter-and-immunomodulator-in-emlabeo-ro-26025.
html.

Lenzi:2011:RSP

[10] Mauro Lenzi. Resuspension of sediment as a possible environmental management method for coastal lagoons and aquaculture ponds. Journal of Aquaculture Research & Development, 2(2):??, ???? 2011. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/resuspension-of-sediment-as-a-possible-environmental-
management-method-for-coastal-lagoons-and-aquaculture-ponds-
26017.html.

Kibenge:2011:ISA

[11] Frederick S. B. Kibenge, Molly J. T. Kibenge, and Elmabrok Masaoud. Infectious Salmon Anaemia Virus (ISAV) ringtest: Validation of the ISAV diagnositic process using virus-spiked fish tissues and ISAV TaqMan(R) real-time RT-PCR. Journal of Aquaculture Research & Development, 2(2):??, ???? 2011. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
infectious-salmon-anaemia-virus-isav-ringtest-validation-of-
the-isav-diagnositic-process-using-virus-spiked-fish-tissues-
and-isav-taqman-real-time-rt-pcr-2155-9546.1000110.pdf.

Suloma:2011:AAM

[12] Ashraf Suloma and Hiroshi Y. Ogata. Arachidonic acid is a major component in gonadal fatty acids of tropical coral reef fish in the Philippines and Japan. Journal of Aquaculture Research & Development, 2(2): ??, ???? 2011. CODEN JARDCR. ISSN 2155-9546. URL https://

www.walshmedicalmedia.com/open-access/arachidonic-acid-is-a-
major-component-in-gonadal-fatty-acids-of-tropical-coral-reef-
fish-in-the-philippines-and-japan-26019.html.

Mohanty:2011:GPD

[13] S. Mohanty, P. K. Choudhury, A. Dash, M. Samanta, and N. K. Maiti. Genotypic and phenotypic diversity of bacillus spp. isolated from freshwater ecosystems. Journal of Aquaculture Research & Development, 2(2):??, ???? 2011. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
genotypic-and-phenotypic-diversity-of-bacillus-spp-isolated-
from-freshwater-ecosystems-26020.html.

Ovissipour:2011:FAA

[14] Mahmoudreza Ovissipour and Barbara Rasco. Fatty acid and amino acid profiles of domestic and wild beluga (Huso huso) roe and impact on fertilization ratio. Journal of Aquaculture Research & Development, 2(3): ??, ???? 2011. CODEN JARDCR. ISSN 2155-9546. URL https:/

/www.walshmedicalmedia.com/open-access/fatty-acid-and-amino-
acid-profiles-of-domestic-and-wild-beluga-huso-huso-roe-and-
impact-on-fertilization-ratio-26081.html.

Strand:2011:GEE

[15] Åsa Strand, Carin Magnhagen, and Anders Alanärä. Growth and energy expenditures of Eurasian perch Perca fluviatilis (Linnaeus) in different temperatures and of different body sizes. Journal of Aquaculture Research & Development, 2(3):??, ???? 2011. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
growth-and-energy-expenditures-of-eurasian-perch-perca-emfluviatilisem-
linnaeus-in-different-temperatures-and-of-differe-26082.html.

Rajesh:2011:LWR

[16] Dahare Rajesh. Length–weight relationship of a riverine fish Barilius barila (Gunther). Journal of Aquaculture Research & Development, 2(3):??, ???? 2011. CODEN JARDCR. ISSN 2155-9546. URL

https://www.walshmedicalmedia.com/open-access/length-weight-
relationship-of-a-riverine-fish-embarilius-barilaem-gunther-
26083.html.

Buonocore:2011:IFA

[17] Francesco Buonocore. Importance of fish antimicrobial peptides for aquaculture and biomedicine. Journal of Aquaculture Research & Development, 2(4):??, ???? 2011. CODEN JARDCR. ISSN 2155-9546. URL

https://www.walshmedicalmedia.com/open-access/importance-of-
fish-antimicrobial-peptides-for-aquaculture-and-biomedicine-
26122.html.

Ssanyu:2011:EPC

[18] Grace A. Ssanyu, Joseph Rasowo, Elmada Auma, and Moses Ndunguru. Evaluation of plankton community structure in fish refugia acting as Oreochromic niloticus propagation and nursery units for rice/ fish trials, Uganda. Journal of Aquaculture Research & Development, 2(4):??, ???? 2011. CODEN JARDCR. ISSN 2155-9546. URL

https://www.walshmedicalmedia.com/open-access/evaluation-of-
plankton-community-structure-in-fish-refugia-acting-as-oreochromic-
niloticus-propagation-and-nursery-units-26123.html.

Huynh:2011:EDS

[19] Huy P. V. Huynh and Dayanthi Nugegoda. Effects of dietary supplements on growth performance and phosphorus waste production of Australian catfish, Tandanus tandanus, fed with diets containing canola meal as fishmeal replacement. Journal of Aquaculture Research & Development, 2(4):??, ???? 2011. CODEN JARDCR. ISSN 2155-9546.

URL https://www.walshmedicalmedia.com/open-access/effects-
of-dietary-supplements-on-growth-performance-and-phosphorus-
waste-production-of-australian-catfish-emtandanus-ta-26184.html.

Ndi:2012:RDP

[20] Olasumbo L. Ndi and Mary D. Barton. Resistance determinants of Pseudomonas species from aquaculture in Australia. Journal of Aquaculture Research & Development, 3(1):??, ???? 2012. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
resistance-determinants-of-pseudomonas-species-from-aquaculture-
in-australia-26236.html.

Afonso:2012:UUC

[21] Luis O. B. Afonso, Zina Richmond, Alexandra A. Eaves, Jon Richard, Laura M. Hawley, and Kyle A. Garver. Use of ultraviolet C (UVC) radiation to inactivate infectious hematopoietic necrosis virus (IHNV) and viral hemorrhagic septicemia virus (VHSV) in fish processing plant effluent. Journal of Aquaculture Research & Development, 3(1):??, ???? 2012. CODEN JARDCR. ISSN 2155-9546. URL https:/

/www.walshmedicalmedia.com/open-access/use-of-ultraviolet-c-
uvc-radiation-to-inactivate-infectious-hematopoietic-necrosis-
virus-ihnv-and-viral-hemorrhagic-sept-26239.html.

Shelar:2012:EDO

[22] G. S. Shelar, H. D. Dhaker, D. I. Pathan, and M. M. Shirdhankar. Effect of different organic manures on the growth of screw vallisneria, Vallisneria spiralis Linne 1753. Journal of Aquaculture Research & Development, 3(1):??, ???? 2012. CODEN JARDCR.

ISSN 2155-9546. URL https://www.walshmedicalmedia.com/open-
access/effect-of-different-organic-manures-on-the-growth-of-
screw-vallisneria-emvallisneria-spiralisem-linne-1753-26289.
html.

Paulraj:2012:ACF

[23] Ambrose Paulraj, José Vitor Lima-Filho, Kareem Altaff, and Silvio Peixoto. Atypical chytridiomyces fungal infection in cultured Macrobrachium rosenbergii in India. Journal of Aquaculture Research & Development, 3(1):??, ???? 2012. CODEN JARDCR.

ISSN 2155-9546. URL https://www.walshmedicalmedia.com/open-
access/atypical-chytridiomyces-fungal-infection-in-cultured-
macrobrachium-rosenbergii-in-india-2155-9546.1000122.pdf.

Saavedra:2012:ARG

[24] M. J. Saavedra. Antibiotic resistance of the genus Aeromonas spp. Journal of Aquaculture Research & Development, 3(2):??, ???? 2012. CODEN JARDCR. ISSN 2155-9546. URL https://

www.walshmedicalmedia.com/open-access/antibiotic-resistance-
of-the-genus-aeromonas-spp-2155-9546.1000e101.pdf.

Ceesay:2012:BAH

[25] Abdoulie Ceesay, Mariana Nor Shamsudin, Norfarrah Mohammed Alipiah, and Intan Safinar Ismail. Biodiversity and adaptability of Holothuria leucospilota and Stichopus japonicus sea cucumber species in artificial environment. Journal of Aquaculture Research & Development, 3(2):??, ???? 2012. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
biodiversity-and-adaptability-of-holothuria-leucospilota-and-
stichopus-japonicus-sea-cucumber-species-in-artificial-environment-
2155-9546.1000123.pdf.

Ogunremi:2012:EAR

[26] J. B. Ogunremi, E. O. Faturoti, and O. I. Oladele. Effects of aquaculture researchers’ job characteristics on linkage activities in Nigeria. Journal of Aquaculture Research & Development, 3(2):??, ???? 2012. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/effects-of-aquaculture-researchers-job-characteristics-
on-linkage-activities-in-nigeria-26318.html.

Plaipetch:2012:UYF

[27] Pichet Plaipetch and Amararatne Yakupitiyage. Use of yeast-fermented canola meal to replace fishmeal in the diet of Asian sea bass Lates calcarifer (Bloch, 1790). Journal of Aquaculture Research & Development, 3(2):??, ???? 2012. CODEN JARDCR. ISSN 2155-9546. URL https://

www.walshmedicalmedia.com/open-access/use-of-yeastfermented-
canola-meal-to-replace-fishmeal-in-the-diet-of-asian-sea-bass-
emlates-calcariferem-bloch-1790-26319.html.

Jensen:2012:ACB

[28] Nathan R. Jensen, Jon J. Amberg, James A. Luoma, Liza R. Walleser, and Mark P. Gaikowski. Assessing consumption of bioactive micro-particles by filter-feeding Asian carp. Journal of Aquaculture Research & Development, 3(2):??, ???? 2012. CODEN JARDCR. ISSN 2155-9546. URL https://

www.walshmedicalmedia.com/open-access/assessing-consumption-
of-bioactive-micro-particles-by-filter-feeding-asian-carp-2155-
9546.1000126.pdf.

Ottesen:2012:DLB

[29] O. H. Ottesen, E. Dunaevskaya, and J. D’Arcy. Development of Labrus bergylta (Ascanius 1767) larvae from hatching to metamorphosis. Journal of Aquaculture Research & Development, 3(3):??, ???? 2012. CODEN JARDCR. ISSN 2155-9546. URL https://

www.walshmedicalmedia.com/open-access/development-of-labrus-
bergylta-ascanius-1767-larvae-from-hatching-to-metamorphosis-
26321.html.

Suarez:2012:EPL

[30] J. A. Suárez, G. Gaxiola, R. Posso, G. Garcia, G. Alanis, A. Suárez, J. Faillace, and G. Cuzon. Energy partitioning by Litopenaeus vannamei (Boone, 1931) given plant protein diets. Journal of Aquaculture Research & Development, 3(3):??, ???? 2012. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
energy-partitioning-by-emlitopenaeus-vannameiem-boone-1931-given-
plant-protein-diets-26322.html.

Chowdhary:2012:EPR

[31] Shipra Chowdhary, P. P. Srivastava, Suman Mishra, R. Dayal, A. K. Yadav, and J. K. Jena. Evaluation of partial replacement of dietary animal protein from plant protein blended with glucosamine on growth and body indices of Asian catfish (Clarias batrachus) fingerlings. Journal of Aquaculture Research & Development, 3(3):??, ???? 2012. CODEN JARDCR. ISSN 2155-9546. URL https://

www.walshmedicalmedia.com/open-access/evaluation-of-partial-
replacement-of-dietary-animal-protein-from-plant-protein-blended-
with-glucosamine-on-growth-and-body-indices-of-asian-catfish-
clarias-batrachus-fingerlings-2155-9546.1000129.pdf.

Allayie:2012:PEA

[32] Sartaj Ahmad Allayie, S. Hemalatha, C. Elanchezhiyan, V. Manoharan, N. Silambarasan, K. Balasubramanian, N. Raaman, and C. Siveraj. Protective efficacy of aqueous extract of Naringi crenulata leaves against shampoo (endocrine disruptor) induced toxicity in Labeo rohita. Journal of Aquaculture Research & Development, 3(3):??, ???? 2012. CODEN JARDCR. ISSN 2155-9546. URL https:

//www.walshmedicalmedia.com/open-access/protective-efficacy-
of-aqueous-extract-of-naringi-crenulata-leaves-against-shampoo-
endocrine-disruptor-induced-toxicity-in-labeo-rohita-2155-9546.
1000130.pdf.

Dias:2012:ARP

[33] Carla Dias, Vânia Mota, António Martinez-Murcia, and Maria JoséSaavedra. Antimicrobial resistance patterns of Aeromonas spp. isolated from ornamental fish. Journal of Aquaculture Research & Development, 3 (3):??, ???? 2012. CODEN JARDCR. ISSN 2155-9546. URL

https://www.walshmedicalmedia.com/open-access/antimicrobial-
resistance-patterns-of-aeromonas-spp-isolated-from-ornamental-
fish-2155-9546.1000131.pdf.

Beaven:2012:AMP

[34] Utete Beaven and Victor Muposhi. Aspects of a monosex population of Oreochromis niloticus fingerlings produced using 17-α methyl testosterone hormone. Journal of Aquaculture Research & Development, 3(3):

??, ???? 2012. CODEN JARDCR. ISSN 2155-9546. URL https:/

/www.walshmedicalmedia.com/open-access/aspects-of-a-monosex-
population-of-oreochromis-niloticus-fingerlings-produced-using-
17-methyl-testosterone-hormone-2155-9546.1000132.pdf.

Olumuji:2012:IBA

[35] O. K. Olumuji and M. K. Mustapha. Induced breeding of African mud catfish, Clarias gariepinus (Burchell 1822), using different doses of normal saline diluted ovaprim. Journal of Aquaculture Research & Development, 3(4):??, ???? 2012. CODEN JARDCR. ISSN 2155-9546. URL https:

//www.walshmedicalmedia.com/open-access/induced-breeding-of-
african-mud-catfish-emclarias-gariepinusem-burchell-1822-using-
different-doses-of-normal-saline-dilu-26435.html.

Srivastava:2012:BLR

[36] Prem Prakash Srivastava, Sudhir Raizada, Rajesh Dayal, Shipra Chowdhary, Wazir Singh Lakra, Akhilesh Kumar Yadav, Priya Sharma, and Jyoti Gupta. Breeding and larval rearing of Asian catfish, Clarias batrachus (Linnaeus, 1758) on live and artificial feed. Journal of Aquaculture Research & Development, 3(4):??, ???? 2012. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
breeding-and-larval-rearing-of-asian-catfish-emclarias-batrachusem-
linnaeus-1758-on-live-and-artificial-feed-26434.html.

Imanpoor:2012:LTE

[37] Hanaee Kashani Z. Mohammad Reza Imanpoor, Mohammad Reza Imanpoor, Ali Shabani, and Saeed Gorgin. Long-term effects and interactions of different levels of dietary vitamin C and E and highly unsaturated fatty acid on sperm parameters in goldfish (Carassius auratus gibelio). Journal of Aquaculture Research & Development, 3(4): ??, ???? 2012. CODEN JARDCR. ISSN 2155-9546. URL https:/

/www.walshmedicalmedia.com/open-access/longterm-effects-and-
interactions-of-different-levels-of-dietary-vitamin-c-and-e-
and-highly-unsaturated-fatty-acid-on-sp-26433.html.

Leyton:2012:DPS

[38] Yanett Leyton, Jorge Borquez, José Darias, Mercedes Cueto, A. R. Díaz-Marrero, and Carlos Riquelme. Diketopiperazines produced by an [sic] Bacillus species inhibits Vibrio parahaemolyticus. Journal of Aquaculture Research & Development, 3(4):1–5, ???? 2012. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/diketopiperazines-produced-by-an-embacillus-
speciesem-inhibits-emvibrio-parahaemolyticusem-26479.html.

Chambers:2012:CGS

[39] Michael Chambers, Jon Bunker, Win Watson III, and W. H. Howell. Comparative growth and survival of juvenile Atlantic cod (Gadus morhua) cultured in copper and nylon net pens. Journal of Aquaculture Research & Development, 3(5):??, ???? 2012. CODEN JARDCR.

ISSN 2155-9546. URL https://www.walshmedicalmedia.com/open-
access/comparative-growth-and-survival-of-juvenile-atlantic-
cod-emgadus-morhuaemcultured-in-copper-and-nylon-net-pens-26469.
html.

Yadav:2012:GRA

[40] Akhilesh Kumar Yadav, Prem Prakash Srivastava, P. Shrivastava, Shipra Chowdhary, Rajesh Dayal, and Joy Krushna Jena. Growth responses of animal and plant origin dietary lipids on the survival, growth and feed efficiency of Asian catfish, Clarias batrachus (Linnaeus, 1758) grow-out. Journal of Aquaculture Research & Development, 3(5):??, ???? 2012. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/growth-responses-of-animal-and-plant-origin-
dietary-lipids-on-the-survivalgrowth-and-feed-efficiency-of-
asian-catfish-cl-26472.html.

Suantika:2012:UZW

[41] Gede Suantika, Dea Indriani Astuti, Rifki R. Arief, Malendra Rusni, and Osman R. Turendro. Use of zero water discharge technology through the application of nitrifying bacteria and textile vertical substrate in grow-out phase of Macrobrachium rosenbergii De Man. Journal of Aquaculture Research & Development, 3(5):??, ???? 2012. CODEN JARDCR. ISSN 2155-

9546. URL https://www.walshmedicalmedia.com/open-access/use-
of-zero-water-discharge-technology-through-the-application-of-
nitrifying-bacteria-and-textile-vertical-substrate-in--26470.
html.

Figueiredo:2012:CDD

[42] Joana Figueiredo, Junda Lin, Justin Anto, and Luís Narciso. The consumption of DHA during embryogenesis as an indicative of the need to supply DHA during early larval development: a review. Journal of Aquaculture Research & Development, 3(5):??, ???? 2012. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/the-consumption-of-dha-during-embryogenesis-
as-an-indicative-of-the-need-to-supply-dha-during-early-larval-
development-a-26471.html.

Rahman:2012:ETC

[43] Muhammad Mizanur Rahman, Mostafa Shamsuzzaman, Saad Mahmood, Subrata Sarker, and Faruk Alam. Economics of tilapia culture in watershed pond in Bangladesh. Journal of Aquaculture Research & Development, 3(5):??, ???? 2012. CODEN JARDCR. ISSN 2155-9546.

URL https://www.walshmedicalmedia.com/open-access/economics-
of-tilapia-culture-in-watershed-pond-in-bangladesh-26487.html.

Dunaevskaya:2012:OBW

[44] Evgenia Dunaevskaya, Anil B. Amin, and Oddvar H. Ottesen. Organogenesis of ballan wrasse Labrus bergylta (Ascanius 1767) larvae. Journal of Aquaculture Research & Development, 3(5):1–6, ???? 2012. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/organogenesis-of-ballan-wrasse-emlabrus-bergyltaem-
emascanius-1767em-larvae-26478.html.

Mancuso:2012:PEB

[45] Monique Mancuso. Photobacteriosis exchange between wild and farmed fish in the Mediterranean area. Journal of Aquaculture Research & Development, 3(6):??, ???? 2012. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
photobacteriosis-exchange-between-wild-and-farmed-fish-in-the-
mediterranean-area-26621.html.

Nizza:2012:WFO

[46] Antonino Nizza. What future for organic aquaculture? Journal of Aquaculture Research & Development, 3(6):??, ???? 2012. CODEN JARDCR.

ISSN 2155-9546. URL https://www.walshmedicalmedia.com/open-
access/what-future-for-organic-aquaculture-2155-9546.1000e103.
pdf.

Singh:2012:EFE

[47] Soibam Khogen Singh, Upasana Mishra, Sibnarayan Dam Roy, N. K. Chadha, and G. Venkateshwarlu. Effect of feeding enriched formulated diet and live feed on growth, survival and fatty acid profile of deccan mahseer, Tor khudree (Sykes) first feeding fry. Journal of Aquaculture Research & Development, 3(6):??, ???? 2012. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
effect-of-feeding-enriched-formulated-diet-and-live-feed-on-
growth-survival-and-fatty-acid-profile-of-deccan-mahseer-emt-
26480.html.

Preetha:2012:EDN

[48] V. V. Preetha, A. Belayneh, A. Palavesam, G. Immanuel, and Z. Tadesse. Effect of dietary nutrient sources on nitrogen and phosphorus loading from culture of tiger shrimp (Penaeus monodon). Journal of Aquaculture Research & Development, 3(6):??, ???? 2012. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
effect-of-dietary-nutrient-sources-on-nitrogen-and-phosphorus-
loading-from-culture-of-tiger-shrimp-empenaeus-monodonem-26473.
html.

Sarkar:2012:ITA

[49] Agniswar Sarkar, Mousumi Saha, and Pranab Roy. Identification and typing of Aeromonas hydrophila through 16S rDNA-PCR fingerprinting. Journal of Aquaculture Research & Development, 3(6):??, ???? 2012. CODEN JARDCR. ISSN 2155-9546. URL https://www.

walshmedicalmedia.com/open-access/identification-and-typing-
of-emaeromonas-hydrophilaem-through-16s-rdnapcr-fingerprinting-
26604.html.

Elkamel:2012:INT

[50] Ahmad A. Elkamel and Gamal M. Mosaad. Immunomodulation of Nile tilapia, Oreochromis niloticus, by Nigella sativa and Bacillus subtilis. Journal of Aquaculture Research & Development, 3(6):??, ???? 2012. CODEN JARDCR. ISSN 2155-9546. URL https:

//www.walshmedicalmedia.com/open-access/immunomodulation-of-
nile-tilapia-emoreochromis-niloticusem-by-nigella-sativa-and-
embacillus-subtilisem-26486.html.

Sivakamavalli:2012:PII

[51] Jeyachandran Sivakamavalli, Perumal Rajakumaran, and Baskaralingam Vaseeharan. Prophenoloxidase and immune indices of Indian white shrimp Fenneropenaeus indicus. Journal of Aquaculture Research & Development, 3(6):??, ???? 2012. CODEN JARDCR. ISSN 2155-9546. URL https:/

/www.walshmedicalmedia.com/open-access/prophenoloxidase-and-
immune-indices-of-indian-white-shrimp-fenneropenaeus-indicus-
26489.html.

Hasan:2012:OFE

[52] B. M. A. Hasan, B. Guha, and S. Datta. Optimization of feeding efficiency in semi-intensive farming system for sustainable and cost effective production of Penaeus monodon Fabricius. Journal of Aquaculture Research & Development, 3(6):??, ???? 2012. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
optimization-of-feeding-efficiency-in-semiintensive-farming-
system-for-sustainable-and-cost-effective-production-of-empe-
26488.html.

Ovissipour:2012:SCC

[53] Mahmoudreza Ovissipour and Barbara Rasco. Sturgeon: Conservation of Caspian Sea stocks. Journal of Aquaculture Research & Development, 3 (7):??, ???? 2012. CODEN JARDCR. ISSN 2155-9546. URL https://

www.walshmedicalmedia.com/open-access/sturgeon-conservation-
of-caspian-sea-stocks-2155-9546.1000e104.pdf.

Monteiro:2012:PAV

[54] Maria Lúcia G. Monteiro, Eliane T. Mársico, Izolda Martins Viriato, José Marcelino Lima de Souza, and Carlos A. Conte Júnior. Preparation of added value byproducts from the waste material of tilapia (Oreochromis niloticus) processing. Journal of Aquaculture Research & Development, 3(7):??, ???? 2012. CODEN JARDCR.

ISSN 2155-9546. URL https://www.walshmedicalmedia.com/open-
access/preparation-of-added-value-byproducts-from-the-waste-
material-of-tilapia-oreochromis-niloticus-processing-26573.html.

Gaber:2012:UCM

[55] Magdy M. Gaber, Magdy M. Elhalfawy, and Amal M. Ramadan. Utilization of cottonseed meal supplemented with iron for detoxification of gossypol in Nile tilapia, broodstock and their impact on the hatchability of their progenies. Journal of Aquaculture Research & Development, 3(7):??, ???? 2012. CODEN JARDCR. ISSN 2155-9546. URL https://www.

walshmedicalmedia.com/open-access/utilization-of-cottonseed-
meal-supplemented-with-iron-for-detoxification-of-gossypol-in-
nile-tilapia-broodstock-and-thei-26574.html.

Gaber:2012:ESD

[56] M. M. Gaber, E. A. Omar, M. Abdel-Rahim, A. M. Nour, M. A. Zaki, and T. M. Srour. Effects of stocking density and water exchange rates on growth performance of tiger shrimp, Penaeus semisulcatus cultured in earthen ponds. Journal of Aquaculture Research & Development, 3(7):??, ???? 2012. CODEN JARDCR.

ISSN 2155-9546. URL https://www.walshmedicalmedia.com/open-
access/effects-of-stocking-density-and-water-exchange-rates-
on-growth-performance-of-tiger-shrimp-penaeus-semisulcatus-cultured-
26605.html.

Intriago:2012:NFR

[57] P. Intriago, J. Espinoza, J. Cabrera, A. Sanchez, and A. Navarrete. Nitrogen flow in a recirculating operation of Litopenaeus vannamei maturation in Ecuador. Journal of Aquaculture Research & Development, 3(8):??, ???? 2012. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
nitrogen-flow-in-a-recirculating-operation-of-emlitopenaeus-
vannameiem-maturation-in-ecuador-26582.html.

Bag:2012:AWP

[58] Mukti Pada Bag, Subhash Chandra Mahapatra, and Pavuluri Srinivasa Rao. Aquatic weed as potential feed for Mozambique tilapia, Oreochromis mossambicus. Journal of Aquaculture Research & Development, 3(8):??, ???? 2012. CODEN JARDCR. ISSN 2155-9546.

URL https://www.walshmedicalmedia.com/open-access/aquatic-
weed-as-potential-feed-for-mozambique-tilapia-emoreochromis-
mossambicusem-26633.html.

Kefi:2012:GRS

[59] A. S. Kefi, J. Kang’ombe, D. Kassam, and C. Katongo. Growth, reproduction and sex ratios in Oreochromis andersonii (Castelnau, 1861) fed with varying levels of 17α-methyl testosterone. Journal of Aquaculture Research & Development, 3(8):??, ???? 2012. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
growth-reproduction-and-sex-ratios-in-emoreochromis-andersoniiem-
castelnau-1861-fed-with-varying-levels-of-17945methyl-t-26625.
html.

Asusena:2012:EEA

[60] Aguiñaga-Cruz Jazmín Asusena, Sainz-Hernández Juan Carlos, Fierro-Coronado Jesús Arturo, and Diarte-Plata Genaro. The effects of eyestalk ablation on the reproductive and immune function of female Macrobrachium americanum. Journal of Aquaculture Research & Development, 3(8):??, ???? 2012. CODEN JARDCR. ISSN 2155-

9546. URL https://www.walshmedicalmedia.com/open-access/the-
effects-of-eyestalk-ablation-on-the-reproductive-and-immune-
function-of-female-emmacrobrachium-americanumem-26623.html.

Adloo:2012:EFE

[61] Mohammad Nabi Adloo, Abbas Matinfar, and Iman Sourinezhad. Effects of feeding enriched Artemia fransiscana with HUFA, vitamin C and E on growth performance, survival and stress resistance of yellowfin seabream larvae. Journal of Aquaculture Research & Development, 3(8):??, ???? 2012. CODEN JARDCR. ISSN 2155-9546.

URL https://www.walshmedicalmedia.com/open-access/effects-
of-feeding-enriched-artemia-fransiscana-with-hufa-vitamin-c-
and-e-on-growth-performance-survival-and-stress-resi-26624.html.

Hosseinzadeh:2013:SCS

[62] Mahboubeh Hosseinzadeh, Mohammad Reza Imanpoor, Ali Shabani, and Hamed Nekoubin. Seasonal changes in serum calcium and 17β-estradiol levels in Persian sturgeon, Acipenser persicus. Journal of Aquaculture Research & Development, 4(1):??, ???? 2013. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
seasonal-changes-in-serum-calcium-and-17946estradiol-levels-
in-persian-sturgeon-emacipenser-persicusem-26663.html.

Mustapha:2013:EED

[63] Aba Mustapha, Belghyti Driss, and Benabid Mohamed. Effects of extruded diets with different energy levels on body composition of fat content in different parts of dorsal, ventral of fillet of rainbow trout (Oncorhynchus mykiss). Journal of Aquaculture Research & Development, 4(1):??, ???? 2013. CODEN JARDCR. ISSN 2155-9546. URL https://www.

walshmedicalmedia.com/open-access/effects-of-extruded-diets-
with-different-energy-levels-on-body-composition-of-fat-content-
in-different-parts-of-dorsal-v-26665.html.

Varadharajan:2013:FFH

[64] D. Varadharajan and N. Pushparajan. Food and feeding habits of aquaculture candidate a potential crustacean of Pacific white shrimp Litopenaeus vannamei, South East Coast of India. Journal of Aquaculture Research & Development, 4(1):??, ???? 2013. CODEN JARDCR. ISSN 2155-9546.

URL https://www.walshmedicalmedia.com/open-access/food-and-
feeding-habits-of-aquaculture-candidate-a-potential-crustacean-
of-pacific-white-shrimp-emlitopenaeus-vannameiem-26664.html.

He:2013:ICP

[65] Lin He, Jing Xie, Qing Li, Yunlong Zhao, Yaping Wang, and Qun Wang. Isolation and characterization of polymorphic microsatellite loci in the redclaw crayfish, Cherax quadricarinatus. Journal of Aquaculture Research & Development, 4(1):??, ???? 2013. CODEN JARDCR. ISSN 2155-9546.

URL https://www.walshmedicalmedia.com/open-access/isolation-
and-characterization-of-polymorphic-microsatellite-loci-in-the-
redclaw-crayfish-emcherax-quadricarinatusem-26666.html.

Bojorquez-Mascareno:2013:END

[66] E. I. Bojórquez-Mascareño and M. F. Soto-Jiménez. Effects of natural diet on growth on white-leg shrimp Litopenaeus vannamei under experimental mesocosms emulating an intensive culture system. Journal of Aquaculture Research & Development, 4(1):??, ???? 2013. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
effects-of-natural-diet-on-growth-on-white-leg-shrimp-litopenaeus-
vannamei-under-experimental-mesocosms-emulating-an-intensive-
culture-system-2155-9546.1000163.pdf.

Buonocore:2013:MBD

[67] Francesco Buonocore. Marine biotechnology: Developments and perspectives. Journal of Aquaculture Research & Development, 4(2):??, ???? 2013. CODEN JARDCR. ISSN 2155-9546. URL https:/

/www.walshmedicalmedia.com/open-access/marine-biotechnology-
developments-and-perspectives-2155-9546.1000e105.pdf.

Mohanta:2013:EDA

[68] Kedar Nath Mohanta, Sankaran Subramanian, and Veeratayya Sidweerayya Korikanthimath. Evaluation of different animal protein sources in formulating the diets for blue gourami, Trichogaster trichopterus fingerlings. Journal of Aquaculture Research & Development, 4(2):??, ???? 2013. CODEN JARDCR. ISSN 2155-9546. URL

https://www.walshmedicalmedia.com/open-access/evaluation-of-
different-animal-protein-sources-in-formulating-the-diets-for-
blue-gourami-trichogaster-trichopterus-finge-26693.html.

Pushparajan:2013:FFS

[69] N. Pushparajan, D. Varadharajan, and P. Soundarapandian. Feasibility of fish and shellfish curries stored in indigenous polymer coated tin free steel cans. Journal of Aquaculture Research & Development, 4(2): ??, ???? 2013. CODEN JARDCR. ISSN 2155-9546. URL https:

//www.walshmedicalmedia.com/open-access/feasibility-of-fish-
and-shellfish-curries-stored-in-indigenous-polymer-coated-tin-
free-steel-cans-26694.html.

Ozovehe:2013:GPH

[70] Bello Nuhu Ozovehe. Growth performance, haematological indices and some biochemical enzymes of juveniles Clarias gariepinus (Burchell 1822) fed varying levels of Moringa oleifera leaf meal diet. Journal of Aquaculture Research & Development, 4(2):??, ???? 2013. CODEN JARDCR.

ISSN 2155-9546. URL https://www.walshmedicalmedia.com/open-

access/growth-performance-haematological-indices-and-some-biochemical-
enzymes-of-juveniles-emclarias-gariepinusem-burchell-1822-26695.
html.

Nekoubin:2013:EDT

[71] Hamed Nekoubin and Mohammad Sudagar. Effect of different types of plants (Lemna sp., Azolla filiculoides and alfalfa) and artificial diet (with two protein levels) on growth performance, survival rate, biochemical parameters and body composition of grass carp (Ctenopharyngodon idella). Journal of Aquaculture Research & Development, 4(2):1–6, ???? 2013. CODEN JARDCR. ISSN 2155-9546. URL https://www.

walshmedicalmedia.com/open-access/effect-of-different-types-
of-plants-emlemnaem-sp-emazolla-filiculoidesem-and-emalfalfaem-
and-artificial-diet-with-two-pr-26696.html.

Magnesen:2013:EDF

[72] Thorolf Magnesen, Anita Jacobsen, and Malebo Hellen Moepi. Effect of different filter methods on seawater quality at a marine scallop hatchery. Journal of Aquaculture Research & Development, 4(2): ??, ???? 2013. CODEN JARDCR. ISSN 2155-9546. URL https:

//www.walshmedicalmedia.com/open-access/effect-of-different-
filter-methods-on-seawater-quality-at-a-marine-scallop-hatchery-
26697.html.

Benhaim:2013:UCU

[73] David Benhaïm, Marie-Laure Bégout, and Béatrice Chatain. Unfamiliar congener used as a visual attractor in wild caught and domesticated sea bass (Dicentrarchus labrax) placed in a T-maze. Journal of Aquaculture Research & Development, 4(2):??, ???? 2013. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
unfamiliar-congener-used-as-a-visual-attractor-in-wild-caught-
and-domesticated-sea-bass-emdicentrarchus-labraxem-placed--26699.
html.

Kalantarian:2013:EDL

[74] S. H. Kalantarian, G. H. Rafiee, M. Farhangi, and Mojazi Amiri B. Effect of different levels of dietary calcium and potassium on growth indices, biochemical composition and some whole body minerals in rainbow trout (Oncorhynchus mykiss) fingerlings. Journal of Aquaculture Research & Development, 4(3):??, ???? 2013. CODEN JARDCR. ISSN 2155-9546.

URL https://www.walshmedicalmedia.com/open-access/effect-of-
different-levels-of-dietary-calcium-and-potassium-on-growth-
indices-biochemical-composition-and-some-whole-bod-26777.html.

Keysomi:2013:ECR

[75] M. M. Esfandani Keysomi, M. Sudagar, and Nasirir Asl. Effect of citalopram on reducing transportation stress in rainbow trout (Oncorhynchus mykiss). Journal of Aquaculture Research & Development, 4(3):1–3, ???? 2013. CODEN JARDCR. ISSN 2155-9546. URL https:/

/www.walshmedicalmedia.com/open-access/effect-of-citalopram-
on-reducing-transportation-stress-in-rainbow-troutemoncorhynchus-
mykissem-26778.html.

Keivanloo:2013:FSV

[76] Saeide Keivanloo and Mohammad Sudagar. Feasibility studies on vitrification of Persian sturgeon (Acipenser persicus) embryos. Journal of Aquaculture Research & Development, 4(3):??, ???? 2013. CODEN JARDCR.

ISSN 2155-9546. URL https://www.walshmedicalmedia.com/open-
access/feasibility-studies-on-vitrification-of-persian-sturgeon-
emacipenser-persicusemembryos-26779.html.

Bbalali:2013:RBN

[77] Saeed Bbalali, Seyed Abbas Hoseini, Rasool Ghorbani, and Hamideh Kordi. Relationships between nutrients and chlorophyll a concentration in the International Alma Gol Wetland, Iran. Journal of Aquaculture Research & Development, 4(3):??, ???? 2013. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
relationships-between-nutrients-and-chlorophyll-a-concentration-
in-the-international-alma-gol-wetland-iran-26783.html.

Rodrigues:2013:CBA

[78] Bruna Leal Rodrigues, Thiago Silveira Alvares, Marion Pereira da Costa, Guilherme Sicca Lopes Sampaio, César Aquiles Lázaro de la Torre, Eliane Teixeira Mársico, and Carlos Adam Conte Júnior. Concentration of biogenic amines in rainbow trout (Oncorhynchus mykiss) preserved in ice and its relationship with physicochemical parameters of quality. Journal of Aquaculture Research & Development, 4(3):??, ???? 2013. CODEN JARDCR. ISSN 2155-9546. URL https://www.

walshmedicalmedia.com/open-access/concentration-of-biogenic-
amines-in-rainbow-trout-emoncorhynchus-mykissempreserved-in-
ice-and-its-relationship-with-phys-26784.html.

Soheil:2013:EFA

[79] Lameeihassankiadeh Soheil, Khara Hossein, Nezamibaloochi Shabanali, Borani Mohammad, Mohammadalikhani Mehdi, and Abbasian Firouz. The effects of folic acid treatment on biometric and blood parameters of fingerling rainbow trout fishes (Oncorhynchus mykiss). Journal of Aquaculture Research & Development, 4(3):??, ???? 2013. CODEN JARDCR.

ISSN 2155-9546. URL https://www.walshmedicalmedia.com/open-
access/the-effects-of-folic-acid-treatment-on-biometric-and-
blood-parameters-of-fingerling-rainbow-trout-fishes-emoncorhynchus-
-26785.html.

Akinpelu:2013:GAP

[80] O. M. Akinpelu, A. A. Ayeloja, F. O. A. George, G. L. Adebisi, W. A. Jimoh, and S. D. Idowu. Gender analysis of processing activities among commercial catfish processors within Ibadan Metropolis, Oyo State South-Western Nigeria. Journal of Aquaculture Research & Development, 4(3):??, ???? 2013. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
gender-analysis-of-processing-activities-among-commercial-catfish-
processors-within-ibadan-metropolis-oyo-state-southwes-26788.
html.

Aramli:2013:CSS

[81] Mohammad Sadegh Aramli, Mohammad Reza Kalbassi, and Rajab Mohammad Nazari. Comparative study of sex steroid levels of Persian sturgeon, Acipenser persicus males in responding negative and positive to LHRH-A_{2} hormone. Journal of Aquaculture Research & Development, 4(3):??, ???? 2013. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
comparative-study-of-sex-steroid-levels-of-persian-sturgeon-
emacipenser-persicusem-males-in-responding-negative-and-posi-
26786.html.

Lenzi:2013:FA

[82] Mauro Lenzi. The future of aquaculture. Journal of Aquaculture Research & Development, 4(4):??, ???? 2013. CODEN JARDCR. ISSN 2155-

9546. URL https://www.walshmedicalmedia.com/open-access/the-
future-of-aquaculture-2155-9546.1000e106.pdf.

Srilatha:2013:DFF

[83] G. Srilatha, P. Mayavu, D. Varadharajan, and K. Chamundeeswari. Distribution of fin-fish eggs and larvae from Point Calimere and Muthupettai, South East Coast of India. Journal of Aquaculture Research & Development, 4(4):??, ???? 2013. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
distribution-of-finfish-eggs-and-larvae-from-point-calimere-
and-muthupettai-south-east-coast-of-india-27047.html.

Leyton:2013:OAD

[84] Yanett Leyton and Carlos Riquelme. Oleic acid and diketopiperazines produced by marine bacteria reduce the load of the pathogen Vibrio parahaemolyticus in Argopecten purpuratus. Journal of Aquaculture Research & Development, 4(4):??, ???? 2013. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
oleic-acid-and-diketopiperazines-produced-by-marine-bacteria-
reduce-the-load-of-the-pathogen-emvibrio-parahaemolyticusem-
26867.html.

Azarin:2013:EDV

[85] Hajar Azarin, Mohammad Reza Imanpoor, and Mohammad Pourdehghani. Effect of 17α, 20β-dihydroxyprogesterone on in vitro oocyte maturation in Persian sturgeon (Acipenser persicus) and sterlet (Acipenser ruthenus). Journal of Aquaculture Research & Development, 4(4):1==4, ???? 2013. CODEN JARDCR. ISSN 2155-9546.

URL https://www.walshmedicalmedia.com/open-access/effect-of-
17945-20946dihydroxyprogesterone-on-emin-vitroem-oocyte-maturation-
in-persian-sturgeon-emacipenser-persicusem--26868.html.

Manoharan:2013:LWR

[86] J. Manoharan, A. Gopalakrishnan, D. Varadharajan, C. Udayakumar, and S. Priyadharsini. Length–weight relationship of crescent perch Terapon jarbua (Forsskal) from Parangipettai Coast, South East Coast of India. Journal of Aquaculture Research & Development, 4(4):??, ???? 2013. CODEN JARDCR. ISSN 2155-9546. URL https://www.

walshmedicalmedia.com/open-access/lengthweight-relationship-
of-crescent-perch-terapon-jarbua-forsskalfrom-parangipettai-
coast-south-east-coast-of-india-26869.html.

Orire:2013:GBC

[87] A. M. Orire, S. O. Omotoyinbo, and S. O. E. Sadiku. The growth and body composition of Clarias gariepinus fingerlings fed combined different sources of lipid. Journal of Aquaculture Research & Development, 4 (4):??, ???? 2013. CODEN JARDCR. ISSN 2155-9546. URL https:

//www.walshmedicalmedia.com/open-access/the-growth-and-body-
composition-of-emclarias-gariepinusem-fingerlings-fed-combined-
different-sources-of-lipid-26982.html.

Srivastava:2013:SVE

[88] Pallavi Srivastava and Ajay Singh. Study of in-vivo effects caused by metabolites (1,2,4-trizole alanine) of steroid-inhibitor fungicide on aquatic life (fish). Journal of Aquaculture Research & Development, 4(4):??, ???? 2013. CODEN JARDCR. ISSN 2155-9546.

URL https://www.walshmedicalmedia.com/open-access/study-of-
invivo-effects-caused-by-metabolites-124trizole-alanine-of-steroidinhibitor-
fungicide-on-aquatic-life-fish-26983.html.

Priyadharsini:2013:FFH

[89] S. Priyadharsini, A. Subramaniyan, J. Manoharan, and D. Varadharajan. Food and feeding habits of red lionfish Pterois volitans from Cuddalore Coast, South East Coast of India. Journal of Aquaculture Research & Development, 4(4):??, ???? 2013. CODEN JARDCR. ISSN 2155-9546.

URL https://www.walshmedicalmedia.com/open-access/food-and-
feeding-habits-of-red-lionfish-empterois-volitansem-from-cuddalore-
coast-south-east-coast-of-india-27048.html.

Azzouzi:2013:CRH

[90] N. Azzouzi, M. M. Ennaji, M. N. Benchekroun, M. Rakotomanga, J. F. Baroillier, and F. Galibert. Construction of radiation hybrid panels for two major aquaculture species: Sturgeon and oyster. Journal of Aquaculture Research & Development, 4(4):??, ???? 2013. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/construction-of-radiation-hybrid-panels-for-
two-major-aquaculture-species-sturgeon-and-oyster-26984.html.

Dhara:2013:CBA

[91] Kishore Dhara and Nimai Chandra Saha. Controlled breeding of Asian catfish Clarias batrachus using pituitary gland extracts and ovaprim at different temperatures, latency periods and their early development. Journal of Aquaculture Research & Development, 4(4): ??, ???? 2013. CODEN JARDCR. ISSN 2155-9546. URL https:

//www.walshmedicalmedia.com/open-access/controlled-breeding-
of-asian-catfish-emclarias-batrachusem-using-pituitary-gland-
extracts-and-ovaprim-at-different-tempe-27049.html.

Coelho:2013:EDS

[92] Ana Rita Coelho, Maria Teresa Dinis, and Joaquim Reis. Effect of diet and stocking densities on life history traits of Clea helena (Philippi 1847) reared in captivity. Journal of Aquaculture Research & Development, 4(5):??, ???? 2013. CODEN JARDCR. ISSN 2155-9546.

URL https://www.walshmedicalmedia.com/open-access/effect-of-
diet-and-stocking-densities-on-life-history-traits-of-emclea-
helenaem-philippi-1847-reared-in-captivity-27664.html.

Cuong:2013:PEC

[93] Do Bien Cuong, Vu Kim Dung, Nguyen Thi Thu Hien, and Dang Thi Thu. Prebiotic evaluation of copra-derived mannooligosaccharides in white-leg shrimps. Journal of Aquaculture Research & Development, 4 (5):??, ???? 2013. CODEN JARDCR. ISSN 2155-9546. URL https:/

/www.walshmedicalmedia.com/open-access/prebiotic-evaluation-
of-copraderived-mannooligosaccharides-in-whiteleg-shrimps-27665.
html.

Kpundeh:2013:SDC

[94] Mathew D. Kpundeh, Pao Xu, Hong Yang, Jun Qiang, and Jie He. Stocking densities and chronic zero culture water exchange stress effects on biological performances, hematological and serum biochemical indices of GIFT tilapia juveniles (Oreochromis niloticus). Journal of Aquaculture Research & Development, 4(5):??, ???? 2013. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/stocking-densities-and-chronic-zero-culture-
water-exchange-stress-effects-on-biological-performances-hematological-
and-s-27666.html.

Wang:2013:LSD

[95] Xiaodan Wang, Erchao Li, Zequan Xiong, Ke Chen, Na Yu, Zhenyu Du, and Liqiao Chen. Low salinity decreases the tolerance to two pesticides, beta-cypermethrin and acephate, of white-leg shrimp, Litopenaeus vannamei. Journal of Aquaculture Research & Development, 4(5):??, ???? 2013. CODEN JARDCR. ISSN 2155-

9546. URL https://www.walshmedicalmedia.com/open-access/low-
salinity-decreases-the-tolerance-to-two-pesticides-betacypermethrin-
and-acephate-of-whiteleg-shrimp-emlitopenaeus-va-27667.html.

Mancuso:2013:FES

[96] Monique Mancuso, Renata Zaccone, Francesca Carella, Paola Maiolino, and Gionata De Vico. First episode of shell disease syndrome in Carcinus aestuarii (Crustacea: Decapoda: Portunidae) in the Volturno River. Journal of Aquaculture Research & Development, 4 (5):??, ???? 2013. CODEN JARDCR. ISSN 2155-9546. URL

https://www.walshmedicalmedia.com/open-access/first-episode-
of-shell-disease-syndrome-in-emcarcinus-aestuariiem-crustacea-
decapoda-portunidae-in-the-volturno-river-27668.html.

Khalafalla:2013:NVD

[97] Malik M. Khalafalla. Nutritive value of diets containing Digeston-1 as a feed additive for Nile tilapia (Oreochromis niloticus) fingerlings. Journal of Aquaculture Research & Development, 4(5):??, ???? 2013. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/nutritive-value-of-diets-containing-digeston1-
as-a-feed-additive-for-nile-tilapia-emoreochromis-niloticusem-
fingerlings-27669.html.

Workagegn:2013:EDI

[98] Kassaye Balkew Workagegn, Elias Dadebo Ababbo, and Bishaw Tadele Tossa. The effect of dietary inclusion of Jatropha curcas kernel meal on growth performance, feed utilization efficiency and survival rate of juvenile Nile tilapia. Journal of Aquaculture Research & Development, 4(5): ??, ???? 2013. CODEN JARDCR. ISSN 2155-9546. URL https://

www.walshmedicalmedia.com/open-access/the-effect-of-dietary-
inclusion-of-emjatropha-curcasem-kernel-meal-on-growth-performance-
feed-utilization-efficiency-and-27670.html.

Suantika:2013:UIP

[99] Gede Suantika, Pingkan Aditiawati, Dea Indriani Astuti, and Zarah Fazri Khotimah. The use of indigenous probiotic Halomonas aquamarina and Shewanella algae for white shrimp (Litopenaeus vannamei Boone) hatchery productivity in zero water discharge system. Journal of Aquaculture Research & Development, 4(5):??, ???? 2013. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/the-use-of-indigenous-probiotic-emhalomonas-
aquamarinaem-and-emshewanella-algaeem-for-white-shrimp-emlitopenaeus-
vanname-27671.html.

deOliveira:2013:CEN

[100] Flávia Martins Franco de Oliveira and Maria Cristina Basílio Crispim. Compost extract as a nutrient source for algal cultures. Journal of Aquaculture Research & Development, 4(5):??, ???? 2013. CODEN JARDCR.

ISSN 2155-9546. URL https://www.walshmedicalmedia.com/open-
access/compost-extract-as-a-nutrient-source-for-algal-cultures-
27672.html.

Varadharajan:2013:SAA

[101] D. Varadharajan and P. Soundarapandian. Science of advance in aqua farming study of food and feeding habits of jumbo tiger shrimp, Penaeus monodon (Fabricius, 1798) from Parangipettai, South East Coast of India. Journal of Aquaculture Research & Development, 4(6):??, ???? 2013. CODEN JARDCR. ISSN 2155-9546.

URL https://www.walshmedicalmedia.com/open-access/science-
of-advance-in-aqua-farming-study-of-food-and-feeding-habits-
of-jumbo-tiger-shrimp-empenaeus-monodonem-fabricius--27776.html.

Kritsanapuntu:2013:EDP

[102] Sirusa Kritsanapuntu, Nilnaj Chaitanawisuti, and Wannanee Santaweesuk. Effects of dietary partial replacement of tuna oil by corn oil in formulated diets for growth performance and proximate composition of juvenile spotted babylon Babylonia areolata under hatchery conditions. Journal of Aquaculture Research & Development, 4(6):??, ???? 2013. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/effects-of-dietary-partial-replacement-of-tuna-
oil-by-corn-oil-in-formulated-diets-for-growth-performance-and-
proximate--27777.html.

Chamundeeswari:2013:ICN

[103] K. Chamundeeswari, S. Saranya, S. Shanker, D. Varadharajan, D. Varadharajan, and T. Balasubramanian. Incidental catch and new distributional records of asteroids by bottom trawling activities from Southeast Coast of India. Journal of Aquaculture Research & Development, 4(6): ??, ???? 2013. CODEN JARDCR. ISSN 2155-9546. URL https:/

/www.walshmedicalmedia.com/open-access/incidental-catch-and-
new-distributional-records-of-asteroids-by-bottom-trawling-activities-
from-southeast-coast-of-india-27778.html.

Thilagavathi:2013:DDM

[104] B. Thilagavathi, D. Varadharajan, A. Babu, J. Manoharan, S. Vijayalakshmi, and T. Balasubramanian. Distribution and diversity of macrobenthos in different mangrove ecosystems of Tamil Nadu Coast, India. Journal of Aquaculture Research & Development, 4(6):??, ???? 2013. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/distribution-and-diversity-of-macrobenthos-in-
different-mangrove-ecosystems-of-tamil-nadu-coast-india-27779.
html.

R:2013:EAH

[105] Mary Lini R, Pani Prasad Kurcheti, Gireesh Babu, and C. S. Purushothaman. Effect of Aeromonas hydrophila infection on caspase-3 expression and activity in rohu, Labeo rohita. Journal of Aquaculture Research & Development, 4(6):??, ???? 2013. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
effect-of-emaeromonas-hydrophilaem-infection-on-caspase3-expression-
and-activity-in-rohu-labeo-rohita-27780.html.

Priyadharsini:2013:RBH

[106] S. Priyadharsini, J. Manoharan, D. Varadharajan, and A. Subramaniyan. Reproductive biology and histological study of red lionfish Pterois volitans from Cuddalore, South East Coast of India. Journal of Aquaculture Research & Development, 4(6):??, ???? 2013. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
reproductive-biology-and-histological-study-of-red-lionfish-
pterois-volitans-from-cuddalore-south-east-coast-of-india-27781.
html.

Mancuso:2013:FWA

[107] Monique Mancuso. Fish welfare in aquaculture. Journal of Aquaculture Research & Development, 4(6):??, ???? 2013. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
fish-welfare-in-aquaculture-27782.html.

Lu:2014:DOM

[108] X. Lu, M. J. Talbott, J. P. V. Eenennaam, M. A. H. Webb, S. I. Doroshov, M. Ovissipour, and B. Rasco. Determining ovarian maturity in farmed sturgeon 1 (Acipenser transmontanus) for caviar production using Fourier Transform Infrared Spectroscopy (FT-IR). Journal of Aquaculture Research & Development, 5(1):1–6, ???? 2014. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
determining-ovarian-maturity-in-farmed-sturgeon-1-acipenser-
transmontanus-for-caviar-production-using-fourier-transform-
-27856.html.

Bolarinwa:2014:LWR

[109] Josef Bamidele Bolarinwa and Q. Popoola. Length–weight relationships of some economic fishes of Ibeshe Waterside, Lagos Lagoon, Nigeria. Journal of Aquaculture Research & Development, 5(1):??, ???? 2014. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/lengthweight-relationships-of-some-economic-
fishes-of-ibeshe-waterside-lagos-lagoon-nigeria-27857.html.

Soundarapandian:2014:AMS

[110] P. Soundarapandian, G. K. Dinakaran, and D. Varadharajan. Alternative mating strategies in male morphotypes of the prawn Macrobrachium idellaidella (Hilgendorf, 1898). Journal of Aquaculture Research & Development, 5(1):??, ???? 2014. CODEN JARDCR.

ISSN 2155-9546. URL https://www.walshmedicalmedia.com/open-
access/alternative-mating-strategies-in-male-morphotypes-of-
the-prawn-macrobrachiumidellaidella-hilgendorf-1898-27858.html.

Paritova:2014:ICZ

[111] Assel Paritova, Nurzhan Sarsembayeva, Bożena Łozowicka, Amangeldy Maulanov, Gulnur Kuzembekova, Aida Abzhalieva, and Piotr Kaczyński.

The influence of Chankanay zeolites as feed additives on the chemical, biochemical and histological profile of the rainbow trout (Oncorhynchus mykiss). Journal of Aquaculture Research & Development, 5(1):1–8, ???? 2014. CODEN JARDCR. ISSN 2155-9546. URL

https://www.walshmedicalmedia.com/open-access/the-influence-
of-chankanay-zeolites-as-feed-additives-on-the-chemical-biochemical-
and-histological-profile-of-the-rainbo-27859.html.

ElBaky:2014:ISP

[112] H. Abd El Baky, K. F. Hanaa El Baz, and S. A. El-Latife. Induction of sulfated polysaccharides in Spirulina platensis as response to nitrogen concentration and its biological evaluation. Journal of Aquaculture Research & Development, 5(1):??, ???? 2014. CODEN JARDCR. ISSN 2155-9546.

URL https://www.walshmedicalmedia.com/open-access/induction-
of-sulfated-polysaccharides-in-spirulina-platensis-as-response-
to-nitrogen-concentration-and-its-biological-ev-27860.html.

AMohamed:2014:ROB

[113] S. M. AMohamed, I. H. Magdy, A. M. Olfat, A. T. Ebtsam, and S. Y. Nesreen. Role of oligosaccharides as biological additives in cultured Oreochromis niloticus. Journal of Aquaculture Research & Development, 5(1):??, ???? 2014. CODEN JARDCR. ISSN 2155-9546.

URL https://www.walshmedicalmedia.com/open-access/role-of-
oligosaccharides-as-biological-additives-in-cultured-oreochromis-
niloticus-27861.html.

Langeland:2014:DEA

[114] Markus Langeland, Jan Erik Lindberg, and Torbjörn Lundh. Digestive enzyme activity in Eurasian perch (Perca fluviatilis) and Arctic charr (Salvelinus alpinus). Journal of Aquaculture Research & Development, 5(1):??, ???? 2014. CODEN JARDCR. ISSN 2155-9546.

URL https://www.walshmedicalmedia.com/open-access/digestive-
enzyme-activity-in-eurasian-perch-perca-fluviatilis-and-arctic-
charr-salvelinus-alpinus-27862.html.

Abro:2014:EDS

[115] Rani Abro, Torbjörn Lundh, and Jan Erik Lindberg. Effect of dietary starch inclusion rate on digestibility and amylase activity in Arctic charr (Salvelinus alpinus) and Eurasian perch (Perca fluviatilis). Journal of Aquaculture Research & Development, 5(1):??, ???? 2014. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/effect-of-dietary-starch-inclusion-rate-on-digestibility-
and-amylase-activity-in-arctic-charr-salvelinus-alpinus-and-
eur-27863.html.

Sudhakar:2014:MSP

[116] S. Sudhakar, P. Soundarapandian, and D. Varadharajan. Mass seed production of Macrobrachium idae (Heller, 1862). Journal of Aquaculture Research & Development, 5(1):??, ???? 2014. CODEN JARDCR. ISSN 2155-

9546. URL https://www.walshmedicalmedia.com/open-access/mass-
seed-production-of-macrobrachium-idae-heller-1862-27864.html.

Mancuso:2014:AA

[117] Monique Mancuso. Aquaculture advancement. Journal of Aquaculture Research & Development, 5(1):??, ???? 2014. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
aquaculture-advancement-27865.html.

Thirunavukkarasu:2014:PCC

[118] K. Thirunavukkarasu, P. Soundarapandian, D. Varadharajan, and B. Gunalan. Phytoplankton composition and community structure of Kottakudi and Nari Backwaters, South East of Tamil Nadu. Journal of Aquaculture Research & Development, 5(2):??, ???? 2014. CODEN JARDCR.

ISSN 2155-9546. URL https://www.walshmedicalmedia.com/open-
access/phytoplankton-composition-and-community-structure-of-
kottakudi-and-nari-backwaters-south-east-of-tamil-nadu-28144.
html.

Luin:2014:SMG

[119] Marianne Luin, Ching Fui Fui, and Shigeharu Senoo. Sexual maturation and gonad development in tiger grouper (Epinephelus fuscoguttatus) × giant grouper (E. Lanceolatus) hybrid. Journal of Aquaculture Research & Development, 5(2):??, ???? 2014. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
sexual-maturation-and-gonad-development-in-tiger-grouperemepinephelus-
fuscoguttatusem-x-giant-grouper-eme-lanceolatusem--27959.html.

Andi:2014:LSA

[120] G. T. Andi, Dahlifa, Ratnawati, Mardiana, and P. A. AndiRezki. Land suitability analysis of tiger shrimp aquaculture (Penaeusmonodon. Fab.) in the coastal area of Labakkang District South Sulawesi — Indonesia. Journal of Aquaculture Research & Development, 5(2):1–7, ???? 2014. CODEN JARDCR. ISSN 2155-9546. URL https://www.

walshmedicalmedia.com/open-access/land-suitability-analysis-
of-tiger-shrimp-aquaculture-penaeusmonodon-fab-in-the-coastal-
area-of-labakkang-district-south-28145.html.

Laith:2014:AHA

[121] A. R. Laith and M. Najiah. Aeromonas hydrophila: Antimicrobial susceptibility and histopathology of isolates from diseased catfish, Clarias gariepinus (Burchell). Journal of Aquaculture Research & Development, 5(2):??, ???? 2014. CODEN JARDCR. ISSN 2155-9546. URL https:/

/www.walshmedicalmedia.com/open-access/aeromonas-hydrophila-
antimicrobial-susceptibility-and-histopathology-of-isolates-
from-diseased-catfish-clarias-gariepinu-28146.html.

Bottari:2014:EDA

[122] T. Bottari, B. Busalacchi, A. Profeta, M. Mancuso, D. Giordano, and P. Rinelli. Elasmobranch distribution and assemblages in the Southern Tyrrhenian Sea (Central Mediterranean). Journal of Aquaculture Research & Development, 5(2):??, ???? 2014. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
elasmobranch-distribution-and-assemblages-in-the-southern-tyrrhenian-
sea-central-mediterranean-28147.html.

Varadharajan:2014:PCM

[123] D. Varadharajan and P. Soundarapandian. Proximate composition and mineral contents of freshwater crab Spiralothelphusa hydrodroma (Herbst, 1794) from Parangipettai, South East Coast of India. Journal of Aquaculture Research & Development, 5(2):??, ???? 2014. CODEN JARDCR. ISSN 2155-9546. URL https://www.

walshmedicalmedia.com/open-access/proximate-composition-and-
mineral-contents-of-freshwater-crab-spiralothelphusa-hydrodroma-
herbst-1794-from-parangipettai-28148.html.

Roncarati:2014:CGP

[124] Alessandra Roncarati, Oliviero Mordenti, Luca Stocchi, and Paolo Melotti. Comparison of growth performance of ‘common catfish Ameiurus melas, Rafinesque 1820’ reared in pond and in recirculating aquaculture system. Journal of Aquaculture Research & Development, 5(2): 1–6, ???? 2014. CODEN JARDCR. ISSN 2155-9546. URL https:/

/www.walshmedicalmedia.com/open-access/comparison-of-growth-
performance-of-1951621948364194732common-catfish-ameiurus-melas-
rafinesque182019516219483641948482--28143.html.

Mahmoud:2014:IMM

[125] Nisreen E. Mahmoud, Mohga F. M. Badawy, and M. M. Fahmy. Investigations on mass mortalities among Oreochromis niloticus at Mariotteya Stream, Egypt: Parasitic infestation and environmental pollution impacts. Journal of Aquaculture Research & Development, 5(2):

??, ???? 2014. CODEN JARDCR. ISSN 2155-9546. URL https://www.

walshmedicalmedia.com/open-access/investigations-on-mass-mortalities-
among-oreochromis-niloticus-at-mariotteya-stream-egypt-parasitic-
infestation-and-envi-27966.html.

Carreno-Leon:2014:GMP

[126] Diana Carreño-León, Ilie Racotta-Dimitrov, Ramón Casillas-Hernández, Armando Monge-Quevedo, Lucía Ocampo-Victoria, José Naranjo-Páramo, and Humberto Villarreal. Growth, metabolic and physiological response of juvenile Cherax quadricarinatus fed different available nutritional substrates. Journal of Aquaculture Research & Development, 5(2): 1–7, ???? 2014. CODEN JARDCR. ISSN 2155-9546. URL https:/

/www.walshmedicalmedia.com/open-access/growth-metabolic-and-
physiological-response-of-juvenile-cheraxquadricarinatus-fed-
different-available-nutritional-substr-28149.html.

Roohi:2014:ESC

[127] Zahra Roohi and Mohammad Reza Imanpoor. Effects of spearmint (/- carvon) oil and methyl salicylate oil emulsion on anesthesia of common carp (Cyprinus carpio L., 1758). Journal of Aquaculture Research & Development, 5(2):??, ???? 2014. CODEN JARDCR. ISSN 2155-9546.

URL https://www.walshmedicalmedia.com/open-access/effects-
of-spearmint-carvon-oil-and-methyl-salicylate-oil-emulsion-on-
anesthesia-of-common-carp-cyprinus-carpio-l-1758-28150.html.

Ronald:2014:ESD

[128] Ntanzi Ronald, Bwanika Gladys, and Eriku Gasper. The effects of stocking density on the growth and survival of Nile tilapia (Oreochromis niloticus) fry at Son Fish Farm, Uganda. Journal of Aquaculture Research & Development, 5(2):??, ???? 2014. CODEN JARDCR. ISSN 2155-

9546. URL https://www.walshmedicalmedia.com/open-access/the-
effects-of-stocking-density-on-the-growth-and-survival-of-nile-
tilapia-oreochromis-niloticus-fry-at-son-fish-farm-ug-28261.
html.

Ovissipour:2014:FOP

[129] Mahmoudreza Ovissipour and Barbara Rasco. Fish oil: Perspectives on a sustainable multi-use resource. Journal of Aquaculture Research & Development, 5(3):??, ???? 2014. CODEN JARDCR. ISSN 2155-9546.

URL https://www.walshmedicalmedia.com/open-access/fish-oil-
perspectives-on-a-sustainable-multiuse-resource-2155-9546.1000e109.
pdf.

Baky:2014:LID

[130] Hanaa H. Abd El Baky, Gamal S. El-Baroty, and Abderrahim Bouaid. Lipid induction in Dunaliella salina culture aerated with various levels CO_{2} and its biodiesel production. Journal of Aquaculture Research & Development, 5(3):??, ???? 2014. CODEN JARDCR.

ISSN 2155-9546. URL https://www.walshmedicalmedia.com/open-
access/lipid-induction-in-dunaliella-salina-culture-aerated-
with-various-levels-co2-and-its-biodiesel-production-28262.html.

Sanmukh:2014:BCD

[131] Swapnil Sanmukh, Benedict Bruno, Udhaya Ramakrishnan, Krishna Khairnar, Sandhya Swaminathan, and Waman Paunikar. Bioactive compounds derived from microalgae showing antimicrobial activities. Journal of Aquaculture Research & Development, 5(3):??, ???? 2014. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/bioactive-compounds-derived-from-microalgae-
showing-antimicrobial-activities-28264.html.

O:2014:DPE

[132] Ashaolu Michael O. Development and performance evaluation of a motorized fish smoking kiln. Journal of Aquaculture Research & Development, 5(3):??, ???? 2014. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
development-and-performance-evaluation-of-a-motorized-fish-smoking-
kiln-28225.html.

Wotton:2014:CHS

[133] Heather J. Wotton, Sara L. Purcell, Jennifer M. Covello, Ben F. Koop, and Mark D. Fast. Comparison of host selection and gene expression of adult Lepeophtheirus salmonis and Salmo salar during a cohabitation of initially infected and uninfected fish. Journal of Aquaculture Research & Development, 5(3):??, ???? 2014. CODEN JARDCR. ISSN 2155-9546. URL

https://www.walshmedicalmedia.com/open-access/comparison-of-
host-selection-and-gene-expression-of-adult-lepeophtheirus-salmonis-
and-salmo-salar-during-a-cohabitation--28226.html.

Jufaili:2014:BEF

[134] Sarah Al Jufaili, Vladimir Machkevskyi, Sami Al Sulaimi, Redha Bait Faraj, and Nashwa Al Mazrooei. Biological and ecological features of Poecilancistrum sp. plerocercoid (Cestoda: Trypanorhyncha) infection of Arabian Sea meager Argyrosomus heinii (Steindachner, 1907). Journal of Aquaculture Research & Development, 5(3):1–7, ????

2014. CODEN JARDCR. ISSN 2155-9546. URL https://www.

walshmedicalmedia.com/open-access/biological-and-ecological-
features-of-poecilancistrum-sp-plerocercoid-cestoda-trypanorhyncha-
infection-of-arabian-sea-me-28227.html.

Chen:2014:RAR

[135] Yihong Chen, Xiaoyun Li, and Jianguo He. Recent advances in researches on shrimp immune pathway involved in white spot syndrome virus genes regulation. Journal of Aquaculture Research & Development, 5(3):??, ???? 2014. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
recent-advances-in-researches-on-shrimp-immune-pathway-involved-
in-white-spot-syndrome-virus-genes-regulation-28228.html.

Ren:2014:MCE

[136] Gang Ren, Wen-Ying Shen, Wei-Fen Li, and Yao rong Zhu. Molecular characterization and expression pattern of a liver-expressed antimicrobial peptide 2 (LEAP-2) gene in yellow catfish (Pelteobagrus fulvidraco). Journal of Aquaculture Research & Development, 5(3):??, ???? 2014. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/molecular-characterization-and-expression-pattern-
of-a-liverexpressed-antimicrobial-peptide-2-leap2-gene-in-yellow-
catfi-28229.html.

Geraylou:2014:FAO

[137] Zahra Geraylou, Eugene Rurangwa, Tom Van De Wiele, Christophe M. Courtin, Jan A. Delcour, Johan Buyse, and Frans Ollevier. Fermentation of arabinoxylan-oligosaccharides, oligofructose and their monomeric sugars by hindgut bacteria from Siberian sturgeon and African catfish in batch culture in vitro. Journal of Aquaculture Research & Development, 5(3):??, ???? 2014. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
fermentation-of-arabinoxylanoligosaccharides-oligofructose-and-
their-monomeric-sugars-by-hindgut-bacteria-from-siberian--28249.
html.

Kang:2014:GDC

[138] Jung-Ha Kang, Eun-Soo Noh, Jae-Hyun Lim, Hyeung-Kyun Han, Bong-Seok Kim, and Sang-Ku Lim. Genetic differentiation of Carassius auratus and C. cuvieri by the cytochrome C oxidase I gene analysis. Journal of Aquaculture Research & Development, 5(3):??, ???? 2014. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/genetic-differentiation-of-emcarassius-auratusem-

andem-c-cuvieriem-by-the-cytochrome-c-oxidase-i-gene-analysis-
28250.html.

Yadollahvand:2014:HCI

[139] Reza Yadollahvand and Haji Gholi Kami. Habitat changes and its impacts on the Caspian pond turtle (Mauremys caspica) population in the Golestan and Mazandaran Provinces of Iran. Journal of Aquaculture Research & Development, 5(3):??, ???? 2014. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
habitat-changes-and-its-impacts-on-the-caspian-pond-turtle-mauremys-
caspica-population-in-the-golestan-and-mazandaran-pr-28251.html.

Amare:2014:PIP

[140] A. Amare, A. Alemayehu, and A. Aylate. Prevalence of internal parasitic helminthes infected Oreochromis niloticus (Nile tilapia), Clarias gariepinus (African catfish) and Cyprinus carpio (common carp) in Lake Lugo (Hayke), Northeast Ethiopia. Journal of Aquaculture Research & Development, 5(3):??, ???? 2014. CODEN JARDCR.

ISSN 2155-9546. URL https://www.walshmedicalmedia.com/open-
access/prevalence-of-internal-parasitic-helminthes-infected-
oreochromis-niloticus-nile-tilapia-clarias-gariepinus-african-
catfi-28252.html.

Hixson:2014:FNC

[141] Stefanie M. Hixson. Fish nutrition and current issues in aquaculture: The balance in providing safe and nutritious seafood, in an environmentally sustainable manner. Journal of Aquaculture Research & Development, 5(3):??, ???? 2014. CODEN JARDCR.

ISSN 2155-9546. URL https://www.walshmedicalmedia.com/open-
access/fish-nutrition-and-current-issues-in-aquaculture-the-
balance-in-providing-safe-and-nutritious-seafood-in-an-environmenta-
28253.html.

Workagegn:2014:GPN

[142] Kassaye Balkew Workagegn, Elias Dadebo Ababboa, Girma Tilahun Yimer, and Tigist Ashagre Amare. Growth performance of the Nile tilapia (Oreochromis niloticus L.) fed different types of diets formulated from varieties of feed ingredients. Journal of Aquaculture Research & Development, 5(3):??, ???? 2014. CODEN JARDCR. ISSN 2155-9546. URL https://

www.walshmedicalmedia.com/open-access/growth-performance-of-
the-nile-tilapia-oreochromis-niloticus-l-fed-different-types-
of-diets-formulated-from-varieties-of-28254.html.

Qamer:2014:CBG

[143] Mahwish Qamer, Farkhanda Asad, and Nimra Tahir. Changes in body growth of Labeo rohita in relation to dietary carbohydrate content and protein levels using gelatinized and non-gelatinized corn. Journal of Aquaculture Research & Development, 5(4):??, ???? 2014. CODEN JARDCR. ISSN 2155-9546. URL https://www.

walshmedicalmedia.com/open-access/changes-in-body-growth-of-
labeo-rohita-in-relation-to-dietary-carbohydrate-content-and-
protein-levels-using-gelatinized--28392.html.

Valenzuela-Munoz:2014:MCT

[144] Valentina Valenzuela-Muñoz, Gustavo Nuñez-Acuña, and Cristian Gallardo-Escárate. Molecular characterization and transcription analysis of p-glycoprotein gene from the salmon louse Caligus rogercresseyi. Journal of Aquaculture Research & Development, 5(4):??, ???? 2014. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/molecular-characterization-and-transcription-
analysis-of-pglycoprotein-gene-from-the-salmon-louse-caligus-
rogercresseyi-28255.html.

Lin:2014:LET

[145] Yi-Pei Lin, Wen-Jen Chung, Jen-Leih Wu, and Shao-Yang Hu. The liver-enriched transcription factors HNF-1α, HNF-3β, and C/EBPβ contribute to the growth hormone-induced transcription of the progranulin a gene in zebrafish (Danio rerio). Journal of Aquaculture Research & Development, 5(4):1–7, ???? 2014. CODEN JARDCR. ISSN 2155-

9546. URL https://www.walshmedicalmedia.com/open-access/the-
liverenriched-transcription-factors-hnf1195381194177-hnf3195381194178-
and-c-ebp195381194178-contribute-to-the-growth-28256.html.

Davidson:2014:GPF

[146] J. W. Davidson, P. B. Kenney, M. Manor, C. M. Good, G. M. Weber, A. Aussanasuwannakul, P. J. Turk, C. Welsh, and S. T. Summerfelt. Growth performance, fillet quality, and reproductive maturity of rainbow trout (Oncorhynchus mykiss) cultured to 5 kilograms within freshwater recirculating systems. Journal of Aquaculture Research & Development, 5(4):??, ???? 2014. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
growth-performance-fillet-quality-and-reproductive-maturity-
of-rainbow-trout-oncorhynchus-mykiss-cultured-to-5-kilograms-
28292.html.

Rha:2014:NCB

[147] Sung-Ju Rha, Jae-Kwon Cho, Seon-Jae Kim, Wook-Min Park, Tai-Sun Shin, and Jae-Ho Hwang. Nutritional characteristics of black rockfish (Sebastes schlegeli) fed a diet of fish skin. Journal of Aquaculture Research & Development, 5(4):??, ???? 2014. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
nutritional-characteristics-of-black-rockfish-sebastes-schlegeli-
fed-a-diet-of-fish-skin-28291.html.

Al-Faragi:2014:EPG

[148] Jamal K. Al-Faragi. The efficacy of prebiotic (β-glucan) as a feed additive against toxicity of aflatoxin B1 in common carp, Cyprinus carpio L. Journal of Aquaculture Research & Development, 5(4):??, ???? 2014. CODEN JARDCR. ISSN 2155-9546. URL https://www.

walshmedicalmedia.com/open-access/the-efficacy-of-prebiotic-
223glucan-as-a-feed-additive-against-toxicity-of-aflatoxin-b1-
in-common-carp-cyprinus-carpio-l-28293.html.

Sakthivel:2014:EMD

[149] Alagarsamy Sakthivel, Periyasamy Selvakumar, and Ayyaru Gopalakrishnan. Effect of mineral deposition on shrimp Litopenaeus vannamei in high alkaline water of Pennar River, Andhra Pradesh of Southeast Coast of India. Journal of Aquaculture Research & Development, 5(4):??, ???? 2014. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
effect-of-mineral-deposition-on-shrimp-litopenaeus-vannamei-
in-high-alkaline-water-of-pennar-river-andhra-pradesh-of-sou-
28340.html.

Eissa:2014:COB

[150] Alaa Eldin Eissa and Mohamed Faisal. Clinical outbreaks of bacterial kidney disease (BKD) in hatchery-raised brook trout (Salvelinus fontinalis) (Mitchill, 1814): Lessons learned. Journal of Aquaculture Research & Development, 5(4):??, ???? 2014. CODEN JARDCR. ISSN 2155-9546.

URL https://www.walshmedicalmedia.com/open-access/clinical-
outbreaks-of-bacterial-kidney-disease-bkd-in-hatcheryraised-
brook-trout-salvelinus-fontinalis-mitchill-1814-les-28341.html.

Meetei:2014:IFR

[151] L. I. Meetei, A. S. Ninawe, and R. Chakrabarti. Influence of feeding regimes on the digestive enzyme profile and ultrastructure of digestive tract of first feeding Catla catla (Hamilton) larvae. Journal of Aquaculture Research & Development, 5(4):??, ???? 2014. CODEN JARDCR.

ISSN 2155-9546. URL https://www.walshmedicalmedia.com/open-
access/influence-of-feeding-regimes-on-the-digestive-enzyme-
profile-and-ultrastructure-of-digestive-tract-of-first-feeding-
catla-catla-hamilton-larvae-2155-9546.1000243.pdf.

Estim:2014:WQR

[152] Abentin Estim and Saleem Mustafa. Water quality remediation using geotextile in fish hatchery systems. Journal of Aquaculture Research & Development, 5(4):??, ???? 2014. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
water-quality-remediation-using-geotextile-in-fish-hatchery-
systems-28342.html.

Haghshenas:2014:PPF

[153] Mehrdad Haghshenas, Hedayat Hosseini, Kooshan Nayebzadeh, Amin Mosavi Khanghah, Behnoosh Shabkoohi Kakesh, and Rozita Komeily Fonood. Production of prebiotic functional shrimp nuggets using β-glucan and reduction of oil absorption by carboxymethyl cellulose: Impacts on sensory and physical properties. Journal of Aquaculture Research & Development, 5(4):??, ???? 2014. CODEN JARDCR. ISSN 2155-9546. URL

https://www.walshmedicalmedia.com/open-access/production-of-
prebiotic-functional-shrimp-nuggets-using-223glucan-and-reduction-
of-oil-absorption-by-carboxymethyl-cellu-28343.html.

Wani:2014:FRA

[154] M. A. Wani and S. P. S. Dutta. First record of an anomalous Catla catla (Ham. Buch) in fresh water fish ponds of Gurdaspur District, Punjab (India). Journal of Aquaculture Research & Development, 5(4):??, ???? 2014. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
first-record-of-an-anomalous-catla-catla-ham-buch-in-fresh-water-
fish-ponds-of-gurdaspur-district-punjab-india-28344.html.

Aly:2014:AUA

[155] Salah Mesalhy Aly and Aqel Albutti. Antimicrobials use in aquaculture and their public health impact. Journal of Aquaculture Research & Development, 5(4):??, ???? 2014. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
antimicrobials-use-in-aquaculture-and-their-public-health-impact-
28346.html.

Dubey:2014:MLS

[156] S. K. Dubey, R. K. Trivedi, S. K. Rout, B. K. Chand, and A. Choudhury. Median lethal salinity (MLS_{96h}) of two small indigenous fish species Amblypharyngodon mola and Pethia ticto from Indian Sundarban. Journal of Aquaculture Research & Development, 5(5):??, ???? 2014. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/median-lethal-salinity-mls96-h-of-two-small-
indigenous-fish-species-amblypharyngodon-mola-and-pethia-ticto-
from-indian-s-28393.html.

Asad:2014:AND

[157] Farkhanda Asad, Mahwish Qamer, and Nimra Tahir. Apparent nutrient digestibility assessment and influence of gelatinized and non-gelatinized corn starch based diet in Labeo rohita. Journal of Aquaculture Research & Development, 5(5):??, ???? 2014. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
apparent-nutrient-digestibility-assessment-and-influence-of-
gelatinized-and-nongelatinized-corn-starch-based-diet-in-lab-
28394.html.

Chi:2014:PMG

[158] Jing Ruei Chi, Chang-Wen Huang, Jen Leih Wu, and Shao Yang Hu. Prolactin I microsatellite as genetic markers for characterization of five Oreochromis tilapia species and two Oreochromis niloticus strains. Journal of Aquaculture Research & Development, 5(5):1–7, ???? 2014. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/prolactin-i-microsatellite-as-genetic-markers-
for-characterization-of-five-oreochromis-tilapia-species-and-
two-oreochrom-28395.html.

Omobepade:2014:TEA

[159] B. P. Omobepade, O. T. Adebayo, and T. T. Amos. Technical efficiency of aquaculturists in Ekiti State, Nigeria. Journal of Aquaculture Research & Development, 5(5):??, ???? 2014. CODEN JARDCR. ISSN 2155-9546.

URL https://www.walshmedicalmedia.com/open-access/technical-
efficiency-of-aquaculturists-in-ekiti-state-nigeria-28396.html.

Soliman:2014:NAC

[160] Soha A. Soliman. New aspect in cartilage growth “The Invasive Interstitial Type”. Journal of Aquaculture Research & Development, 5 (5):1–4, ???? 2014. CODEN JARDCR. ISSN 2155-9546. URL

https://www.walshmedicalmedia.com/open-access/new-aspect-in-
cartilage-growth-1951621948364194339the-invasive-interstitial-
type1951621948364194157-28397.html.

P:2014:SUR

[161] Rama Nisha P, Elezabeth Mary A, M. Uthayasiva, and S. Arularasan. Seaweed Ulva reticulata a potential feed supplement for growth, colouration and disease resistance in fresh water ornamental gold fish, Carassius auratus. Journal of Aquaculture Research & Development, 5(5):??, ???? 2014. CODEN JARDCR. ISSN 2155-9546. URL https://www.

walshmedicalmedia.com/open-access/seaweed-ulva-reticulata-a-
potential-feed-supplement-for-growth-colouration-and-disease-
resistance-in-fresh-water-ornamen-28398.html.

Gunarto:2014:CMC

[162] Gunarto and Andi Parenrengi. Crablet of mangrove crab, Scylla olivacea rearing at the different salinity regimes. Journal of Aquaculture Research & Development, 5(5):??, ???? 2014. CODEN JARDCR. ISSN 2155-9546.

URL https://www.walshmedicalmedia.com/open-access/crablet-
of-mangrove-crab-scylla-olivacea-rearing-at-the-different-salinity-
regimes-28399.html.

Hua:2014:RBE

[163] Nhan Thai Hua and Harry Ako. Reproductive biology and effect of arachidonic acid level in broodstock diet on final maturation of the Hawaiian limpet Cellana sandwicensis. Journal of Aquaculture Research & Development, 5(5):??, ???? 2014. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
reproductive-biology-and-effect-of-arachidonic-acid-level-in-
broodstock-diet-on-final-maturation-of-the-hawaiian-limpet--
28400.html.

Fattah:2014:PEM

[164] Muhammad Hattah Fattah, Muh. Saenong, Asbar, and St Rahbiah Busaeri. Production of endemic microcrustacean Phronima suppa (Phronima sp) to subtitute Artemia salina in tiger prawn cultivation. Journal of Aquaculture Research & Development, 5(5):??, ???? 2014. CODEN JARDCR. ISSN 2155-9546. URL https://

www.walshmedicalmedia.com/open-access/production-of-endemic-
microcrustacean-phronima-suppa-phronima-sp-to-subtitute-artemia-
salina-in-tiger-prawn-cultivation-28401.html.

Ponnusamy:2014:HMC

[165] K. Ponnusamy, P. Sivaperumal, M. Suresh, S. Arularasan, S. Munilkumar, and A. K. Pal. Heavy metal concentration from biologically important edible species of bivalves (Perna viridis and Modiolus metcalfei) from Vellar Estuary, South East Coast of India. Journal of Aquaculture Research & Development, 5(5):??, ???? 2014. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/heavy-metal-concentration-from-biologically-
important-edible-species-of-bivalves-perna-viridis-and-modiolus-
metcalfei-fr-28402.html.

Kayhan:2014:AOA

[166] Muhammet Hayati Kayhan and Murtaza Olmez. Aquaculture and organic aquaculture in Turkey. Journal of Aquaculture Research & Development, 5(5):??, ???? 2014. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
aquaculture-and-organic-aquaculture-in-turkey-28403.html.

Good:2014:IWE

[167] Christopher Good, John Davidson, Ryan L. Earley, Elizabeth Lee, and Steven Summerfelt. The impact of water exchange rate and treatment processes on water-borne hormones in recirculation aquaculture systems containing sexually maturing Atlantic salmon Salmo salar. Journal of Aquaculture Research & Development, 5(5):??, ???? 2014. CODEN JARDCR. ISSN 2155-9546. URL https:

//www.walshmedicalmedia.com/open-access/the-impact-of-water-
exchange-rate-and-treatment-processes-on-waterborne-hormones-
in-recirculation-aquaculture-systems-co-28404.html.

Yadollahvand:2014:ADB

[168] Reza Yadollahvand and Behzad Rahnama. The age determination of black pomfret (Parastromateus niger), based on otolith cross sections in Iranian Coast of Oman Sea. Journal of Aquaculture Research & Development, 5(5):??, ???? 2014. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
the-age-determination-of-black-pomfret-parastromateus-niger-
based-on-otolith-cross-sections-in-iranian-coast-of-oman-sea-
28405.html.

Marchand:2014:SPB

[169] C. Marchand, N. Molnar, J. Deborde, L. D. Patrona, and T. Meziane. Seasonal pattern of the biogeochemical properties of mangrove sediments receiving shrimp farm effluents (New Caledonia). Journal of Aquaculture Research & Development, 5(5):??, ???? 2014. CODEN JARDCR.

ISSN 2155-9546. URL https://www.walshmedicalmedia.com/open-
access/seasonal-pattern-of-the-biogeochemical-properties-of-
mangrove-sediments-receiving-shrimp-farm-effluents-new-caledonia-
28406.html.

Saoud:2014:EAC

[170] Patrick Saoud I, S. Naamani, J. Ghanawi, and N. Nasser. Effects of acute and chronic nitrite exposure on rabbitfish Siganus rivulatus growth, hematological parameters, and gill histology. Journal of Aquaculture Research & Development, 5(6):??, ???? 2014. CODEN JARDCR. ISSN 2155-9546.

URL https://www.walshmedicalmedia.com/open-access/effects-
of-acute-and-chronic-nitrite-exposure-on-rabbitfish-siganus-
rivulatus-growth-hematological-parameters-and-gill-h-28521.html.

Cheng:2014:EPW

[171] A. C. Cheng, C. F. Lee, Y. Y. Chen, and J. C. Chen. The effect of packing on water quality parameters, survival and NNV load of Epinephelus coioides fry after simulated transport. Journal of Aquaculture Research & Development, 5(6):??, ???? 2014. CODEN JARDCR. ISSN 2155-

9546. URL https://www.walshmedicalmedia.com/open-access/the-
effect-of-packing-on-water-quality-parameters-survival-and-nnv-
load-of-epinephelus-coioides-fry-after-simulated-tran-28522.
html.

Mzengereza:2014:NVL

[172] K. Mzengereza, O. V. Msiska, F. Kapute, J. Kang’ombe, W. Singini, and A. Kamangira. Nutritional value of locally available plants with potential for diets of Tilapia rendalli in pond aquaculture in NkhataBay, Malawi. Journal of Aquaculture Research & Development, 5(6):??, ???? 2014. CODEN JARDCR. ISSN 2155-9546. URL https:/

/www.walshmedicalmedia.com/open-access/nutritional-value-of-
locally-available-plants-with-potential-for-diets-of-tilapia-
rendalli-in-pond-aquaculture-in-nkhata-28529.html.

Mesa-Rodriguez:2014:SDM

[173] A. Mesa-Rodríguez, C. M. Hernández-Cruz, J. A. Socorro, H. Fernández-Palacios, M. S. Izquierdo, and J. Roo. Skeletal development and mineralization pattern of the vertebral column, dorsal, anal and caudal fin complex in Seriola rivoliana (Valenciennes, 1833) larvae. Journal of Aquaculture Research & Development, 5(6):??, ???? 2014. CODEN JARDCR. ISSN 2155-9546. URL https:/

/www.walshmedicalmedia.com/open-access/skeletal-development-
and-mineralization-pattern-of-the-vertebral-column-dorsal-anal-
and-caudal-fin-complex-in-seriola-ri-28528.html.

Tessema:2014:APC

[174] A. Tessema, A. Mohammed, T. Birhanu, and T. Negu. Assessment of physico-chemical water quality of Bira Dam, Bati Wereda, Amhara Region, Ethiopia. Journal of Aquaculture Research & Development, 5(6):??, ???? 2014. CODEN JARDCR. ISSN 2155-9546. URL

https://www.walshmedicalmedia.com/open-access/assessment-of-
physicochemical-water-quality-of-bira-dam-bati-wereda-amhara-
region-ethiopia-28523.html.

Chen:2014:WSL

[175] Y. Y. Chen, J. C. Chen, Y. C. Lin, Yeh, K. P. Chao, and C. S. Lee. White shrimp Litopenaeus vannamei that have received petalonia binghamiae extract activate immunity, increase immune response and resistance against Vibrio alginolyticus. Journal of Aquaculture Research & Development, 5(6):??, ???? 2014. CODEN JARDCR.

ISSN 2155-9546. URL https://www.walshmedicalmedia.com/open-
access/white-shrimp-litopenaeus-vannamei-that-have-received-
petalonia-binghamiae-extract-activate-immunity-increase-immune-
resp-28524.html.

Quaiyum:2014:EBC

[176] M. A. Quaiyum, R. Jahan, N. Jahan, T. Akhter, and Islam M. Sadiqul. Effects of bamboo charcoal added feed on reduction of ammonia and growth of Pangasius hypophthalmus. Journal of Aquaculture Research & Development, 5(6):??, ???? 2014. CODEN JARDCR. ISSN 2155-9546.

URL https://www.walshmedicalmedia.com/open-access/effects-
of-bamboo-charcoal-added-feed-on-reduction-of-ammonia-and-growth-
of-pangasius-hypophthalmus-28525.html.

Saha:2014:IEC

[177] M. Saha, A. Sarkar, and B. Bandhophadhyay. Introduction to establish the comparative analysis of 16S rRNA gene sequences with amoA and nxrA for nitrifying bacteria isolated from East Kolkata Wetland: an International Ramsar Site. Journal of Aquaculture Research & Development, 5(6):??, ???? 2014. CODEN JARDCR. ISSN 2155-9546. URL https://www.

walshmedicalmedia.com/open-access/introduction-to-establish-
the-comparative-analysis-of-16s-rrna-gene-sequences-with-amoa-
and-nxra-for-nitrifying-bacteria-28526.html.

Buric:2014:SER

[178] M. Buric, J. Bláhovec, and J. Kouril. A simple and effective recirculating hatchery for salmonids. Journal of Aquaculture Research & Development, 5(6):??, ???? 2014. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
a-simple-and-effective-recirculating-hatchery-for-salmonids-
28527.html.

Carani:2014:EGR

[179] Fernanda Regina Carani, Bruno Oliveira da Silva Duran, Warlen Pereira Piedade, Fernanda Antunes Alves da Costa, Vera Maria Fonseca de Almeida-Val, and Maeli Dal-Pai-Silva. Expression of growthrelated factors in skeletal muscle of pirarucu (Arapaima gigas) during growth. Journal of Aquaculture Research & Development, 5 (6):??, ???? 2014. CODEN JARDCR. ISSN 2155-9546. URL

https://www.walshmedicalmedia.com/open-access/expression-of-
growthrelated-factors-in-skeletal-muscle-of-pirarucuarapaima-
gigas-during-growth-28641.html.

Kasnir:2014:WQP

[180] Muhammad Kasnir, Harlina, and Rosmiati. Water quality parameter analysis for the feasibility of shrimp culture in Takalar Regency, Indonesia. Journal of Aquaculture Research & Development, 5(6):??, ???? 2014. CODEN JARDCR. ISSN 2155-9546. URL

https://www.walshmedicalmedia.com/open-access/water-quality-
parameter-analysis-for-the-feasibility-of-shrimp-culture-in-
takalar-regency-indonesia-28642.html.

Zainuddin:2014:EDC

[181] Zainuddin, Haryati, and Siti Aslamyah. Effect of dietary carbohydrate levels and feeding frequencies on growth and carbohydrate digestibility by white shrimp Litopenaeus vannamei under laboratory conditions. Journal of Aquaculture Research & Development, 5(6):??, ???? 2014. CODEN JARDCR. ISSN 2155-9546.

URL https://www.walshmedicalmedia.com/open-access/effect-of-
dietary-carbohydrate-levels-and-feeding-frequencies-on-growth-
and-carbohydrate-digestibility-by-white-shrimp-l-28643.html.

Yang:2014:EMB

[182] Chih-Chiu Yang, Shiu-Nan Chen, Chung-Lun Lu, Sherwin Chen, Kam-Chiu Lai, and Wen-Liang Liao. Effect of mushroom beta glucan (MBG) on immune and haemocyte response in Pacific white shrimp (Litopenaeus vannamei). Journal of Aquaculture Research & Development, 5(6):??, ???? 2014. CODEN JARDCR. ISSN 2155-9546.

URL https://www.walshmedicalmedia.com/open-access/effect-of-
mushroom-beta-glucan-mbg-on-immune-and-haemocyte-response-in-
pacific-white-shrimp-litopenaeus-vannamei-28644.html.

Cipriano:2014:ETM

[183] A. Cipriano, G. Burnell, S. Culloty, and S. Long. Evaluation of 3 tagging methods in marking sea urchin, Paracentrotuslividus, populations under both laboratory and field conditions. Journal of Aquaculture Research & Development, 5(6):1–7, ???? 2014. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
evaluation-of-3-tagging-methods-in-marking-sea-urchinparacentrotuslividus-
populations-under-both-laboratory-and-field-co-28645.html.

Rao:2014:CBH

[184] M. Vishwas Rao and T. T. Ajith Kumar. Captive breeding and hatchery production of mouth brooding jewel cardinal perch, Pterapogon kauderni, (Koumanns, 1933) using brackish water: The role of live prey and green water enrichment in juvenile production. Journal of Aquaculture Research & Development, 5(7):??, ???? 2014. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/captive-breeding-and-hatchery-production-of-
mouth-brooding-jewel-cardinal-perch-pterapogon-kauderni-2155-
9546-5-277.pdf.

Junaidi:2014:LDO

[185] Junaidi, Hafrijal Syandri, and Azrita. Loading and distribution of organic materials in Maninjau Lake West Sumatra Province-Indonesia. Journal of Aquaculture Research & Development, 5(7):??, ???? 2014. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/loading-and-distribution-of-organic-materials-
in-maninjau-lake-west-sumatra-provinceindonesia-28674.html.

Chatterjee:2014:SMV

[186] N. R. Chatterjee, D. Sahoo, and C. Chetri. Study on the monthly variation in hydro biological condition and its relation to fish production of a sewage fed Bheri System at Suburban Kolkata. Journal of Aquaculture Research & Development, 5(7):??, ???? 2014. CODEN JARDCR. ISSN 2155-9546.

URL https://www.walshmedicalmedia.com/open-access/study-on-
the-monthly-variation-in-hydro-biological-condition-and-its-
relation-to-fish-production-of-a-sewage-fed-bheri-s-28675.html.

Soundarapandian:2014:ETE

[187] P. Soundarapandian, G. K. Dinakaran, and D. Varadharajan. Effect of temperatures on the embryonic development, morphometrics and survival of Macrobrachium idella Idella (Hilgendorf, 1898). Journal of Aquaculture Research & Development, 5(7):1–6, ???? 2014. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
effect-of-temperatures-on-the-embryonic-development-morphometrics-
and-survival-of-macrobrachium-idella-idella-hilgendorf-28676.
html.

Bailung:2014:SIB

[188] Bhenila Bailung and S. P. Biswas. Successful induced breeding of a Bagrid catfish, Mystus dibrugarensis in captive condition. Journal of Aquaculture Research & Development, 5(7):1–3, ???? 2014. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
successful-induced-breeding-of-a-bagrid-catfish-mystus-dibrugarensis-
in-captive-condition-28677.html.

Olusegun:2014:HRS

[189] Akinwole A. Olusegun and Oguntuga O. Adedayo. Haematological responses, serum biochemistry and histology of Clarias gariepinus (Burchell, 1822) exposed to sublethal concentrations of cold water fresh root bark extracts of Plumbago zeylanica (leadwort). Journal of Aquaculture Research & Development, 5(7):??, ???? 2014. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/haematological-responses-serum-biochemistry-
and-histology-of-clarias-gariepinus-burchell-1822-exposed-to-
sublethal-conce-28678.html.

Chandanapalli:2014:DDA

[190] Suresh Babu Chandanapalli, Sreenivasa Reddy E, and Rajya Lakshmi D. Design and deployment of aqua monitoring system using wireless sensor networks and IAR-Kick. Journal of Aquaculture Research & Development, 5(7):??, ???? 2014. CODEN JARDCR. ISSN 2155-9546. URL https://

www.walshmedicalmedia.com/open-access/design-and-deployment-
of-aqua-monitoring-system-using-wireless-sensor-networks-and-
iarkick-28679.html.

ElBaky:2014:CAA

[191] H. S. Abd El Baky, G. S. El-Baroty, A. E. Ibrahim, and F. K. El Baz. Cytotoxicity, antioxidants and antimicrobial activities of lipids extracted from some marine algae. Journal of Aquaculture Research & Development, 5(7):??, ???? 2014. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
cytotoxicity-antioxidants-and-antimicrobial-activities-of-lipids-
extracted-from-some-marine-algae-28680.html.

Hwang:2014:VCC

[192] Jae-Ho Hwang, Nam-Gil Kim, Hee-Chul Woo, Sung-Ju Rha, Seon-Jae Kim, and Tai-Sun Shin. Variation in the chemical composition of Saccharina japonica with harvest area and culture period. Journal of Aquaculture Research & Development, 5(7):??, ????

2014. CODEN JARDCR. ISSN 2155-9546. URL https://www.

walshmedicalmedia.com/open-access/variation-in-the-chemical-
composition-of-saccharina-japonica-with-harvest-area-and-culture-
period-28681.html.

Soundarapandian:2014:AAP

[193] P. Soundarapandian, D. Varadharajan, C. Sivasubramanian, and A. S. Irin Kumari. Amino acid profiles of ridged swimming crab, Charybdis natator Herbst. Journal of Aquaculture Research & Development, 5(7): ??, ???? 2014. CODEN JARDCR. ISSN 2155-9546. URL https:

//www.walshmedicalmedia.com/open-access/amino-acid-profiles-
of-ridged-swimming-crab-charybdis-natator-herbst-28682.html.

Soundarapandian:2014:AAH

[194] P. Soundarapandian, Shyamalendu Roy, and D. Varadharajan. Antioxidant activity in hard and soft shell crabs of Charybdis lucifera (Fabricius, 1798). Journal of Aquaculture Research & Development, 5(7): ??, ???? 2014. CODEN JARDCR. ISSN 2155-9546. URL https:/

/www.walshmedicalmedia.com/open-access/antioxidant-activity-
in-hard-and-soft-shell-crabs-of-charybdis-luciferafabricius-
1798-28683.html.

JafariZadeh:2014:DIH

[195] Mahsa JafariZadeh, Rahim Peyghan, and Shadi Eftekhar Manavi. The detection of Ichthyophonus hoferi in naturally infected fresh water ornamental fishes. Journal of Aquaculture Research & Development, 5 (7):??, ???? 2014. CODEN JARDCR. ISSN 2155-9546. URL

https://www.walshmedicalmedia.com/open-access/the-detection-
of-ichthyophonus-hoferi-in-naturally-infected-fresh-water-ornamental-
fishes-28684.html.

Gupta:2015:RSS

[196] Sandipan Gupta. Review on Sperata seenghala (Sykes, 1839), a freshwater catfish of Indian Subcontinent. Journal of Aquaculture Research & Development, 6(1):??, ???? 2015. CODEN JARDCR. ISSN 2155-9546.

URL https://www.walshmedicalmedia.com/open-access/review-on-
sperata-seenghala-sykes-1839-a-freshwater-catfish-of-indian-
subcontinent-28685.html.

Budiati:2015:MQC

[197] Titik Budiati, Gulam Rusul, Wan Nadiah Wan-Abdullah, Rosma Ahmad, and Yahya Mat Arip. Microbiological quality of catfish (Clarias gariepinus) and tilapia (Tilapia mossambica) obtained from wet markets and ponds in Malaysia. Journal of Aquaculture Research & Development, 6(1):??, ???? 2015. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
microbiological-quality-of-catfish-clarias-gariepinus-and-tilapia-
tilapia-mossambica-obtained-from-wet-markets-and-ponds-28830.
html.

Lenzi:2015:WCD

[198] Mauro Lenzi. What can be done about massive macroalgal blooms? Journal of Aquaculture Research & Development, 6(1):??, ???? 2015. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/what-can-be-done-about-massive-macroalgal-blooms-
2155-9546-6-292.pdf.

Rasco:2015:EWA

[199] Barbara Rasco. Electrolyzed water applications in aquaculture and the seafood industry. Journal of Aquaculture Research & Development, 6(1): ??, ???? 2015. CODEN JARDCR. ISSN 2155-9546. URL https://www.

walshmedicalmedia.com/open-access/electrolyzed-water-applications-
in-aquaculture-and-the-seafood-industry-2155-9546-6-293.pdf.

Sayeda:2015:ESD

[200] M. A. Sayeda, M. Y. Eman, M. K. Amany, B. I. Taghreed, and T. A. Wafaa. Effect of El-Sail drain wastewater on Nile tilapia (Oreochromis niloticus) from River Nile at Aswan, Egypt. Journal of Aquaculture Research & Development, 6(1):??, ???? 2015. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
effect-of-elsail-drain-wastewater-on-nile-tilapia-oreochromis-
niloticusfrom-river-nile-at-aswan-egypt-28833.html.

Charoenwattanasak:2015:DAG

[201] Siripavee Charoenwattanasak, Rakpong Petkham, Aruneepong Srisathapom, and Pornthep Niamphithak. The development and application of genetic markers for giant freshwater prawns Macrobrachium rosenbergii by microsatellites. Journal of Aquaculture Research & Development, 6(1): ??, ???? 2015. CODEN JARDCR. ISSN 2155-9546. URL https:

//www.walshmedicalmedia.com/open-access/the-development-and-
application-of-genetic-markers-for-giant-freshwater-prawns-macrobrachium-
rosenbergii-by-microsatelli-28834.html.

Sarkar:2015:RFC

[202] Uttam Kumar Sarkar, Jyoti Sharma, and Bijoy Kali Mahapatra. A review on the fish communities in the Indian reservoirs and enhancement of fisheries and aquatic environment. Journal of Aquaculture Research & Development, 6(1):??, ???? 2015. CODEN JARDCR. ISSN 2155-

9546. URL https://www.walshmedicalmedia.com/open-access/a-
review-on-the-fish-communities-in-the-indian-reservoirs-and-
enhancement-of-fisheries-and-aquatic-environment-28835.html.

Uddowla:2015:CEC

[203] Md. Hasan Uddowla, Ah Ran Kim, Won gyu Park, and Hyun-Woo Kim. cDNAs encoding chitin synthase from shrimp (Pandalopsis japonica): Molecular characterization and expression analysis. Journal of Aquaculture Research & Development, 6(1):??, ???? 2015. CODEN JARDCR. ISSN 2155-9546. URL https://

www.walshmedicalmedia.com/open-access/cdnas-encoding-chitin-
synthase-from-shrimp-pandalopsis-japonica-molecular-characterization-
and-expression-analysis-28836.html.

Elgaml:2015:PES

[204] Shimaa A. Elgaml, Ryad Khalil, Emad A. Hashish, and Abdelhakeem El-Murr. Protective effects of selenium and alpha-tocopherol against lead-induced hepatic and renal toxicity in Oreochromis niloticus. Journal of Aquaculture Research & Development, 6(1):??, ???? 2015. CODEN JARDCR. ISSN 2155-9546. URL https://

www.walshmedicalmedia.com/open-access/protective-effects-of-
selenium-and-alphatocopherol-against-leadinduced-hepatic-and-
renal-toxicity-in-oreochromis-nilotic-28837.html.

Jaafar:2015:DDE

[205] Rzgar M. Jaafar, Jakob Skov, Per W. Kania, and Kurt Buchmann. Dose dependent effects of dietary immunostimulants on rainbow trout immune parameters and susceptibility to the parasite Ichthyophthirius multifiliis. Journal of Aquaculture Research & Development, 6(1):??, ???? 2015. CODEN JARDCR. ISSN 2155-9546. URL https://www.

walshmedicalmedia.com/open-access/dose-dependent-effects-of-
dietary-immunostimulants-on-rainbow-trout-immune-parameters-
and-susceptibility-to-the-parasite-26163.html.

Takahashi:2015:IEU

[206] Megumu Takahashi, Koji Mikami, Hiroyuki Mizuta, and Naotsune Saga. Identification and efficient utilization of antibiotics for the development of a stable transformation system in Porphyra yezoensis (Bangiales, Rhodophyta). Journal of Aquaculture Research & Development, 6(1):??, ???? 2015. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
identification-and-efficient-utilization-of-antibiotics-for-

the-development-of-a-stable-transformation-system-in-emporph-
26505.html.

Conti:2015:DOO

[207] G. Oliveri Conti, C. Ledda, M. Zuccarello, M. Fiore, R. Fallico, S. Sciacca, and M. Ferrante. Detection of Ostreopsis ovata from Eastern Sicily Coast using SEM microscopy coupled to the Vibrio fischeri test. Journal of Aquaculture Research & Development, 6(1):??, ???? 2015. CODEN JARDCR. ISSN 2155-9546.

URL https://www.walshmedicalmedia.com/open-access/detection-
of-emostreopsis-ovataem-from-eastern-sicily-coast-using-sem-
microscopy-coupled-to-the-emvibrio-fischeriem-test-26182.html.

Ninan:2015:ECM

[208] George Ninan, Lalitha K. V, Zynudheen A. A, and Jose Joseph. Effect of chilling on microbiological, biochemical and sensory attributes of whole aquacultured rainbow trout (Oncorhynchus mykiss Walbaum, 1792). Journal of Aquaculture Research & Development, 6(1):??, ???? 2015. CODEN JARDCR. ISSN 2155-9546. URL https://

www.walshmedicalmedia.com/open-access/effect-of-chilling-on-
microbiological-biochemical-and-sensory-attributes-of-whole-
aquacultured-rainbow-trout-oncorhynchu-26183.html.

Leung:2015:GES

[209] Christelle Leung, Pierre Magnan, and Bernard Angers. Genetic evidence for sympatric populations of yellow perch (Perca flavescens) in Lake Saint-Pierre (Canada): the crucial first step in developing a fishery management plan. Journal of Aquaculture Research & Development, 6(1):??, ???? 2015. CODEN JARDCR.

ISSN 2155-9546. URL https://www.walshmedicalmedia.com/open-
access/genetic-evidence-for-sympatric-populations-of-yellow-
perch-perca-flavescens-in-lake-saintpierre-canada-the-crucial-
first-26222.html.

Sanz:2015:GDE

[210] A. Sanz, M. Furné, M. C. Hidalgo, A. Domezain, and M. García-Gallego. Growth and digestive enzymatic profile of Acipenser naccarii and Oncorhynchus mykiss fed on different dietary macronutrient levels. A comparative study. Journal of Aquaculture Research & Development, 6(2): ??, ???? 2015. CODEN JARDCR. ISSN 2155-9546. URL https:/

/www.walshmedicalmedia.com/open-access/growth-and-digestive-
enzymatic-profile-of-acipenser-naccarii-and-oncorhynchus-mykiss-
fed-on-different-dietary-macronutri-28851.html.

Islam:2015:DPL

[211] M. A. Islam, M. Asadujjaman, S. Biswas, M. Manirujjaman, M. Rahman, M. A. Hossain, A. M. M. Uddin, M. Asaduzzaman, M. S. Rahman, and S. Munira. Determination of protein, lipid and carbohydrate contents of conventional and non-conventional feed items used in carp polyculture pond. Journal of Aquaculture Research & Development, 6(2):??, ???? 2015. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
determination-of-protein-lipid-and-carbohydrate-contents-of-
conventional-and-nonconventional-feed-items-used-in-carp-pol-
28852.html.

Girao:2015:LSF

[212] P. R. N. Vieira-Girão, I. R. C. B. Rocha, M. Gazzieno, P. R. N. Vieira, H. M. R. Lucena, F. H. F. Costa, and Rádis-Baptista. Low salinity facilitates the replication of infectious myonecrosis virus and viral co-infection in the shrimp Litopenaeus vannamei. Journal of Aquaculture Research & Development, 6(2):1–6, ???? 2015. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/low-salinity-facilitates-the-replication-of-
infectious-myonecrosis-virus-and-viral-coinfection-in-the-shrimp-
litopenaeus-28853.html.

Rasco:2015:HHI

[213] Barbara Rasco, Kenny Down, and Mahmoudreza Ovissipour. Humane harvesting initiative: The influence of harvest and post-harvest handling practices on fish welfare and product quality. Journal of Aquaculture Research & Development, 6(2):??, ???? 2015. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
humane-harvesting-initiative-the-influence-of-harvest-and-postharvest-
handling-practices-on-fish-welfare-and-product-quality-2155-
9546-6-303.pdf.

Mude:2015:GCW

[214] Jagadish Naik Mude and Danya Babu Ravuru. Growth of cultured white leg shrimp Litopenaeus vannamei (Boone, 1931) of brackish water culture system in winter season with artificial diet. Journal of Aquaculture Research & Development, 6(2):??, ???? 2015. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
growth-of-cultured-white-leg-shrimp-litopenaeus-vannamei-boone-
1931of-brackish-water-culture-system-in-winter-season-wit-28855.
html.

Moosavi:2015:EDL

[215] Mahsa Javadi Moosavi and Vali-Allah Jafari Shamushaki. Effects of different levels of copper sulfate on growth and reproductive performances in guppy (P. reticulate). Journal of Aquaculture Research & Development, 6(2):??, ???? 2015. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
effects-of-different-levels-of-copper-sulfate-on-growth-and-
reproductive-performances-in-guppy-p-reticulate-28856.html.

Tunde:2015:EAC

[216] Adeniyi Bashir Tunde, M. P. Kuton, Ayegbokiki Adedayo Oladipo, and Lawal Hakeem Olasunkanmi. Economic analyze of costs and return of fish farming in Saki-East Local Government Area of Oyo State, Nigeria. Journal of Aquaculture Research & Development, 6(2):??, ???? 2015. CODEN JARDCR. ISSN 2155-9546. URL https://www.

walshmedicalmedia.com/open-access/economic-analyze-of-costs-
and-return-of-fish-farming-in-sakieast-local-government-area-
of-oyo-state-nigeria-28857.html.

Aripin:2015:EEM

[217] Siti-Ariza Aripin, Orapint Jintasataporn, and Ruangvit Yoonpundh. Effects of exogenous melatonin in Clarias macrocephalus male broodstock first puberty stage. Journal of Aquaculture Research & Development, 6 (2):??, ???? 2015. CODEN JARDCR. ISSN 2155-9546. URL https:/

/www.walshmedicalmedia.com/open-access/effects-of-exogenous-
melatonin-in-clarias-macrocephalus-male-broodstock-first-puberty-
stage-28858.html.

Bekele:2015:PIP

[218] Jossy Bekele and Daniel Hussien. Prevalence of internal parasites of Oreochromis niloticus and Clarias gariepinus fish species in Lake Ziway, Ethiopia. Journal of Aquaculture Research & Development, 6 (2):??, ???? 2015. CODEN JARDCR. ISSN 2155-9546. URL

https://www.walshmedicalmedia.com/open-access/prevalence-of-
internal-parasites-of-oreochromis-niloticus-and-clarias-gariepinus-
fish-species-in-lake-ziway-ethiopia-28859.html.

Azrita:2015:EPS

[219] Azrita, Yuneidi Basri, and Hafrijal Syandri. EA preliminary study on domestication of bluespotted snakehead (Channa lucius, Channidae) in concrete tank. Journal of Aquaculture Research & Development, 6(2):??, ???? 2015. CODEN JARDCR. ISSN 2155-

9546. URL https://www.walshmedicalmedia.com/open-access/ea-
preliminary-study-on-domestication-of-bluespotted-snakehead-
channa-lucius-channidae-in-concrete-tank-28860.html.

Biswas:2015:CMO

[220] S. P. Biswas, Santosh Kumar Singh A, and J. N. Das. Conservation and management of ornamental fish resources of North East India. Journal of Aquaculture Research & Development, 6(3):??, ???? 2015. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/conservation-and-management-of-ornamental-fish-
resources-of-north-east-india-28916.html.

Petersen:2015:RCB

[221] Elizabeth H. Petersen, Brett D. Glencross, Nguyen Van Tien, Le Anh Tuan, Vu An Tuan, and Truong Ha Phuong. Recent changes in the bioeconomic of finfish mariculture in Vietnam. Journal of Aquaculture Research & Development, 6(3):??, ???? 2015. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
recent-changes-in-the-bioeconomic-of-finfish-mariculture-in-
vietnam-28917.html.

Khalafalla:2015:ESU

[222] Malik M. Khalafalla and Abd elaziz M. A. El-Hais. Evaluation of seaweeds Ulva rigida and Pterocladia capillaceaas dietary supplements in Nile tilapia fingerlings. Journal of Aquaculture Research & Development, 6(3):??, ???? 2015. CODEN JARDCR. ISSN 2155-9546. URL

https://www.walshmedicalmedia.com/open-access/evaluation-of-
seaweeds-ulva-rigida-and-pterocladia-capillaceaas-dietary-supplements-
in-nile-tilapia-fingerlings-28918.html.

Mahdy:2015:AMI

[223] Olfat A. Mahdy and Nadia M. T. Abu El Ezz. Additional morphological information on young female Caligus kuwaitensis (Copepoda Siphonostomatoida) from Egyptian marine water fish (Pagrus pagrus). Journal of Aquaculture Research & Development, 6(3):??, ???? 2015. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/additional-morphological-information-on-young-
female-caligus-kuwaitensiscopepoda-siphonostomatoida-from-egyptian-
marine--28909.html.

Green:2015:ESL

[224] Bartholomew W. Green and Kevin K. Schrader. Effect of stocking large channel catfish in a biofloc technology production system on production and incidence of common microbial off-flavor compounds. Journal of Aquaculture Research & Development, 6(3):??, ???? 2015. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/effect-of-stocking-large-channel-catfish-in-
a-biofloc-technology-production-system-on-production-and-incidence-
of-common-28910.html.

Das:2015:ESD

[225] Basanta Kumar Das and Jyotirmayee Pradhan. Effects of supplementation diet containing Microcystis aeruginosa on haematological and biochemical changes in Labeo rohita infected with Aeromonas hydrophila. Journal of Aquaculture Research & Development, 6(3):??, ???? 2015. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/effects-of-supplementation-diet-containing-microcystis-
aeruginosa-on-haematological-and-biochemical-changes-in-labeo-
roh-28913.html.

Reshid:2015:SCT

[226] Mohammed Reshid, Marshet Adugna, Yisehak Tsegaye Redda, Nesibu Awol, and Awot Teklu. A study of Clinostomum (trematode) and Contracaecum (nematode) parasites affecting Oreochromis niloticus in Small Abaya Lake, Silite Zone, Ethiopia. Journal of Aquaculture Research & Development, 6(3):??, ???? 2015. CODEN JARDCR. ISSN 2155-9546.

URL https://www.walshmedicalmedia.com/open-access/a-study-
of-clinostomum-trematode-and-contracaecum-nematode-parasites-
affecting-oreochromis-niloticus-in-small-abaya-lake-28912.html.

Iromo:2015:ODT

[227] Heppi Iromo, M. Zairin Junior, M. Agus Suprayudi, and Wasmen Manalu. The optimum dose of thyroxine hormone supplementation in broodstock mud crab (Scylla serrata) to accelerate ovarian maturation. Journal of Aquaculture Research & Development, 6(3):??, ???? 2015. CODEN JARDCR. ISSN 2155-9546. URL https:

//www.walshmedicalmedia.com/open-access/the-optimum-dose-of-
thyroxine-hormone-supplementation-in-broodstock-mud-crab-scylla-
serrata-to-accelerate-ovarian-matura-28911.html.

Khater:2015:EFR

[228] El-Sayed G. Khater and Samir A. Ali. Effect of flow rate and length of gully on lettuce plants in aquaponic and hydroponic systems. Journal of Aquaculture Research & Development, 6(3):??, ???? 2015. CODEN JARDCR. ISSN 2155-9546. URL https:

//www.walshmedicalmedia.com/open-access/effect-of-flow-rate-
and-length-of-gully-on-lettuce-plants-in-aquaponic-and-hydroponic-
systems-28914.html.

Zimba:2015:SAO

[229] Paul V. Zimba and Casey C. Grimm. Statistical approaches to optimize detection of MIB off-flavor in aquaculture raised channel catfish. Journal of Aquaculture Research & Development, 6(3):??, ???? 2015. CODEN JARDCR. ISSN 2155-9546. URL https://www.

walshmedicalmedia.com/open-access/statistical-approaches-to-
optimize-detection-of-mib-offflavor-in-aquaculture-raised-channel-
catfish-28952.html.

Ibrahim:2015:EDF

[230] E. B. Ibrahim, A. K. El-Tarabily, A. A. Kassab, M. E. Abdel-Fattah, and N. M. Rasheed. Evaluation of date fiber as feed ingredient for Nile tilapia Oreochromis niloticus fingerlings. Journal of Aquaculture Research & Development, 6(3):??, ???? 2015. CODEN JARDCR.

ISSN 2155-9546. URL https://www.walshmedicalmedia.com/open-
access/evaluation-of-date-fiber-as-feed-ingredient-for-nile-
tilapia-oreochromis-niloticus-fingerlings-28915.html.

Grace:2015:EGM

[231] Olaleye Ibukun Grace. Effects of grasshopper meal in the diet of Clarias gariepinus fingerlings. Journal of Aquaculture Research & Development, 6(4):??, ???? 2015. CODEN JARDCR. ISSN 2155-9546.

URL https://www.walshmedicalmedia.com/open-access/effects-
of-grasshopper-meal-in-the-diet-of-clarias-gariepinus-fingerlings-
29070.html.

Eisapour:2015:CRM

[232] Mina Eisapour, Seed Jafar Seyfabadi, and Behnam Daghooghi. Comparative radular morphology in some intertidal gastropods along Hormozgan Province, Iran. Journal of Aquaculture Research & Development, 6 (4):??, ???? 2015. CODEN JARDCR. ISSN 2155-9546. URL https:

//www.walshmedicalmedia.com/open-access/comparative-radular-
morphology-in-some-intertidal-gastropods-along-hormozgan-province-
iran-29071.html.

Kritsanapuntu:2015:UTC

[233] Sirusa Kritsanapuntu and Nilnaj Chaitanawisuti. Use of tuna-cooking liquid effluent as a dietary protein and lipid source replacing fishmeal in formulated diets for growing hatchery-reared juvenile spotted babylon (Babylonia areolata). Journal of Aquaculture Research & Development, 6(4):??, ???? 2015. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
use-of-tunacooking-liquid-effluent-as-a-dietary-protein-and-

lipid-source-replacing-fishmeal-in-formulated-diets-for-grow-
29072.html.

Kritsanapuntu:2015:RFP

[234] Sirusa Kritsanapuntu and Nilnaj Chaitanawisuti. Replacement of fishmeal by poultry by-product meal in formulated diets for growing hatcheryreared juvenile spotted babylon (Babylonia areolata). Journal of Aquaculture Research & Development, 6(4):1–6, ???? 2015. CODEN JARDCR.

ISSN 2155-9546. URL https://www.walshmedicalmedia.com/open-
access/replacement-of-fishmeal-by-poultry-byproduct-meal-in-
formulated-diets-for-growing-hatchery19516219483641948220reared-
juv-29073.html.

Dada:2015:ITO

[235] Adekunle Ayokanmi Dada. Improvement of tilapia (Oreochromis niloticus Linnaeus, 1758) growth performance fed three commercial feed additives in diets. Journal of Aquaculture Research & Development, 6(4):??, ???? 2015. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
improvement-of-tilapia-oreochromis-niloticus-linnaeus-1758-growth-
performance-fed-three-commercial-feed-additives-in-die-29074.
html.

Hasan:2015:GVA

[236] Innifa Hasan and Mrigendra Mohan Goswami. Genetic variation among cat fish (Mystus vittatus) population assessed by randomly amplified polymorphic (RAPD) markers from Assam, India. Journal of Aquaculture Research & Development, 6(4):??, ???? 2015. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
genetic-variation-among-cat-fish-mystus-vittatus-population-
assessed-by-randomly-amplified-polymorphic-rapd-markers-from-
29075.html.

Yogeshwari:2015:PDL

[237] Govintharaj Yogeshwari, Chandrasekar Jagruthi, Jesu Arockiaraj, and Ramasamy Harikrishnan. Poly D, L-lactide-co-glycolic acid (PLGA)- encapsulated CpG-oligonucleotide (ODN) on immune response in Cyprinus carpio against Aeromonas hydrophila. Journal of Aquaculture Research & Development, 6(4):??, ???? 2015. CODEN JARDCR. ISSN 2155-

9546. URL https://www.walshmedicalmedia.com/open-access/poly-
d-llactidecoglycolic-acid-plgaencapsulated-cpgoligonucleotide-
odn-on-immune-response-in-cyprinus-carpio-against-aer-29076.
html.

Authman:2015:UFB

[238] M. M. N. Authman, M. S. Zaki, E. A. Khallaf, and H. H. Abbas. Use of fish as bio-indicator of the effects of heavy metals pollution. Journal of Aquaculture Research & Development, 6(4):??, ???? 2015. CODEN JARDCR.

ISSN 2155-9546. URL https://www.walshmedicalmedia.com/open-
access/use-of-fish-as-bioindicator-of-the-effects-of-heavy-metals-
pollution-29077.html.

Ballesteros:2015:GDA

[239] Natalia Ballesteros, Néstor Aguirre, Julio Coll, Sara I. Pérez-Prieto, and Sylvia Rodríguez Saint-Jean. Gene2Path: a data analysis tool to study fish gene pathways by automatic search of orthologous genes. Journal of Aquaculture Research & Development, 6(4):??, ???? 2015. CODEN JARDCR. ISSN 2155-9546. URL https://www.

walshmedicalmedia.com/open-access/gene2path-a-data-analysis-
tool-to-study-fish-gene-pathways-by-automatic-search-of-orthologous-
genes-29078.html.

Kusumaningrum:2015:CEC

[240] Indah Kusumaningrum. Coast environment conservation by home scale grouper culture in Pecaron–Situbondo. Journal of Aquaculture Research & Development, 6(4):??, ???? 2015. CODEN JARDCR.

ISSN 2155-9546. URL https://www.walshmedicalmedia.com/open-
access/coast-environment-conservation-by-home-scale-grouper-
culture-in-pecaron8211situbondo-29079.html.

Claudia:2015:EDW

[241] Carmona-Osalde Claudia, Puerto-Novelo Enrique, and Miguel Rodriguez-Serna. Effect of decreasing water levels on the spawning rate and egg count of female crayfish Procambarus (Austrocambarus) llamasi (Villalobos, 1955). Journal of Aquaculture Research & Development, 6(4): 1–4, ???? 2015. CODEN JARDCR. ISSN 2155-9546. URL https:/

/www.walshmedicalmedia.com/open-access/effect-of-decreasing-
water-levels-on-the-spawning-rate-and-egg-count-of-female-crayfish-
procambarus-austrocambarus-llama-29080.html.

Zabetakis:2015:UNR

[242] Ioannis Zabetakis. The urgent need to re-think outside the “Omega-3 Pufas” box. Journal of Aquaculture Research & Development, 6(5): ??, ???? 2015. CODEN JARDCR. ISSN 2155-9546. URL https://www.

walshmedicalmedia.com/open-access/the-urgent-need-to-rethink-
outside-the-8220omega3-pufas8221-box-29147.html.

Geremew:2015:DSC

[243] Akewake Geremew, Abebe Getahun, and Krishen Rana. Digestibility of soybean cake, Niger seed cake and linseed cake in juvenile Nile tilapia, Oreochromis niloticus L. Journal of Aquaculture Research & Development, 6(5):??, ???? 2015. CODEN JARDCR. ISSN 2155-9546. URL

https://www.walshmedicalmedia.com/open-access/digestibility-
of-soybean-cake-niger-seed-cake-and-linseed-cake-in-juvenile-
nile-tilapia-oreochromis-niloticus-l-29148.html.

Fedorovykh:2015:ELC

[244] J. V. Fedorovykh, S. V. Ponomarev, J. M. Bakaneva, N. M. Bakanev, J. V. Sergeeva, A. A. Bakhareva, J. N. Grozesku, and V. I. Egorova. The effect of lipid composition in diets on ovicell generating of the Russian sturgeon females. Journal of Aquaculture Research & Development, 6 (5):??, ???? 2015. CODEN JARDCR. ISSN 2155-9546. URL https:

//www.walshmedicalmedia.com/open-access/the-effect-of-lipid-
composition-in-diets-on-ovicell-generating-of-the-russian-sturgeon-
females-29149.html.

Lenzi:2015:CLB

[245] Mauro Lenzi. A closer link between politics and the scientific community, as well as global coordination are essential to tackle environmental collapse. Journal of Aquaculture Research & Development, 6(5): ??, ???? 2015. CODEN JARDCR. ISSN 2155-9546. URL https://

www.walshmedicalmedia.com/open-access/a-closer-link-between-
politics-and-the-scientific-community-as-well-as-global-coordination-
are-essential-to-tackle-environmental-collapse-2155-9546-1000335.
pdf.

Rehman:2015:IFF

[246] Muhammad Hafeez ur Rehman, Khalid Javed Iqbal, Farzana Abbas, Mirza Muhammad Haroon Mushtaq, Fayyaz Rasool, and Shakeela Parveen. Influence of feeding frequency on growth performance and body indices of goldfish (Carrassius auratus). Journal of Aquaculture Research & Development, 6(5):??, ???? 2015. CODEN JARDCR. ISSN 2155-9546.

URL https://www.walshmedicalmedia.com/open-access/influence-
of-feeding-frequency-on-growth-performance-and-body-indices-
of-goldfish-carrassius-auratus-29151.html.

Kurnia:2015:MCR

[247] A. Kurnia, S. Satoh, Y. Haga, H. Kudo, M. Nakada, H. Matsumura, Y. Watanabe, and S. Adachi. Muscle coloration of rainbow trout with astaxanthin sources from marine bacteria and synthetic astaxanthin. Journal of Aquaculture Research & Development, 6(5):??, ???? 2015. CODEN JARDCR. ISSN 2155-9546. URL https:/

/www.walshmedicalmedia.com/open-access/muscle-coloration-of-
rainbow-trout-with-astaxanthin-sources-from-marine-bacteria-
and-synthetic-astaxanthin-29152.html.

Bellos:2015:ETS

[248] Georgios Bellos, Panagiotis Angelidis, and Helen Miliou. Effect of temperature and seasonality principal epizootiological risk factor on vibriosis and photobacteriosis outbreaks for European sea bass in Greece (1998–2013). Journal of Aquaculture Research & Development, 6(5):??, ???? 2015. CODEN JARDCR. ISSN 2155-9546.

URL https://www.walshmedicalmedia.com/open-access/effect-of-
temperature-and-seasonality-principal-epizootiological-risk-
factor-on-vibriosis-and-photobacteriosis-outbreaks-29153.html.

Ramirez:2015:MME

[249] Besay Ramírez, Leonor Ortega, Daniel Montero, Fernando Tuya, and Ricardo Haroun. Monitoring a massive escape of European sea bass (Dicentrarchus labrax) at an oceanic island: Potential species establishment. Journal of Aquaculture Research & Development, 6(5):??, ???? 2015. CODEN JARDCR. ISSN 2155-9546. URL https:/

/www.walshmedicalmedia.com/open-access/monitoring-a-massive-
escape-of-european-sea-bass-dicentrarchus-labrax-at-an-oceanic-
island-potential-species-establishme-29154.html.

Hartviksen:2015:PPE

[250] Mali Hartviksen, Jose L. Gonzalez Vecino, Anu Kettunen, Reidar Myklebust, Kari Ruohonen, Simon Wadsworth, and Einar Ringø. Probiotic and pathogen ex-vivo exposure of Atlantic salmon (Salmo salar L.) intestine from fish fed four different protein sources. Journal of Aquaculture Research & Development, 6(5):1–8, ???? 2015. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
probiotic-and-pathogen-exvivo-exposure-of-atlantic-salmon-salmo-
salar-l-intestine-from-fish-fed-four-different-protein-s-29155.
html.

Saka:2015:GDF

[251] Babatunde Akeem Saka and Olanike kudirat Adeyemo. Gonad development in the female Nigerian Clarias gariepinus Burchell 1822. Journal of Aquaculture Research & Development, 6(6):??, ???? 2015. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.

com/open-access/gonad-development-in-the-female-nigerian-clarias-
gariepinus-burchell-1822-29370.html.

Drawany:2015:BSR

[252] El-Drawany. On the biology of Siganus rivulatus inhabits Bitter Lakes in Egypt. Journal of Aquaculture Research & Development, 6(6):??, ???? 2015. CODEN JARDCR. ISSN 2155-9546. URL https://www.

walshmedicalmedia.com/open-access/on-the-biology-of-siganus-
rivulatus-inhabits-bitter-lakes-in-egypt-29371.html.

Ferosekhan:2015:ELD

[253] S. Ferosekhan, S. K. Sahoo, S. S. Giri, A. Saha, and M. Paramanik. Embryonic and larval development of yellow tail catfish, Pangasius pangasius. Journal of Aquaculture Research & Development, 6(6):??, ???? 2015. CODEN JARDCR. ISSN 2155-9546. URL https:/

/www.walshmedicalmedia.com/open-access/embryonic-and-larval-
development-of-yellow-tail-catfish-pangasius-pangasius-29372.
html.

deSouza:2015:MEW

[254] Gabriel Marcos Domingues de Souza, Lucienne Garcia Pretto-Giordano, Gislayne Trindade Vilas-Bôas, Túlio Oliveira de Carvalho, Ângela Teresa Silva-Souza, Mauro Caetano Filho, Ronaldos Tamanini, and Laurival Antônio Vilas-Boas. Microbiological evaluation of water and fillets in the production chain of Nile tilapia (Oreochromis niloticus). Journal of Aquaculture Research & Development, 6(6):??, ???? 2015. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/microbiological-evaluation-of-water-and-fillets-
in-the-production-chain-of-nile-tilapia-oreochromis-niloticus-
29373.html.

Mokhtar:2015:CSO

[255] Doaa M. Mokhtar. Comparative structural organization of skin in red-tail shark (Epalzeorhynchos bicolor) and guppy (Poecilia reticulata). Journal of Aquaculture Research & Development, 6(6):??, ???? 2015. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/comparative-structural-organization-of-skin-
in-redtail-sharkepalzeorhynchos-bicolor-and-guppy-poecilia-reticulata-
29374.html.

Elgendy:2015:CEP

[256] M. Y. Elgendy, M. Abdelsalam, M. Moustafa, A. M. Kenawy, and A. Seida. Caligus elongatus and Photobacterium damselae subsp piscicida concomitant infections affecting broodstock European seabass, Dicentrarchus labrax, with special reference to histopathological responses. Journal of Aquaculture Research & Development, 6(7):??, ???? 2015. CODEN JARDCR. ISSN 2155-9546. URL https://

www.walshmedicalmedia.com/open-access/caligus-elongatus-and-
photobacterium-damselae-subsp-piscicida-concomitant-infections-
affecting-broodstock-european-seaba-29375.html.

Aripin:2015:EZA

[257] Siti-Ariza Aripin, Orapint Jintasataporn, and Ruangvit Yoonpundh. Effects of zinc amino acid in walking catfish (Clarias macrocephalus) female broodstock first sexual maturation. Journal of Aquaculture Research & Development, 6(7):??, ???? 2015. CODEN JARDCR. ISSN 2155-9546.

URL https://www.walshmedicalmedia.com/open-access/effects-
of-zinc-amino-acid-in-walking-catfish-clarias-macrocephalus-
female-broodstock-first-sexual-maturation-29376.html.

El-Drawany:2015:GFF

[258] M. A. El-Drawany and W. G. Elnagar. Growth, food and feeding habits of Bagrus bayad and Bagrus docmac inhibiting Muess Channel, Sharkia Province, Egypt. Journal of Aquaculture Research & Development, 6(7):??, ???? 2015. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
growth-food-and-feeding-habits-of-bagrus-bayad-and-bagrus-docmac-
inhibiting-muess-channel-sharkia-province-egypt-29377.html.

Swetha:2015:RCF

[259] C. H. Swetha, B. P. Girish, and P. S. Reddy. Reproductive cycle and fecundity in natural population of edible freshwater crab, Oziothelphusa senex senex (Fabricius, 1798) (Decapoda: Brachyura). Journal of Aquaculture Research & Development, 6(7):??, ???? 2015. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/reproductive-cycle-and-fecundity-in-natural-
population-of-edible-freshwater-crab-oziothelphusa-senex-senex-
fabricius-179-29378.html.

Owhonda:2015:WSV

[260] K. N. Owhonda. Wet season variations of some physicochemical parameters of the brackish water farm, Buguma, Niger Delta, Nigeria. Journal of Aquaculture Research & Development, 6(7):??, ???? 2015. CODEN JARDCR. ISSN 2155-9546. URL https://

www.walshmedicalmedia.com/open-access/wet-season-variations-
of-some-physicochemical-parameters-of-the-brackish-water-farm-
buguma-niger-delta-nigeria-29379.html.

Islam:2015:SSA

[261] Md. Robiul Islam, Mamnur Rashid M, Md. Hashmi Sakib, and Most. Waheda Rahman Ansary. Serological studies of Aeromonas hydrophila in Bangladesh. Journal of Aquaculture Research & Development, 6(7): ??, ???? 2015. CODEN JARDCR. ISSN 2155-9546. URL https:

//www.walshmedicalmedia.com/open-access/serological-studies-
of-aeromonas-hydrophila-in-bangladesh-29380.html.

Adam:2015:SEE

[262] Bownik Adam and Stepniewska Zofia. Subchronic effects of ectoine on survival, growth and physiological parameters of Daphnia magna. Journal of Aquaculture Research & Development, 6(8):??, ???? 2015. CODEN JARDCR. ISSN 2155-9546. URL https://

www.walshmedicalmedia.com/open-access/subchronic-effects-of-
ectoine-on-survival-growth-and-physiological-parameters-of-daphnia-
magna-29466.html.

Khalafalla:2015:BTB

[263] Malik M. Khalafalla and B. El-Sayed. Biological treatment of broad bean hulls and its evaluation through tilapia fingerlings (Oreochromis niloticus) feeding. Journal of Aquaculture Research & Development, 6(8):??, ???? 2015. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
biological-treatment-of-broad-bean-hulls-and-its-evaluation-
through-tilapia-fingerlings-oreochromis-niloticus-feeding-29467.
html.

Mojekwu:2015:ATM

[264] T. O. Mojekwu and C. I. Anumudu. Advanced techniques for morphometric analysis in fish. Journal of Aquaculture Research & Development, 6 (8):??, ???? 2015. CODEN JARDCR. ISSN 2155-9546. URL https:

//www.walshmedicalmedia.com/open-access/advanced-techniques-
for-morphometric-analysis-in-fish-29468.html.

CasteloBrancoRocha:2015:TEA

[265] Ítalo Régis Castelo Branco Rocha, Tadeu Dote Sá, Rommel Rocha de Sousa, Gutemberg Costa de Lima, José Renato de Oliveira César, and Francisco Hiran Farias Costa. Technical and environmental analysis of shrimp farming in the Coreau River Estuary, Ceara State, Brazil. Journal of Aquaculture Research & Development, 6 (8):??, ???? 2015. CODEN JARDCR. ISSN 2155-9546. URL

https://www.walshmedicalmedia.com/open-access/technical-and-
environmental-analysis-of-shrimp-farming-in-the-coreau-river-
estuary-ceara-state-brazil-29469.html.

Reddy:2015:CRA

[266] P. Sreenivasula Reddy and M. Srilatha. 13-cis-retinoic acid-induced hyperglycemia in the fresh water edible crab, Oziothelphusa senex Senex is mediated by triggering release of hyperglycemic hormone from eyestalks. Journal of Aquaculture Research & Development, 6(8):??, ???? 2015. CODEN JARDCR. ISSN 2155-9546. URL https://www.

walshmedicalmedia.com/open-access/13cisretinoic-acidinduced-
hyperglycemia-in-the-fresh-water-edible-crab-oziothelphusa-senex-
senex-is-mediated-by-triggeri-29470.html.

Moniruzzaman:2015:ESD

[267] Mohammad Moniruzzaman, Kazi Belal Uddin, Sanjib Basak, Yahia Mahmud, Muhammad Zaher, and Sungchul C. Bai. Effects of stocking density on growth, body composition, yield and economic returns of monosex tilapia (Oreochromis niloticus L.) under cage culture system in Kaptai Lake of Bangladesh. Journal of Aquaculture Research & Development, 6 (8):??, ???? 2015. CODEN JARDCR. ISSN 2155-9546. URL https:

//www.walshmedicalmedia.com/open-access/effects-of-stocking-
density-on-growth-body-composition-yield-and-economic-returns-
of-monosex-tilapia-oreochromis-nilotic-29471.html.

ElShafei:2015:SHM

[268] Hussien M. El Shafei. Some heavy metals concentration in water, muscles and gills of Tilapia niloticus as biological indicator of Manzala Lake pollution. Journal of Aquaculture Research & Development, 6(9):??, ???? 2015. CODEN JARDCR. ISSN 2155-

9546. URL https://www.walshmedicalmedia.com/open-access/some-
heavy-metals-concentration-in-water-muscles-and-gills-of-tilapia-
niloticus-as-biological-indicator-of-manzala-lake--29472.html.

Suantika:2015:PZW

[269] G. Suantika, G. Lumbantoruan, H. Muhammad, F. F. N. Azizah, and P. Aditiawati. Performance of zero water discharge (ZWD) system with nitrifying bacteria and microalgae Chaetoceros calcitrans components in super intensive white shrimp (Litopenaeus vannamei) culture. Journal of Aquaculture Research & Development, 6(9):??, ???? 2015. CODEN JARDCR. ISSN 2155-9546. URL https://www.

walshmedicalmedia.com/open-access/performance-of-zero-water-

discharge-zwd-system-with-nitrifying-bacteria-and-microalgae-
chaetoceros-calcitrans-components-29473.html.

Om:2015:MCG

[270] A. D. Om, S. Sharif, S. Jasmani, Y. Y. Sung, and A. A. Bolong. Molecular characteristic of giant grouper (Epinephelus lanceolatus) Vitellogenin. Journal of Aquaculture Research & Development, 6(9):??, ???? 2015. CODEN JARDCR. ISSN 2155-9546.

URL https://www.walshmedicalmedia.com/open-access/molecular-
characteristic-of-giant-grouper-epinephelus-lanceolatusvitellogenin-
29475.html.

Orisasona:2015:GMU

[271] O. Orisasona and E. K. Ajani. The growth and mineral utilization of clarias gariepinus fingerlings fed phytase-supplemented toasted lima bean (Phaseolus lunatus) diets. Journal of Aquaculture Research & Development, 6(9):??, ???? 2015. CODEN JARDCR. ISSN 2155-

9546. URL https://www.walshmedicalmedia.com/open-access/the-
growth-and-mineral-utilization-of-clarias-gariepinus-fingerlings-
fed-phytasesupplemented-toasted-lima-bean-phaseolus-29477.html.

Miah:2015:PSC

[272] Md. Nasir Uddin Miah, Md. Mostafa Shamsuzzaman, Ahmed Harun-Al-Rashid, and Partho Protim Barman. Present status of coastal fisheries in Sitakunda Coast with special reference on climate change and fish catch. Journal of Aquaculture Research & Development, 6(9):??, ???? 2015. CODEN JARDCR. ISSN 2155-9546. URL https://www.

walshmedicalmedia.com/open-access/present-status-of-coastal-
fisheries-in-sitakunda-coast-with-special-reference-on-climate-
change-and-fish-catch-29478.html.

Pretto-Giordano:2015:SIU

[273] Lucienne Garcia Pretto-Giordano, Josiane Aniele Scarpassa, André Rocha Barbosa, Carla Suzuki Altrão, Carolina Galdino Gumiero Ribeiro, and Laurival Antônio Vilas-Boas. Streptococcus iniae: an unusual important pathogen fish in Brazil. Journal of Aquaculture Research & Development, 6(9):??, ???? 2015. CODEN JARDCR. ISSN 2155-9546. URL https:

//www.walshmedicalmedia.com/open-access/streptococcus-iniae-
an-unusual-important-pathogen-fish-in-brazil-29479.html.

Schrader:2015:AAM

[274] Kevin K. Schrader, Mark T. Hamann, James D. McChesney, Douglas L. Rodenburg, and Mohamed A. Ibrahim. Antibacterial activities of metabolites from Platanus occidentalis (American sycamore) against fish pathogenic bacteria. Journal of Aquaculture Research & Development, 6(10):??, ???? 2015. CODEN JARDCR. ISSN 2155-9546. URL

https://www.walshmedicalmedia.com/open-access/antibacterial-
activities-of-metabolites-from-platanus-occidentalis-american-
sycamore-against-fish-pathogenic-bacteria-29670.html.

Karthik:2015:ANI

[275] R. Karthik, K. Ramalingam, D. Yuvaraj, M. C. Vanitha, and R. Muthezhilan. Attenuation of negative impacts by micro algae and enriched Artemia salina on Penaeus monodon and Litopenaeus vannamei larval culture. Journal of Aquaculture Research & Development, 6(10):??, ???? 2015. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
attenuation-of-negative-impacts-by-micro-algae-and-enriched-
artemia-salina-on-penaeus-monodon-and-litopenaeus-vannamei-l-
29671.html.

Chanda:2015:TRF

[276] Angsuman Chanda. A taxonomic research on the first record of the genus Miyadiella kubo, 1949 by Kunju from Indian water. Journal of Aquaculture Research & Development, 6(10):??, ???? 2015. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/a-taxonomic-research-on-the-first-record-of-
the-genus-miyadiella-kubo-1949-by-kunju-from-indian-water-29672.
html.

Carmona-Osalde:2015:TPS

[277] Claudia Carmona-Osalde, Miguel Rodríguez-Serna, Héctor Hernández-Moreno, and J. L. Arredondo-Figueroa. Total and partial substitution of dietary fish oil with palm oil to juvenile crayfish, Procambarus llamasi. Journal of Aquaculture Research & Development, 6(10):??, ???? 2015. CODEN JARDCR. ISSN 2155-9546.

URL https://www.walshmedicalmedia.com/open-access/total-and-
partial-substitution-of-dietary-fish-oil-with-palm-oil-to-juvenile-
crayfish-procambarus-llamasi-29673.html.

Syandri:2015:PSF

[278] Hafrijal Syandri, Azrita, Niagara, and Junaidi. Preliminary study on the feeding schedule of laboratory reared of bonylip barb larva, Osteochilus vittatus Cyprinidae. Journal of Aquaculture Research & Development, 6 (10):??, ???? 2015. CODEN JARDCR. ISSN 2155-9546. URL https:/

/www.walshmedicalmedia.com/open-access/preliminary-study-on-

the-feeding-schedule-of-laboratory-reared-of-bonylip-barb-larva-
osteochilus-vittatus-cyprinidae-29674.html.

Kumar:2015:SHC

[279] Muneesh Kumar, Parvinder Kumar, and Sangeeta Devi. To study the histopathological changes in the gills of Clarias batrachus, an air breathing teleost after short term exposure of copper sulphate. Journal of Aquaculture Research & Development, 6(10):??, ???? 2015. CODEN JARDCR.

ISSN 2155-9546. URL https://www.walshmedicalmedia.com/open-
access/to-study-the-histopathological-changes-in-the-gills-of-
clarias-batrachus-an-air-breathing-teleost-after-short-term-
expos-29675.html.

Shamsuzzaman:2015:PSM

[280] M. M. Shamsuzzaman, A. H. A. Rashid, M. A. A. Mamun, S. K. Mazumder, and M. A. Haque. Present status of marine puffer fishes in Bangladesh. Journal of Aquaculture Research & Development, 6(10):??, ???? 2015. CODEN JARDCR. ISSN 2155-9546.

URL https://www.walshmedicalmedia.com/open-access/present-
status-of-marine-puffer-fishes-in-bangladesh-29676.html.

Sadeq:2015:PPA

[281] Qays Mohammed Sadeq and Wan Ismail Bin Wan Yusoff. Porosity and permeability analysis from well logs and core in fracture, Vugy and intercrystalline carbonate reservoirs. Journal of Aquaculture Research & Development, 6(10):1–5, ???? 2015. CODEN JARDCR. ISSN 2155-9546.

URL https://www.walshmedicalmedia.com/open-access/porosity-
and-permeability-analysis-from-well-logs-and-core-in-fracture-
vugy-and-intercrystalline-carbonate-reservoirs-29677.html.

Harlina:2015:PSK

[282] Harlina Harlina, Arief Prajitno, Eddy Suprayitno, Happy Nursyam, and Rosmiati. Potential study of kopasanda (Chromolaena odorata L.) leaves as antibacterial against Vibrio harveyi, disease causative agent of tiger shrimp (Penaeus monodon Fabricius) post larvae. Journal of Aquaculture Research & Development, 6(10):??, ???? 2015. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
potential-study-of-kopasanda-chromolaena-odorata-l-leaves-as-
antibacterial-against-vibrio-harveyi-disease-causative-agen-
29678.html.

Campuzano:2015:MAE

[283] F. J. Campuzano, J. M. Gutiérrez, T. Senabre, M. D. Mateus, A. Perán, A. Belmonte, V. Aliaga, and R. Neves. A modelling approach to estimate the environmental and productive carrying capacity for a Mediterranean coastal marine culture park. Journal of Aquaculture Research & Development, 6(11):??, ???? 2015. CODEN JARDCR. ISSN 2155-

9546. URL https://www.walshmedicalmedia.com/open-access/a-
modelling-approach-to-estimate-the-environmental-and-productive-
carrying-capacity-for-a-mediterranean-coastal-marine-c-29679.
html.

Mokhtar:2015:LSEa

[284] Doaa M. Mokhtar, Enas A. Abd-Elhafez, and A. H. Hassan. Light and scanning electron microscopic studies on the intestine of grass carp (Ctenopharyngodon idella): I — anterior intestine. Journal of Aquaculture Research & Development, 6(11):??, ???? 2015. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
light-and-scanning-electron-microscopic-studies-on-the-intestine-
of-grass-carp-ctenopharyngodon-idella-ianterior-intesti-29680.
html.

Mancuso:2015:SSP

[285] Monique Mancuso. Scorpaena scrofa: a promising aquaculture candidate for Sicilian aquaculture. Journal of Aquaculture Research & Development, 6(11):??, ???? 2015. CODEN JARDCR. ISSN 2155-9546.

URL https://www.walshmedicalmedia.com/open-access/scorpaena-
scrofa-a-promising-aquaculture-candidate-for-sicilian-aquaculture-
29681.html.

Singh:2015:AEE

[286] Narendra Singh, Ampee Tasung, Sonal Tripathi, Pathik Baldev Patel, Ajeet Mulchand Bafna, and Rattan Govind Patil. Aquaculture effluent: Effect on yield, nutrient content and uptake in Salicornia bachiata Roxb. Journal of Aquaculture Research & Development, 6(11): ??, ???? 2015. CODEN JARDCR. ISSN 2155-9546. URL https:/

/www.walshmedicalmedia.com/open-access/aquaculture-effluent-
effect-on-yield-nutrient-content-and-uptake-in-salicornia-bachiata-
roxb-29682.html.

Aryani:2015:EDP

[287] Netti Aryani and Indra Suharman. Effect of dietary protein level on the reproductive performance of female of green catfish (Hemibagrus nemurus Bagridae). Journal of Aquaculture Research & Development, 6(11):1–5, ???? 2015. CODEN JARDCR. ISSN 2155-9546. URL https://www.

walshmedicalmedia.com/open-access/effect-of-dietary-protein-
level-on-the-reproductive-performance-of-female-of-green-catfish-
hemibagrus-nemurus-bagridae--29683.html.

Bekeh:2015:GHI

[288] Ada Fidelis Bekeh, Ayotunde Ezekiel Olatunji, and William Kinsley Bassey. Gonado-hepatosomatic indexes of Clarias gariepinus sub-adult exposed to artrazine, Cocos nucifera water and phyllanthus muelarianus extract. Journal of Aquaculture Research & Development, 6(11): ??, ???? 2015. CODEN JARDCR. ISSN 2155-9546. URL https:

//www.walshmedicalmedia.com/open-access/gonadohepatosomatic-
indexes-of-clarias-gariepinus-subadult-exposed-to-artrazine-
cocos-nucifera-water-and-phyllanthus-mue-30338.html.

Iromo:2015:SDT

[289] H. Iromo, M. Z. Junior, M. S. Agus, and W. Manalu. Suplementation doses thyroxine hormone of broodstock mud crab (Scylla serrata) during ovarian maturation. Journal of Aquaculture Research & Development, 6(12):??, ???? 2015. CODEN JARDCR.

ISSN 2155-9546. URL https://www.walshmedicalmedia.com/open-
access/suplementation-doses-thyroxine-hormone-of-broodstock-
mud-crab-scylla-serrata-during-ovarian-maturation-29684.html.

Mokhtar:2015:LSEb

[290] D. M. Mokhtar, E. A. Abd-Elhafez, and A. H. Hassan. Light and scanning electron microscopic studies on the intestine of grass carp (Ctenopharyngodon idella): II — posterior intestine. Journal of Aquaculture Research & Development, 6(12):??, ???? 2015. CODEN JARDCR. ISSN 2155-9546.

URL https://www.walshmedicalmedia.com/open-access/light-and-
scanning-electron-microscopic-studies-on-the-intestine-of-grass-
carp-ctenopharyngodon-idella-iiposterior-intes-29685.html.

Mandal:2015:SCC

[291] Basudev Mandal, Sourabh Kumar Dubey, Dona Bhattacharya, and Bimal Kinkar Chand. Significance of cross-contamination on bacteriological quality of black tiger shrimp (Penaeus monodon Fabricius 1798) for export trade produced in fish processing plant. Journal of Aquaculture Research & Development, 6(12):??, ???? 2015. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
significance-of-crosscontamination-on-bacteriological-quality-
of-black-tiger-shrimp-penaeus-monodon-fabricius-1798-for-e-29686.
html.

Grasteau:2015:EGC

[292] Alexandra Grasteau, Thomas Guiraud, Patrick Daniel, Ségolène Calvez, Valérie Chesneau, and Michel Le Hénaff. Evaluation of glutaraldehyde, chloramine-T, bronopol, incimaxx aquatic(R) and hydrogen peroxide as biocides against Flavobacterium psychrophilum for sanitization of rainbow trout eyed eggs. Journal of Aquaculture Research & Development, 6(12):??, ???? 2015. CODEN JARDCR. ISSN 2155-9546. URL

https://www.walshmedicalmedia.com/open-access/evaluation-of-
glutaraldehyde-chloraminet-bronopol-incimaxx-aquatic1958218194174-
and-hydrogen-peroxide-as-biocides-agains-29687.html.

Varadharajan:2015:BAP

[293] D. Varadharajan and P. Soundarapandian. Biodiversity and abundance of phytoplankton from Muthupettai Mangrove Region, South East Coast of India. Journal of Aquaculture Research & Development, 6(12):??, ???? 2015. CODEN JARDCR. ISSN 2155-9546. URL https://www.

walshmedicalmedia.com/open-access/biodiversity-and-abundance-
of-phytoplankton-from-muthupettai-mangrove-region-south-east-
coast-of-india-29688.html.

Mohideen:2015:EPM

[294] Meeran Mohideen and M. A. Haniffa. Effect of probiotic on microbiological and haematological responsiveness of cat fish (Heteropnuestes fossilis) challenged with bacteria Aeromonas hydrophila and fungi Aphanomyces invadans. Journal of Aquaculture Research & Development, 6(12): ??, ???? 2015. CODEN JARDCR. ISSN 2155-9546. URL https:

//www.walshmedicalmedia.com/open-access/effect-of-probiotic-
on-microbiological-and-haematological-responsiveness-of-cat-
fish-heteropnuestes-fossilis-challenged--29689.html.

Chatterji:2015:SGJ

[295] Anil Chatterji, Siddhartha Pati, and B. P. Dash. A study on the growth of juveniles of tiger prawn, Penaeus monodon (Fabricius) under different photoperiods. Journal of Aquaculture Research & Development, 6(12): ??, ???? 2015. CODEN JARDCR. ISSN 2155-9546. URL https://

www.walshmedicalmedia.com/open-access/a-study-on-the-growth-
of-juveniles-of-tiger-prawn-penaeus-monodon-fabricius-under-
different-photoperiods-29690.html.

Adelodun:2015:PYA

[296] O. B. Adelodun. Participation of youth in aquaculture. Journal of Aquaculture Research & Development, 6(12):??, ???? 2015. CODEN JARDCR.

ISSN 2155-9546. URL https://www.walshmedicalmedia.com/open-
access/participation-of-youth-in-aquaculture-29691.html.

Asefa:2016:LST

[297] Wondimagegne Asefa and Tarekegn Beranu. Levels of some trace metals in fishes tissues, water and sediment at Tendaho Water Reservoir, Afar Region, Ethiopia. Journal of Aquaculture Research & Development, 7(1): ??, ???? 2016. CODEN JARDCR. ISSN 2155-9546. URL https:/

/www.walshmedicalmedia.com/open-access/levels-of-some-trace-
metals-in-fishes-tissues-water-and-sediment-at-tendaho-water-
reservoir-afar-region-ethiopia-29768.html.

Mzengereza:2016:EDS

[298] K. Mzengereza and J. Kang’ombe. Effect of dietary salt (sodium chloride) supplementation on growth, survival and feed utilization of Oreochromis shiranus (Trewavas, 1941). Journal of Aquaculture Research & Development, 7(1):??, ???? 2016. CODEN JARDCR. ISSN 2155-9546.

URL https://www.walshmedicalmedia.com/open-access/effect-of-
dietary-salt-sodium-chloride-supplementation-on-growth-survival-
and-feed-utilization-of-emoreochromis-shiranus-29769.html.

Yones:2016:EFM

[299] A. M. M. Yones and A. A. Metwalli. Effects of fish meal substitution with poultry by-product meal on growth performance, nutrients utilization and blood contents of juvenile Nile tilapia (Oreochromis niloticus). Journal of Aquaculture Research & Development, 7(1):??, ???? 2016. CODEN JARDCR. ISSN 2155-9546. URL https:/

/www.walshmedicalmedia.com/open-access/effects-of-fish-meal-
substitution-with-poultry-byproduct-meal-on-growth-performance-
nutrients-utilization-and-blood-cont-29770.html.

Ambas:2016:SIM

[300] I. Ambas, R. Fotedar, and N. Buller. Survival and immunity of marron Cherax cainii (Austin, 2002) fed Bacillus mycoides supplemented diet under simulated transport. Journal of Aquaculture Research & Development, 7(1):??, ???? 2016. CODEN JARDCR. ISSN 2155-9546. URL https://

www.walshmedicalmedia.com/open-access/survival-and-immunity-
of-marron-emcherax-cainiiem-austin-2002-fed-embacillus-mycoidesem-
supplemented-diet-under-simulate-29771.html.

Syandri:2016:SSF

[301] H. Syandri, Elfiondri, Junaidi, and Azrita. Social status of the fish-farmers of floating-net-cages in Lake Maninjau, Indonesia. Journal of Aquaculture Research & Development, 7(1):??, ???? 2016. CODEN JARDCR.

ISSN 2155-9546. URL https://www.walshmedicalmedia.com/open-

access/social-status-of-the-fishfarmers-of-floatingnetcages-
in-lake-maninjau-indonesia-29772.html.

Qambrani:2016:RBG

[302] G. R. Qambrani, A. N. Soomro, Z. A. Palh, W. A. Baloch, S. Tabasum, K. H. Lashari, and M. A. Qureshi. Reproductive biology of Glossogobius giuris (Hamilton), in Manchar Lake Sindh, Pakistan. Journal of Aquaculture Research & Development, 7(1):??, ???? 2016. CODEN JARDCR. ISSN 2155-9546. URL https:/

/www.walshmedicalmedia.com/open-access/reproductive-biology-
of-emglossogobius-giurisem-hamilton-in-manchar-lake-sindh-pakistan-
29756.html.

Andrewartha:2016:ASS

[303] S. J. Andrewartha, N. G. Elliott, J. W. McCulloch, and P. B. Frappell. Aquaculture sentinels: Smart-farming with biosensor equipped stock. Journal of Aquaculture Research & Development, 7(1):??, ???? 2016. CODEN JARDCR. ISSN 2155-9546. URL https://

www.walshmedicalmedia.com/open-access/aquaculture-sentinels-
smartfarming-with-biosensor-equipped-stock-29773.html.

Humphries:2016:PGM

[304] Fran Humphries. Patenting genetic material in aquaculture: a red herring or an emerging issue to tackle? Journal of Aquaculture Research & Development, 7(1):??, ???? 2016. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/patenting-genetic-material-in-aquaculture-a-
red-herring-or-an-emerging-issue-to-tackle-29774.html.

Mussabekov:2016:CAH

[305] A. T. Mussabekov, S. N. Borovikov, Zh. A. Suranshiyev, and A. S. Shamshidin. Comparative analysis of Holstein, Black-Motley, Angler, Simmental bulls semen. Journal of Aquaculture Research & Development, 7(2):??, ???? 2016. CODEN JARDCR. ISSN 2155-9546. URL https:/

/www.walshmedicalmedia.com/open-access/comparative-analysis-
of-holstein-blackmotley-angler-simmental-bulls-semen-29933.html.

Mzengereza:2016:AND

[306] K. Mzengereza, W. Singini, O. V. Msiska, F. Kapute, J. Kang’ombe, and A. Kamangira. Apparent nutrient digestibility of plant based diets by Tilapia rendalli (Boulenger, 1896). Journal of Aquaculture Research & Development, 7(2):??, ???? 2016. CODEN JARDCR.

ISSN 2155-9546. URL https://www.walshmedicalmedia.com/open-
access/apparent-nutrient-digestibility-of-plant-based-diets-
by-emtilapia-rendalliem-boulenger-1896-29934.html.

Sadeq:2016:CFM

[307] Qays Mohammed Sadeq, Swapan Kumar Bhattacharya, and Wan Ismail. Conceptual fracture modelling for carbonate reservoir in Bai Hassan Oil Field Northern Iraq. Journal of Aquaculture Research & Development, 7 (2):??, ???? 2016. CODEN JARDCR. ISSN 2155-9546. URL https://

www.walshmedicalmedia.com/open-access/emconceptual-fracture-
modelling-for-carbonate-reservoir-in-bai-hassan-oil-field-northern-
iraqem-30053.html.

Agbabiaka:2016:CSN

[308] L. A. Agbabiaka, O. A. Kuforiji, and C. C. Egobuike. Comparative studies on the nutrients, sensory and storage qualities of moon-fish (Citharinus citharus Geoffery Saint-Hilaire 1809) pre-treated with extracts from two spices. Journal of Aquaculture Research & Development, 7(2):??, ???? 2016. CODEN JARDCR. ISSN 2155-9546. URL https:

//www.walshmedicalmedia.com/open-access/comparative-studies-
on-the-nutrients-sensory-and-storage-qualities-of-moonfish-emcitharinus-
citharusem-geoffery-sainthil-30054.html.

Agbabiaka:2016:SME

[309] L. A. Agbabiaka, O. A. Kuforiji, and O. E. Ndumnigwe. Storage and microbial evaluation of black pepper pre-treated oven-dried moon fish (Citharinus citharus Geoffery Saint-Hilaire 1809). Journal of Aquaculture Research & Development, 7(2):??, ???? 2016. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/emstorage-and-microbial-evaluation-of-black-
pepper-pretreated-oven-dried-moon-fish-citharinus-citharus-geoffery-
sainthil-30055.html.

Gupta:2016:PPH

[310] Sandipan Gupta. Pangasius pangasius (Hamilton, 1822), a threatened fish of Indian subcontinent. Journal of Aquaculture Research & Development, 7(2):??, ???? 2016. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
empangasius-pangasiusem-hamilton-1822-a-threatened-fish-of-indian-
subcontinent-30056.html.

Pai:2016:DCE

[311] I. K. Pai, Maryem Shaikh Altaf, and K. N. Mohanta. Development of cost effective nutritionally balanced food for freshwater ornamental fish black molly (Poecilia latipinna). Journal of Aquaculture Research & Development, 7(2):??, ???? 2016. CODEN JARDCR.

ISSN 2155-9546. URL https://www.walshmedicalmedia.com/open-
access/development-of-cost-effective-nutritionally-balanced-
food-for-freshwater-ornamental-fish-black-molly-empoecilia-latipinn-
30057.html.

Dutta:2016:RAS

[312] S. P. S. Dutta. Record of an abnormal Schizothoraichthys esocinus (Heckel) from the Himalayan River Chenab Draining Jammu Region of the J&K State, India. Journal of Aquaculture Research & Development, 7(2):??, ???? 2016. CODEN JARDCR. ISSN 2155-9546.

URL https://www.walshmedicalmedia.com/open-access/record-of-
an-abnormal-emschizothoraichthys-esocinusem-heckel-from-the-
himalayan-river-chenab-draining-jammu-region-of-th-30059.html.

Akpoilih:2016:ODC

[313] B. U. Akpoilih, E. K. Ajani, and B. O. Omitoyin. Optimum dietary Ca/P ratio and phytase for growth and bone mineralization in juvenile Clarias gariepinus fed soya bean-based diet. Journal of Aquaculture Research & Development, 7(2):??, ???? 2016. CODEN JARDCR.

ISSN 2155-9546. URL https://www.walshmedicalmedia.com/open-
access/optimum-dietary-cap-ratio-and-phytase-for-growth-and-
bone-mineralization-in-juvenile-emclarias-gariepinusem-fed-soya-
bea-30060.html.

Bogdan:2016:TSS

[314] Ana-Maria Bogdan and Suren Kulshreshtha. Temporal and spatial scale community level water use in Saskatchewan. Journal of Aquaculture Research & Development, 7(2):??, ???? 2016. CODEN JARDCR.

ISSN 2155-9546. URL https://www.walshmedicalmedia.com/open-
access/temporal-and-spatial-scale-community-level-water-use-
in-saskatchewan-30061.html.

Salencia:2016:BAS

[315] H. R. Salência, J. L. P. Mouriño, G. S. Ferreira, R. F. Arantes, M. Ubert, M. Ubert, K. R. Lapa, and W. Q. Seiffert. A bioaugmentation agent in super intensive marine shrimp farming system with zero water exchange. Journal of Aquaculture Research & Development, 7(2):??, ???? 2016. CODEN JARDCR. ISSN 2155-

9546. URL https://www.walshmedicalmedia.com/open-access/a-
bioaugmentation-agent-in-super-intensive-marine-shrimp-farming-
system-with-zero-water-exchange-30062.html.

Ogunlela:2016:DPE

[316] A. O. Ogunlela and A. A. Adebayo. Development and performance evaluation of an automatic fish feeder. Journal of Aquaculture Research & Development, 7(2):??, ???? 2016. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
development-and-performance-evaluation-of-an-automatic-fish-
feeder-30063.html.

Yardimci:2016:ACT

[317] R. E. Yardimci and Gülsen Timur. Antigenic characterisation of Tenacibaculum maritimum isolates from sea bass (Dicentrarchus labrax, L.) farmed on the Aegean Sea Coasts of Turkey. Journal of Aquaculture Research & Development, 7(2):??, ???? 2016. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
antigenic-characterisation-of-emtenacibaculum-maritimumem-isolates-
from-sea-bass-emdicentrarchus-labraxem-l-farmed-on-th-30064.
html.

Kusuma:2016:CTI

[318] H. S. Kusuma, A. F. P. Putra, and M. Mahfud. Comparison of two isolation methods for essential oils from orange peel (Citrus auranticum L.) as a growth promoter for fish: Microwave steam distillation and conventional steam distillation. Journal of Aquaculture Research & Development, 7(2):??, ???? 2016. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
comparison-of-two-isolation-methods-for-essential-oils-from-
orange-peel-emcitrus-auranticumem-l-as-a-growth-promoter-for-
30065.html.

Valdez:2016:ADF

[319] A. S. M. Valdez and T. R. Castillo. Abundance and distribution of freshwater eels in Pangi River, Maitum, Sarangani Province. Journal of Aquaculture Research & Development, 7(2):??, ???? 2016. CODEN JARDCR.

ISSN 2155-9546. URL https://www.walshmedicalmedia.com/open-
access/abundance-and-distribution-of-freshwater-eels-in-pangi-
river-maitum-sarangani-province-30066.html.

Siddik:2016:SIC

[320] M. Siddik, M. Chaklader, M. Hanif, M. Islam, M. Sharker, and M. Rahman. Stock identification of critically endangered olive barb, Puntius sarana (Hamilton, 1822) with emphasis on management implications. Journal of Aquaculture Research & Development, 7(2):??, ???? 2016. CODEN JARDCR. ISSN 2155-9546. URL https:/

/www.walshmedicalmedia.com/open-access/stock-identification-
of-critically-endangered-olive-barb-empuntius-saranaem-hamilton-
1822-with-emphasis-on-management-im-30074.html.

Nguyen:2016:ABP

[321] Nguyen, Thi Thanh Thuy, and Quang Thanh. An analysis of brand personality on brand loyalty in frozen seafood supermarkets in Ho Chi Minh City. Journal of Aquaculture Research & Development, 7(3): ??, ???? 2016. CODEN JARDCR. ISSN 2155-9546. URL https:/

/www.walshmedicalmedia.com/open-access/an-analysis-of-brand-
personality-on-brand-loyalty-in-frozen-seafood-supermarkets-
in-ho-chi-minh-city-30117.html.

Oramary:2016:FCC

[322] Shokri Omar Mustafa Oramary, Dasthal Mamy Ibraheem Koramarky, Sarmad Ali Salih, and Ayiad Ahmed Mustafa. Feeding common carp fish (Cyprinus carpio) on natural foods (algae, phytoplankton, zooplankton and others) on Tigris River in Mosul Dam/Duhok, Kurdistan Region of Iraq. Journal of Aquaculture Research & Development, 7(3): ??, ???? 2016. CODEN JARDCR. ISSN 2155-9546. URL https:

//www.walshmedicalmedia.com/open-access/feeding-common-carp-
fish-emcyprinus-carpioem-on-natural-foods-algaephytoplankton-
zooplankton-and-others-on-tigris-river--30118.html.

Li:2016:EDC

[323] Tieliang Li, Chuan He, Zhihong Ma, Wei Xing, Na Jiang, Wentong Li, Xiangjun Sun, and Lin Luo. Effects of different carotenoids on pigmentation of blood parrot (Cichlasoma synspilum × Cichlasoma citrinellum). Journal of Aquaculture Research & Development, 7(3):??, ???? 2016. CODEN JARDCR. ISSN 2155-9546. URL https:/

/www.walshmedicalmedia.com/open-access/effects-of-different-
carotenoids-on-pigmentation-of-blood-parrot-emcichlasoma-synspilumem-
215-emcichlasoma-citrinellumem-30119.html.

Li:2016:EDD

[324] Weifeng Li, Xiaoyi Wu, Senda Lu, Shuntian Jiang, Yuan Luo, Mingjuan Wu, and Jun Wang. Effects of different dietary carbohydrate/lipid ratios on growth, feed utilization and body composition of early giant grouper Epinephelus lanceolatus juveniles. Journal of Aquaculture Research & Development, 7(3):??, ???? 2016. CODEN JARDCR. ISSN 2155-9546.

URL https://www.walshmedicalmedia.com/open-access/effects-
of-different-dietary-carbohydratelipid-ratios-on-growth-feed-
utilization-and-body-composition-of-early-giant-gro-30122.html.

Sayed:2016:OSI

[325] Alaa El-Din H. Sayed and Nasser S. Abou Khalil. Oxidative stress induction in monosex Nile tilapia (Oreochromis niloticus, Linnaeus, 1758): a field study on the side effects of methyltestosterone. Journal of Aquaculture Research & Development, 7(3):??, ???? 2016. CODEN JARDCR.

ISSN 2155-9546. URL https://www.walshmedicalmedia.com/open-
access/oxidative-stress-induction-in-monosex-nile-tilapia-emoreochromis-
niloticus-linnaeusem-1758-a-field-study-on-the-side-eff-30107.
html.

Chaklader:2016:MPA

[326] M. R. Chaklader, M. A. B. Siddik, Ashfaqun Nahar, M. A. Hanif, M. J. Alam, and Sultan Mahmud. Morphometric parameters and allometric growth in paradise threadfin Polynemus paradiseus (Linnaeus, 1758) from a coastal river of Bangladesh. Journal of Aquaculture Research & Development, 7(3):??, ???? 2016. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
morphometric-parameters-and-allometric-growth-in-paradise-threadfin-
empolynemus-paradiseusem-linnaeus-1758-from-a-coasta-30108.html.

Jo:2016:PAT

[327] Jihoon Jo, Chunsik Park, Munhwan Kim, and Chungoo Park. Phylogenetic analysis of the three color variations of the sea cucumber Apostichopus japonicus. Journal of Aquaculture Research & Development, 7(3): ??, ???? 2016. CODEN JARDCR. ISSN 2155-9546. URL https://

www.walshmedicalmedia.com/open-access/phylogenetic-analysis-
of-the-three-color-variations-of-the-sea-cucumber-emapostichopus-
japonicusem-30120.html.

Guy:2016:EMF

[328] Jeffrey A. Guy, D. A. Stephen, and Smith. Effects of meal frequency on growth performance and feed efficiency of two yearold mulloway (Argyrosomus japonicus; pisces: Sciaenidae) reared in tanks. Journal of Aquaculture Research & Development, 7(3):??, ???? 2016. CODEN JARDCR. ISSN 2155-9546. URL https://www.

walshmedicalmedia.com/open-access/effects-of-meal-frequency-
on-growth-performance-and-feed-efficiency-of-two-yearold-mulloway-
emargyrosomus-japonicusem-pi-30123.html.

Ali:2016:IFD

[329] A. Hyder Ali, A. Jawahar Ali, M. Saiyad Musthafa, M. S. Arun Kumar, Mohamed Saquib Naveed, Mehrajuddin War, and K. Altaff. Impact of formulated diets on the growth and survival of ornamental fish Pterophyllum scalare (angel fish). Journal of Aquaculture Research & Development, 7(4):??, ???? 2016. CODEN JARDCR. ISSN 2155-9546. URL https:/

/www.walshmedicalmedia.com/open-access/impact-of-formulated-
diets-on-the-growth-and-survival-of-ornamental-fish-empterophyllum-
scalareem-angel-fish-30160.html.

Hassaan:2016:EEO

[330] Mohamed S. Hassaan and Magdy A. Soltan. Evaluation of essential oil of fennel and garlic separately or combined with Bacillus licheniformis on the growth, feeding behaviour, hemato-biochemical indices of Oreochromis niloticus (L.) fry. Journal of Aquaculture Research & Development, 7(4):??, ???? 2016. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
evaluation-of-essential-oil-of-fennel-and-garlic-separately-
or-combined-with-embacillus-licheniformisem-on-the-growth-fe-
30161.html.

Banu:2016:GFP

[331] Rubia Banu and Annie Christianus. Giant freshwater prawn Macrobrachium rosenbergii farming: a review on its current status and prospective in Malaysia. Journal of Aquaculture Research & Development, 7(4):??, ???? 2016. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
giant-freshwater-prawn-emmacrobrachium-rosenbergiiem-farming-
a-review-on-its-current-status-and-prospective-in-malaysia-30162.
html.

Hosamani:2016:NIE

[332] Neelima Hosamani, Ramachandra Reddy Pamuru, and Sreenivasula Reddy Pamanji. Natural and induced (eyestalk ablation) molt cycle in freshwater rice field crab Oziothelphusa senex Senex. Journal of Aquaculture Research & Development, 7(4):??, ???? 2016. CODEN JARDCR.

ISSN 2155-9546. URL https://www.walshmedicalmedia.com/open-
access/natural-and-induced-eyestalk-ablation-molt-cycle-in-freshwater-
rice-field-crab-emoziothelphusa-senex-senexem-30175.html.

Manaf:2016:EVT

[333] Sharifah Raina Manaf, Hassan Mohd. Daud, Abdul Razak Alimon, Noordin Mohamed Mustapha, Ruhil Hayati Hamdan, Kumari Geetha Muniandy, Nora Faten Afifah Mohamed, Rashidah Razak, and Nur Hidayahanum Hamid. The effects of Vitex trifolia, Strobilanthes crispus and Aloe vera herbal-mixed dietary supplementation on growth performance and disease resistance in red hybrid tilapia (Oreochromis sp.). Journal of Aquaculture Research & Development, 7(4):??, ???? 2016. CODEN JARDCR.

ISSN 2155-9546. URL https://www.walshmedicalmedia.com/open-
access/the-effects-of-emvitex-trifolia-strobilanthes-crispusem-
and-emaloe-veraem-herbalmixed-dietary-supplementation-on-growth-
-30185.html.

Beruat:2016:SSC

[334] Agmy Beruat, A. N. Bambang, and Ambaryanto. Status of seagrass community in coastal area in the Kei Besar District of North-East, South-East Maluku Regency. Journal of Aquaculture Research & Development, 7(5):??, ???? 2016. CODEN JARDCR. ISSN 2155-9546.

URL https://www.walshmedicalmedia.com/open-access/status-of-
seagrass-community-in-coastal-area-in-the-kei-besar-district-
of-northeast-southeast-maluku-regency-30225.html.

El-Gawad:2016:EAA

[335] Eman A. Abd El-Gawad, Ashraf M. Abd El-latif, and Ramy M. Shourbela. Enhancement of antioxidant activity, non-specific immunity and growth performance of Nile tilapia, Oreochromis niloticus by dietary fructooligosaccharide. Journal of Aquaculture Research & Development, 7(5):??, ???? 2016. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
enhancement-of-antioxidant-activity-nonspecific-immunity-and-
growth-performance-of-emnile-tilapiaem-emoreochromis-niloti-
30231.html.

Huicab-Pech:2016:CSB

[336] Z. G. Huicab-Pech, C. Landeros-Sánchez, M. R. Castañeda-Chávez, F. Lango-Reynoso, C. J. López-Collado, and D. E. Platas Rosado. Current state of bacteria pathogenicity and their relationship with host and environment in tilapia Oreochromis niloticus. Journal of Aquaculture Research & Development, 7(5):??, ???? 2016. CODEN JARDCR. ISSN 2155-9546.

URL https://www.walshmedicalmedia.com/open-access/current-
state-of-bacteria-pathogenicity-and-their-relationship-with-
host-andenvironment-in-tilapia-oreochromis-niloticus-2155-9546-
1000428.pdf.

Gupta:2016:FFH

[337] Sandipan Gupta and Samir Banerjee. Food, feeding habit and reproductive biology of tire-track spiny eel (Mastacembelus armatus): a review. Journal of Aquaculture Research & Development, 7(5):??, ???? 2016. CODEN JARDCR. ISSN 2155-9546. URL https://www.

walshmedicalmedia.com/open-access/food-feeding-habit-and-reproductive-
biology-of-tiretrack-spiny-eel-emmastacembelus-armatusem-a-review-
30254.html.

Mairapetyan:2016:PBI

[338] Stepan Mairapetyan, Vardan Mamikonyan, Juletta Alexanyan, Anahit Tovmasyan, Mahsa Daryadar, and Bella Stepanian. Productivity, biochemical indices and antioxidant activity of peppermint (Mentha piperita L.) and basil (Ocimum basilicum L.) in conditions of hydroponics. Journal of Aquaculture Research & Development, 7(6):??, ???? 2016. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/productivity-biochemical-indices-and-antioxidant-
activity-of-peppermintmentha-piperita-l-and-basil-ocimum-basilicum-
l-in-condition-2155-9546-1000430.pdf.

Mapenzi:2016:RSG

[339] Levinus Leonard Mapenzi and Aviti John Mmochi. Role of salinity on growth performance of Oreochromis niloticus [female-sign] and Oreochromis urolepis urolepis [male-sign] hybrids. Journal of Aquaculture Research & Development, 7(6):1–5, ???? 2016. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
role-of-salinity-on-growth-performance-of-emoreochromis-niloticusem9792-
and-emoreochromis-urolepisem-emurolepisem9794-hy-30288.html.

Kefi:2016:SVT

[340] Ferdaous Jaafar Kefi, Anwar Mleiki, Jihen Maâtoug Béjaoui, and Najoua Trigui El Menif. Seasonal variations of trace metal concentrations in the soft tissue of Lithophaga lithophaga collected from the Bizerte Bay (Northern Tunisia, Mediterranean Sea). Journal of Aquaculture Research & Development, 7(6):??, ???? 2016. CODEN JARDCR.

ISSN 2155-9546. URL https://www.walshmedicalmedia.com/open-
access/seasonal-variations-of-trace-metal-concentrations-in-
the-soft-tissue-of-emlithophaga-lithophagaem-collected-from-
the-biz-30320.html.

Lenzi:2016:ITA

[341] Mauro Lenzi. Insights on the trends in aquaculture research. Journal of Aquaculture Research & Development, 7(7):??, ???? 2016. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/insights-on-the-trends-in-aquaculture-research-
2155-9546-1000e111.pdf.

Lenzi:2016:TON

[342] Mauro Lenzi. Tilapia Oreochromis niloticus — a choice of aqua farming. Journal of Aquaculture Research & Development, 7(7):??, ???? 2016. CODEN JARDCR. ISSN 2155-9546. URL https:

//www.walshmedicalmedia.com/open-access/tilapia-oreochromis-
niloticus--a-choice-of-aqua-farming-2155-9546-1000e110.pdf.

Abu-Elala:2016:EDS

[343] Nermeen M. Abu-Elala, Mona K. Galal, Reham M. Abd-Elsalam, Omnia Mohey-Elsaeed, and Naela M. Ragaa. Effects of dietary supplementation of Spirulina platensis and garlic on the growth performance and expression levels of immune-related genes in Nile tilapia (Oreochromis niloticus). Journal of Aquaculture Research & Development, 7(7):??, ???? 2016. CODEN JARDCR. ISSN 2155-9546.

URL https://www.walshmedicalmedia.com/open-access/effects-
of-dietary-supplementation-of-emspirulina-platensisem-and-garlic-
on-the-growth-performance-and-expression-levels-30401.html.

Bejaoui:2016:FAV

[344] Jihen Maâtoug Béjaoui, Ferdaous Jaafar Kefi, Anwar Mleiki, and Najoua Trigui El Menif. Foot abnormalities in Venericardia antiquata and Venus verrucosa from the Bizerte Lagoon Complex (Northern Tunisia): Hydrodynamics and sediment texture inductions. Journal of Aquaculture Research & Development, 7(7):??, ???? 2016. CODEN JARDCR. ISSN 2155-9546. URL https://

www.walshmedicalmedia.com/open-access/foot-abnormalities-in-
emvenericardia-antiquataem-and-emvenus-verrucosaem-from-the-
bizerte-lagoon-complex-northern-tunisi-30404.html.

Gorji:2016:GPE

[345] Abdolvahab Ebrahimpour Gorji, Hossein Rahmani, and Ghodrat Rahimi Miyanji. Growth parameters evaluation and identification of growth hormone receptor gene polymorphisms in various strains of rainbow trout Oncorhynchus mykiss. Journal of Aquaculture Research & Development, 7(7):??, ???? 2016. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
growth-parameters-evaluation-and-identification-of-growth-hormone-
receptor-gene-polymorphisms-in-various-strains-of-rain-30405.
html.

Samanta:2016:HUA

[346] Palas Samanta, Sandipan Pal, Aloke Kumar Mukherjee, Tarakeshwar Senapati, and Apurba Ratan Ghosh. Histopathological and ultrastructural alterations in Anabas testudineus exposed to glyphosate-based herbicide, Excel Mera 71 under field and laboratory conditions. Journal of Aquaculture Research & Development, 7(7):??, ???? 2016. CODEN JARDCR.

ISSN 2155-9546. URL https://www.walshmedicalmedia.com/open-
access/histopathological-and-ultrastructural-alterations-in-
anabas-testudineus-exposed-to-glyphosatebased-herbicide-excel-
mera--30413.html.

Chow:2016:EDC

[347] Edwin Pei Yong Chow, Kah Heng Liong, and Elke Schoeters. The effect of dietary carotenoids of different forms: Microemulsified and nonmicroemulsified on the growth performance, pigmentation and hematological parameters in hybrid catfish (Clarias macrocephalus × Clarias gariepinus). Journal of Aquaculture Research & Development, 7(7):1– 6, ???? 2016. CODEN JARDCR. ISSN 2155-9546. URL https://

www.walshmedicalmedia.com/open-access/the-effect-of-dietary-
carotenoids-of-different-forms-microemulsified-and-nonmicroemulsified-
on-the-growth-performance-pi-30416.html.

Kumary:2016:CEB

[348] KS Anila Kumary. Comparative ecology of backwater and mangrove environments of Kayamkulam Lake, Kerala. Journal of Aquaculture Research & Development, 7(8):??, ???? 2016. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
comparative-ecology-of-backwater-and-mangrove-environments-of-
kayamkulam-lake-kerala-30428.html.

Nakkina:2016:SGR

[349] Mounika Nakkina. Study of growth rate in Nile tilapia (Oreochromis niloticus). Journal of Aquaculture Research & Development, 7(8):??, ???? 2016. CODEN JARDCR. ISSN 2155-9546. URL https:/

/www.walshmedicalmedia.com/open-access/study-of-growth-rate-
in-nile-tilapia-emoreochromis-niloticusem-30440.html.

Mosha:2016:EOI

[350] Sebastian S. Mosha, Jeremiah Kang’ombe, Wilson Jere, and Nazael Madalla. Effect of organic and inorganic fertilizers on natural food composition and performance of African catfish (Clarias gariepinus) fry produced under artificial propagation. Journal of Aquaculture Research & Development, 7(8):??, ???? 2016. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
effect-of-organic-and-inorganic-fertilizers-on-natural-food-
composition-and-performance-of-african-catfish-clarias-garie-
30454.html.

Tsygankov:2016:KIP

[351] Vasiliy Yu Tsygankov. The Kuril Islands as a potential region for aquaculture: Organochlorine pesticides in pink and chum salmon. Journal of Aquaculture Research & Development, 7(8):??, ???? 2016. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/the-kuril-islands-as-a-potential-region-for-
aquaculture-organochlorine-pesticides-in-pink-and-chum-salmon-
30453.html.

Adebayo:2016:TRY

[352] IA Adebayo and BJ Akin-Obasola. Toxicity of rotenone-yielding plant extracts on fish fry and selected predatory aquatic organisms. Journal of Aquaculture Research & Development, 7(8):??, ???? 2016. CODEN JARDCR. ISSN 2155-9546. URL https:/

/www.walshmedicalmedia.com/open-access/toxicity-of-rotenone-
yielding-plant-extracts-on-fish-fry-and-selected-predatory-aquatic-
organisms-30486.html.

Anjugam:2016:ACP

[353] Mahalingam Anjugam, Arokiadhas Iswarya, Thiruselvam Indumathi, Baskaralingam Vaseeharan, Raman Pachaiappan, Narayanan Gopi, and Palaniyandi Velusamy. Antibiofilm competency of Portunus pelagicus haemolymph and identification of its bioactive compounds. Journal of Aquaculture Research & Development, 7(8):??, ???? 2016. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/antibiofilm-competency-of-portunus-pelagicus-
haemolymph-and-identification-of-its-bioactive-compounds-30490.
html.

Patra:2016:RDF

[354] Avijit Patra, Sudeshna Sarker, Sayani Banerjee, Harresh Adikesavalu, Debadyuti Biswas, and Thangapalam Jawahar Abraham. Rapid detection of Flavobacterium columnare infection in fish by speciesspecific polymerase chain reaction. Journal of Aquaculture Research & Development, 7(9):??, ???? 2016. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
rapid-detection-of-flavobacterium-columnare-infection-in-fish-
by-speciesspecific-polymerase-chain-reaction-30495.html.

Tohamy:2016:PIG

[355] Hossam G. Tohamy and Ramy M. Shourbela. Pathological investigations on galilee tilapia (Sarotherodon galilaeus) following chronic exposure to cadmium chloride. Journal of Aquaculture Research & Development, 7(9):??, ???? 2016. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
pathological-investigations-on-galilee-tilapia-sarotherodon-
galilaeus-following-chronic-exposure-to-cadmium-chloride-30496.
html.

Zhang:2016:PTE

[356] J. Y. Zhang, X. H. Liu, Y. L. Zhao, F. L. Wei, and C. Z. Li. Pilot test of elemental fingerprint in otolith core as a natural biological tag to discriminate among Gymnocypris przewalskii stocks in Lake Qinghai in Qinghai-Tibet Plateau, China. Journal of Aquaculture Research & Development, 7(9):??, ???? 2016. CODEN JARDCR. ISSN 2155-9546. URL

https://www.walshmedicalmedia.com/open-access/pilot-test-of-
elemental-fingerprint-in-otolith-core-as-a-natural-biological-
tag-to-discriminate-among-gymnocypris-przewa-30497.html.

Olivares:2016:NES

[357] Pamela Olivares, Rodrigo Sanchez, Erico Carmona, Allisson Astuya, Hector Herrera, and Jorge Parodi. Nucleotides and effect over starving condition on fish SHK-1 cells model. Journal of Aquaculture Research & Development, 7(9):??, ???? 2016. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
nucleotides-and-effect-over-starving-condition-on-fish-shk1-
cells-model-30500.html.

Hasan:2016:ICC

[358] Mehedi Mahmudul Hasan and Shuva Bhowmik. Impact of climate change on the socio-economics of aquaculture in the District of Noakhali, Bangladesh. Journal of Aquaculture Research & Development, 7(9):??, ???? 2016. CODEN JARDCR. ISSN 2155-9546.

URL https://www.walshmedicalmedia.com/open-access/impact-of-
climate-change-on-the-socioeconomics-of-aquaculture-in-the-district-
of-noakhali-bangladesh-30503.html.

Chen:2016:NVA

[359] Xu Chen, Qiang-Qiang Liu, Jia-Jun Xie, Yun-Qiang Zhang, and Jin Niu. Nutritional value and apparent digestibility for dry matter, protein, energy and essential amino acid in ten selected feedstuffs for juvenile Penaeus monodon. Journal of Aquaculture Research & Development, 7(10):??, ???? 2016. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
nutritional-value-and-apparent-digestibility-for-dry-matter-
protein-energy-and-essential-amino-acid-in-ten-selected-feed-
30581.html.

Mortensen:2016:ETE

[360] Atle Mortensen, Oyvind J. Hansen, and Velmurugu Puvanendran. Evaluation of three external marking methods of farmed Atlantic salmon for the future use of differentiating it from wild Atlantic salmon. Journal of Aquaculture Research & Development, 7(10):??, ???? 2016. CODEN JARDCR. ISSN 2155-9546. URL https:

//www.walshmedicalmedia.com/open-access/evaluation-of-three-
external-marking-methods-of-farmed-atlantic-salmon-for-the-future-
use-of-differentiating-it-from-wil-30582.html.

Soultani:2016:AFL

[361] Georgia Soultani, Irini F. Strati, Panagiotis Zoumpoulakis, Sofia Miniadis-Meimaroglou, and Vassilia J. Sinanoglou. Assessment of functional lipid constituents of red (Aristaeomorpha foliacea) and pink (Parapenaeus longirostris) shrimps. Journal of Aquaculture Research & Development, 7(10):??, ???? 2016. CODEN JARDCR. ISSN 2155-9546. URL

https://www.walshmedicalmedia.com/open-access/assessment-of-
functional-lipid-constituents-of-red-aristaeomorpha-foliacea-
and-pink-parapenaeus-longirostris-shrimps-30583.html.

Sanchez:2016:FFS

[362] Fabiola Chong Sánchez, Martha Enríquez Díaz, Imelda Martínez Morales, and Dalila Aldana Aranda. Formulated feed for Strombus pugilis (Mollusca, Gastropoda) allowed effective gonad maturity. Journal of Aquaculture Research & Development, 7(10):??, ???? 2016. CODEN JARDCR. ISSN 2155-9546. URL https:

//www.walshmedicalmedia.com/open-access/formulated-feed-for-
strombus-pugilis-mollusca-gastropoda-allowed-effective-gonad-
maturity-30589.html.

Mosbahi:2016:IEC

[363] Nawfel Mosbahi, Jean-Philippe Pezy, Jean-Claude Dauvin, and Lassad Neifar. Immediate effect of clam harvesting on intertidal benthic communities in the mudflat zones of Kneiss Islands (Central Mediterranean Sea). Journal of Aquaculture Research & Development, 7(11):??, ???? 2016. CODEN JARDCR. ISSN 2155-9546.

URL https://www.walshmedicalmedia.com/open-access/immediate-
effect-of-clam-harvesting-on-intertidal-benthic-communities-
in-the-mudflat-zones-of-kneiss-islands-central-med-30610.html.

Saei:2016:EDL

[364] Mohsen Mohamadi Saei, Kiarash Beiranvand, Hadis Mansouri Taee, and Hamed Nekoubin. Effects of different levels of BioAcid ultra on growth performance, survival, hematologichal and biochemical parameters of fingerlings rainbow trout (Oncorhynchus mykiss). Journal of Aquaculture Research & Development, 7(11):??, ???? 2016. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
effects-of-different-levels-of-bioacid-ultra-on-growth-performance-
survival-hematologichal-and-biochemical-parameters-of-30632.
html.

Yang:2016:TCE

[365] Lin Yang, Ting-Ting Lin, Dong Zhang, and Xin Liu. Time course effect of low salinity on the plasma osmotic pressure, ion concentrations and Na^{+} /K^{+}-ATPase activity in the gill of juvenile lined seahorse, Hippocampus erectus. Journal of Aquaculture Research & Development, 7(11):??, ???? 2016. CODEN JARDCR. ISSN 2155-9546. URL https://www.

walshmedicalmedia.com/open-access/time-course-effect-of-low-
salinity-on-the-plasma-osmotic-pressure-ion-concentrations-and-
nakatpase-activity-in-the-gill--30633.html.

Ostrensky:2016:UCM

[366] Antonio Ostrensky, Giorgi Dal Pont, Gisela Geraldine Castilho Westphal, and Ana Silvia Pedrazzani. Use of clove, mint and camphor essential oils on confinement of clown anemonefish Amphiprion ocellaris (Cuvier 1830): Anesthetic effects and influence on water quality. Journal of Aquaculture Research & Development, 7(11):??, ???? 2016. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
use-of-clove-mint-and-camphor-essential-oils-on-confinement-
of-clown-anemonefish-amphiprion-ocellaris-cuvier-1830-anesth-
30634.html.

Adikesavalu:2016:PPP

[367] Harresh Adikesavalu, Pradipta Paul, Siddhartha N. Joardar, and Jawahar T. Abraham. Polypeptide profiling of pangas catfish (Pangasius pangasius) serum globulin protein fraction and development of anti-pangas serum globulin-HRPO immunoconjugate for rapid detection of bacterial infection. Journal of Aquaculture Research & Development, 7(11):??, ???? 2016. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
polypeptide-profiling-of-pangas-catfish-pangasius-pangasius-
serum-globulin-protein-fraction-and-development-of-antipanga-
30635.html.

Aly:2016:AAC

[368] Hadir A. Aly, Mohamed M. Abdel Rahim, Ayman M. Lotfy, Basem S. Abdelaty, and Ghada M. Sallam. The applicability of activated carbon, natural zeolites, and probiotics (EM(R)) and its effects on ammonia removal efficiency and fry performance of European seabass Dicentrarchus labrax. Journal of Aquaculture Research & Development, 7(11): ??, ???? 2016. CODEN JARDCR. ISSN 2155-9546. URL https:/

/www.walshmedicalmedia.com/open-access/the-applicability-of-
activated-carbon-natural-zeolites-and-probiotics-emsup174sup-
and-its-effects-on-ammonia-removal-eff-30638.html.

Lenzi:2016:ENJ

[369] Mauro Lenzi. Editor note: Journal of Aquaculture Research & Development vol 7 issue 12. Journal of Aquaculture Research & Development, 7 (12):??, ???? 2016. CODEN JARDCR. ISSN 2155-9546. URL https:

//www.walshmedicalmedia.com/open-access/editor-note-journal-
of-aquaculture-research--development-vol-7-issue12-2155-9546-
1000e112.pdf.

Akinwole:2016:GPA

[370] A. O. Akinwole, A. B. Dauda, and O. A. Ololade. Growth performance of African catfish (Clarias gariepinus) juveniles reared in wastewater treated with alum and Moringa oleifera seed. Journal of Aquaculture Research & Development, 7(12):??, ???? 2016. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
growth-performance-of-african-catfish-clarias-gariepinus-juveniles-
reared-in-wastewater-treated-with-alum-and-moringa-ol-30657.
html.

Siddik:2016:BLC

[371] M. A. B. Siddik, M. A. Islam, M. A. Hanif, M. R. Chaklader, and R. Kleindienst. Barramundi, Lates calcarifer (Bloch, 1790): a new dimension to the fish farming in coastal Bangladesh. Journal of Aquaculture Research & Development, 7(12):??, ???? 2016. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
barramundi-lates-calcarifer-bloch-1790-a-new-dimension-to-the-
fish-farming-in-coastal-bangladesh-30673.html.

Khallaf:2016:CSW

[372] Elsayed A. Khallaf, M. Galal, M. N. Mohammad Authman, and R. A. Zaid. A comparative study on the water quality of two Nilotic canals in the Delta of Egypt. Journal of Aquaculture Research & Development, 7 (12):??, ???? 2016. CODEN JARDCR. ISSN 2155-9546. URL https:

//www.walshmedicalmedia.com/open-access/a-comparative-study-
on-the-water-quality-of-two-nilotic-canals-in-the-delta-of-egypt-
30699.html.

Hou:2016:CPE

[373] Dongwei Hou, Shenzheng Zeng, Jian Liu, Muting Yan, Shaoping Weng, Jianguo He Jianguo He, and Zhijian Huang. Characterization of prokaryotic and eukaryotic microbial community in Pacific white shrimp ponds. Journal of Aquaculture Research & Development, 7(12): ??, ???? 2016. CODEN JARDCR. ISSN 2155-9546. URL https:

//www.walshmedicalmedia.com/open-access/characterization-of-
prokaryotic-and-eukaryotic-microbial-community-in-pacific-white-
shrimp-ponds-30702.html.

Barros:2016:VDI

[374] Inês Barros, Susana Mendes, Domitilia Rosa, Ricardo Serrão Santos, and Raul Bettencourt. Vibrio diabolicus immunomodulatory effects on Bathymodiolus azoricus during long-term acclimatization at atmospheric pressure. Journal of Aquaculture Research & Development, 7(12):??, ???? 2016. CODEN JARDCR. ISSN 2155-9546. URL https://www.

walshmedicalmedia.com/open-access/vibrio-diabolicus-immunomodulatory-
effects-on-bathymodiolus-azoricus-during-longterm-acclimatization-
at-atmospheric-pres-30718.html.

Alkobaby:2017:ATC

[375] Akram I. Alkobaby and Rasha K. Abd El-Wahed. The acute toxicity of copper to Nile tilapia (Oreochromis niloticus) fingerlings and its effects on gill and liver histology. Journal of Aquaculture Research & Development, 8(1):??, ???? 2017. CODEN JARDCR. ISSN 2155-9546. URL https://

www.walshmedicalmedia.com/open-access/the-acute-toxicity-of-
copper-to-nile-tilapia-oreochromis-niloticus-fingerlings-and-
its-effects-on-gill-and-liver-histolo-30739.html.

Banu:2017:RPM

[376] Husne Banu and Kurcheti Pani Prasad. Role of plasmids in microbiology. Journal of Aquaculture Research & Development, 8(1): ??, ???? 2017. CODEN JARDCR. ISSN 2155-9546. URL https:

//www.walshmedicalmedia.com/open-access/role-of-plasmids-in-
microbiology-30789.html.

Tripathi:2017:PAS

[377] Manoj Kumar Tripathi, Maheep Kumar, S. Deepali, Ravi Kumar Asthana, and Subhasha Nigam. Proteomic analysis of sensitive and resistant isolates of Escherichia coli in understanding target(s) of a cyanobacterial biomolecule hapalindole-T. Journal of Aquaculture Research & Development, 8(1):??, ???? 2017. CODEN JARDCR. ISSN 2155-9546.

URL https://www.walshmedicalmedia.com/open-access/proteomic-
analysis-of-sensitive-and-resistant-isolates-of-escherichia-
coli-in-understanding-targets-of-a-cyanobacterial--30799.html.

Lenzi:2017:ENJ

[378] Mauro Lenzi. Editor note: Journal of Aquaculture Research & Development volume 8, issue 2. Journal of Aquaculture Research & Development, 8(2):??, ???? 2017. CODEN JARDCR. ISSN 2155-9546. URL https:

//www.walshmedicalmedia.com/open-access/editor-note-journal-
of-aquaculture-research--development-volume-8-issue-2-2155-9546-
1000e113.pdf.

Omitoyin:2017:BTF

[379] Bamidele Oluwarotimi Omitoyin, Emmanuel Kolawole Ajani, Oluwabusayo Israel Okeleye, Benjamin Uzezi Akpoilih, and Adeniyi Adewale Ogunjobi. Biological treatments of fish farm effluent and its reuse in the culture of Nile tilapia (Oreochromis niloticus). Journal of Aquaculture Research & Development, 8(2):??, ???? 2017. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/biological-treatments-of-fish-farm-effluent-
and-its-reuse-in-the-culture-of-nile-tilapia-oreochromis-niloticus-
30870.html.

Nahar:2017:SSC

[380] Ashfaqun Nahar, Md Reaz Chaklader, Muhammad Abu Bakar Siddik, Ilham Ilham, Hung Duc Pham, and Sukham Munilkumar. Stock structure of the critically endangered Clupisoma garua (Hamilton, 1822): an investigation based on discriminant analysis approach. Journal of Aquaculture Research & Development, 8(2):??, ???? 2017. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
stock-structure-of-the-critically-endangered-clupisoma-garua-
hamilton-1822-an-investigation-based-on-discriminant-analys-
30871.html.

Rahman:2017:BBT

[381] Md. Anisur Rahman, Tayfa Ahmed, Md. Mehedi Hasan Pramanik, R. Flura, Md. Monjurul Hasan, Md. Golam Sajed Riar, Khandaker Rashidul Hasan, Masud Hossain Khan, and Yahia Mahmud. On-board breeding trial of hilsa (Tenualosa ilisha, Ham. 1822) and testing of larval rearing in Bangladesh. Journal of Aquaculture Research & Development, 8(2):??, ???? 2017. CODEN JARDCR.

ISSN 2155-9546. URL https://www.walshmedicalmedia.com/open-
access/onboard-breeding-trial-of-hilsa-tenualosa-ilisha-ham-
1822-and-testing-of-larval-rearing-in-bangladesh-30879.html.

Al-Taee:2017:VSI

[382] Asaad M. R. Al-Taee, Najem R. Khamees, and Nadia A. H. Al-Shammari. Vibrio species isolated from farmed fish in Basra City in Iraq. Journal of Aquaculture Research & Development, 8(2):??, ???? 2017. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/vibrio-species-isolated-from-farmed-fish-in-
basra-city-in-iraq-30882.html.

El-Shafei:2017:ALS

[383] Hussien M. El-Shafei. Alterations in the leucocytes and serum biochemistry in grey mullet (Mugil cephalus L.) fingerlings exposed to sub lethal doses of lead for different exposure periods. Journal of Aquaculture Research & Development, 8(3):??, ???? 2017. CODEN JARDCR.

ISSN 2155-9546. URL https://www.walshmedicalmedia.com/open-
access/alterations-in-the-leucocytes-and-serum-biochemistry-
in-grey-mu-llet-mugil-cephalus-l-fingerlings-exposed-to-sub-
lethal--30919.html.

Dinesh:2017:DTN

[384] R. Dinesh, Chandra Prakash, N. K. Chadha, Nalini Poojary, and Sherry Abraham. Does tobacco (Nicotiana tabacum) leaf dust save the life of rohu (Labeo rohita) fingerlings during transport? Journal of Aquaculture Research & Development, 8(3):??, ???? 2017. CODEN JARDCR. ISSN 2155-

9546. URL https://www.walshmedicalmedia.com/open-access/does-
tobacco-nicotiana-tabacum-leaf-dust-save-the-life-of-rohu-labeo-
rohita-fingerlings-during-transport-30920.html.

Franklyn:2017:TAR

[385] Awonfor Franklyn, Yajun Wang, Na Yu, Jianbo Wang, Qijun Le, Xiaohuan Cao, and Huakun Zheng. Transcriptome analysis reveals lamino acids as olfactory stimulant in the large yellow croaker (Larimichthys crocea). Journal of Aquaculture Research & Development, 8 (3):??, ???? 2017. CODEN JARDCR. ISSN 2155-9546. URL

https://www.walshmedicalmedia.com/open-access/transcriptome-
analysis-reveals-lamino-acids-as-olfactory-stimulant-in-the-
large-yellow-croaker-larimichthys-crocea-30927.html.

Sallam:2017:RTB

[386] Ghada R. Sallam, Walied A. Fayed, Mohamed A. El-Absawy, Hadir A. Ali, and Zeinab A. El-Greisy. Red tilapia broodstocks and larval production under different water salinities without acclimation. Journal of Aquaculture Research & Development, 8(3):??, ???? 2017. CODEN JARDCR.

ISSN 2155-9546. URL https://www.walshmedicalmedia.com/open-
access/red-tilapia-broodstocks-and-larval-production-under-different-
water-salinities-without-acclimation-30928.html.

ChavesFabrini:2017:SPD

[387] Bruno Chaves Fabrini, Weslley Fernandes Braga, Estefânia Souza Andrade, Daniela Aparecida de Jesus Paula, Renan Rosa Paulino, Adriano Carvalho Costa, Luciano Vilela Paiva, Fabrício Lelis da Silva, and Luis David Solis Murgas. Sulfated polysaccharides in diets for Nile tilapia (Oreochromis niloticus) in the initial growth phase. Journal of Aquaculture Research & Development, 8(4):1–6, ???? 2017. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
sulfated-polysaccharides-in-diets-for-nile-tilapia-oreochromis-
niloticus-in-the-initial-growth-phase-30962.html.

A:2017:NAI

[388] Ignacio Agudo-Padrón A. New additions to inventory of marine mollusc species from Santa Catarina State/SC, Central Southern Brazil. Journal of Aquaculture Research & Development, 8(4): ??, ???? 2017. CODEN JARDCR. ISSN 2155-9546. URL

https://www.walshmedicalmedia.com/open-access/new-additions-
to-inventory-of-marine-mollusc-species-from-santa-catarina-statesc-
central-southern-brazil-30972.html.

Alemayehu:2017:EFF

[389] Tewodros Abate Alemayehu and Ababe Getahun. Effect of feeding frequency on growth performance and survival of Nile tilapia (Oreochromis niloticus L. 1758) in a cage culture system in Lake Hora-Arsedi, Ethiopia. Journal of Aquaculture Research & Development, 8(4):??, ???? 2017. CODEN JARDCR. ISSN 2155-9546.

URL https://www.walshmedicalmedia.com/open-access/effect-of-
feeding-frequency-on-growth-performance-and-survival-of-nile-
tilapia-oreochromis-niloticus-l-1758-in-a-cage-cu-30979.html.

Chin-Long:2017:SGP

[390] K. Y. Chin-Long and Gilles Le Moullac. Shell growth performance of hatchery produced Pinctada margaritifera: Family effect and relation with cultured pearl weight. Journal of Aquaculture Research & Development, 8(4):??, ???? 2017. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
shell-growth-performance-of-hatchery-produced-pinctada-margaritifera-
family-effect-and-relation-with-cultured-pearl-weig-30973.html.

Kitanovski:2017:ESL

[391] Vladimir Dimche Kitanovski, Dessislava Borislavova Vlahova-Vangelova, Stefan Georgiev Dragoev, Hristo Nikolov Nikolov, and Dessislav Kostadinov Balev. Extension the shelf-life of fresh golden rainbow trout via ultrafast air or cryogenic carbon dioxide super chilling. Journal of Aquaculture Research & Development, 8(4):??, ???? 2017. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
extension-the-shelflife-of-fresh-golden-rainbow-trout-via-ultrafast-
air-or-cryogenic-carbon-dioxide-super-chilling-30978.html.

Abdel-Aziz:2017:EUF

[392] Mohamed F. A. Abdel-Aziz and Mohammed A. Ragab. Effect of use fresh macro algae (seaweed) Ulva fasciata and Enteromorpha flaxusa with or without artificial feed on growth performance and feed utilization of rabbitfish (Siganus rivulatus) fry. Journal of Aquaculture Research & Development, 8(4):??, ???? 2017. CODEN JARDCR. ISSN 2155-9546. URL

https://www.walshmedicalmedia.com/open-access/effect-of-use-
fresh-macro-algae-seaweed-ulva-fasciata-and-enteromorpha-flaxusa-
with-or-without-artificial-feed-on-growth-30982.html.

Pramanik:2017:GNS

[393] Md. Mehedi Hasan Pramanik, Md. Anisur Rahman, Tayfa Ahmed, Flura, Md. Monjurul Hasan, Masud Hossain Khan, and Yahia Mahmud. Gill net selectivity of hilsa (Tenualosa ilisha) in the Meghna River Estuary of Bangladesh. Journal of Aquaculture Research & Development, 8(4): ??, ???? 2017. CODEN JARDCR. ISSN 2155-9546. URL https:/

/www.walshmedicalmedia.com/open-access/gill-net-selectivity-
of-hilsa-tenualosa-ilisha-in-the-meghna-river-estuary-of-bangladesh-
30983.html.

Thirumal:2017:HPM

[394] Yasodha Thirumal and Suresh Laavu. HPLC profile of medicinal plant extracts and its application in aquaculture. Journal of Aquaculture Research & Development, 8(5):??, ???? 2017. CODEN JARDCR. ISSN 2155-

9546. URL https://www.walshmedicalmedia.com/open-access/hplc-
profile-of-medicinal-plant-extracts-and-its-application-in-aquaculture-
31035.html.

Nyunja:2017:SSD

[395] Cynthia Nyunja, Joyce Maina, Joshua Amimo, Felix Kibegwa, David Harper, and Joseph Jung’a. Stock structure delineation of the African catfish (Clarius gariepinus) in selected populations in Kenya using mitochondrial DNA (Dloop) variability. Journal of Aquaculture Research & Development, 8(5):??, ???? 2017. CODEN JARDCR.

ISSN 2155-9546. URL https://www.walshmedicalmedia.com/open-
access/stock-structure-delineation-of-the-african-catfish-clarius-
gariepinus-in-selected-populations-in-kenya-using-mitochondri-
31038.html.

Blessy:2017:EAO

[396] G. Blessy, C. Ajan, T. Citarasu, and Michael Babu M. Effect of algal oil incorporated diet on growth biochemical and immunological response in ornamental fish Daniorerio rerio. Journal of Aquaculture Research & Development, 8(5):??, ???? 2017. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
effect-of-algal-oil-incorporated-diet-on-growth-biochemical-
and-immunological-response-in-ornamental-fish-daniorerio-rer-
31036.html.

El-Feky:2017:EZE

[397] Mohamed M. M. El-Feky. Effect of zooplankton and environmental parameters on African catfish Clarias gariepinus (Burchell, 1822) in Egypt. Journal of Aquaculture Research & Development, 8(5):??, ???? 2017. CODEN JARDCR. ISSN 2155-9546. URL https://www.

walshmedicalmedia.com/open-access/effect-of-zooplankton-and-
environmental-parameters-on-african-catfish-clarias-gariepinus-
burchell-1822-in-egypt-31052.html.

Julinta:2017:HWH

[398] R. B. Julinta, T. J. Abraham, Anwesha Roy, Jasmine Singha, Gadadhar Dash, T. S. Nagesh, and P. K. Patil. Histopathology and wound healing in oxytetracycline treated Oreochromis niloticus (L.) against Aeromonas hydrophila intramuscular challenge. Journal of Aquaculture Research & Development, 8(5):??, ???? 2017. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
histopathology-and-wound-healing-in-oxytetracycline-treated-
oreochromis-niloticus-l-against-aeromonas-hydrophila-intramu-
31053.html.

Rahman:2017:IAT

[399] Md Anisur Rahman, Md Mehedi Hasan Pramanik, Flura, Md Monjurul Hasan, Tayfa Ahmed, Masud Hossain Khan, and Yahia Mahmud. Impact assessment of twenty-two days fishing ban in the major spawning grounds of Tenualosa ilisha (Hamilton, 1822) on its spawning success in Bangladesh. Journal of Aquaculture Research & Development, 8(6):

??, ???? 2017. CODEN JARDCR. ISSN 2155-9546. URL https:/

/www.walshmedicalmedia.com/open-access/impact-assessment-of-
twentytwo-days-fishing-ban-in-the-major-spawning-grounds-of-
tenualosa-ilisha-hamilton-1822-on-its-s-31080.html.

Prameswari:2017:RAA

[400] K. Prameswari, M. Hemalatha, B. Kishori, and P. Sreenivasula Reddy. Role of arachidonic acid and COX inhibitors in the regulation of reproduction in freshwater crab Oziothelphusa senex senex. Journal of Aquaculture Research & Development, 8(6):??, ???? 2017. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/role-of-arachidonic-acid-and-cox-inhibitors-
in-the-regulation-of-reproduction-in-freshwater-crab-oziothelphusa-
senex-sen-31087.html.

Jalal:2017:EIS

[401] Valiallahi Jalal. Environmental impact of shrimp culture at Gwatr culture site in Chabahar, Sistan-Baluchestan Province. Journal of Aquaculture Research & Development, 8(6):??, ???? 2017. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
environmental-impact-of-shrimp-culture-at-gwatr-culture-site-
in-chabahar-sistanbaluchestan-province-31081.html.

Paul:2017:MSR

[402] Thankam Theresa Paul, Rani Palaniswamy, S. Manoharan, Usha Unnithan, and U. K. Sarkar. Management strategies for reservoirs fisheries. Journal of Aquaculture Research & Development, 8(6):??, ???? 2017. CODEN JARDCR. ISSN 2155-9546. URL https://www.

walshmedicalmedia.com/open-access/management-strategies-for-
reservoirs-fisheries-31090.html.

Ayari:2017:GMS

[403] Tahani El Ayari, Anwar Mleiki, and Najoua Trigui El Menif. Genitalia malformations in Stramonita haemastoma (Gastropoda: Muricidae) from Atlantic and Mediterranean Coast. Journal of Aquaculture Research & Development, 8(6):??, ???? 2017. CODEN JARDCR. ISSN 2155-9546.

URL https://www.walshmedicalmedia.com/open-access/genitalia-
malformations-in-stramonita-haemastoma-gastropoda-muricidae-
from-atlantic-and-mediterranean-coast-31089.html.

Chowdhury:2017:DPB

[404] Istiaq Ahmad Chowdhury, Jewel Das, and Nani Gopal Das. Development of probiotics based culture system of Macrobrachium rosenbergii using different stocking densities. Journal of Aquaculture Research & Development, 8(6):??, ???? 2017. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
development-of-probiotics-based-culture-system-of-macrobrachium-
rosenbergii-using-different-stocking-densities-31101.html.

Plamoottil:2017:TNM

[405] Mathews Plamoottil. Taxonomic notes on Mystus species of Northern Kerala. Journal of Aquaculture Research & Development, 8(6):??, ???? 2017. CODEN JARDCR. ISSN 2155-9546. URL https://www.

walshmedicalmedia.com/open-access/taxonomic-notes-on-mystus-
species-of-northern-kerala-31102.html.

Aboyadak:2017:NSR

[406] Ibrahim M. Aboyadak, Nadia GM Ali, Ashraf MAS Goda, Walaa Saad, and Asmaa ME Salam. Non-selectivity of R-S media for Aeromonas hydrophila and TCBS media for Vibrio species isolated from diseased Oreochromis niloticus. Journal of Aquaculture Research & Development, 8 (7):1–5, ???? 2017. CODEN JARDCR. ISSN 2155-9546. URL https:/

/www.walshmedicalmedia.com/open-access/nonselectivity-of-rs-
media-for-aeromonas-hydrophila-and-tcbs-media-for-vibrio-species-
isolated-from-diseased-oreochromis-31122.html.

Mchunu:2017:FSF

[407] Ntobeko Mchunu, Gareth Lagerwall, and Adian Senzanje. Food sovereignty for food security, aquaponics system as a potential method: a review. Journal of Aquaculture Research & Development, 8(7):??, ???? 2017. CODEN JARDCR. ISSN 2155-9546. URL https://www.

walshmedicalmedia.com/open-access/food-sovereignty-for-food-
security-aquaponics-system-as-a-potential-method-a-review-31132.
html.

Mekkawy:2017:CMS

[408] W. Mekkawy, B. K. Barman, A. H. M. Kohinoor, and J. A. H. Benzie. Characterization of mono-sex Nile tilapia (Oreochromis niloticus) hatcheries in Bangladesh. Journal of Aquaculture Research & Development, 8(7):??, ???? 2017. CODEN JARDCR.

ISSN 2155-9546. URL https://www.walshmedicalmedia.com/open-
access/characterization-of-monosex-nile-tilapia-oreochromis-
niloticus-hatcheries-inbangladesh-2155-9546-1000498.pdf.

El-Barbary:2017:PGC

[409] Manal I. El-Barbary and Ahmed M. Hal. Phenotypic and genotypic characterization of some Pseudomonas sp. associated with Burkholderia cepacia isolated from various infected fishes. Journal of Aquaculture Research & Development, 8(7):??, ???? 2017. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
phenotypic-and-genotypic-characterization-of-some-pseudomonas-
sp-associated-with-burkholderia-cepacia-isolated-from-vari-31144.
html.

Abdelsalam:2017:PRA

[410] Mohamed Abdelsalam. Potential role of anaerobic bacteria as fish pathogens. Journal of Aquaculture Research & Development, 8(7):??, ???? 2017. CODEN JARDCR. ISSN 2155-9546. URL https://www.

walshmedicalmedia.com/open-access/potential-role-of-anaerobic-
bacteria-as-fish-pathogens-2155-9546-1000500.pdf.

Redman:2017:AUU

[411] Natalie Redman, Christopher Good, and Brian J. Vinci. Assessing the utility of ultraviolet irradiation to reduce bacterial biofilms in fish hatchery well water supplies. Journal of Aquaculture Research & Development, 8(7):??, ???? 2017. CODEN JARDCR. ISSN 2155-9546. URL https://

www.walshmedicalmedia.com/open-access/assessing-the-utility-
of-ultraviolet-irradiation-to-reduce-bacterial-biofilms-in-fish-
hatchery-well-water-supplies-31149.html.

Oedjoe:2017:CNC

[412] Marcelien Dj Ratoe Oedjoe. Composition of nutritional content of sea cucumbers (Holothuroidea) in Mania Waters, Sabu Raijua Regency, East Nusa Tenggara. Journal of Aquaculture Research & Development, 8(7):??, ???? 2017. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
composition-of-nutritional-content-of-sea-cucumbers-holothuroidea-
in-mania-waters-sabu-raijua-regency-east-nusa-tenggara-31147.
html.

Mustami:2017:FPC

[413] Muhammad Khalifah Mustami. The formation of polyploidy on Cyprinus carpio Linn Punten race by heat shocking temperature. Journal of Aquaculture Research & Development, 8(8):1–4, ???? 2017. CODEN JARDCR.

ISSN 2155-9546. URL https://www.walshmedicalmedia.com/open-
access/the-formation-of-polyploidy-on-cyprinus-carpio-linn-punten-
race-by-heatshocking-temperature-2155-9546-1000503.pdf.

Bashirichelkasari:2017:PCE

[414] Nasim Bashirichelkasari and Reza Yadollahvandmiandoab. Placobdella costata an ectoparasite for Mauremys caspica in North of Iran. Journal of Aquaculture Research & Development, 8(9):??, ???? 2017. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/emplacobdella-costata-eman-ectoparasite-for-
emmauremys-caspicaem-in-north-of-iran-31224.html.

Khater:2017:EWT

[415] El-Sayed G. Khater, Samir A. Ali, and Waheed E. Mohamed. Effect of water temperature on masculinization and growth of Nile tilapia fish. Journal of Aquaculture Research & Development, 8(9):??, ???? 2017. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/effect-of-water-temperature-on-masculinization-
and-growth-of-nile-tilapia-fish-31235.html.

Ogundiran:2017:NCH

[416] M. A. Ogundiran. Nutritional compositions and heavy metal accumulation potentials of selected fish species from Ijede River, Lagos, Nigeria. Journal of Aquaculture Research & Development, 8(9):??, ???? 2017. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/nutritional-compositions-and-heavy-metal-accumulation-
potentials-of-selected-fish-species-from-ijede-river-lagos-nigeria-
31250.html.

Pathmanandakumar:2017:ECM

[417] Vyddiyaratnam Pathmanandakumar. The effectiveness of co-management practices: The case of small-scale fisheries in Sri Lanka. Journal of Aquaculture Research & Development, 8(9):??, ???? 2017. CODEN JARDCR.

ISSN 2155-9546. URL https://www.walshmedicalmedia.com/open-
access/the-effectiveness-of-comanagement-practices-the-case-
of-smallscale-fisheries-in-sri-lanka-31261.html.

Kumar:2017:ADS

[418] Vikash Kumar and Suvra Roy. Aquaculture drugs: Sources, active ingredients, pharmaceutic preparations and methods of administration. Journal of Aquaculture Research & Development, 8(9):??, ???? 2017. CODEN JARDCR. ISSN 2155-9546. URL https://www.

walshmedicalmedia.com/open-access/aquaculture-drugs-sources-
active-ingredients-pharmaceutic-preparations-and-methods-of-
administration-31262.html.

Kumar:2017:RPR

[419] Amrendra Kumar, Sudhanshu Shekhar, and A. Saravanakumar. Role of PmRab7 regulation in WSSV infection and functional validation of small molecule as PmRab7 GTPase inhibitor. Journal of Aquaculture Research & Development, 8(9):1–5, ???? 2017. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/role-of-pmrab7-regulation-in-wssv-infection-
and-functional-validation-of-small-molecule-as-pmrab7-gtpase-
inhibitor-31285.html.

Olanrewaju:2017:PCS

[420] A. N. Olanrewaju, E. K. Ajani, and O. K. Kareem. Physicochemical status of Eleyele Reservoir, Ibadan, Nigeria. Journal of Aquaculture Research & Development, 8(9):??, ???? 2017. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/physicochemical-status-of-eleyele-reservoir-
ibadan-nigeria-31284.html.

Iqbal:2017:ESI

[421] Sonia Iqbal, Usman Atique, Muhammad Sharif Mughal, Noor Khan, Muhammad Sultan Haider, Khalid Javed Iqbal, and Muhammad Akmal. Effect of selenium incorporated in feed on the hematological profile of tilapia (Oreochromis niloticus). Journal of Aquaculture Research & Development, 8(10):??, ???? 2017. CODEN JARDCR. ISSN 2155-9546.

URL https://www.walshmedicalmedia.com/open-access/effect-of-
selenium-incorporated-in-feed-on-the-hematological-profile-of-
tilapia-emoreochromis-niloticusem-31323.html.

Wan:2017:SDS

[422] Yu Wan, Jia He, and Yang Bai. Spatial distribution of sediment bacterial communities in eutrophicated Meiliang Bay: Correlation with environmental factors. Journal of Aquaculture Research & Development, 8(10): ??, ???? 2017. CODEN JARDCR. ISSN 2155-9546. URL https:/

/www.walshmedicalmedia.com/open-access/spatial-distribution-
of-sediment-bacterial-communities-in-eutrophicated-meiliang-
bay-correlation-with-environmental-fact-31326.html.

Sarker:2017:PSS

[423] Subrata Sarker, Shyamal Chandra Basak, Jahid Hasan, Md. Solaiman Hossain, Muhammad Mizanur Rahman, and Md. Ahsanul Islam. Production in small scale aquaculture farm: a success story from Bangladesh. Journal of Aquaculture Research & Development, 8(10):??, ???? 2017. CODEN JARDCR. ISSN 2155-9546. URL https://www.

walshmedicalmedia.com/open-access/production-in-small-scale-
aquaculture-farm-a-success-story-from-bangladesh-31324.html.

Chen:2017:REI

[424] J. C. Chen, Y. C. Lin, Y. Y. Chen, and C. L. Huang. Up-regulated expressions of immune parameters, ppa, propo, sod, and hsp70 in white shrimp Litopenaeus vannamei reared at unfavorable low salinities. Journal of Aquaculture Research & Development, 8(11):??, ???? 2017. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/upregulated-expressions-of-immune-parameters-
ppa-propo-sod-and-hsp70-in-white-shrimp-emlitopenaeus-vannameiem-
reared-at--31381.html.

Al-Ngada:2017:EDS

[425] R. S. Al-Ngada, A. M. Abdelwahab, and S. M. El-Bahr. Effect of dietary supplementation of green tea (Camellia sinensis) on growth, body composition and serum biochemistry of the Asian seabass, Lates calcarifer fingerlings. Journal of Aquaculture Research & Development, 8(11):??, ???? 2017. CODEN JARDCR. ISSN 2155-9546.

URL https://www.walshmedicalmedia.com/open-access/effect-of-
dietary-supplementation-of-green-tea-camellia-sinensis-on-growth-
body-composition-and-serum-biochemistry-of-th-31391.html.

Shen:2017:MPP

[426] Hui Shen, Ge Jiang, Xihe Wan, Xianping Fan, Yi Qiao, Wenjun Shi, Hui Li, and Libao Wang. Multiple pathogens prevalent in shrimp Penaeus vannamei cultured from greenhouse ponds in Jiangsu Province of China. Journal of Aquaculture Research & Development, 8(12):??, ???? 2017. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
multiple-pathogens-prevalent-in-shrimp-empenaeus-vannameiem-
cultured-from-greenhouse-ponds-in-jiangsu-province-of-china-
31325.html.

Chen:2018:GPL

[427] Mingrui Chen, Yinglong Wu, Qiuping Yan, Zhenzhen Lv, Mei He, Limei Feng, Jiaqi Zhao, and Zhengxin Duan. Growth performance, lipid deposition and hepatic lipid metabolism related gene expression in Schizothorax prenanti fed with dietary acidolysis-oxidized konjac glucomannan supplementation. Journal of Aquaculture Research & Development, 9(1):1–8, ???? 2018. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
growth-performance-lipid-deposition-and-hepatic-lipid-metabolism-
related-gene-expression-in-schizothorax-prenanti-fed-wi-31424.
html.

Worku:2018:IWH

[428] Melese Worku and Samuel Sahile. Impact of water hyacinth, Eichhornia crassipes (Martius) (Pontederiaceae) in Lake Tana Ethiopia: a review. Journal of Aquaculture Research & Development, 9(1):??, ???? 2018. CODEN JARDCR. ISSN 2155-9546. URL https://www.

walshmedicalmedia.com/open-access/impact-of-water-hyacinth-eichhornia-
crassipes-martius-pontederiaceae-in-lake-tana-ethiopia-a-review-
31425.html.

Agidie:2018:CWH

[429] Aklilu Agidie, Samuel Sahle, Adugnaw Admas, and Mehari Alebachew. Controlling water hyacinth, Eichhorni acrassipes (Mart.) Solms using some selected eco-friendly chemicals. Journal of Aquaculture Research & Development, 9(1):??, ???? 2018. CODEN JARDCR.

ISSN 2155-9546. URL https://www.walshmedicalmedia.com/open-
access/controlling-water-hyacinth-eichhorni-acrassipes-mart-
solms-using-some-selected-ecofriendly-chemicals-31430.html.

Kattakdad:2018:PCD

[430] Supalug Kattakdad, Orapint Jintasataporn, Wanchai Worawattanamateekul, and Srinoy Chumkam. pH characterization of digestive enzyme and in vitro digestibility of red bee shrimp Caridina cantonensis (Decapoda: Atyidae). Journal of Aquaculture Research & Development, 9(2): 1–6, ???? 2018. CODEN JARDCR. ISSN 2155-9546. URL https:

//www.walshmedicalmedia.com/open-access/ph-characterization-
of-digestive-enzyme-andem-in-vitroem-digestibility-of-red-bee-
shrimp-emcaridina-cantonensisem-decapo-31478.html.

Suantika:2018:ERS

[431] Gede Suantika, Magdalena Lenny Situmorang, Abdul Khakim, Indra Wibowo, Pingkan Aditiawati, Shrikumar Suryanarayan, Sri Sailaja Nori, Sawan Kumar, and Ferisca Putri. Effect of red seaweed Kappaphycus alvarezii on growth, survival, and disease resistance of Pacific white shrimp Litopenaeus vannamei against Vibrio harveyi in the nursery phase. Journal of Aquaculture Research & Development, 9(2):??, ???? 2018. CODEN JARDCR. ISSN 2155-9546. URL

https://www.walshmedicalmedia.com/open-access/effect-of-red-
seaweedem-kappaphycus-alvareziiem-on-growth-survival-and-disease-
resistance-of-pacific-white-shrimp-emlito-31479.html.

Purkait:2018:IFP

[432] Soumyadip Purkait, Thangapalam Jawahar Abraham, Sutanu Karmakar, Biswadeep Dey, and Anwesha Roy. Inhibition of fish pathogenic Aeromonas hydrophila and Edwardsiella tarda by Centella asiatica in-vitro. Journal of Aquaculture Research & Development, 9(2): 1–5, ???? 2018. CODEN JARDCR. ISSN 2155-9546. URL

https://www.walshmedicalmedia.com/open-access/inhibition-of-

fish-pathogenic-emaeromonas-hydrophilaem-and-emedwardsiella-
tardaem-by-emcentella-asiatica-invitroem-31480.html.

Gupta:2018:RFR

[433] Sandipan Gupta. A review on feeding and reproductive biology of Ompok pabda with an emphasis on its conservation. Journal of Aquaculture Research & Development, 9(2):??, ???? 2018. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
a-review-on-feeding-and-reproductive-biology-of-ompok-pabda-
with-an-emphasis-on-its-conservation-31490.html.

Moshaand:2018:EDC

[434] Sebastian S. Moshaand and Frank T. Mlingi. Effects of different catfish pituitary gland extract dosages on eggs and hatchlings quantity of African catfish, Clarias gariepinus at a constant latency period. Journal of Aquaculture Research & Development, 9(3):??, ???? 2018. CODEN JARDCR.

ISSN 2155-9546. URL https://www.walshmedicalmedia.com/open-
access/effects-of-different-catfish-pituitary-gland-extract-
dosages-on-eggs-and-hatchlings-quantity-of-african-catfish-clarias-
-31515.html.

Mulyadi:2018:DEF

[435] Yeyes Mulyadi, Kriyo Sambodho, Nur Syahroni, Muhammad Zikra, and Winda Amalia Herdianti. Development of eco-friendly aquaculture design for lobster cultivation in Indonesia. Journal of Aquaculture Research & Development, 9(3):??, ???? 2018. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
development-of-ecofriendly-aquaculture-design-for-lobster-cultivation-
in-indonesia-31516.html.

Mosha:2018:RCS

[436] Sebastian S. Mosha. Recent comparative studies on the performance and survival rate of African catfish (Clarias gariepinus) larval produced under natural and synthetics hormones: a review. Journal of Aquaculture Research & Development, 9(3):??, ???? 2018. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
recent-comparative-studies-on-the-performance-and-survival-rate-
of-african-catfish-clarias-gariepinus-larval-produced-un-31517.
html.

Khalil:2018:EPU

[437] Hala Saber Khalil, Abdallah Tageldein Mansour, Ashraf Mohamed Abdelsamee Goda, Ahmed Kamel El-Hammady, and Eglal Ali Omar. Effect of poly-unsaturated fatty acids fortification on growth performance, survival, fatty acid composition and antioxidant balance of meagre, Argyrosomus regius larvae. Journal of Aquaculture Research & Development, 9(3):??, ???? 2018. CODEN JARDCR. ISSN 2155-9546. URL https://www.

walshmedicalmedia.com/open-access/effect-of-polyunsaturated-
fatty-acids-fortification-on-growth-performance-survival-fatty-
acid-composition-and-antioxidan-31529.html.

Suantika:2018:AIR

[438] Gede Suantika, Magdalena Lenny Situmorang, Adani Nurfathurahmi, Intan Taufik, Pingkan Aditiawati, Nasukha Yusuf, and Rizkiyanti Aulia. Application of indoor recirculation aquaculture system for white shrimp (Litopenaeus vannamei) growout super-intensive culture at low salinity condition. Journal of Aquaculture Research & Development, 9(3): ??, ???? 2018. CODEN JARDCR. ISSN 2155-9546. URL https://

www.walshmedicalmedia.com/open-access/application-of-indoor-
recirculation-aquaculture-system-for-white-shrimp-litopenaeus-
vannamei-growout-superintensive-cult-31530.html.

Al-Sheraa:2018:MQT

[439] Amal S. Al-Sheraa. Microbial quality of three imported fresh locally produced marine fishes in Al-Faw City, Basrah, Iraq. Journal of Aquaculture Research & Development, 9(4):??, ???? 2018. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
microbial-quality-of-three-imported-fresh-locally-produced-marine-
fishes-in-alfaw-city-basrah-iraq-2155-9546-1000531.pdf.

Phiri:2018:TET

[440] Francis Phiri and Xinhua Yuan. Technical efficiency of tilapia production in Malawi and China: Application of stochastic frontier production approach. Journal of Aquaculture Research & Development, 9(4): ??, ???? 2018. CODEN JARDCR. ISSN 2155-9546. URL https:/

/www.walshmedicalmedia.com/open-access/technical-efficiency-
of-tilapia-production-in-malawi-and-china-application-of-stochastic-
frontier-production-approach-31570.html.

Oh:2018:IPR

[441] Sueng-Bo Oh, Chi-Hoon Lee, and Young-Don Lee. Induction of puberty in red spotted grouper, Epinephelus akaara by water temperature. Journal of Aquaculture Research & Development, 9(5):??, ???? 2018. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/induction-of-puberty-in-red-spotted-grouper-
epinephelus-akaara-by-water-temperature-31600.html.

Georges:2018:ESM

[442] Fonkwa Georges, Lekeufack Folefack Guy Benoît, Tchuinkam Timoléon, Ishtiyaq Ahmad, and Tchoumboue Joseph. Effect of season on Myxosporean infections in Oreochromis niloticus Linnaeus, 1758 (Cichlidae) at MAPE Dam in Adamawa, Cameroon. Journal of Aquaculture Research & Development, 9(5):??, ???? 2018. CODEN JARDCR. ISSN 2155-9546.

URL https://www.walshmedicalmedia.com/open-access/effect-of-
season-on-emmyxosporeanem-infections-in-emoreochromis-niloticusem-
linnaeus-1758-cichlidae-at-mape-dam-in-adama-31594.html.

Niu:2018:SFB

[443] Bingshan Niu, Guangyao Li, Fang Peng, Jing Wu, Long Zhang, and Zhenbo Li. Survey of fish behavior analysis by computer vision. Journal of Aquaculture Research & Development, 9(5):??, ???? 2018. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/survey-of-fish-behavior-analysis-by-computer-
vision-31595.html.

Phiri:2018:EPT

[444] Francis Phiri and Xinhua Yuan. Economic profitability of tilapia production in Malawi and China. Journal of Aquaculture Research & Development, 9(5):??, ???? 2018. CODEN JARDCR. ISSN 2155-9546.

URL https://www.walshmedicalmedia.com/open-access/economic-
profitability-of-tilapia-production-in-malawi-and-china-31596.
html.

Echem:2018:ECC

[445] Roldan T. Echem, Herbert M. Barba, Guangyao Li, Fang Peng, and Nikka Joy C. Buenaventura. Endoparasites in Chanos chanos (Forsskal, 1775) from the wetlands of Zamboanga City, Western Mindanao, Philippines. Journal of Aquaculture Research & Development, 9(5):??, ???? 2018. CODEN JARDCR. ISSN 2155-9546. URL https://www.

walshmedicalmedia.com/open-access/endoparasites-in-emchanos-
chanosem-forsskal-1775-from-the-wetlands-of-zamboanga-city-western-
mindanao-philippines-31597.html.

Islam:2018:CBA

[446] Seikh Razibul Islam, Md. Aminul Islam, Mohsina Mannan, Mahadi Hasan Osman, Md. Shaheed Reza, Md. Naim Uddin, and Md. Khalilur Rahman. Comparative bacteriological analysis of giant freshwater prawn (Macrobrachium rosenbergii) cultured in southwestern coastal areas of Bangladesh. Journal of Aquaculture Research & Development, 9(6):??, ???? 2018. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
comparative-bacteriological-analysis-of-giant-freshwater-prawn-
emmacrobrachium-rosenbergiiem-cultured-in-southwestern-co-31624.
html.

Novriadi:2018:EDL

[447] Romi Novriadi and Gabriel Proano. Effects of different light intensities and nutrient source on growth rate and crude fat content in Chlorella vulgaris. Journal of Aquaculture Research & Development, 9 (6):??, ???? 2018. CODEN JARDCR. ISSN 2155-9546. URL https:/

/www.walshmedicalmedia.com/open-access/effects-of-different-
light-intensities-and-nutrient-source-on-growth-rate-and-crude-
fat-content-in-emchlorella-vulgarise-31625.html.

Goncalves:2018:FHD

[448] Rui A. Gonçalves, Sofia Engrola, Cláudia Aragão, Simon Mackenzie, Gerlinde Bichl, Tibor Czabany, and Dian Schatzmayr. Fate of [^{3}H]- deoxynivalenol in rainbow trout (Oncorhynchus mykiss) juveniles: Tissue distribution and excretion. Journal of Aquaculture Research & Development, 9(7):??, ???? 2018. CODEN JARDCR. ISSN 2155-9546.

URL https://www.walshmedicalmedia.com/open-access/fate-of-
sup3suphdeoxynivalenol-in-rainbow-trout-emoncorhynchus-mykissem-
juveniles-tissue-distribution-and-excretion-31643.html.

Rana:2018:HFR

[449] Muhammad Enayet Ullah Rana, Abdus Salam, K. M. Shahriar Nazrul, and Md. Monjurul Hasan. Hilsa fishers of Ramgati, Lakshmipur, Bangladesh: an overview of socio-economic and livelihood context. Journal of Aquaculture Research & Development, 9(7): ??, ???? 2018. CODEN JARDCR. ISSN 2155-9546. URL

https://www.walshmedicalmedia.com/open-access/hilsa-fishers-
of-ramgati-lakshmipur-bangladesh-an-overview-of-socio-economic-
and-livelihood-context-31644.html.

Ram:2018:CRP

[450] E. J. Ram. Cleaning regime for Pinctada margaritifera (Linneaus, 1758) in Savusavu Bay, Fiji. Journal of Aquaculture Research & Development, 9 (7):??, ???? 2018. CODEN JARDCR. ISSN 2155-9546. URL https:

//www.walshmedicalmedia.com/open-access/cleaning-regime-for-
empinctada-margaritiferaem-linneaus-1758-in-savusavu-bay-fiji-
31651.html.

Miah:2018:AGD

[451] Md. Faruque Miah, Saeed Anwar, Md. Hazrat Ali, Niamul Naser M, and Kawser Ahmed. Assessment of genetic diversity of freshwater mud eel (Monopterus cuchia) using RAPD and RFLP markers. Journal of Aquaculture Research & Development, 9(7):??, ???? 2018. CODEN JARDCR. ISSN 2155-9546. URL https://

www.walshmedicalmedia.com/open-access/assessment-of-genetic-
diversity-of-freshwater-mud-eel-monopterus-cuchia-using-rapd-
and-rflp-markers-31668.html.

Mosha:2018:RSA

[452] Sebastian S. Mosha. A review on significance of azolla meal as a protein plant source in finfish culture. Journal of Aquaculture Research & Development, 9(7):??, ???? 2018. CODEN JARDCR. ISSN 2155-9546.

URL https://www.walshmedicalmedia.com/open-access/a-review-
on-significance-of-azolla-meal-as-a-protein-plant-source-in-
finfish-culture-31669.html.

Amenyogbe:2018:RGA

[453] Eric Amenyogbe, Gang Chen, Zhongliang Wang, Mingde Lin, Xiaoying Lu, Denicia Atujona, and Emmanuel D. Abarike. A review of ghanas aquaculture industry. Journal of Aquaculture Research & Development, 9(8):??, ???? 2018. CODEN JARDCR. ISSN 2155-

9546. URL https://www.walshmedicalmedia.com/open-access/a-
review-of-ghanas-aquaculture-industry-31693.html.

Ram:2018:GSP

[454] E. J. Ram and C. Morris. Growth and survival of Pinctada margaritifera mother of pearl oyster cocultured with Kappaphycus alvarezii seaweeds. Journal of Aquaculture Research & Development, 9(8):??, ???? 2018. CODEN JARDCR. ISSN 2155-9546. URL https:

//www.walshmedicalmedia.com/open-access/growth-and-survival-
of-pinctada-margaritifera-mother-of-pearl-oyster-cocultured-
with-kappaphycus-alvarezii-seaweeds-31694.html.

Guo:2018:QDL

[455] Quanyou Guo, Thanhhoa Truonghuynh, Chaojun Jiang, Luming Wang, Baoguo Li, and Xiaoliang Xing. Quality differences of large yellow croaker (Pseudosciaena crocea) cultured in deep-water sea cages of two China regions. Journal of Aquaculture Research & Development, 9(9): ??, ???? 2018. CODEN JARDCR. ISSN 2155-9546. URL https:

//www.walshmedicalmedia.com/open-access/quality-differences-

of-large-yellow-croaker-pseudosciaena-crocea-cultured-in-deepwater-
sea-cages-of-two-china-regions-31707.html.

Emeish:2018:AIA

[456] Walaa F. A. Emeish, Hams M. A. Mohamed, and Ahmad A. Elkamel. Aeromonas infections in African sharptooth catfish. Journal of Aquaculture Research & Development, 9(9):??, ???? 2018. CODEN JARDCR.

ISSN 2155-9546. URL https://www.walshmedicalmedia.com/open-
access/aeromonas-infections-in-african-sharptooth-catfish-31735.
html.

Khalila:2018:DSS

[457] Hala Saber Khalila, Walied Mohamed Fayed, Abdallah Tageldein Mansour, Tarek Mohamed Srour, Eglal Ali Omar, Shawky Ibrahim Darwish, and Abdel Aziz Moussa Nour. Dietary supplementation of spirulina, Arthrospira platensis, with plant protein sources and their effects on growth, feed utilization and histological changes in Nile tilapia, Oreochromis niloticus. Journal of Aquaculture Research & Development, 9(10):1–9, ???? 2018. CODEN JARDCR. ISSN 2155-9546.

URL https://www.walshmedicalmedia.com/open-access/dietary-
supplementation-of-spirulina-emarthrospira-platensisem-with-
plant-protein-sources-and-their-effects-on-growth-fe-31765.html.

Elkhodary:2018:NVE

[458] Gihan M. Elkhodary, Nahla E. Omran, Amal A. Hassan, Salwa A. El-Saidy, and Mohamed H. Mona. Nutritional value of the Egyptian freshwater bivalve Spathopsis rubens arcuata under the effect of depuration. Journal of Aquaculture Research & Development, 9(10): ??, ???? 2018. CODEN JARDCR. ISSN 2155-9546. URL https:/

/www.walshmedicalmedia.com/open-access/nutritional-value-of-
the-egyptian-freshwater-bivalve-emspathopsis-rubensem-arcuata-
under-the-effect-of-depuration-31779.html.

Eirna-Liza:2018:DPC

[459] N. Eirna-Liza, Hasliza Abu Hassim, Chong Chou Min, Fadhil Syukri, and Murni Karim. The duration of protection conferred by garlic on African catfish (Clarias gariepinus) against Aeromonas hydrophila. Journal of Aquaculture Research & Development, 9(10):??, ???? 2018. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/the-duration-of-protection-conferred-by-garlic-
on-african-catfish-emclarias-gariepinusem-against-emaeromonas-
hydrophilae-31780.html.

Singh:2018:EPF

[460] Gajender Singh, Anita Bhatnagar, Kalla Alok, and Singh Ashneel Ajay. Enzymatic profiling and feeding preferences of catla: Catla catla, rohu: Labeo rohita and mrigala: Cirrhinus mrigala in rural polyculture ponds. Journal of Aquaculture Research & Development, 9(10): 1–8, ???? 2018. CODEN JARDCR. ISSN 2155-9546. URL https:

//www.walshmedicalmedia.com/open-access/enzymatic-profiling-
and-feeding-preferences-of-catla-emcatla-catlaem-rohu-emlabeo-
rohitaem-and-mrigala-emcirrhinus-mriga-31781.html.

Oyebamiji:2018:LWR

[461] O. Fafioye Oyebamiji, A. Asiru Raheem, and Y. Oladunjoye Rasheed. Length–weight relationship, abundance and sex ratio of the giant river prawn Macrobrachium vollenhovenii (Herklots, 1857) from River Osun, Southwestern Nigeria. Journal of Aquaculture Research & Development, 9(10):1–5, ???? 2018. CODEN JARDCR.

ISSN 2155-9546. URL https://www.walshmedicalmedia.com/open-
access/lengthweight-relationship-abundance-and-sex-ratio-of-
the-giant-river-prawnemmacrobrachium-vollenhoveniiem-herklots-
1857--31782.html.

Anto:2018:PCC

[462] Junianto Anto, Iskandar, and Achmad Rizal. Physico-chemical characteristics and levels of preference for drinking collagen drinks the result of extracts from Nilem fish skins. Journal of Aquaculture Research & Development, 9(11):1–4, ???? 2018. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
physicochemical-characteristics-and-levels-of-preference-for-
drinking-collagen-drinks-the-result-of-extracts-from-nilem--
31808.html.

Limon-Morales:2018:SRC

[463] M. C. Limon-Morales, H. Hernandez-Moreno, C. Carmona-Osalde, and M. Rodriguez-Serna. Study of the reproduction of Cambarellus montezumae (Saussure, 1857) under different sex relations. Journal of Aquaculture Research & Development, 9(11):??, ???? 2018. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
study-of-the-reproduction-of-emcambarellus-montezumaeem-saussure-
1857-under-different-sex-relations-31809.html.

Pramono:2018:LWR

[464] Taufik Budhi Pramono, Diana Arfiati, Maheno Sri, and Uun Yanuhar. Length–weight relationship and condition factor senggaringan fish (Mystus singaringan) from Klawing River, Central Java, Indonesia. Journal of Aquaculture Research & Development, 9(12):??, ???? 2018. CODEN JARDCR. ISSN 2155-9546. URL https://www.

walshmedicalmedia.com/open-access/lengthweight-relationship-
and-condition-factor-senggaringan-fish-mystus-singaringan-from-
klawing-river-central-java-indo-31810.html.

Parandavar:2018:EGP

[465] Hossein Parandavar, Mizanur Rahman, Phoo War War, and Chang-Hoon Kim. Effects of growth on polychaete rockworm, Marphysa sanguinea integrated culture with olive flounder, Paralichythys olivaceus in flow through system. Journal of Aquaculture Research & Development, 9(12): ??, ???? 2018. CODEN JARDCR. ISSN 2155-9546. URL https:/

/www.walshmedicalmedia.com/open-access/effects-of-growth-on-
polychaete-rockworm-emmarphysa-sanguineaem-integrated-culture-
with-olive-flounder-emparalichythys-o-31812.html.

Imran:2019:SEAa

[466] Siti Zaharah Imran and Lee Seong Wei. Selected essential amino acid enhancement by Bacillus cereus from solid state fermentation of soy pulp. Journal of Aquaculture Research & Development, 10(1):??, ???? 2019. CODEN JARDCR. ISSN 2155-9546.

URL https://www.walshmedicalmedia.com/open-access/selected-
essential-amino-acid-enhancement-by-embacillus-cereusem-from-
solid-state-fermentation-of-soy-pulp-31820.html.

Wahidi:2019:CBS

[467] Budi Rianto Wahidi, Uun Yanuhar, Mohamad Fadjar, and Sri Andayani. Comparison between single-round PCR assay and IHC for detection of koi herpes-virus (KHV) in Oreochromis niloticus. Journal of Aquaculture Research & Development, 10(1):??, ???? 2019. CODEN JARDCR.

ISSN 2155-9546. URL https://www.walshmedicalmedia.com/open-
access/comparison-between-singleround-pcr-assay-and-ihc-for-
detection-of-koi-herpesvirus-khv-in-emoreochromis-niloticusem-
31821.html.

Allah:2019:ECC

[468] Mohammed M. Abd-Allah, Ashraf A. Ramadan, Noha M. Said, Ibrahim H. Ibrahim, and Esam A. Abdel-karim. Effects of cadmium chloride and glyphosate on antioxidants as biochemical biomarkers in Nile tilapia. Journal of Aquaculture Research & Development, 10(1):??, ???? 2019. CODEN JARDCR. ISSN 2155-9546.

URL https://www.walshmedicalmedia.com/open-access/effects-

of-cadmium-chloride-and-glyphosate-on-antioxidants-as-biochemical-
biomarkers-in-nile-tilapia-31822.html.

Thu:2019:SEG

[469] Em Vo Thi Thu, Mizanur Rahman, War War Phoo, and Chang-Hoon Kim. Salinity effects on growth and survival of the polychaete rockworm Marphysa sanguinea (Montagu, 1813) juveniles and adults. Journal of Aquaculture Research & Development, 10(2):??, ???? 2019. CODEN JARDCR. ISSN 2155-9546. URL https:

//www.walshmedicalmedia.com/open-access/salinity-effects-on-
growth-and-survival-of-the-polychaete-rockworm-emmarphysa-sanguineaem-
montagu-1813-juveniles-and-adu-31823.html.

Rahman:2019:IPG

[470] Zinia Rahman, Al Mamun, Istiyak Ahmad, and Ibrahim Rashid. Influence of probiotics on the growth performance of sex reversed Nile tilapia (Oreochromis niloticus, Linnaeus, 1758) fry. Journal of Aquaculture Research & Development, 10(2):??, ???? 2019. CODEN JARDCR.

ISSN 2155-9546. URL https://www.walshmedicalmedia.com/open-
access/influence-of-probiotics-on-the-growth-performance-of-
sex-reversed-nile-tilapia-emoreochromis-niloticusem-linnaeus-
1758-f-31850.html.

Agbeko:2019:TSD

[471] Etornyo Agbeko, Daniel Adjei-Boateng, Nelson W. Agbo, and Thomas K. Agyemang. Trophic status and development of Aquaculture Management Areas (AMAs) for the two major reservoirs: Tono and Vea, in the Upper East Region of Ghana. Journal of Aquaculture Research & Development, 10(2):??, ???? 2019. CODEN JARDCR. ISSN 2155-9546.

URL https://www.walshmedicalmedia.com/open-access/trophic-
status-and-development-of-aquaculture-management-areas-amas-
for-the-two-major-reservoirs-tono-and-vea-in-the-upp.pdf.

Orire:2019:ECP

[472] A. M. Orire and G. I. Emine. Effects of crude protein levels and binders on feed buoyancy. Journal of Aquaculture Research & Development, 10(3):??, ???? 2019. CODEN JARDCR. ISSN 2155-9546.

URL https://www.walshmedicalmedia.com/open-access/effects-
of-crude-protein-levels-and-binders-on-feed-buoyancy.pdf.

Imran:2019:SEAb

[473] Siti Zaharah Imran and Lee Seong Wei. Selected essential amino acid enhancement by Bacillus cereus from solid state fermentation of soy pulp through carbon concentration and fermentation period. Journal of Aquaculture Research & Development, 10(3):??, ???? 2019. CODEN JARDCR.

ISSN 2155-9546. URL https://www.walshmedicalmedia.com/open-
access/selected-essential-amino-acid-enhancement-by-bacillus-
cereus-from-solid-state-fermentation-of-soy-pulp-through-carbon-
co.pdf.

Maulu:2019:RST

[474] Sahya Maulu, Brian Pelekelo Munganga, Oliver Jolezya Hasimuna, Lloyd Haninga Haambiya, and Bornwell Seemani. A review of the science and technology developments in Zambia’s aquaculture industry. Journal of Aquaculture Research & Development, 10(4):??, ???? 2019. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/a-review-of-the-science-and-technology-developments-
in-zambias-aquaculture-industry.pdf.

Jiboye:2019:IEDa

[475] Joy Oluwatomi Jiboye, Christopher Ogolo Ikporukpo, and Charles Olufisayo Olatubara. Impact of environmental degradation on livelihoods in the coastal areas of South West Nigeria. Journal of Aquaculture Research & Development, 10(4):??, ???? 2019. CODEN JARDCR. ISSN 2155-

9546. URL https://www.walshmedicalmedia.com/abstract/impact-
of-environmental-degradation-on-livelihoods-in-the-coastal-areas-
of-south-west-nigeria-31911.html.

Aimerance:2019:ETF

[476] Kenfack Donhachi Aimérance, Efole Ewoukem Thomas, Nana Towa Algrient, Zebaze Toguet Serge Hubert, and J. Tchoumboue. Effects of the type of fertilizer on the taxonomic composition and zooplankton distribution in ponds of Batie in Western Cameroon. Journal of Aquaculture Research & Development, 10(5):??, ???? 2019. CODEN JARDCR. ISSN 2155-9546.

URL https://www.walshmedicalmedia.com/open-access/effects-
of-the-type-of-fertilizer-on-the-taxonomic-composition-and-zooplankton-
distribution-in-ponds-of-batie-in-western.pdf.

Jiboye:2019:IEDb

[477] Joy Oluwatomi Jiboye, Christopher Ogolo Ikporukpo, and Charles Olufisayo Olatubara. Impact of environmental degradation on livelihoods in the coastal areas of South West Nigeria. Journal of Aquaculture Research & Development, 10(6):??, ???? 2019. CODEN JARDCR. ISSN 2155-9546.

URL https://www.walshmedicalmedia.com/open-access/impact-of-
environmental-degradation-on-livelihoods-in-the-coastal-areas-
of-south-west-nigeria.pdf.

Ogundiran:2019:CSN

[478] M. A. Ogundiran and T. A. Ayandiran. Comparative studies on the nutritional fitness of some selected cichlids sampled from OBA Reservoir, South Western Nigeria. Journal of Aquaculture Research & Development, 10(7):??, ???? 2019. CODEN JARDCR. ISSN 2155-9546. URL https:

//www.walshmedicalmedia.com/open-access/comparative-studies-
on-the-nutritional-fitness-of-some-selected-cichlids-sampled-
from-oba-reservoir-south-western-nigeri.pdf.

Lee:2019:EAC

[479] Bong Joo Lee, So Young Kim, Hyo Rin Kang, Hyun Jeong Hwang, Mi So Seong, Ye Eun Jeong, and Jae Hun Cheong. Effect of aquafeed containing soybean meal on lipid metabolism and immune gene expression in the liver of olive flounder. Journal of Aquaculture Research & Development, 10(7):??, ???? 2019. CODEN JARDCR. ISSN 2155-9546.

URL https://www.walshmedicalmedia.com/open-access/effect-of-
aquafeed-containing-soybean-meal-on-lipid-metabolism-and-immune-
gene-expression-in-the-liver-of-olive-flounder.pdf.

Gyan:2019:ERF

[480] Watson Ray Gyan, Stephen Ayiku, and Qihui Yang. Effects of replacing fishmeal with soybean products in fish and crustaceans performance. Journal of Aquaculture Research & Development, 10(8):??, ???? 2019. CODEN JARDCR. ISSN 2155-9546. URL https:/

/www.walshmedicalmedia.com/open-access/effects-of-replacing-
fishmeal-with-soybean-products-in-fish-and-crustaceans-performance-
44334.html.

Ram:2019:PMF

[481] E. J. Ram and A. Kalla. Pinctada margaritifera farming in Fiji. Journal of Aquaculture Research & Development, 10(9):??, ???? 2019. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/empinctada-margaritiferaem-farming-in-fiji-44707.
html.

Wright:2019:CPT

[482] Samuel James Wright. Customer perceptions towards aquatic pets and how that influences farming/catching, trading and caring for these animals in the aquatic trade industry from both fish farms and wild caught populations. Journal of Aquaculture Research & Development, 10(10): ??, ???? 2019. CODEN JARDCR. ISSN 2155-9546. URL https:/

/www.walshmedicalmedia.com/open-access/customer-perceptions-

towards-aquatic-pets-and-how-that-influences-farmingcatching-
trading-and-caring-for-these-animals-in-the-aqua-44708.html.

Sansuwan:2019:EDZ

[483] Kanokwan Sansuwan, El-Orapint Jintasataporn, and Srinoy Chumkam. Effects of dietary zinc amino acid complex and zinc sulfate on growth performance, digestive enzyme activity and immune response in Asian seabass (Lates calcarifer). Journal of Aquaculture Research & Development, 10(11):??, ???? 2019. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
effects-of-dietary-zinc-amino-acid-complex-and-zinc-sulfate-
on-growth-performance-digestive-enzyme-activity-and-immune-response-
in-44333.html.

Tanjung:2019:SRE

[484] Livia Rossila Tanjung. Survival rate and early development of giant gourami Padang strain Osphronemus gouramy (Perciformes: Osphronemidae). Journal of Aquaculture Research & Development, 10 (12):??, ???? 2019. CODEN JARDCR. ISSN 2155-9546. URL

https://www.walshmedicalmedia.com/open-access/survival-rate-
and-early-development-of-giant-gourami-padang-strain-emosphronemus-
gouramyem-perciformes-osphronemidae-45346.html.

Sun:2019:EFC

[485] Ying Sun, Xing Zhao, Haiyan Liu, and Zhencai Yang. Effect of fiber content in practical diet on feed utilization and antioxidant capacity of loach, Misgurnus anguillicaudatus. Journal of Aquaculture Research & Development, 10(12):??, ???? 2019. CODEN JARDCR. ISSN 2155-9546.

URL https://www.walshmedicalmedia.com/open-access/effect-of-
fiber-content-in-practical-diet-on-feed-utilization-and-antioxidant-
capacity-of-loach-emmisgurnus-anguillicaudatusem-45459.html.

Sahoo:2020:IID

[486] P. K. Sahoo, A. Paul, M. K. Sahoo, S. Pattanayak, P. Rajesh Kumar, and B. K. Das. Incidences of infectious diseases in freshwater aquaculture farms of Eastern India: a passive surveillance based study from 2014–2018. Journal of Aquaculture Research & Development, 11(1):??, ???? 2020. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
incidences-of-infectious-diseases-in-freshwater-aquaculture-
farms-of-eastern-india-a-passive-surveillance-based-study-from-
2014201-52836.html.

Panagiota:2020:EOV

[487] Panagiotaki Panagiota, Ntantali Olga, Karatzinos Theodoros, Antoniou Zoe, and Fleris Georgios. Effect of Origanum vulgare, Eugenia aromatica and Cinnamomum zeylanicum essential oils as welfare promoters in gilthead seabream during slaughter. Journal of Aquaculture Research & Development, 11(2):??, ???? 2020. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
effect-of-origanum-vulgare-eugenia-aromatica-and-cinnamomum-
zeylanicum-essential-oils-as-welfare-promoters-in-gilthead-s.
pdf.

Paruntu:2020:ERE

[488] Carolus P. Paruntu, Suria Darwisito, Antonius P. Rumengan, Defny S. Wewengkang, and Henki Rotinsulu. The effects of rabbitfish existence in polyculture system and feed type against the growth performance of bigeye trevally in floating net cage. Journal of Aquaculture Research & Development, 11(2):??, ???? 2020. CODEN JARDCR. ISSN 2155-

9546. URL https://www.walshmedicalmedia.com/open-access/the-
effects-of-rabbitfish-existence-in-polyculture-system-and-feed-
type-against-the-growth-performance-of-bigeye-trevall.pdf.

Abiyu:2020:PIN

[489] Muluken Abiyu, Gebrekrustos Mekonnen, and Kidanu Hailay. Prevalence of internal nematode parasites of Nile tilapia (Oreochromis niloticus) fish species caught from southwestern part of Lake Tana, Central Gondar, Ethiopia. Journal of Aquaculture Research & Development, 11(3):??, ???? 2020. CODEN JARDCR. ISSN 2155-9546. URL

https://www.walshmedicalmedia.com/open-access/prevalence-of-
internal-nematode-parasites-of-nile-tilapia-oreochromis-niloticus-
fish-species-caught-from-southwestern-pa.pdf.

Tantu:2020:VSC

[490] Andi Gusti Tantu, Muhammad Ishak, and Suryawati Salam. Vaname shrimp cultivation (Litopenaeus vannamei) on high stocking densities in controlled ponds. Journal of Aquaculture Research & Development, 11(3):1–6, ???? 2020. CODEN JARDCR. ISSN 2155-9546. URL

https://www.walshmedicalmedia.com/open-access/vaname-shrimp-
cultivation-litopenaeus-vannamei-on-high-stocking-densities-
in-controlled-ponds.pdf.

Hasimuna:2020:AHM

[491] Oliver J. Hasimuna, Sahya Maulu, and Joseph Mphande. Aquaculture health management practices in Zambia: Status, challenges and proposed biosecurity measures. Journal of Aquaculture Research & Development, 11(4):??, ???? 2020. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
aquaculture-health-management-practices-in-zambia-status-challenges-
and-proposed-biosecurity-measures.pdf.

Chellappan:2020:DAT

[492] Ajan Chellappan, Praba Thangamani, Shyni Markose, Selvaraj Thangaswamy, Uma Ganapathi, Citarasu Thavasimuthu, and Michael Babu Mariavincent. Development of alternative technology for the long-term storage of microalgal stock culture in fish hatcheries. Journal of Aquaculture Research & Development, 11(4):??, ???? 2020. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
development-of-alternative-technology-for-the-longterm-storage-
of-microalgal-stock-culture-in-fish-hatcheries.pdf.

Kalima:2020:GPM

[493] Sauti Kalima, W. L. Jere, and Alexander Shula Kefi. Growth performance of monosex and mixed sex of Oreochromis tanganicae (Günther, 1894) raised in semi concreate ponds. Journal of Aquaculture Research & Development, 11(4):??, ???? 2020. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/growth-performance-of-monosex-and-mixed-sex-
of-oreochromis-tanganicae-gnther-1894-raised-in-semi-concreate-
ponds.pdf.

Tzohari:2020:IUP

[494] Nir Tzohari. The importance of using a precise and intelligence data analyzing in the aquaculture industry before and after the COVID19. Journal of Aquaculture Research & Development, 11(4):??, ???? 2020. CODEN JARDCR. ISSN 2155-9546. URL https://www.

walshmedicalmedia.com/open-access/the-importance-of-using-a-
precise-and-intelligence-data-analyzing-in-the-aquaculture-industry-
before-and-after-the-covid.pdf.

Saleh:2020:EPT

[495] Hamed H. E. Saleh. Effect of partial or total replacement of fish meal by plant protein sources on productive performance of gilthead sea bream (Sparus aurata L.) fry under the water environment of Qaroun Lake. Journal of Aquaculture Research & Development, 11(5): ??, ???? 2020. CODEN JARDCR. ISSN 2155-9546. URL https:/

/www.walshmedicalmedia.com/open-access/effect-of-partial-or-

total-replacement-of-fish-meal-by-plant-protein--sources-on-
productive-performance-of-gilthead-sea-.pdf.

Shachi:2020:QPD

[496] Kumari Shachi, Sanjeev Kumar, N. K. Dubey, and Usha Dubey. Qualitative plankton diversity of two carp culture ageing ponds of LNMU Campus Darbhanga, India. Journal of Aquaculture Research & Development, 11 (5):??, ???? 2020. CODEN JARDCR. ISSN 2155-9546. URL https:/

/www.walshmedicalmedia.com/open-access/qualitative-plankton-
diversity-of-two-carp-culture-ageing-ponds-of--lnmu-campus-darbhanga-
india.pdf.

Hirayama:2020:LTC

[497] Shin Hirayama, Shuichi Tashiro, Kohei Inoue, Kazuya Urata, Masafumi Iima, and Yasuyuki Ikegami. Long-term culture characteristics of sterile Ulva spp. (Chlorophyta). Journal of Aquaculture Research & Development, 11(5):??, ???? 2020. CODEN JARDCR.

ISSN 2155-9546. URL https://www.walshmedicalmedia.com/open-
access/longterm-culture-characteristics-of-sterile-ulva-spp-
chlorophyta-53743.html.

Hoseini:2020:SAD

[498] Zia Hoseini. Shrimp aquaculture’s development perspectives reviewing on improving strategies and management plans in Ecuadorian shrimp industry. Journal of Aquaculture Research & Development, 11(6):??, ???? 2020. CODEN JARDCR. ISSN 2155-9546. URL https://www.

walshmedicalmedia.com/open-access/shrimp-aquaculturersquos-development-
perspectives-reviewing-on-improving-strategies-and-management-
plans-in-ecuadorian-shrimp-indu-54385.html.

Heneash:2020:MWQ

[499] Ahmed M. M. Heneash and Ahmed E. Alprol. Monitoring of water quality and zooplankton community in presence of different dietary levels of commercial wood charcoal of red tilapia. Journal of Aquaculture Research & Development, 11(6):??, ???? 2020. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/monitoring-of-water-quality-and-zooplankton-
community-in-presence-of-different-dietary-levels-of-commercial-
wood-charcoal-of-red-t-54386.html.

Sunny:2020:ALS

[500] Atiqur Rahman Sunny, Kazi Mohammad Masum, Nusrat Islam, Mizanur Rahman, Arifur Rahman, Jahurul Islam, Saidur Rahman, Khandaker Jafor Ahmed, and Shamsul Haque Prodhan. Analyzing livelihood sustainability of climate vulnerable fishers: Insight from Bangladesh. Journal of Aquaculture Research & Development, 11(6): ??, ???? 2020. CODEN JARDCR. ISSN 2155-9546. URL https:/

/www.walshmedicalmedia.com/open-access/analyzing-livelihood-
sustainability-of-climate-vulnerable-fishers-insight-from-bangladesh-
54387.html.

Mehdi:2020:NMS

[501] Shahmoradi Mehdi, Dehvari Aslam, Motaghi Mohammad Mehdi, Hossein Rameshi, and Tadayon Mohammad Hassan. Novobiocin: The most sensitive antibiotic against Vibrionaceae isolated from Pinctada margaritifera (black-lipped pearl oyster) in the South of Iran. Journal of Aquaculture Research & Development, 11(7):1–4, ???? 2020. CODEN JARDCR. ISSN 2155-9546. URL https:

//www.walshmedicalmedia.com/open-access/novobiocin-the-most-
sensitive-antibiotic-against-vibrionaceae-isolated-from-pinctada-
margaritifera-blacklipped-pearl-oys.pdf.

Trigo:2020:NNM

[502] Juan E. Trigo and Meritxell Mondéjar. New natural method for the elimination of salmon farms parasite copepods. Journal of Aquaculture Research & Development, 11(7):??, ???? 2020. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
new-natural-method-for-the-elimination-of-salmon-farms-parasite-
copepods-55534.html.

Deomedes:2020:APP

[503] Lameck Deomedes. Assessment of pond productivity in African catfish farms. Journal of Aquaculture Research & Development, 11 (7):??, ???? 2020. CODEN JARDCR. ISSN 2155-9546. URL

https://www.walshmedicalmedia.com/open-access/assessment-of-
pond-productivity-in-african-catfish-farms-55535.html.

Setyaningrum:2020:CPA

[504] Ervina Wahyu Setyaningrum, Endang Dewi Masithah, Mega Yuniartik, Agustina Tri-Kusuma Dewi, and Maghdalena Putri Nugrahani. Comparison of plankton abundance, water condition, organism growth performance of vaname shrimp (Litopenaeus vannameiei) on intensive and extensive culture systems in Banyuwangi District. Journal of Aquaculture Research & Development, 11(7): ??, ???? 2020. CODEN JARDCR. ISSN 2155-9546. URL

https://www.walshmedicalmedia.com/open-access/comparison-of-

plankton-abundance-water-condition-organism-growth-performance-
of-vaname-shrimp-litopenaeus-vannameiei-on-.pdf.

Tagarao:2020:LWR

[505] Sheena M. Tagarao, Chennie L. Solania, Joycelyn C. Jumawan, Shirlamaine G. Masangcay, and Laurence B. Calagui. Length–weight relationship (LWR), gonadosomatic index (GSI) and fecundity of Johnius borneensis (Bleeker, 1850) from Lower Agusan River basin, Butuan City, Philippines. Journal of Aquaculture Research & Development, 11(7):??, ???? 2020. CODEN JARDCR. ISSN 2155-9546. URL https://www.

walshmedicalmedia.com/open-access/lengthweight-relationship-
lwr-gonadosomatic-index-gsi-and-fecundity-of-johnius-borneensis-
bleeker-1850-from-lower-agusan.pdf.

Harlina:2020:AAC

[506] Harlina Harlina, Andi Gusti Tantu, Rosmiati Rosmiati, and Kamarudin Kamaruddin. The application of asterales (Chromolaena odorata L.) leaves on the feed for disease prevention in tiger shrimp cultivation (Penaeus monodon Fabricius) in pond. Journal of Aquaculture Research & Development, 11(8):??, ???? 2020. CODEN JARDCR. ISSN 2155-

9546. URL https://www.walshmedicalmedia.com/open-access/the-
application-of-asterales-chromolaena-odorata-l-leaves-on-the-
feed-for-disease-prevention-in-tiger-shrimp-cultivation.pdf.

Malini:2020:PNE

[507] H. Malini, A. Kalarani, V. Vinodha, and R. M. Inbaraj. Presence of neurosteroids and expression of neuropeptides in the brain of Indian salmon, Eleutheronema tetradactylum. Journal of Aquaculture Research & Development, 11(8):??, ???? 2020. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/presence-of-neurosteroids-and-expression-of-
neuropeptides-in-the-brain-of-indian-salmon-eleutheronema-tetradactylum.
pdf.

Zanna:2020:ATC

[508] Babagana Zanna, Abba I. Tijjani, and Mohammed Musa. Aquaculture: a tool for community sustainable development under the background of cultural and economic consideration in Nigeria. Journal of Aquaculture Research & Development, 11(8):??, ???? 2020. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
aquaculture-a-tool-for-community-sustainable-development-under-
the-background-of-cultural-and-economic-consideration-in-.pdf.

Adeli:2020:CPF

[509] Afshin Adeli and Masoomeh Hassannejad. Consumption preferences of fishery products among the young people of the Iran (case study: Gorgan City). Journal of Aquaculture Research & Development, 11(8):??, ???? 2020. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
consumption-preferences-of-fishery-products-among-the-young-
people-of-the-iran-case-study-gorgan-city.pdf.

Reyes:2020:GPS

[510] Guillermo Galindo Reyes. Genotoxicity by PAHs in shrimp (Litopenaeus vannamei) and its impact on the aquaculture of two coastal ecosystems of the Gulf of California, Mexico. Journal of Aquaculture Research & Development, 11(8):1–8, ???? 2020. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
genotoxicity-by-pahs-in-shrimp-litopenaeus-vannamei-and-its-
impact-on-the-aquaculture-of-two-coastal-ecosystems-of-the-g.
pdf.

Aro:2020:LWR

[511] O. O. Aro, F. I. Adesoun, I. T. Omoniyi, W. O. Abdul, A. O. Bashir, O. A. Amosu, and O. Egunjobi. Length–weight relationship, condition factor and natural diet of Hyperopisus bebe occidentalis and Brycinus macrolepidotus in Ikere Gorge, South-West, Nigeria. Journal of Aquaculture Research & Development, 11(9):??, ???? 2020. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
lengthweight-relationship-condition-factor-and-natural-diet-
of-hyperopisus-bebe-occidentalis-and-brycinus-macrolepidotus-
in-ikere--58371.html.

Farhaduzzaman:2020:PSD

[512] A. M. Farhaduzzaman, Md. Abu Hanif, Md. Suzan Khan, Mahadi Hasan Osman, Md. Neamul Hasan Shovon, Md. Khalilur Rahman, and Shahida Binte Ahmed. Perfect stocking density ensures best production and economic returns in floating cage aquaculture system. Journal of Aquaculture Research & Development, 11(9):??, ???? 2020. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/perfect-stocking-density-ensures-best-production-
and-economic-returns-in-floating-cage-aquaculture-system.pdf.

Holeh:2020:IAM

[513] G. M. Holeh, J. O. Ochiewo, S. Tsuma, and D. O. Mirera. Impact of aquaculture and mariculture information dissemination to the local coastal communities in

Kenya. Journal of Aquaculture Research & Development, 11(9):??, ???? 2020. CODEN JARDCR. ISSN 2155-9546. URL https://www.

walshmedicalmedia.com/open-access/impact-of-aquaculture-and-
mariculture-information-dissemination-to-the-local-coastal-communities-
in-kenya-58375.html.

Khalil:2020:ESN

[514] Maryyum Khalil, Summia Perveen, Hafiz Muhammad Arslan-Amin, Iqra Anwar, and Mahrukh Butt. The effectiveness of silver nanoparticles for the treatment of abdominal dropsy in Labeo rohita. Journal of Aquaculture Research & Development, 11(9):??, ???? 2020. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
the-effectiveness-of-silver-nanoparticles-for-the-treatment-
of-abdominal-dropsy-in-labeo-rohita.pdf.

Ahamed:2020:SEL

[515] G. S. Ahamed, M. T. Alam, M. A. Hussain, and S. Sultana. Socio economics and livelihoods of beel fishermen: Cases from North-Western Bangladesh. Journal of Aquaculture Research & Development, 11(9): 1–10, ???? 2020. CODEN JARDCR. ISSN 2155-9546. URL https:

//www.walshmedicalmedia.com/open-access/socio-economics-and-
livelihoods-of-beel-fishermen-cases-from-northwestern-bangladesh.
pdf.

Thiruvengadam:2020:DSW

[516] Vaitheeswaran Thiruvengadam. Deep sea water decapod, Munidopsis species (Crustacea: Decapoda: Munidopsidae) off Thoothukudi Coast of Gulf of Mannar collected in 2019–2020: Primary documentation of 53 species, India (08^{o} 31.912’N 78^{o} 25.327’E) (318 M). Journal of Aquaculture Research & Development, 11(10):1–3, ???? 2020. CODEN JARDCR.

ISSN 2155-9546. URL https://www.walshmedicalmedia.com/open-
access/deep-sea-water-decapod-munidopsis-species-crustacea-decapoda-
munidopsidae-off-thoothukudi-coast-of-gulf-of-mannar-collec.
pdf.

Prithi:2020:SEF

[517] C. Prithi, V. Vinodha, A. Kalarani, and R. M. Inbaraj. Study of endocrine factors influencing the reproductive stages of Indian salmon, Eleutheronema tetradactylum. Journal of Aquaculture Research & Development, 11(10):??, ???? 2020. CODEN JARDCR. ISSN 2155-9546.

URL https://www.walshmedicalmedia.com/open-access/study-of-
endocrine-factors-influencing-the-reproductive-stages-of-indian-
salmon-eleutheronema-tetradactylum.pdf.

Egbal:2020:PRP

[518] O. A. Egbal, B. A. El-Nouman, B. M. Siham, and M. A. Mona. Partial replacement of plant protein by fish meal in the diets of Nile tilapia, Oreochromis niloticus (Linnaeus, 1758). Journal of Aquaculture Research & Development, 11(10):??, ???? 2020. CODEN JARDCR.

ISSN 2155-9546. URL https://www.walshmedicalmedia.com/open-
access/partial-replacement-of-plant-protein-by-fish-meal-in-
the-diets-of-nile-tilapia-oreochromis-niloticus-linnaeus-1758.
pdf.

Bedassa:2020:DSH

[519] Wakessa Miheretu Bedassa. Determination of some heavy metals in Nile tilapia (Oreochromis niloticus) fish species in the water of Lake Hawassa: Review. Journal of Aquaculture Research & Development, 11 (10):??, ???? 2020. CODEN JARDCR. ISSN 2155-9546. URL https://

www.walshmedicalmedia.com/open-access/determination-of-some-
heavy-metals-in-nile-tilapia-oreochromis-niloticus-fish-species-
in-the-water-of-lake-hawassa-revie.pdf.

Adebayo:2020:GPS

[520] I. A. Adebayo, B. W. Obe, and T. Jegede. Growth performance and survivability of Gymnarchus niloticus (Cuvier, 1892) juveniles fed both live and formulated diets in earthen ponds. Journal of Aquaculture Research & Development, 11(11):??, ???? 2020. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
growth-performance-and-survivability-of-gymnarchus-niloticus-
cuvier-1892-juveniles-fed-both-live-and-formulated-diets-in.
pdf.

Kurcheti:2020:SNA

[521] Pani Prasad Kurcheti, M. Dhayanath, Abisha Juliet Mary, and Hina Alim. Silver nanoparticles (AgNPs) — potential applications in aquatic animal health management. Journal of Aquaculture Research & Development, 11 (11):??, ???? 2020. CODEN JARDCR. ISSN 2155-9546.

Kurcheti:2020:MAP

[522] Pani Prasad Kurcheti, Abisha Juliet Mary, and M. Dhayanath. Marine antimicrobial peptides (AMPs) — a promising natural alternative to antibiotics. Journal of Aquaculture Research & Development, 11(11):??, ???? 2020. CODEN JARDCR. ISSN 2155-

9546. URL https://www.walshmedicalmedia.com/abstract/marine-
antimicrobial-peptides-amps-a-promising-natural-alternative-
to-antibiotics-60127.html.

Holeh:2020:SES

[523] G. M. Holeh, E. W. Magondu, J. M. Njiru, S. Tsuma, A. Salim, A. M. Muriuki, A. Fulanda, J. Kilonzo, O. Ochola, S. Ndirangu, M. S. Zamu, G. Athoni, and J. Luyesi. Social economic survey and feasibility study to initiate cage fish farming in Kenyan coastal creeks. Journal of Aquaculture Research & Development, 11(11):??, ???? 2020. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
social-economic-survey-and-feasibility-study-to-initiate-cage-
fish-farming-in-kenyan-coastal-creeks.pdf.

Ullah:2020:CDA

[524] Md. Akram Ullah, Mohammad Abu Naeem, Amir Hossain, Abdulla-Al-Asif, and Md. Robiul Hasan. Categorization and distribution of aqua-chemicals used in coastal farming of south-eastern part of Bangladesh. Journal of Aquaculture Research & Development, 11(11): ??, ???? 2020. CODEN JARDCR. ISSN 2155-9546. URL https://www.

walshmedicalmedia.com/open-access/categorization-and-distribution-
of-aquachemicals-used-in-coastal-farming-of-southeastern-part-
of-bangladesh.pdf.

Sivakumar:2020:BDW

[525] J. Sivakumar, P. Harinath Reddy, and S. Surya Bhaskar Rao. Basic deviations of water quality parameters of shrimp (L. vannamei) culture ponds at Kongodu, Mogalipalem, and Gorripudi areas of East Godavari District, Andhra Pradesh, India. Journal of Aquaculture Research & Development, 11(12):??, ???? 2020. CODEN JARDCR. ISSN 2155-9546. URL https:

//www.walshmedicalmedia.com/open-access/basic-deviations-of-
water-quality-parameters-of-shrimp-l-vannamei-culture-ponds-
at-kongodu-mogalipalem-and-gorripudi-are.pdf.

Al-Hisnawi:2021:DSB

[526] Ali A. Abid Al-Hisnawi and Doaa Ali Beiwi. Dietary supplementation of Bacillus subtilis as probiotic affects haemato-immunological parameters of common carp (Cyprinus carpio). Journal of Aquaculture Research & Development, 12(1):??, ???? 2021. CODEN JARDCR. ISSN 2155-9546.

URL https://www.walshmedicalmedia.com/open-access/dietary-
supplementation-of-bacillus-subtilis-as-probiotic-affects-haematoimmunological-
parameters-of-common-carp-cyprinu.pdf.

Sime:2021:SPB

[527] Berhanu Sime and Shimelis Tesfaye. Stability performance of bread wheat (Triticum aestivum L.) genotype for yield and yield components in Oromia, Ethiopia. Journal of Aquaculture Research & Development, 12(1):1–8, ???? 2021. CODEN JARDCR. ISSN 2155-9546.

URL https://www.walshmedicalmedia.com/open-access/stability-
performance-of-bread-wheat-triticum-aestivum-l-genotype-for-
yield-and-yield-components-in-oromia-ethiopia.pdf.

Athman:2021:FCS

[528] S. H. Athman, G. M. Hole, C. Magori, S. Ndirangu, and M. S. Zamu. Fish cage site selection at Kibuyuni in Kwale County, Kenya: Tidal variations, waves height, current speed and direction status. Journal of Aquaculture Research & Development, 12(1):??, ???? 2021. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
fish-cage-site-selection-at-kibuyuni-in-kwale-county-kenya--
tidal-variations-waves-height-current-speed-and-direction-st.
pdf.

Rahman:2021:EPBa

[529] Ashikur Rahman, Shoaibe Hossain Talukder Shefat, Mohammed Anas Chowdhury, and Saif Uddin Khan. Effects of probiotic bacillus on growth performance, immune response and disease resistance in aquaculture. Journal of Aquaculture Research & Development, 12(1):1–6, ???? 2021. CODEN JARDCR. ISSN 2155-9546.

URL https://www.walshmedicalmedia.com/open-access/effects-
of-probiotic-bacillus-on-growth-performance-immune-response-
and-disease-resistance-in-aquaculture.pdf.

Sine:2021:SSF

[530] Sehanu Sine. Small-scale fisheries will provide sustainable benefits to the fishing communities in Sri Lanka? Journal of Aquaculture Research & Development, 12(2):??, ???? 2021. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
smallscale-fisheries-will-provide-sustainable-benefits-to-the-
fishing-communities-in-sri-lanka.pdf.

Abu:2021:DAA

[531] Akram Abu. Drug administration in aquaculture. Journal of Aquaculture Research & Development, 12(2):??, ???? 2021. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
drug-administration-in-aquaculture.pdf.

Wanchal:2021:AAS

[532] Suplang Wanchal. Aquaculture applications in seafood industry. Journal of Aquaculture Research & Development, 12(2):??, ???? 2021. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/aquaculture-applications-in-seafood-industry.
pdf.

Tilg:2021:IPM

[533] Frank Tilg. Importance of plasmids in microbiology. Journal of Aquaculture Research & Development, 12(2):??, ???? 2021. CODEN JARDCR.

ISSN 2155-9546. URL https://www.walshmedicalmedia.com/open-
access/importance-of-plasmids-in-microbiology.pdf.

Lui:2021:BEA

[534] Xin Lui. Brief editorial about fish health and growth. Journal of Aquaculture Research & Development, 12(2):??, ???? 2021. CODEN JARDCR.

ISSN 2155-9546. URL https://www.walshmedicalmedia.com/open-
access/brief-editorial-about-fish-health-and-growth.pdf.

Anirudhan:2021:ADC

[535] Anupa Anirudhan, Okomoda Victor Tosin, Mohd Effendy Wahid, and Yeong Yik Sung. Alternative disease control methods in shrimp aquaculture: a review. Journal of Aquaculture Research & Development, 12 (3):??, ???? 2021. CODEN JARDCR. ISSN 2155-9546. URL https:

//www.walshmedicalmedia.com/open-access/alternative-disease-
control-methods-in-shrimp-aquaculture-a-reviews.pdf.

Delgado:2021:SRC

[536] E. Delgado, D. Valles-Rosales, G. A. Pámanes-Carrasco, P. Cooke, N. C. Flores, and D. Reyes-Jáquez. Structural, rheological and calorimetric properties of an extruded shrimp feed using glandless cottonseed meal as a protein source. Journal of Aquaculture Research & Development, 12(3):1–6, ???? 2021. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
structural-rheological-and-calorimetric-properties-of-an-extruded-
shrimp-feed-using-glandless-cottonseed-meal-as-a-prote.pdf.

Workagegn:2021:ASIa

[537] Kassaye Belkew Workagegn. Aeromonas septicemia infection in cultured Nile tilapia, Oreochromis niloticus L. Journal of Aquaculture Research & Development, 12(3):1–5, ???? 2021. CODEN JARDCR. ISSN 2155-9546.

URL https://www.walshmedicalmedia.com/open-access/aeromonas-
septicemia-infection-in-cultured-nile-tilapia-oreochromis-niloticus-
l.pdf.

Mishra:2021:AWT

[538] Bimlendu Kumar Mishra, Monowar Alam Khalid, and Shyam Narayan Labh. Assessment of water temperature on growth performance and protein profile of rainbow trout Oncorhynchus mykiss (Walbaum, 1792). Journal of Aquaculture Research & Development, 12 (3):??, ???? 2021. CODEN JARDCR. ISSN 2155-9546. URL

https://www.walshmedicalmedia.com/open-access/assessment-of-
water-temperature-on-growth-performance-and-protein-profile-
of-rainbow-trout-oncorhynchus-my-kiss-walbaum-.pdf.

Ajit:2021:FND

[539] Sri Lakshmi Ajit. Fish is a natural diabetic animal why? Journal of Aquaculture Research & Development, 12(3):??, ???? 2021. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/fish-is-a-natural-diabetic-animal-why.pdf.

Mathew:2021:IPS

[540] Anu Mathew. Identification of protease synthesizing Micrococcus yunnanensis from the gut of mud crab Scylla serrata. Journal of Aquaculture Research & Development, 12(4):??, ???? 2021. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
identification-of-protease-synthesizing-micrococcus-yunnanensis-
from-the-gutof-mud-crab-scylla-serrata.pdf.

Ong:2021:BCP

[541] Rachel Ong. Breeding crayfish in paddy fields: a BIG economic boost for farmers. Journal of Aquaculture Research & Development, 12(4): ??, ???? 2021. CODEN JARDCR. ISSN 2155-9546. URL https:/

/www.walshmedicalmedia.com/open-access/breeding-crayfish-in-
paddy-fields-a-big-economic-boost-for-farmers.pdf.

Sunny:2021:ULC

[542] Atiqur Rahman Sunny, Shamsul Haque Prodhan, Md. Ashrafuzzaman, Golam Shakil Ahamed, Sharif Ahmed Sazzad, Mahmudul Hasan Mithun, K. M. Nadim Haider, and Md Tariqul Alam. Understanding livelihood characteristics and vulnerabilities of small-scale fishers in Coastal Bangladesh. Journal of Aquaculture Research & Development, 12(4):??, ???? 2021. CODEN JARDCR. ISSN 2155-9546. URL

https://www.walshmedicalmedia.com/open-access/understanding-
livelihood-characteristics-and-vulnerabilities-of-smallscale-
fishers-in-coastal-bangladesh.pdf.

Adi:2021:EFU

[543] Bayu Prakoso Adi. The efficiency of feed utilization, growth, and survival rate performance of saline tilapia (Oreochromisniloticus) by the additional of papain enzyme in artificial feed. Journal of Aquaculture Research & Development, 12(4):??, ???? 2021. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
the-efficiency-of-feed-utilization-growth-and-survival-rate-
performance-ofsaline-tilapia-oreochromisniloticus-by-the-add.
pdf.

Workagegn:2021:ASIb

[544] Kassaye Balkew Workagegn, Bekele Lema, P. Natarajan, and L. Prabadevi. Aeromonas spp. infection in farmed Nile tilapia, Oreochromis niloticus. Journal of Aquaculture Research & Development, 12(4):1– 6, ???? 2021. CODEN JARDCR. ISSN 2155-9546. URL

https://www.walshmedicalmedia.com/open-access/aeromonas-spp-
infection-in-farmed-nile-tilapia-oreochromis-niloticus.pdf.

Ajit:2021:GAD

[545] Sri Lakshmi Ajit. Genomics and aquaculture developments overview. Journal of Aquaculture Research & Development, 12(5):??, ???? 2021.

CODEN JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/genomics-and-aquaculture-developments-overview.
pdf.

Ajit:2021:PEM

[546] Sri Lakshmi Ajit. Plastic and its effect on marine environment. Journal of Aquaculture Research & Development, 12(5):??, ???? 2021. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/plastic-and-its-effect-on-marine-environment.
pdf.

Ajit:2021:MPA

[547] Sri Lakshmi Ajit. Marine protected area and fisheries. Journal of Aquaculture Research & Development, 12(5):??, ???? 2021. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/marine-protected-area-and-fisheries.pdf.

Ajit:2021:RFI

[548] Sri Lakshmi Ajit. The role of fish in improving nutrition and health outcomes. Journal of Aquaculture Research & Development, 12(5): ??, ???? 2021. CODEN JARDCR. ISSN 2155-9546. URL https:

//www.walshmedicalmedia.com/open-access/the-role-of-fish-in-
improving-nutrition-and-health-outcomes.pdf.

Kiwewa:2021:PSL

[549] Bernard Kiwewa, Rosemary Nalwanga, and Charles K. Twesigye. The performance of small and large leaf plants in aquaponics. Journal of Aquaculture Research & Development, 12(5):??, ???? 2021. CODEN JARDCR.

ISSN 2155-9546. URL https://www.walshmedicalmedia.com/open-
access/the-performance-of-small-and-large-leaf-plants-in-aquaponics.
pdf.

Hernandez:2021:CIS

[550] Luis Remedios Hernandez. The controverted issue of small-scale rural aquaculture in the world invitation to debate. Journal of Aquaculture Research & Development, 12(5):??, ???? 2021. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
the-controverted-issue-of-smallscale-rural-aquaculture-in-the-
world-invitation-to-debate.pdf.

Rahman:2021:EPBb

[551] Ashikur Rahman, Shoaibe Hossain Talukder Shefat, Mohammed Anas Chowdhury, and Saif Uddin Khan. Effects of probiotic bacillus on fish growth performance, immune response and disease resistance in aquaculture. Journal of Aquaculture Research & Development, 12(6):1–10, ???? 2021. CODEN JARDCR. ISSN 2155-9546. URL https:/

/www.walshmedicalmedia.com/open-access/effects-of-probiotic-
bacillus-on-fish-growth-performance-immune-response-and-disease-
resistance-in-aquaculture-80131.html.

Chari:2021:DZS

[552] T. Jagadeeshwara Chari, B. Srishylam, and A. V. Rajashekhar. Diversity of zooplankton in Shanigaram Reservoir, Siddipet District, Telangana, India. Journal of Aquaculture Research & Development, 12(6):??, ???? 2021. CODEN JARDCR. ISSN 2155-9546.

URL https://www.walshmedicalmedia.com/open-access/diversity-

of-zooplankton-in-shanigaram-reservoir-siddipet-district-telangana-
india-80134.html.

Gutema:2021:GRN

[553] B. Gutema, A. Yimer, and F. Hailemichael. Growth response of Nile tilapia (Oreochromis niloticus) fries fed on silage-based diets. Journal of Aquaculture Research & Development, 12(6):??, ???? 2021. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/growth-response-of-nile-tilapia-emoreochromis-
niloticusem-fries-fed-on-silagebased-diets-80135.html.

Bernard:2021:PAT

[554] Kiweewa Bernard. Performance of aquaponics technologies in the Albertine Region, Western Uganda. Journal of Aquaculture Research & Development, 12(6):??, ???? 2021. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
performance-of-aquaponics-technologies-in-the-albertine-region-
western-uganda-80240.html.

Wijekoon:2021:EGT

[555] M. Wijekoon, C. C. Parrish, and A. Mansour. Effect of growth temperature on muscle lipid class and fatty acid composition in adult steelhead trout (Oncorhynchus mykiss) fed commercial diets with different ω6 to ω3 fatty acid ratios. Journal of Aquaculture Research & Development, 12(6):??, ???? 2021. CODEN JARDCR. ISSN 2155-9546.

URL https://www.walshmedicalmedia.com/open-access/effect-of-
growth-temperature-on-muscle-lipid-class-and-fatty-acid-composition-
in-adult-steelhead-trout-emoncorhynchus-mykissem-fed-80891.html.

Nilamsari:2021:GPS

[556] Kiki Wahyu Nilamsari, Johannes Hutabarat, Diana Rachmawati, and Bayu Prakoso Adi. Growth performance and survival rates of milkfish (Chanos chanos) fed diets supplemented with pineapple extract. Journal of Aquaculture Research & Development, 12(7):??, ???? 2021. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/growth-performance-and-survival-rates-of-milkfish-
emchanos-chanosem-fed-diets-supplemented-with-pineapple-extract-
82724.html.

Eshete:2021:EPR

[557] Teshome Belay Eshete, Kassaye Balkew Workagegn, and Natarajan Pavanasam. Effects of partial replacement of fishmeal by locally available ingredients on growth performance and feed utilization efficiency of Nile tilapia, Oreochromis niloticus. Journal of Aquaculture Research & Development, 12(7):??, ???? 2021. CODEN JARDCR. ISSN 2155-9546.

URL https://www.walshmedicalmedia.com/open-access/effects-
of-partial-replacement-of-fishmeal-by-locally-available-ingredients-
on-growth-performance-and-feed-utilization-efficiency--82428.
html.

KedirJema:2021:GYR

[558] KedirJema and Mohammed Aliyi. Growth and yield response of food barley (Hordeum vulgare L.) varieties to nitrogen fertilizer rates Inkofele District, Southeastern Oromia. Journal of Aquaculture Research & Development, 12(7):??, ???? 2021. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
growth-and-yield-response-of-food-barley-emhordeum-vulgareem-
l-varieties-to-nitrogen-fertilizer-rates-inkofele-district-southeaste-
82429.html.

Choudhary:2021:MPV

[559] Sanju Choudhary and Shailesh Godika. Morphological and pathogenic variability in different isolates of Fusarium oxysporum f. sp. cumini. Journal of Aquaculture Research & Development, 12(7):??, ???? 2021. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/morphological-and-pathogenic-variability-in-
different-isolates-of-emfusarium-oxysporumem-f-sp-emcuminiem-
82430.html.

LizanneAC:2021:FDU

[560] Maria Lizanne AC. Formulation of diet using conventional and non-conventional protein sources. Journal of Aquaculture Research & Development, 12(7):??, ???? 2021. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
formulation-of-diet-using-conventional-and-nonconventional-protein-
sources-82431.html.

vatika:2021:UBW

[561] Sushma vatika. Use of brackish water. Journal of Aquaculture Research & Development, 12(8):??, ???? 2021. CODEN JARDCR. ISSN 2155-

9546. URL https://www.walshmedicalmedia.com/open-access/use-
of-brackish-water-83972.html.

Bekele:2021:HHC

[562] Begonesh Bekele, P. Natarajan, Kassaye Balkew Workagegn, and Devika Pillai. Hematological and histopathological changes in artificially infected Nile tilapia with vibrio species. Journal of Aquaculture Research & Development, 12(8):??, ???? 2021. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
hematological-and-histopathological-changes-in-artificially-
infected-nile-tilapia-with-emvibrioem-species-83766.html.

Ajay:2021:DCS

[563] V. S. Ajay and Amrutha Krishnan R. Design characteristics and specifications of gill net operated along the lower stretches of Vembanad Wetlands, Kerala, India. Journal of Aquaculture Research & Development, 12(8):??, ???? 2021. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
design-characteristics-and-specifications-of-gill-net-operated-
along-the-lower-stretches-of-vembanad-wetlands-kerala-india-
83767.html.

Tuluka:2021:DFM

[564] Tekalign Tuluka, Ketema Bekele, and Kumilachew Alamerie. Determinants of fish market supply in the case of Lake Hawassa, Sidama National Regional State, Ethiopia. Journal of Aquaculture Research & Development, 12(8):??, ???? 2021. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
determinants-of-fish-market-supply-in-the-case-of-lake-hawassa-
sidama-national-regional-state-ethiopia-83768.html.

Akinloye:2021:HAS

[565] Ogundiran Mathew Akinloye, Fawole Olubanjo Olatunde, and Ajala Olumuyiwa Olasunmibo. Histological aberrations in selected organs of two Clariids species obtained from Asa River, Nigeria. Journal of Aquaculture Research & Development, 12(8):??, ???? 2021. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
evaluation-of-health-status-of-emheterobranchus-longifilisem-
and-emclarias-buthupogonem-obtained-from-asa-river-nigeria-using-
hist-83959.html.

Kotha:2021:NRS

[566] Rakshitha Kotha. Nursery rearing and seed stocking of silver pompano. Journal of Aquaculture Research & Development, 12(9):??, ???? 2021. CODEN JARDCR. ISSN 2155-9546. URL https:

//www.walshmedicalmedia.com/open-access/nursery-rearing-and-
seed-stocking-of-silver-pompano-85209.html.

Kotha:2021:ENP

[567] Rakshitha Kotha. Editorial note on photosynthetic organisms in the ocean. Journal of Aquaculture Research & Development, 12(9):??, ???? 2021. CODEN JARDCR. ISSN 2155-9546. URL https://www.

walshmedicalmedia.com/open-access/editorial-note-on-photosynthetic-
organisms-in-the-ocean-85211.html.

Kotha:2021:NSA

[568] Rakshitha Kotha. Nutrient supplementation to aquatic species and their metabolic growth. Journal of Aquaculture Research & Development, 12(9):??, ???? 2021. CODEN JARDCR. ISSN 2155-9546.

URL https://www.walshmedicalmedia.com/open-access/nutrient-
supplementation-to-aquatic-species-and-their-metabolic-growth-
85215.html.

Kotha:2021:CAH

[569] Rakshitha Kotha. Combination of aquaculture and hydroponics. Journal of Aquaculture Research & Development, 12(9):??, ???? 2021. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/combination-of-aquaculture-and-hydroponics-85216.
html.

Kotha:2021:CPH

[570] Rakshitha Kotha. Commercial phase of hatchery system. Journal of Aquaculture Research & Development, 12(9):??, ???? 2021. CODEN JARDCR.

ISSN 2155-9546. URL https://www.walshmedicalmedia.com/open-
access/commercial-phase-of-hatchery-system-85217.html.

Mathew:2021:ICS

[571] Anu Mathew, Imelda Joseph, and V. Susmitha. Isolation and characterization of siderophore and HCN producing Pseudomonas aeruginosa from Indian pompano Trachinotus mookalee Cuvier, 1832. Journal of Aquaculture Research & Development, 12(10):??, ???? 2021. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
isolation-and-characterization-of-siderophore-and-hcn-producing-
empseudomonas-aeruginosaem-from-indian-pompano-emtrachinotus-
mooka-87442.html.

Ronald:2021:EDP

[572] Komguep N. Ronald, Efole E. Thomas, Defang F. Henry, Nana T. Alrigent, Mube K. Herve, and Tagning Z. P. Davy. Effects of dietary protein level on growth performance of pre-growing catfish Clarias jaensis (Boulanger 1909) fingerlings in captivity. Journal of Aquaculture Research & Development, 12(10):??, ???? 2021. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
effects-of-dietary-protein-level-on-growth-performance-of-pregrowing-
catfish-emclarias-jaensisem-boulanger-1909-in-captivity-87444.
html.

Kotha:2021:FHM

[573] Rakshitha Kotha. Facultative, highly migratory activity of American eel. Journal of Aquaculture Research & Development, 12(10):??, ???? 2021. CODEN JARDCR. ISSN 2155-9546. URL https://www.

walshmedicalmedia.com/open-access/facultative-highly-migratory-
activity-of-american-eel-87521.html.

Kotha:2021:FWM

[574] Rakshitha Kotha. Food web in marine life. Journal of Aquaculture Research & Development, 12(10):??, ???? 2021. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
food-web-in-marine-life-87525.html.

Kotha:2021:FFW

[575] Rakshitha Kotha. Flora and fauna of water bodies. Journal of Aquaculture Research & Development, 12(10):??, ???? 2021. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
flora-and-fauna-of-water-bodies-87533.html.

Kotha:2021:LSH

[576] Rakshitha Kotha. The life of a sea horse. Journal of Aquaculture Research & Development, 12(11):??, ???? 2021. CODEN JARDCR. ISSN 2155-9546.

Kotha:2021:SSU

[577] Rakshitha Kotha. The stages of a sea urchin’s existence. Journal of Aquaculture Research & Development, 12(11):??, ???? 2021. CODEN JARDCR. ISSN 2155-9546.

Kotha:2021:IWB

[578] Rakshitha Kotha. Inland water bodies and the organisms that live there. Journal of Aquaculture Research & Development, 12(11):??, ???? 2021. CODEN JARDCR. ISSN 2155-9546.

Kotha:2021:EGR

[579] Rakshitha Kotha. An editorial on gill respiration in sea life. Journal of Aquaculture Research & Development, 12(11):??, ???? 2021. CODEN JARDCR. ISSN 2155-9546.

Kotha:2021:AOS

[580] Rakshitha Kotha. Aquatic organisms that survive in the temperature. Journal of Aquaculture Research & Development, 12(12):??, ???? 2021. CODEN JARDCR. ISSN 2155-9546. URL https://www.

walshmedicalmedia.com/open-access/aquatic-organisms-that-survive-
in-the-temperature-109271.html.

Kotha:2021:FNA

[581] Rakshitha Kotha. Food and nutrition in aquatic species. Journal of Aquaculture Research & Development, 12(12):??, ???? 2021. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/food-and-nutrition-in-aquatic-species-109272.
html.

Kotha:2021:RPA

[582] Rakshitha Kotha. A report on pathology in aquatic organisms. Journal of Aquaculture Research & Development, 12(12):??, ???? 2021. CODEN JARDCR. ISSN 2155-9546. URL https://

www.walshmedicalmedia.com/open-access/a-report-on-pathology-
in-aquatic-organisms-109273.html.

Kotha:2021:ENF

[583] Rakshitha Kotha. An editorial note on fish farming. Journal of Aquaculture Research & Development, 12(12):??, ???? 2021. CODEN JARDCR.

ISSN 2155-9546. URL https://www.walshmedicalmedia.com/open-
access/an-editorial-note-on-fish-farming-109274.html.

Kotha:2021:ERS

[584] Rakshitha Kotha. Editorial on recreational or sport fishing. Journal of Aquaculture Research & Development, 12(12):??, ???? 2021. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/editorial-on-recreational-or-sport-fishing-109275.
html.

Yang:2022:FSP

[585] Xiaowei Yang. Fish spoilage and preservation techniques. Journal of Aquaculture Research & Development, 13(1):??, ???? 2022. CODEN JARDCR.

ISSN 2155-9546. URL https://www.walshmedicalmedia.com/open-
access/fish-spoilage-and-preservation-techniques-109795.html.

Benneth:2022:BNS

[586] Andrew Benneth. Brief note on seagrass biology. Journal of Aquaculture Research & Development, 13(1):??, ???? 2022. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
brief-note-on-seagrass-biology-109796.html.

Ratan:2022:OSF

[587] Kanti Ratan. Overview on shrimp framing. Journal of Aquaculture Research & Development, 13(1):??, ???? 2022. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
overview-on-shrimp-framing-109798.html.

Stephen:2022:DNH

[588] Carrie Stephen. Detailed note on health benefits of salmon fish. Journal of Aquaculture Research & Development, 13(1):??, ???? 2022. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/detailed-note-on-health-benefits-of-salmon-fish-
109800.html.

James:2022:DNP

[589] Mathew James. Detailed note on plankton. Journal of Aquaculture Research & Development, 13(1):??, ???? 2022. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
detailed-note-on-plankton-110060.html.

Ayari:2022:SAE

[590] Tahani El Ayari, Lazhar Mhadhbi, and Najoua Trigui El Menif. Abdominal swelling and skeletal deformity in gilthead sea bream (Sparus aurata) and European sea bass, (Dicentrarchus labrax) developed under rearing conditions. Journal of Aquaculture Research & Development, 13(2):??, ???? 2022. CODEN JARDCR. ISSN 2155-9546.

URL https://www.walshmedicalmedia.com/open-access/abdominal-
swelling-and-skeletal-deformity-in-gilthead-sea-bream-emsparus-
aurataem-and-european-sea-bass-emdicentrarchus-labraxem-d-110167.
html.

Liu:2022:MEM

[591] Chongguo Liu. Major effects of marine pollution. Journal of Aquaculture Research & Development, 13(2):??, ???? 2022. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
major-effects-of-marine-pollution-110200.html.

Paul:2022:IAF

[592] Sourav Paul. Important advantages of fish hilsa. Journal of Aquaculture Research & Development, 13(2):??, ???? 2022. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
important-advantages-of-fish-hilsa-110250.html.

Zhou:2022:DNA

[593] Xiulin Zhou. Detailed note on algal bloom. Journal of Aquaculture Research & Development, 13(2):??, ???? 2022. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
detailed-note-on-algal-bloom-110251.html.

Gebru:2022:NBE

[594] Teklay Gebru. Evaluation of productivity and profitability of Nile tilapia (O. niloticus) with beetroot (Beta vulgaris) and carrot (Daucus carota). Journal of Aquaculture Research & Development, 13 (2):??, ???? 2022. CODEN JARDCR. ISSN 2155-9546. URL

https://www.walshmedicalmedia.com/open-access/evaluation-of-
productivity-and-profitability-of-nile-tilapia-emo-niloticusem-
with-beetroot-embeta-vulgarisem-and-carrot-emdaucus-c-110252.
html.

Racine:2022:RAP

[595] Grant Racine. Role of aquatic plants in maintaining the biota in water bodies. Journal of Aquaculture Research & Development, 13(3):??, ???? 2022. CODEN JARDCR. ISSN 2155-9546. URL https://www.

walshmedicalmedia.com/open-access/role-of-aquatic-plants-in-
maintaining-the-biota-in-water-bodies-111289.html.

Campbell:2022:SSI

[596] Joshua Campbell. Significance of seaweeds in improving the aquatic and terrestrial environment. Journal of Aquaculture Research & Development, 13(3):??, ???? 2022. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
significance-of-seaweeds-in-improving-the-aquatic-and-terrestrial-
environment-111291.html.

Boxman:2022:RPG

[597] Kevan Boxman. The role of pisciculture in the growth of economy. Journal of Aquaculture Research & Development, 13(3):??, ???? 2022. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/the-role-of-pisciculture-in-the-growth-of-economy-
111292.html.

Langdon:2022:IIM

[598] Frank Langdon. Importance of integrated multi tropic aquaculture. Journal of Aquaculture Research & Development, 13(3):??, ???? 2022. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.

com/open-access/importance-of-integrated-multi-tropic-aquaculture-
111293.html.

Alam:2022:IAF

[599] Henry Alam. Importance of aquariums and fish ponds in aquaculture. Journal of Aquaculture Research & Development, 13(3): ??, ???? 2022. CODEN JARDCR. ISSN 2155-9546. URL

https://www.walshmedicalmedia.com/open-access/importance-of-
aquariums-and-fish-ponds-in-aquaculture-111294.html.

Thomas:2022:ECA

[600] Kieliszek Thomas. Environmental conditions that affect the growth of aquaculture species. Journal of Aquaculture Research & Development, 13(4):??, ???? 2022. CODEN JARDCR. ISSN 2155-9546. URL

https://www.walshmedicalmedia.com/open-access/environmental-
conditions-that-affect-the-growth-of-aquaculture-species-111941.
html.

Huston:2022:EFA

[601] Arthur Huston. Efficacy of freshwater aquaculture in India and future goals. Journal of Aquaculture Research & Development, 13(4):??, ???? 2022. CODEN JARDCR. ISSN 2155-9546. URL https://www.

walshmedicalmedia.com/open-access/efficacy-of-freshwater-aquaculture-
in-india-and-future-goals-111942.html.

Paul:2022:IVM

[602] David Paul. Implications in veterinary medicine and welfare and its effective therapy. Journal of Aquaculture Research & Development, 13(4): ??, ???? 2022. CODEN JARDCR. ISSN 2155-9546. URL https://www.

walshmedicalmedia.com/open-access/implications-in-veterinary-
medicine-and-welfare-and-its-effective-therapy-111943.html.

Mignani:2022:RVM

[603] Helena Mignani. Role of veterinary medicine in both humans and animals. Journal of Aquaculture Research & Development, 13(4):??, ???? 2022. CODEN JARDCR. ISSN 2155-9546. URL https://www.

walshmedicalmedia.com/open-access/role-of-veterinary-medicine-
in-both-humans-and-animals-111944.html.

Mignani:2022:ADN

[604] Helena Mignani. Assessment of dietary and nutritional requirements in aquaculture. Journal of Aquaculture Research & Development, 13(4):??, ???? 2022. CODEN JARDCR. ISSN 2155-9546. URL https://www.

walshmedicalmedia.com/open-access/assessment-of-dietary-and-
nutritional-requirements-in-aquaculture-111945.html.

Jung-Hoon:2022:ERC

[605] Jung-Hoon. Environmentally relevant concentrations of human exposure to the arsenic anhydride and manifested in the proteomics technology. Journal of Aquaculture Research & Development, 13(5):??, ???? 2022. CODEN JARDCR. ISSN 2155-9546. URL https://www.

walshmedicalmedia.com/open-access/environmentally-relevant-concentrations-
of-human-exposure-to-the-arsenic-anhydride-and-manifested-in-
the-proteomics-technology-112177.html.

Garlock:2022:EMT

[606] Paul Garlock. Evaluation of molecular techniques in aquaculture and a dynamic approach to bio analytical sensors. Journal of Aquaculture Research & Development, 13(5):??, ???? 2022. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
evaluation-of-molecular-techniques-in-aquaculture-and-a-dynamic-
approach-to-bio-analytical-sensors-112178.html.

Witzell:2022:SDF

[607] Frederick Witzell. Surveillance of delegate fish aquaculture in aquatic pathology and their ecological transpose. Journal of Aquaculture Research & Development, 13(5):??, ???? 2022. CODEN JARDCR.

ISSN 2155-9546. URL https://www.walshmedicalmedia.com/open-
access/surveillance-of-delegate-fish-aquaculture-in-aquatic-
pathology-and-their-ecological-transpose-112179.html.

Colombo:2022:PCA

[608] Juan Colombo. Prominence of cryobanking in aquatic ecosystems and aquaculture in reproductive course of action. Journal of Aquaculture Research & Development, 13(5):??, ???? 2022. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
prominence-of-cryobanking-in-aquatic-ecosystems-and-aquaculture-
in-reproductive-course-of-action-112180.html.

Canzian:2022:MUA

[609] Denis Canzian. Morphological and ultrastructural adaptations and simplification of methods of pharmaceutical research in zebrafish neurobiology. Journal of Aquaculture Research & Development, 13(5): ??, ???? 2022. CODEN JARDCR. ISSN 2155-9546. URL

https://www.walshmedicalmedia.com/open-access/morphological-
and-ultrastructural-adaptations-and-simplification-of-methods-
of-pharmaceutical-research-in-zebrafish-neurobiology-112181.
html.

Mathew:2022:FNF

[610] Stephen Mathew. Fish nutrition, feed types, and management of fish waste. Journal of Aquaculture Research & Development, 13(6):??, ???? 2022. CODEN JARDCR. ISSN 2155-9546. URL https://www.

walshmedicalmedia.com/open-access/fish-nutrition-feed-types-
and-management-of-fish-waste-113071.html.

Seemani:2022:AEU

[611] Bornwell Seemani. Analyzing the effects of using probiotics in aquaculture. Journal of Aquaculture Research & Development, 13(6):??, ???? 2022. CODEN JARDCR. ISSN 2155-9546. URL https://

www.walshmedicalmedia.com/open-access/analyzing-the-effects-
of-using-probiotics-in-aquaculture-113072.html.

Morris:2022:TEA

[612] Keryyn Morris. Toxic effects of ailments of the fish pathology and its risk of loss due to infectious diseases and breeds. Journal of Aquaculture Research & Development, 13(6):??, ???? 2022. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
toxic-effects-of-ailments-of-the-fish-pathology-and-its-risk-
of-loss-due-to-infectious-diseases-and-breeds-113073.html.

Shang:2022:MRI

[613] Shan Shang. Major role and importance of fish farming. Journal of Aquaculture Research & Development, 13(6):??, ???? 2022. CODEN JARDCR.

ISSN 2155-9546. URL https://www.walshmedicalmedia.com/open-
access/major-role-and-importance-of-fish-farming-113074.html.

Michele:2022:GOF

[614] Cristina Michele. A general overview of fisheries and aquaculture. Journal of Aquaculture Research & Development, 13(6):??, ???? 2022. CODEN JARDCR. ISSN 2155-9546. URL https://

www.walshmedicalmedia.com/open-access/a-general-overview-of-
fisheries-and-aquaculture-113076.html.

Stephan:2022:AAP

[615] Chris Stephan. Approaches of aquatic physiology, and monitoring the environmental changes in fish population. Journal of Aquaculture Research & Development, 13(7):??, ???? 2022. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
approaches-of-aquatic-physiology-and-monitoring-the-environmental-
changes-in-fish-population-113275.html.

Cabral:2022:AIO

[616] Haibin Cabral. Analysis and importance of oceanography and its various branches. Journal of Aquaculture Research & Development, 13(7):??, ???? 2022. CODEN JARDCR. ISSN 2155-9546.

URL https://www.walshmedicalmedia.com/open-access/analysis-
and-importance-of-oceanography-and-its-various-branches-113277.
html.

Eugeny:2022:ACS

[617] James Eugeny. Approaches of cryobiology and science of cryopreservation in reproductive medicine. Journal of Aquaculture Research & Development, 13(7):??, ???? 2022. CODEN JARDCR. ISSN 2155-9546. URL

https://www.walshmedicalmedia.com/open-access/approaches-of-
cryobiology-and-science-of-cryopreservation-in-reproductive-
medicine-113278.html.

Nordlund:2022:MRC

[618] Jorge Nordlund. Major role of coastal aquaculture and characteristics of various fields. Journal of Aquaculture Research & Development, 13(7): ??, ???? 2022. CODEN JARDCR. ISSN 2155-9546. URL https://

www.walshmedicalmedia.com/open-access/major-role-of-coastal-
aquaculture-and-characteristics-of-various-fields-113279.html.

Smith:2022:IFS

[619] Brook Smith. Insights of fisheries science, classification and its objectives. Journal of Aquaculture Research & Development, 13(7):??, ???? 2022. CODEN JARDCR. ISSN 2155-9546. URL https://

www.walshmedicalmedia.com/open-access/insights-of-fisheries-
science-classification-and-its-objectives-113280.html.

Edel:2022:EAN

[620] James Edel. Evaluation of aquatic nutrition of animal production and development of specific diet formulations. Journal of Aquaculture Research & Development, 13(8):??, ???? 2022. CODEN JARDCR.

ISSN 2155-9546. URL https://www.walshmedicalmedia.com/open-
access/evaluation-of-aquatic-nutrition-of-animal-production-
and-development-of-specific-diet-formulations-113857.html.

Goss:2022:MRA

[621] Samuel Goss. Major role of aquatic species and the water shed of aquatic species in an ecosystem. Journal of Aquaculture Research & Development, 13(8):??, ???? 2022. CODEN JARDCR. ISSN 2155-9546. URL

https://www.walshmedicalmedia.com/open-access/major-role-of-
aquatic-species-and-the-water-shed-of-aquatic-species-in-an-
ecosystem-113858.html.

Levy:2022:OPA

[622] Alvaro Levy. Oceanographic processes and approaches of various dimensions in coastal ocean system. Journal of Aquaculture Research & Development, 13(8):??, ???? 2022. CODEN JARDCR. ISSN 2155-9546. URL

https://www.walshmedicalmedia.com/open-access/oceanographic-
processes-and-approaches-of-various-dimensions-in-coastal-ocean-
system-113864.html.

Dias:2022:DCM

[623] Joes Dias. Development of conceptual modern fisheries and aquaculture in many parts of world highlights and its importance. Journal of Aquaculture Research & Development, 13(8):??, ???? 2022. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
development-of-conceptual-modern-fisheries-and-aquaculture-in-
many-parts-of-world-highlights-and-its-importance-113865.html.

Bela:2022:MRM

[624] Manthos Bela. Major role of marine microbiology and its various types of marine bacteria. Journal of Aquaculture Research & Development, 13(8): ??, ???? 2022. CODEN JARDCR. ISSN 2155-9546. URL https:/

/www.walshmedicalmedia.com/open-access/major-role-of-marine-
microbiology-and-its-various-types-of-marine-bacteria-113866.
html.

Jung-Hoon:2022:PAU

[625] Jung-Hoon. Proteomics approach for understanding arsenic toxicity in a fish mode. Journal of Aquaculture Research & Development, 13(9): ??, ???? 2022. CODEN JARDCR. ISSN 2155-9546. URL https:

//www.walshmedicalmedia.com/open-access/proteomics-approach-
for-understanding-arsenic-toxicity-in-a-fish-mode-115217.html.

Lynne:2022:OOE

[626] Yuzhu Lynne. An overview of oceanography and its Earth science. Journal of Aquaculture Research & Development, 13(9):??, ???? 2022. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/an-overview-of-oceanography-and-its-earth-science-
114963.html.

Meher:2022:MRP

[627] Lewis Meher. Major role of proteomics in aquaculture production and analysis of aquatic animal pathology. Journal of Aquaculture Research & Development, 13(9):??, ???? 2022. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
major-role-of-proteomics-in-aquaculture-production-and-analysis-
of-aquatic-animal-pathology-114965.html.

Macbeth:2022:EAE

[628] Peter Macbeth. Evaluation of aquaculture environmental interactions and aquaculture population growth in Chile. Journal of Aquaculture Research & Development, 13(9):??, ???? 2022. CODEN JARDCR.

ISSN 2155-9546. URL https://www.walshmedicalmedia.com/open-
access/evaluation-of-aquaculture-environmental-interactions-
and-aquaculture-population-growth-in-chile-114966.html.

Paul:2022:SOF

[629] David Paul. Significance of oyster farming in aquaculture. Journal of Aquaculture Research & Development, 13(9):??, ???? 2022. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/significance-of-oyster-farming-in-aquaculture-
114429.html.

Mignani:2022:SOF

[630] Helena Mignani. Significance of ornamental fishes in aquaculture. Journal of Aquaculture Research & Development, 13(9):??, ???? 2022. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/significance-of-ornamental-fishes-in-aquaculture-
114430.html.

Tikue:2022:IPE

[631] Teklay Gebru Tikue, Kassaye Balkew Workagegn, P. Natarajan, and Belayneh Daniel. Identification of phytoplankton and evaluation of algal biomass in poultry manure fertilized concrete pond in Center for Aquaculture Research and Education, Hawassa University. Journal of Aquaculture Research & Development, 13(9):??, ???? 2022. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
identification-of-phytoplankton-and-evaluation-of-algal-biomass-
in-poultry-manure-fertilized-concrete-pond-in-center-for-aquacultu-
114431.html.

Moore:2022:BOO

[632] Joseph Moore. Biological oceanography: Overview, perspective and characteristics. Journal of Aquaculture Research & Development, 13(10):??, ???? 2022. CODEN JARDCR. ISSN 2155-9546. URL https://www.

walshmedicalmedia.com/open-access/biological-oceanography-overview-
perspective-and-characteristics-115690.html.

Yan:2022:MCF

[633] Kim Yan. Modelling culture facilities marine aquaculture waters for hydrodynamics and solute transport. Journal of Aquaculture Research & Development, 13(10):??, ???? 2022. CODEN JARDCR. ISSN 2155-9546.

URL https://www.walshmedicalmedia.com/open-access/modelling-
culture-facilities-marine-aquaculture-waters-for-hydrodynamics-
and-solute-transport-115691.html.

Dinter:2022:ECF

[634] Walter Dinter. Eddy characteristics in functional oceanographic analysis tools. Journal of Aquaculture Research & Development, 13(10): ??, ???? 2022. CODEN JARDCR. ISSN 2155-9546. URL https:/

/www.walshmedicalmedia.com/open-access/eddy-characteristics-
in-functional-oceanographic-analysis-tools-115692.html.

Yihua:2022:RAB

[635] Wang Yihua. Regional aquaculture businesses’ efficiency in China’s freshwater aquaculture. Journal of Aquaculture Research & Development, 13 (10):??, ???? 2022. CODEN JARDCR. ISSN 2155-9546. URL https:/

/www.walshmedicalmedia.com/open-access/regional-aquaculture-
businesses-efficiency-in-chinas-freshwater-aquaculture-115693.
html.

Hyman:2022:ANE

[636] Karyn Hyman. Aquaculture’s nutrient efficiency and fish nutrition. Journal of Aquaculture Research & Development, 13(10):??, ???? 2022. CODEN JARDCR. ISSN 2155-9546. URL https://

www.walshmedicalmedia.com/open-access/aquacultures-nutrient-
efficiency-and-fish-nutrition-115694.html.

Hartmann:2022:UMM

[637] Jacob Hartmann. Uses of microalgae in meat production of aquaculture. Journal of Aquaculture Research & Development, 13(11):??, ???? 2022. CODEN JARDCR. ISSN 2155-9546. URL https://

www.walshmedicalmedia.com/open-access/uses-of-microalgae-in-
meat-production-of-aquaculture-116073.html.

Barwin:2022:EDQ

[638] Diana Barwin. Ecosystem diversity and quality in freshwater. Journal of Aquaculture Research & Development, 13(11):??, ???? 2022. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/ecosystem-diversity-and-quality-in-freshwater-
116075.html.

Conroy:2022:RSE

[639] Edward Conroy. Remote sensing estimates for dissolved oxygen at aquaculture operations. Journal of Aquaculture Research & Development, 13(11):??, ???? 2022. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
remote-sensing-estimates-for-dissolved-oxygen-at-aquaculture-
operations-116079.html.

Bishop:2022:FFP

[640] Sherman Bishop. Factors on fish pathology identification in aquaculture. Journal of Aquaculture Research & Development, 13(11):??, ???? 2022. CODEN JARDCR. ISSN 2155-9546. URL https://www.

walshmedicalmedia.com/open-access/factors-on-fish-pathology-
identification-in-aquaculture-116081.html.

Elisio:2022:VAM

[641] Darcy Elisio. Various aspects of marine ecology and their significance. Journal of Aquaculture Research & Development, 13(11):??, ???? 2022. CODEN JARDCR. ISSN 2155-9546. URL https://www.

walshmedicalmedia.com/open-access/various-aspects-of-marine-
ecology-and-their-significance-116082.html.

Castillo:2022:ESF

[642] David Castillo. Extensive shrimp farming culture and operation method. Journal of Aquaculture Research & Development, 13(12):??, ???? 2022. CODEN JARDCR. ISSN 2155-9546. URL https://www.

walshmedicalmedia.com/open-access/extensive-shrimp-farming-culture-
and-operation-method-117508.html.

Bates:2022:ACR

[643] Robert Bates. Advantages of coral reefs for ecosystems and biodiversity. Journal of Aquaculture Research & Development, 13(12):??, ???? 2022. CODEN JARDCR. ISSN 2155-9546. URL https://www.

walshmedicalmedia.com/open-access/advantages-of-coral-reefs-
for-ecosystems-and-biodiversity-117509.html.

Matos:2022:EHS

[644] Dalton Matos. Enhancing habitat and the social effects of mariculture. Journal of Aquaculture Research & Development, 13(12):??, ???? 2022. CODEN JARDCR. ISSN 2155-9546. URL https://www.

walshmedicalmedia.com/open-access/enhancing-habitat-and-the-
social-effects-of-mariculture-117510.html.

Tasleem:2022:CAG

[645] Faisal Tasleem, M. Shahid, Hina Fatima, and Numan Gulzar M. Comparative analysis of genetic diversity in natural population of Labeo rohita and Cirrhinus marigala by using species specific molecular markers. Journal of Aquaculture Research & Development, 13(12):??, ???? 2022. CODEN JARDCR. ISSN 2155-9546. URL https:/

/www.walshmedicalmedia.com/open-access/comparative-analysis-
of-genetic-diversity-in-natural-population-of-labeo-rohita--
and-cirrhinus-marigala-by-using-species.pdf.

Shan:2022:MRI

[646] Shan Shan. Major role and importance of veterinary medicine in animal health. Journal of Aquaculture Research & Development, 13(12):??, ???? 2022. CODEN JARDCR. ISSN 2155-9546. URL https://www.

walshmedicalmedia.com/open-access/major-role-and-importance-
of-veterinary-medicine-in-animal-health-117274.html.

Michele:2022:GSI

[647] Cristina Michele. General study and importance of veterinary medicine: an overview. Journal of Aquaculture Research & Development, 13 (12):??, ???? 2022. CODEN JARDCR. ISSN 2155-9546. URL

https://www.walshmedicalmedia.com/open-access/general-study-
and-importance-of-veterinary-medicine-an-overview-117275.html.

Jung-Hoon:2023:PAU

[648] Jung-Hoon. Proteomics approach for understanding arsenic toxicity in a fish mode. Journal of Aquaculture Research & Development, 14(1): ??, ???? 2023. CODEN JARDCR. ISSN 2155-9546. URL https:

//www.walshmedicalmedia.com/open-access/proteomics-approach-
for-understanding-arsenic-toxicity-in-a-fish-mode-115217.html.

Macbeth:2023:EAE

[649] Peter Macbeth. Evaluation of aquaculture environmental interactions and aquaculture population growth in Chile. Journal of Aquaculture Research & Development, 14(1):??, ???? 2023. CODEN JARDCR.

ISSN 2155-9546. URL https://www.walshmedicalmedia.com/open-
access/evaluation-of-aquaculture-environmental-interactions-
and-aquaculture-population-growth-in-chile-114966.html.

Tasleem:2023:CAG

[650] Faisal Tasleem, M. Shahid, Hina Fatima, and Numan Gulzar M. Comparative analysis of genetic diversity in natural population of Labeo rohita and Cirrhinus marigala by using species specific molecular markers. Journal of Aquaculture Research & Development, 14(1):??, ???? 2023. CODEN JARDCR. ISSN 2155-9546. URL https:/

/www.walshmedicalmedia.com/open-access/comparative-analysis-
of-genetic-diversity-in-natural-population-of-labeo-rohita-and-
cirrhinus-marigala-by-using-species-specific-m-117273.html.

Brossard:2023:AFM

[651] Gerard Brossard. Aquaculture farm management using surrogate models utilizing the Internet of Things and AI. Journal of Aquaculture Research & Development, 14(1):??, ???? 2023. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/aquaculture-farm-management-using-surrogate-
models-utilizing-the-internet-of-things-and-ai-118147.html.

Eugeny:2023:ICC

[652] James Eugeny. Impact of cryopreservation, cryoprotectants for ovarian tissue freezing and advances in cryobiology. Journal of Aquaculture Research & Development, 14(1):??, ???? 2023. CODEN JARDCR. ISSN 2155-9546.

URL https://www.walshmedicalmedia.com/open-access/impact-of-
cryopreservation-cryoprotectants-for-ovarian-tissue-freezing-
and-advances-in-cryobiology-117754.html.

Peretti:2023:OMT

[653] Corrado Peretti. Oceanographic model for tidal stream turbines in three dimensions. Journal of Aquaculture Research & Development, 14(2): ??, ???? 2023. CODEN JARDCR. ISSN 2155-9546. URL https:

//www.walshmedicalmedia.com/open-access/oceanographic-model-
for-tidal-stream-turbines-in-three-dimensions-118390.html.

Rongguo:2023:FRA

[654] Wang Rongguo. Floating raft aquaculture images from hyperspectral remote sensing technique. Journal of Aquaculture Research & Development, 14(2):??, ???? 2023. CODEN JARDCR. ISSN 2155-9546. URL

https://www.walshmedicalmedia.com/open-access/floating-raft-
aquaculture-images-from-hyperspectral-remote-sensing-technique-
118391.html.

Griese:2023:TSM

[655] Michael Griese. Tilapia skin microbiomes in Malawi’s aquaculture interact with pond water. Journal of Aquaculture Research & Development, 14(2):??, ???? 2023. CODEN JARDCR. ISSN 2155-9546.

URL https://www.walshmedicalmedia.com/open-access/tilapia-
skin-microbiomes-in-malawis-aquaculture-interact-with-pond-water-
118392.html.

Gabeira:2023:SSW

[656] Helena Gabeira. Study of surface water quality and microsystems limnology. Journal of Aquaculture Research & Development, 14(2):??, ???? 2023. CODEN JARDCR. ISSN 2155-9546. URL https://www.

walshmedicalmedia.com/open-access/study-of-surface-water-quality-
and-microsystems-limnology-118394.html.

Barthez:2023:ERO

[657] Audrey Barthez. Effects of river and ocean particle matter on the Magdalena River Estuary. Journal of Aquaculture Research & Development, 14(2):??, ???? 2023. CODEN JARDCR. ISSN 2155-9546. URL https:/

/www.walshmedicalmedia.com/open-access/effects-of-river-and-
ocean-particle-matter-on-the-magdalena-river-estuary-118395.
html.

Xue:2023:CLS

[658] Liu Xue. Construction of a large-scale subsurface flow wetland for the functional detection of microbes in marine aquaculture water. Journal of Aquaculture Research & Development, 14(3):??, ???? 2023. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/construction-of-a-largescale-subsurface-flow-
wetland-for-the-functional-detection-of-microbes-in-marine-aquaculture-
water-119655.html.

Li:2023:DCC

[659] Kim Li. Development of common carp donor progeny using goldfish as surrogate spawn stock. Journal of Aquaculture Research & Development, 14(3):??, ???? 2023. CODEN JARDCR. ISSN 2155-9546. URL https://

www.walshmedicalmedia.com/open-access/development-of-common-
carp-donor-progeny-using-goldfish-as-surrogate-spawn-stock-119656.
html.

Suassuna:2023:SUL

[660] Pedro Suassuna. Sustainable utilization of lactic acid bacteria in aquaculture. Journal of Aquaculture Research & Development, 14(3):??, ???? 2023. CODEN JARDCR. ISSN 2155-9546. URL https://www.

walshmedicalmedia.com/open-access/sustainable-utilization-of-
lactic-acid-bacteria-in-aquaculture-119657.html.

Perez:2023:ICF

[661] Jimmy Perez. Intense chemical flocculation and coagulation of aquaculture wastewater. Journal of Aquaculture Research & Development, 14(3):??, ???? 2023. CODEN JARDCR. ISSN 2155-9546.

URL https://www.walshmedicalmedia.com/open-access/intense-
chemical-flocculation-and-coagulation-of-aquaculture-wastewater-
119658.html.

Aseev:2023:IPV

[662] Georgy Aseev. Impact of palaemon varians dietary strain on aquaculture development. Journal of Aquaculture Research & Development, 14(3):??, ???? 2023. CODEN JARDCR. ISSN 2155-9546.

URL https://www.walshmedicalmedia.com/open-access/impact-of-
palaemon-varians-dietary-strain-on-aquaculture-development-119659.
html.

Shu:2023:ECE

[663] Katuo Shu. Exploring the causes and effects of common fish diseases. Journal of Aquaculture Research & Development, 14(3):??, ???? 2023. CODEN JARDCR. ISSN 2155-9546. URL https:/

/www.walshmedicalmedia.com/open-access/exploring-the-causes-
and-effects-of-common-fish-diseases-119660.html.

Omi:2023:ITM

[664] Sako Omi. Impact of technology in modern veterinary medicine. Journal of Aquaculture Research & Development, 14(3):??, ???? 2023. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/impact-of-technology-in-modern-veterinary-medicine-
119661.html.

Virkki:2023:IAF

[665] Troell Virkki. The impact of aquaculture feeds on global food security. Journal of Aquaculture Research & Development, 14(3):??, ???? 2023. CODEN JARDCR. ISSN 2155-9546. URL https://www.

walshmedicalmedia.com/open-access/the-impact-of-aquaculture-
feeds-on-global-food-security-119662.html.

Mathe:2023:ABN

[666] Bernado Mathe. Aspects of blue nitrogen to manage nutrient neutrality. Journal of Aquaculture Research & Development, 14(3):??, ????

2023. CODEN JARDCR. ISSN 2155-9546. URL https://www.

walshmedicalmedia.com/open-access/aspects-of-blue-nitrogen-to-
manage-nutrient-neutrality-119664.html.

Luke:2023:IBC

[667] Maria Luke. Importance of blue carbon mangroves in climatic change. Journal of Aquaculture Research & Development, 14(3):??, ???? 2023. CODEN JARDCR. ISSN 2155-9546. URL https://www.

walshmedicalmedia.com/open-access/importance-of-blue-carbon-
mangroves-in-climatic-change-119665.html.

Tamburini:2023:DOF

[668] Elena Tamburini and Edoardo Turolla. The development of oyster farming in Italy: an innovation opportunity for mollusks farming diversification. Journal of Aquaculture Research & Development, 14(4):??, ???? 2023. CODEN JARDCR. ISSN 2155-9546. URL https://www.

walshmedicalmedia.com/open-access/the-development-of-oyster-
farming-in-italy-an-innovation-opportunity-for-mollusks-farming-
diversification-120320.html.

Wergeland:2023:CST

[669] Furuya Wergeland. Conservation strategies for threatened coral ecosystems. Journal of Aquaculture Research & Development, 14(4):??, ???? 2023. CODEN JARDCR. ISSN 2155-9546. URL https://www.

walshmedicalmedia.com/open-access/conservation-strategies-for-
threatened-coral-ecosystems-120181.html.

Zang:2023:AMM

[670] Dang Zang. Advancements in marine microbial culture techniques. Journal of Aquaculture Research & Development, 14(4):??, ???? 2023. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/advancements-in-marine-microbial-culture-techniques-
120182.html.

Ramil:2023:BEG

[671] Samuel Ramil. Balancing economic growth and environmental sustainability in shrimp aquaculture. Journal of Aquaculture Research & Development, 14(4):??, ???? 2023. CODEN JARDCR. ISSN 2155-9546.

URL https://www.walshmedicalmedia.com/open-access/balancing-
economic-growth-and-environmental-sustainability-in-shrimp-aquaculture-
120183.html.

Leng:2023:EVA

[672] Qiu Leng. Exploring the versatile applications of cultivated microalgae in food, fuel, and pharmaceuticals. Journal of Aquaculture Research & Development, 14(4):??, ???? 2023. CODEN JARDCR. ISSN 2155-9546.

URL https://www.walshmedicalmedia.com/open-access/exploring-
the-versatile-applications-of-cultivated-microalgae-in-food-
fuel-and-pharmaceuticals-120184.html.

Fenice:2023:GIN

[673] Yu Fenice. Genetic improvement and nutrition for high-yield aquaculture systems. Journal of Aquaculture Research & Development, 14(4): ??, ???? 2023. CODEN JARDCR. ISSN 2155-9546. URL https:

//www.walshmedicalmedia.com/open-access/genetic-improvement-
and-nutrition-for-highyield-aquaculture-systems-120185.html.

Shan:2023:EIA

[674] Zhang Shan. Exploring the importance and advancements of oceanography. Journal of Aquaculture Research & Development, 14(4): ??, ???? 2023. CODEN JARDCR. ISSN 2155-9546. URL

https://www.walshmedicalmedia.com/open-access/exploring-the-
importance-and-advancements-of-oceanography-120188.html.

Dou:2023:IHE

[675] Domingues Dou. Investigating the health of ecosystems through eco toxicological research. Journal of Aquaculture Research & Development, 14(4):??, ???? 2023. CODEN JARDCR. ISSN 2155-9546. URL

https://www.walshmedicalmedia.com/open-access/investigating-
the-health-of-ecosystems-through-eco-toxicological-research-
120187.html.

Shuichi:2023:IFI

[676] Malherbe Shuichi. Innovative feed and immunological biotechnology for sustainable aquaculture. Journal of Aquaculture Research & Development, 14(4):??, ???? 2023. CODEN JARDCR. ISSN 2155-9546. URL https:

//www.walshmedicalmedia.com/open-access/innovative-feed-and-
immunological-biotechnology-for-sustainable-aquaculture-120186.
html.

Ergena:2023:ZOH

[677] Alazar Ergena, P Natarajan, and Zufan Bedewi. Effect of ginger (Zingiber officinale) on hematological parameters and resistant to Aeromonas hydrophila infection in Nile tilapia, Oreochromis niloticus L. Journal of Aquaculture Research & Development, 14(4):??, ???? 2023. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/effect-of-ginger-emzingiber-officinaleem-on-
hematological-parameters-and-resistant-to-emaeromonas-hydrophilaem-
infection-in-nile-t-119171.html.

Yan:2023:ESC

[678] Duan Yan. Effective strategies for coastal risk management. Journal of Aquaculture Research & Development, 14(5):??, ???? 2023. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/effective-strategies-for-coastal-risk-management-
120526.html.

Li:2023:CCB

[679] Barboza Li. Cryobiology’s contribution to breeding programs in aquaculture. Journal of Aquaculture Research & Development, 14(5):??, ???? 2023. CODEN JARDCR. ISSN 2155-9546. URL https://www.

walshmedicalmedia.com/open-access/cryobiologys-contribution-
to-breeding-programs-in-aquaculture-120530.html.

Wang:2023:EET

[680] Colson Wang. Enhancing ecosystems through salmon aquaculture. Journal of Aquaculture Research & Development, 14(5):??, ???? 2023. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/enhancing-ecosystems-through-salmon-aquaculture-
120531.html.

Shin:2023:EBS

[681] Duncan Shin. Ecological benefits of seabass aquaculture. Journal of Aquaculture Research & Development, 14(5):??, ???? 2023. CODEN JARDCR.

ISSN 2155-9546. URL https://www.walshmedicalmedia.com/open-
access/ecological-benefits-of-seabass-aquaculture-120532.html.

Li:2023:EPF

[682] Zin Li. Exploring the physiological functions of aquatic animals. Journal of Aquaculture Research & Development, 14(5):??, ???? 2023. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/exploring-the-physiological-functions-of-aquatic-
animals-120534.html.

Liu:2023:TEA

[683] Nguyen Liu. The transformative effects of aquatic restoration. Journal of Aquaculture Research & Development, 14(5):??, ???? 2023. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/the-transformative-effects-of-aquatic-restoration-
120535.html.

Markuszewski:2023:AOS

[684] Piotr Markuszewski. Applications of oceanology in seafloor mapping and marine geology. Journal of Aquaculture Research & Development, 14(5): ??, ???? 2023. CODEN JARDCR. ISSN 2155-9546. URL https://www.

walshmedicalmedia.com/open-access/applications-of-oceanology-
in-seafloor-mapping-and-marine-geology-120537.html.

Sagalevich:2023:CCF

[685] Anatoly Sagalevich. Conservation challenges in fish egg development. Journal of Aquaculture Research & Development, 14(5):??, ???? 2023. CODEN JARDCR. ISSN 2155-9546. URL https://www.

walshmedicalmedia.com/open-access/conservation-challenges-in-
fish-egg-development-120539.html.

Jiao:2023:AFT

[686] Mengyu Jiao. Argo floats and their impact on Baltic Sea research. Journal of Aquaculture Research & Development, 14(5):??, ???? 2023. CODEN JARDCR. ISSN 2155-9546. URL https://

www.walshmedicalmedia.com/open-access/argo-floats-and-their-
impact-on-baltic-sea-research-120541.html.

Chen:2023:STV

[687] Shuyin Chen. Spatial and temporal variability of plankton stocks in the Pacific Ocean. Journal of Aquaculture Research & Development, 14(5): ??, ???? 2023. CODEN JARDCR. ISSN 2155-9546. URL https:/

/www.walshmedicalmedia.com/open-access/spatial-and-temporal-
variability-of-plankton-stocks-in-the-pacific-ocean-120542.html.

Ontario:2023:PAE

[688] Antony Ontario. Probiotics in aquaculture: Enhancing health and sustainability. Journal of Aquaculture Research & Development, 14(6):??, ???? 2023. CODEN JARDCR. ISSN 2155-9546. URL https://www.

walshmedicalmedia.com/open-access/probiotics-in-aquaculture-
enhancing-health-and-sustainability-121140.html.

Hong:2023:IIA

[689] Chen Hong. Influence of intensive aquaculture on antibiotic resistance. Journal of Aquaculture Research & Development, 14(6):??, ????

2023. CODEN JARDCR. ISSN 2155-9546. URL https://www.

walshmedicalmedia.com/open-access/influence-of-intensive-aquaculture-
on-antibiotic-resistance-121148.html.

Hong:2023:EIM

[690] Koo Hong. Exploring the intersection of marine science and biotechnology. Journal of Aquaculture Research & Development, 14(6): ??, ???? 2023. CODEN JARDCR. ISSN 2155-9546. URL

https://www.walshmedicalmedia.com/open-access/exploring-the-
intersection-of-marine-science-and-biotechnology-121149.html.

Coasta:2023:ESS

[691] Ionnis Coasta. Enhancing sustainability in shrimp farming. Journal of Aquaculture Research & Development, 14(6):??, ???? 2023. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/enhancing-sustainability-in-shrimp-farming-121150.
html.

Alanara:2023:ARP

[692] Cripps Alanara. Antibiotic resistance prevention in the shrimp aquaculture. Journal of Aquaculture Research & Development, 14(6):??, ???? 2023. CODEN JARDCR. ISSN 2155-9546. URL https://

www.walshmedicalmedia.com/open-access/antibiotic-resistance-
prevention-in-the-shrimp-aquaculture-121151.html.

Hulata:2023:FSL

[693] Kinkar Hulata. Flexibility in season length restrictions for effective fisheries management. Journal of Aquaculture Research & Development, 14 (6):??, ???? 2023. CODEN JARDCR. ISSN 2155-9546. URL https://

www.walshmedicalmedia.com/open-access/flexibility-in-season-
length-restrictions-for-effective-fisheries-management-121152.
html.

Stouart:2023:SAB

[694] Nakoll Stouart. Sustainability in aquaculture: Biochar-enhanced water treatment for fish farming. Journal of Aquaculture Research & Development, 14(6):??, ???? 2023. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
sustainability-in-aquaculture-biocharenhanced-water-treatment-
for-fish-farming-121153.html.

Troung:2023:EBE

[695] Arnold Troung. Examining biological and economic factors in the management of marine resources. Journal of Aquaculture Research & Development, 14(6):??, ???? 2023. CODEN JARDCR. ISSN 2155-9546.

URL https://www.walshmedicalmedia.com/open-access/examining-
biological-and-economic-factors-in-the-management-of-marine-
resources-121154.html.

Win:2023:FDE

[696] Jinu Win. Function of dicer in the environmental stress response of aquatic species. Journal of Aquaculture Research & Development, 14(6): ??, ???? 2023. CODEN JARDCR. ISSN 2155-9546. URL https:/

/www.walshmedicalmedia.com/open-access/function-of-dicer-in-
the-environmental-stress-response-of-aquatic-species-121155.
html.

Findlay:2023:APD

[697] Jolly Findlay. Analysing the proximity and diversity of marine organisms. Journal of Aquaculture Research & Development, 14(6): ??, ???? 2023. CODEN JARDCR. ISSN 2155-9546. URL

https://www.walshmedicalmedia.com/open-access/analysing-the-
proximity-and-diversity-of-marine-organisms-121156.html.

Mahatara:2023:SRT

[698] Mina Mahatara, Ram Bhajan Mandal, Jay Dev Bista, and Sujan Mishra. Sex reversal of tilapia (Oreochromis niloticus) fry by using methyltestosterone (Mt) treated soybean meal. Journal of Aquaculture Research & Development, 14(7):??, ???? 2023. CODEN JARDCR. ISSN 2155-

9546. URL https://www.walshmedicalmedia.com/open-access/sex-
reversal-of-tilapia-ioreochromis-niloticusi-fry-by-using-methyltestosterone-
mt-treated-soybean-meal-121737.html.

Salele:2023:IRF

[699] Hauwau A. Salele, Nafisa Abdurrasheed, and Akeem Babatunde Dauda. Integrated rice-cum-fish farming in Nigeria: Prospects, status and challenges. Journal of Aquaculture Research & Development, 14(7):??, ???? 2023. CODEN JARDCR. ISSN 2155-9546. URL https://www.

walshmedicalmedia.com/open-access/integrated-ricecumfish-farming-
in-nigeria-prospects-status-and-challenges-121738.html.

Gwladys:2023:SGR

[700] Tsoupou Kuété Suzy Gwladys, Kpoumie Nsangou Amidou, Tonfack Achile Peguy, Komguep Nganyo Ronald, Tchouante Tzukam Christelle G, Vemo Bertin Narcisse, Tasse Taboue Géraud C, Ngueguim Jules Romain, and Efole Ewoukem Thomas. Survival, growth response, chemical and biochemical characteristics of the carcass of Clarias jaensis (Boulenger, 1909) post fingerlings fed various dietary energy. Journal of Aquaculture Research & Development, 14(7):??, ???? 2023. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/survival-growth-response-chemical-and-biochemical-
characteristics-of-the-carcass-of-iclarias-jaensisi-boulenger-
1909-post-fingerli-121739.html.

Xia:2023:VRA

[701] Shaohong Xia. Vital role of aquaculture in sustainable seafood production. Journal of Aquaculture Research & Development, 14(7):??, ???? 2023. CODEN JARDCR. ISSN 2155-9546. URL https://www.

walshmedicalmedia.com/open-access/vital-role-of-aquaculture-
in-sustainable-seafood-production-121740.html.

Shung:2023:CCF

[702] Xan Shung. Coastal communities and fisheries sustainability: The vital connection with fishers. Journal of Aquaculture Research & Development, 14(7):??, ???? 2023. CODEN JARDCR. ISSN 2155-9546. URL https:

//www.walshmedicalmedia.com/open-access/coastal-communities-
and-fisheries-sustainability-the-vital-connection-with-fishers-
121741.html.

Jim:2023:ERT

[703] Seoul Jim. Enhancing research through catch and release angling. Journal of Aquaculture Research & Development, 14(7):??, ???? 2023. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/enhancing-research-through-catch-and-release-
angling-121742.html.

Heupel:2023:RSO

[704] Wei Heupel. The role of seasonal oceanographic simulations in preserving marine biodiversity. Journal of Aquaculture Research & Development, 14(7):??, ???? 2023. CODEN JARDCR. ISSN 2155-9546. URL https:/

/www.walshmedicalmedia.com/open-access/the-role-of-seasonal-
oceanographic-simulations-in-preserving-marine-biodiversity-
121743.html.

Jim:2023:EDH

[705] Sanh Jim. Exploring deepwater hyperpycnites and oceanic processes. Journal of Aquaculture Research & Development, 14(7):??, ???? 2023. CODEN JARDCR. ISSN 2155-9546. URL https:

//www.walshmedicalmedia.com/open-access/exploring-deepwater-
hyperpycnites-and-oceanic-processes-121744.html.

Devanathan:2023:EOF

[706] K. Devanathan and P. Soundarapandian. Experimental organic farming of Litopenaeus vannamei (Boon, 1931) near Vellar Estuary in Cuddalore District, Tamil Nadu. Journal of Aquaculture Research & Development, 14(7):??, ???? 2023. CODEN JARDCR. ISSN 2155-9546. URL https:/

/www.walshmedicalmedia.com/open-access/experimental-organic-
farming-of-emlitopenaeus-vannameiem-boon-1931-near-vellar-estuary-
in-cuddalore-district-tamil-nadu.pdf.

Li:2023:IMPa

[707] Grant Li. Impact of micro plastics on juvenile fishes based upon their species. Journal of Aquaculture Research & Development, 14(7):??, ???? 2023. CODEN JARDCR. ISSN 2155-9546. URL https://www.

walshmedicalmedia.com/open-access/impact-of-micro-plastics-on-
juvenile-fishes-based-upon-their-species.pdf.

Antoniadou:2023:PFA

[708] Eleni Antoniadou, Ioannis T. Karapanagiotidis, Panagiota Panagiotaki, and Eleni Golomazou. Phytogenic feed additive supplemented diets as welfare promoters under acute and chronic stress factors in gilthead seabream. Journal of Aquaculture Research & Development, 14(7):??, ???? 2023. CODEN JARDCR. ISSN

2155-9546. URL https://www.walshmedicalmedia.com/open-access/
phytogenic-feed-additive-supplemented-diets-as-welfare-promoters-
under-acute-and-chronic-stress-factors-in-gilthead-seabream-
121295.html.

Muhammad:2023:EIA

[709] Ukasha Muhammad, Abubakar J. Yaji, Sumayya Bashir Yahya, and I. Yunusa. Effect of imidacloprid in aquaculture: a review. Journal of Aquaculture Research & Development, 14(7):??, ???? 2023. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/effect-of-imidacloprid-in-aquaculture-a-review-
120593.html.

Cook:2023:ESD

[710] David Cook, Jill M. Voorhees, and Michael E. Barnes. Evaluation of stirring during artificial spawning of landlocked fall Chinook salmon. Journal of Aquaculture Research & Development, 14 (8):??, ???? 2023. CODEN JARDCR. ISSN 2155-9546. URL

https://www.walshmedicalmedia.com/open-access/evaluation-of-
stirring-during-artificial-spawning-of-landlocked-fall-chinook-
salmon-122853.html.

Fladamar:2023:RUV

[711] Dogen Fladamar. Role of UAV vision revolution in transforming marine cage culture. Journal of Aquaculture Research & Development, 14(8):??, ???? 2023. CODEN JARDCR. ISSN 2155-9546.

URL https://www.walshmedicalmedia.com/open-access/role-of-
uav-vision-revolution-in-transforming-marine-cage-culture-122854.
html.

Van:2023:IWM

[712] Joe Van. Innovations in water management for a thriving marine culture industry. Journal of Aquaculture Research & Development, 14(8): ??, ???? 2023. CODEN JARDCR. ISSN 2155-9546. URL https:/

/www.walshmedicalmedia.com/open-access/innovations-in-water-
management-for-a-thriving-marine-culture-industry-122855.html.

Li:2023:WQE

[713] Fang Li. Water quality and environmental responsibility in shrimp farming. Journal of Aquaculture Research & Development, 14(8): ??, ???? 2023. CODEN JARDCR. ISSN 2155-9546. URL

https://www.walshmedicalmedia.com/open-access/water-quality-
and-environmental-responsibility-in-shrimp-farming-122856.html.

Qui:2023:RMA

[714] Xini Qui. Role of marine ammonia-oxidizing bacteria in nitrogen cycling. Journal of Aquaculture Research & Development, 14(8):??, ???? 2023. CODEN JARDCR. ISSN 2155-9546. URL https://www.

walshmedicalmedia.com/open-access/role-of-marine-ammoniaoxidizing-
bacteria-in-nitrogen-cycling-122857.html.

Oedjoe:2023:KAT

[715] Marcelien Djublina Ratoe Oedjoe, Ade Y. H. Lukas, and Kiik G Sine. The use of selected seed on the growth, production and carrageenan content (Kappaphycus alvarezii) in Tesabela, Kupang Regency, NTT Province, Indonesia. Journal of Aquaculture Research & Development, 14 (8):??, ???? 2023. CODEN JARDCR. ISSN 2155-9546. URL https:

//www.walshmedicalmedia.com/open-access/the-use-of-selected-
seed-on-the-growth-production-and-carrageenan-content-ikappaphycus-
alvareziii-in-tesabela-kupang-regency-ntt-p-123245.html.

San:2023:ISB

[716] Mercy San. Innovative solutions for bycatch mitigation in small-scale fisheries. Journal of Aquaculture Research & Development, 14(8):??, ???? 2023. CODEN JARDCR. ISSN 2155-9546. URL https:/

/www.walshmedicalmedia.com/open-access/innovative-solutions-
for-bycatch-mitigation-in-smallscale-fisheries-123249.html.

Omosuyi:2023:RPL

[717] Olawunmi Mercy Omosuyi. Dietary effects of processed African palm weevil (Rhynchophorus phoenicis) larvae with phyto-additives on growth performance and nutrient utilization of Clarias gariepinus juveniles. Journal of Aquaculture Research & Development, 14(8):??, ???? 2023. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/dietary-effects-of-processed-african-palm-weevil-
irhynchophorus-phoenicisi-larvae-with-phytoadditives-on-growth-
performance-and-nu-123246.html.

Devanathan:2023:LVB

[718] K. Devanathan and P. Soundarapandian. Experimental organic farming of Litopenaeus vannamei (Boon, 1931) near vellar estuary in Cuddalore District, Tamil Nadu. Journal of Aquaculture Research & Development, 14(8):??, ???? 2023. CODEN JARDCR. ISSN 2155-9546. URL https:/

/www.walshmedicalmedia.com/open-access/experimental-organic-
farming-of-emlitopenaeus-vannameiem-boon-1931-near-vellar-estuary-
in-cuddalore-district-tamil-nadu-122266.html.

Li:2023:IMPb

[719] Grant Li. Impact of micro plastics on juvenile fishes based upon their species. Journal of Aquaculture Research & Development, 14(8):??, ???? 2023. CODEN JARDCR. ISSN 2155-9546. URL https://www.

walshmedicalmedia.com/open-access/impact-of-micro-plastics-on-
juvenile-fishes-based-upon-their-species-122267.html.

Heng:2023:SCI

[720] Gal Heng. Sustainability challenges and innovations in Southeast Asian aquaculture. Journal of Aquaculture Research & Development, 14(9):??, ???? 2023. CODEN JARDCR. ISSN 2155-9546. URL https://www.

walshmedicalmedia.com/open-access/sustainability-challenges-
and-innovations-in-southeast-asian-aquaculture-123998.html.

Mei:2023:NCA

[721] Kned Mei. Navigating challenges in aquaculture and fisheries. Journal of Aquaculture Research & Development, 14(9):??, ???? 2023. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/navigating-challenges-in-aquaculture-and-fisheries-
123759.html.

Joe:2023:UPS

[722] Ganf Joe. Unlocking the potential of small-scale cage aquaculture for food security. Journal of Aquaculture Research & Development, 14(9):??, ???? 2023. CODEN JARDCR. ISSN 2155-9546. URL

https://www.walshmedicalmedia.com/open-access/unlocking-the-
potential-of-smallscale-cage-aquaculture-for-food-security-123760.
html.

Shan:2023:RBA

[723] Gan Shan. Role of biochar applications in sustainable aquaculture in improving water quality. Journal of Aquaculture Research & Development, 14(9):??, ???? 2023. CODEN JARDCR. ISSN 2155-9546.

URL https://www.walshmedicalmedia.com/open-access/role-of-
biochar-applications-in-sustainable-aquaculture-in-improving-
water-quality-123761.html.

Joe:2023:OAD

[724] Nac Joe. Optimizing aquaculture diet formulation through genetic selection. Journal of Aquaculture Research & Development, 14(9):??, ???? 2023. CODEN JARDCR. ISSN 2155-9546. URL https://www.

walshmedicalmedia.com/open-access/optimizing-aquaculture-diet-
formulation-through-genetic-selection-123762.html.

Gendo:2023:SAE

[725] Rladama Gendo. Shrimp aquaculture’s economic resilience through antiviral agents. Journal of Aquaculture Research & Development, 14(9): ??, ???? 2023. CODEN JARDCR. ISSN 2155-9546. URL https:

//www.walshmedicalmedia.com/open-access/shrimp-aquacultures-
economic-resilience-through-antiviral-agents-123763.html.

Saas:2023:EBC

[726] Karia Saas. The environmental benefits of closed-loop aquaculture systems. Journal of Aquaculture Research & Development, 14(9):??, ???? 2023. CODEN JARDCR. ISSN 2155-9546. URL https://www.

walshmedicalmedia.com/open-access/the-environmental-benefits-
of-closedloop-aquaculture-systems-123764.html.

Ngueguim:2023:EFA

[727] Derrick Fabrice Ngueguim, Marc Kenmogne Kouam, Georges Fonkwa, Hermann Bridget Katte, Jacques Nack, and Julius Awah-Ndukum. Endogenous factors associated to anatomo-pathological lesions of cultured fish species in the West Region of Cameroon. Journal of Aquaculture Research & Development, 14(10):??, ???? 2023. CODEN JARDCR.

ISSN 2155-9546. URL https://www.walshmedicalmedia.com/open-
access/endogenous-factors-associated-to-anatomopathological-
lesions-of-cultured-fish-species-in-the-west-region-of-cameroon-
124249.html.

Giongu:2023:PPA

[728] Si Giongu. The rising role of Pichia pastoris in aquaculture enhancement. Journal of Aquaculture Research & Development, 14(10):??, ???? 2023. CODEN JARDCR. ISSN 2155-9546. URL https://www.

walshmedicalmedia.com/open-access/the-rising-role-of-ipichia-
pastorisi-in-aquaculture-enhancement-124251.html.

Sang:2023:RQG

[729] Kleen Sang. Managing Rhamdia quelen growth and reproduction for sustainable fisheries and aquaculture. Journal of Aquaculture Research & Development, 14(10):??, ???? 2023. CODEN JARDCR. ISSN 2155-9546.

URL https://www.walshmedicalmedia.com/open-access/managing-
irhamdia-queleni-growth-and-reproduction-for-sustainable-fisheries-
and-aquaculture-124252.html.

Neil:2023:EPM

[730] Jack Neil. The evolution of predictive models in aquaculture. Journal of Aquaculture Research & Development, 14(10):??, ???? 2023. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/the-evolution-of-predictive-models-in-aquaculture-
124254.html.

Natrah:2023:FCR

[731] Zakaria Natrah. Factors contributing to resistance in shrimp aquaculture. Journal of Aquaculture Research & Development, 14(10):??, ???? 2023. CODEN JARDCR. ISSN 2155-9546. URL https:/

/www.walshmedicalmedia.com/open-access/factors-contributing-
to-resistance-in-shrimp-aquaculture-124255.html.

Hashimoto:2023:BNM

[732] Valeria Hashimoto. Role of conservation of Bigelowiella natans in marine biodiversity. Journal of Aquaculture Research & Development, 14(10): ??, ???? 2023. CODEN JARDCR. ISSN 2155-9546. URL https:/

/www.walshmedicalmedia.com/open-access/role-of-conservation-
of-ibigelowiella-natansi-in-marine-biodiversity-124256.html.

Critser:2023:CIH

[733] Boyle Critser. Cryobiology and its impact on human reproduction. Journal of Aquaculture Research & Development, 14(11):??, ???? 2023. CODEN JARDCR. ISSN 2155-9546. URL https:

//www.walshmedicalmedia.com/open-access/cryobiology-and-its-
impact-on-human-reproduction-124551.html.

Lin:2023:ITP

[734] Jack Lin. Innovative technologies in Pacific fisheries management. Journal of Aquaculture Research & Development, 14(11):??, ???? 2023. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/innovative-technologies-in-pacific-fisheries-
management-124595.html.

Kim:2023:IMP

[735] Groell Kim. Impact of marine pearl culture on coastal ecosystems. Journal of Aquaculture Research & Development, 14(11):??, ???? 2023. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/impact-of-marine-pearl-culture-on-coastal-ecosystems-
124596.html.

Choi:2023:RPM

[736] Zou Choi. Role of phytoplankton in mitigating climate change. Journal of Aquaculture Research & Development, 14(11):??, ???? 2023. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/role-of-phytoplankton-in-mitigating-climate-
change-124599.html.

Sia:2023:ESS

[737] Xhanhong Sia. The economic significance of small-scale fisheries. Journal of Aquaculture Research & Development, 14(11):??, ???? 2023. CODEN JARDCR. ISSN 2155-9546. URL https://www.

walshmedicalmedia.com/open-access/the-economic-significance-
of-smallscale-fisheries-124601.html.

Dao:2023:ERW

[738] Shanbai Dao. Exploring the remarkable world of popular freshwater, coastal, and marine fishes. Journal of Aquaculture Research & Development, 14(11):??, ???? 2023. CODEN JARDCR.

ISSN 2155-9546. URL https://www.walshmedicalmedia.com/open-
access/exploring-the-remarkable-world-of-popular-freshwater-
coastal-and-marine-fishes-124613.html.

Xue:2023:PCR

[739] Kin Xue. Protecting coral reefs and their compartmental biodiversity. Journal of Aquaculture Research & Development, 14(11):??, ???? 2023. CODEN JARDCR. ISSN 2155-9546. URL https://www.

walshmedicalmedia.com/open-access/protecting-coral-reefs-and-
their-compartmental-biodiversity-124616.html.

Miao:2023:MDA

[740] Kengyu Miao. Microbial diversity in aquatic environments. Journal of Aquaculture Research & Development, 14(11):??, ???? 2023. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/microbial-diversity-in-aquatic-environments-
124618.html.

Sang:2023:EDS

[741] Kwan Sang. Exploring the diversity and significance of marine shellfish. Journal of Aquaculture Research & Development, 14(11): ??, ???? 2023. CODEN JARDCR. ISSN 2155-9546. URL

https://www.walshmedicalmedia.com/open-access/exploring-the-
diversity-and-significance-of-marine-shellfish-124620.html.

Liu:2023:ECM

[742] Chang Liu. The economic consequences of marine waste pollution. Journal of Aquaculture Research & Development, 14(11):??, ???? 2023. CODEN JARDCR. ISSN 2155-9546. URL https://www.

walshmedicalmedia.com/open-access/the-economic-consequences-
of-marine-waste-pollution-124621.html.

Rang:2023:IHA

[743] Knach Rang. Impact of human activities marine water quality. Journal of Aquaculture Research & Development, 14(11):??, ???? 2023. CODEN

JARDCR. ISSN 2155-9546. URL https://www.walshmedicalmedia.
com/open-access/impact-of-human-activities-marine-water-quality-
124625.html.